A Bibliography of Publications of Jack J. Dongarra

Jack J. Dongarra

Computer Science Department PO Box 2008, Building 6012
University of Tennessee  Mathematical Science Section
Knoxville, TN 37996-1301  Oak Ridge National Laboratory
USA  Oak Ridge, TN 37821-6367

USA

      E-mail: dongarra@cs.utk.edu, dongarra@msr.epm.ornl.gov
 WWW URL: http://www.netlib.org/utk/people/JackDongarra.html

Nelson H. F. Beebe  Stefano Foresti
University of Utah  Utah Supercomputing Institute
Department of Mathematics, 110 LCB  University of Utah
Salt Lake City, UT 84112-0090  Salt Lake City, UT 84112
USA  USA
Tel: +1 801 581 5254  Tel: +1 801 581 3173

FAX: +1 801 585 5366

E-mail: beebe@math.utah.edu  E-mail: stefano@chpc.utah.edu

beebe@acm.org
beebe@computer.org
WWW URL: https://www.math.utah.edu/~beebe/

14 August 2025 Version 1.296

Abstract

This bibliography records publications of Jack J. Dongarra.

Title word cross-reference

[598]. 3 [828]. H [904]. ILU [876]. LU [254, 329, 382, 487, 225, 300, 343, 391, 395, 396, 433, 492, 523, 751, 822, 785, 810, 838, 886, 788, 790, 814]. QR [722, 723, 745, 146, 180, 797, 656, 676, 254, 329, 382, 487, 49, 391, 392, 754, 759, 785, 809, 810, 838, 709, 408, 559, 690, 710, 734, 771, 413, 847, 859].

-590 [266]. -Factorization [49]. -matrix [904].

’08 [1124].

1 [3, 443]. 1.0 [306, 346]. 1.1 [415]. 1.2 [424, 305, 342]. 10 [493]. 10th [1070]. 11 [1057]. 11th [1057, 1093]. ’12 [1137]. 12th [1097]. 13th [1110, 1142]. 14th [1113, 975, 981, 1052, 1145, 703]. 15th [1126, 1024]. 16th [1059, 1130]. 17th [1072, 1076, 1132]. 18-21 [658]. 18th [1135, 1091]. 1986 [961]. 1988 [967]. 1989 [117, 969]. 19th [1138, 1098]. 1st [992, 959].

2 [16, 592, 437]. 2-5 [999]. 20 [41, 51]. 200 [41, 51]. 2000 [1038, 1051]. 2001 [1042, 1052, 1067]. 2002 [1079, 1064, 1065, 1066]. 2003 [1069, 1073, 1075, 1080, 1081, 1082, 1083]. 2004 [1085, 1086, 1087, 1088, 643, 1092]. 2005 [628, 658]. 2006 [667]. 2007 [703]. 2008 [1124]. 2011 [1136]. 2012 [1137]. 2019 [912]. 2020 [909, 921]. 20th [1141]. 21st [1015]. 29 [934]. 2nd [993, 1056].

3 [654, 80, 691]. 37th [1094]. 3rd [1020, 1073, 1075].

4 [41, 44, 51, 55]. 4th [1017, 1085, 1086, 1087, 1088, 1109, 1051, 1067].

500 [511]. 589 [8]. 590 [266]. 5th [1028, 996, 1011, 1079, 1100, 1101, 1102, 1095].

6000 [135, 162]. 6000-550 [266]. 64-bit [641, 678]. 656 [95]. 679 [131]. 680 [791]. 6th [1105, 1106, 1107, 1096, 1033, 991, 1112].

710 [197]. 7th [1039, 1108, 952, 1117, 1118, 1119, 1120, 1127].

810/20 [41, 51]. 860 [137]. ’88 [960]. 8th [1121, 1122, 1123, 1047, 1116, 1061, 1133, 1134].

’90 [970, 319, 308, 344, 345, 350, 446]. ’91 [973, 977]. ’92 [984]. ’93 [989]. ’94 [994]. ’95 [1009, 407, 471]. 95/NT [407, 471]. ’96 [1006, 1007, 1019, 1010]. ’97 [1022]. ’98 [1035, 1036]. 9th [1128, 1129, 1077, 1060, 1139, 1140].

abstracts [1094]. accelerate [867, 915, 670]. Accelerated [748, 824, 696, 793, 848, 933, 852, 841, 916, 744]. Accelerating [922, 699, 749, 799, 880, 823, 884, 856, 888, 927, 719, 741, 184]. Acceleration [830, 940]. accelerator [936]. Accelerators [722, 781, 814, 717, 743, 934, 839, 735, 742]. Access [479, 90, 399, 533, 91, 325, 376, 461, 916, 565]. ACCT [495]. Accuracy [832, 641, 678, 6, 7, 8, 11, 12, 13, 751, 648, 822, 649]. Accurate [262, 611, 539, 623, 936]. Achieving [641, 678, 751, 822, 886]. ACM [1014, 1104, 1124, 1056, 1073, 1075, 1143, 1005]. ACM/IEEE [1014, 1104, 1005, 1027]. Action [707]. Active [607, 1058, 1050, 564, 564]. ACTS [643]. Adaptability [604]. Adaptable [431]. Adaptation [635]. Adapting [577, 609, 588, 587, 631, 567, 538, 571, 578, 629, 552, 642]. Adaptive [892, 460, 562]. adaptivity [639]. Addressing [60]. Administration [596]. Advanced [26, 69, 196, 261, 1001, 1021, 990, 1041, 1053, 101, 971, 70, 113]. Advances [1113, 1135, 1039, 1132, 1060, 597, 634, 1126, 1110, 1130, 1138, 1028, 1017, 1047, 1097, 1033, 1070, 1093]. Affecting [268]. agent [425]. agent-based [425]. Aggregated [762, 763]. Ahead [646]. AI [1143, 896, 938]. Aid [90, 143, 123, 91, 109, 119, 213]. Aims [301]. Alamos [948, 1062]. Albuquerque [977]. Alexandria [992]. Algebra [917, 721, 106, 127, 128, 129, 183, 674, 279, 280, 108, 322, 508, 544, 545, 776, 653, 804, 16, 577, 192, 193, 259, 294, 297, 298, 299, 331, 547, 68, 628, 15, 18, 22, 25, 28, 29, 44, 46, 55, 57, 58, 75, 95, 96, 97, 104, 131, 135, 141, 162, 163, 175, 200, 203, 226, 232, 238, 268, 306, 310, 339, 346, 432, 434, 464, 465, 494, 549, 550, 781, 823, 837, 897, 761, 939, 1115, 646, 713, 767, 619, 440, 473, 498, 242, 718, 696, 743, 445, 446, 414, 447, 848, 918]. algebra [148, 215, 130, 861, 937, 369, 702, 578, 194, 253, 255, 295, 330, 332, 333, 579, 112, 629, 54, 80, 133, 164, 166, 201, 230, 231, 271, 314, 388, 490, 731, 914, 838, 786, 736, 827, 620, 843, 742, 744, 671, 720, 972, 115]. algebraic [480]. Algorithm [125, 248, 482, 701, 832, 679, 33, 39, 43, 70, 76, 94, 168, 169, 233, 757, 783, 689, 787, 812, 408, 559, 497, 274, 317, 275, 444, 476, 413, 773, 797, 818, 282, 52, 77, 206, 336, 392, 583, 782, 883, 645, 688, 733, 668, 8, 95, 131, 197]. Algorithm-Based [832, 679, 274, 317, 701, 757, 783]. Algorithmic [244, 318, 675, 456, 805, 779, 439, 472, 845]. Algorithms [1143, 653, 462, 16, 704, 222, 381, 195, 834, 15, 18, 22, 25, 44, 45, 55, 56, 137, 143, 464, 488, 494, 966, 123, 1020, 760, 811, 761, 840, 1034, 958, 714, 693, 766, 767, 768, 497, 903, 718, 699, 110, 702, 963, 257, 296, 964, 629, 54, 119, 213, 262, 490, 493, 614, 754, 807, 913, 838, 786, 738, 816, 962, 972, 223, 258, 1034]. Algorithms-By-Tiles [693, 714]. AlgoWiki [845]. Alignment [624, 640]. alloys [890]. Alto [1042]. Always [1144]. amount [916]. AMS [1058]. Amsterdam [1141, 1064, 1065, 1066]. Analysis [872, 90, 419, 821, 167, 340, 341, 394, 761, 786, 1137, 1136, 530, 534, 637, 979, 937, 91, 996, 975, 981, 764, 901, 669, 620, 672, 1131]. Analyzing [64, 84]. anatomy [883].

Annapolis [1010]. Annealing [594, 574]. Annotated [529]. Annual [993, 1058, 1094, 1050]. API [566, 568]. Appl [934]. Application [507, 484, 580, 405, 406, 976, 530, 1025, 846, 517, 622, 829]. application-level [622]. Application-Specific [1025]. Applications [146, 502, 503, 681, 240, 1019, 552, 588, 966, 436, 496, 528, 590, 632, 1109, 1048, 605, 849, 746, 180, 963, 1030, 161, 1001, 616, 899, 1021, 900, 1099, 650, 672]. Applied [937, 675, 834, 965, 311, 1009, 338, 528, 951, 840, 1024, 1013, 1067, 1095, 1112, 1127, 1133, 1134, 1139, 1140, 1142, 1145, 701, 1108, 1116]. Applying [420, 589]. Approach [922, 244, 776, 40, 492, 523, 318, 843]. Approaching [55]. Approximate [874]. Approximation [846]. April [993, 1072, 982, 1003, 1011, 1022, 1059, 1074, 1076, 1090, 1091, 1098, 1012, 1000, 1064, 1065, 1066, 978, 1041, 1053]. Arbitrary [935]. Architecture [502, 69, 407, 791, 1046, 196, 261, 712]. Architectures [745, 674, 653, 656, 676, 224, 18, 26, 39, 74, 172, 211, 232, 269, 270, 1020, 709, 761, 408, 559, 1023, 813, 714, 693, 694, 737, 766, 767, 768, 770, 1025, 444, 476, 413, 891, 906, 933, 181, 750, 852, 677, 702, 101, 212, 914, 786, 738, 816, 962, 971]. Areas [1143]. Argonne [950, 113, 41, 38]. Arithmetic [952, 648, 927, 918, 867, 885, 887, 649]. Array [385, 428, 429, 430, 386, 387, 828]. Art [950, 1108, 1116, 890, 1078, 697]. Articles [932]. Asia [1022, 1092]. Aspect [589]. Aspect-Orient [589]. Aspects [967]. Assessing [777, 855, 820]. assessment [1016]. assisted [817]. Asynchronous [747, 748, 967, 795, 645, 905]. Athens [959, 1032]. Atlanta [1100, 1101, 1102, 455]. ATLAS [524, 447, 500, 541]. atmospheric [988]. August [1008, 1009, 1089, 999, 1004, 1077, 1050, 1061, 961, 1040, 1052, 1024, 1013, 967, 808].

Austin [288]. Australia [1020, 1080, 1081, 1082, 1083]. Austria [1077, 1060, 961, 1138]. Auto [716]. Auto-tuning [716]. Automated [821, 500, 541]. Automatic [579, 547, 587, 529, 178, 145, 537, 600, 672, 179, 495]. Automatically [499, 414, 447]. Autotuned [745]. Autotuner [943]. Autotuning [765, 789, 791, 831, 931]. autotuning-based [931]. Available [202]. Avoiding [819, 771]. Award [909, 1143]. Aware [762, 763, 826, 817]. Azzam [934].

B [937]. back [864, 827]. balancing [910]. Balatonfüred [1039]. Baltimore [1063]. Band [694, 737, 670]. Banded [449, 450, 451, 381, 35, 36, 61, 83]. Barbara [1003]. Barcelona [1033]. Barrett [1146]. Based [185, 186, 675, 832, 679, 1069, 303, 274, 317, 828, 849, 934, 830, 149, 701, 425, 757, 783, 589, 616, 826, 839, 764, 856, 888, 813, 931, 564, 719, 741, 601, 622]. Basic [917, 148, 509, 297, 331, 28, 46, 75, 96, 97, 131, 238, 330, 80, 166, 508, 544, 545, 259, 95, 104, 115, 141, 175, 549, 550]. Basis [340, 341, 394, 936]. batch [860, 880]. Batched [873, 917, 834, 839, 840, 886, 857, 894, 841, 934]. batches [908]. Baton [1128, 1129]. Beach [1049]. Beautiful [1114]. been [262]. Beijing [1117, 1118, 1119, 1120]. Belgium [1012]. Bell [111, 142, 161, 174]. Benchmark [99, 307, 516, 551, 612, 685, 708, 896, 647, 100, 136, 352, 515, 853, 882, 584]. Benchmarking [72, 160, 971]. Benchmarks [985, 273]. Benefit [907]. Berlin [955, 1056, 1024]. Berry [1146]. Better [721, 926]. bi [880]. bi-diagonalization [880]. Bialystok [1142]. Biannual [511]. Bidiagonal [294, 737, 766, 255, 295, 884, 816, 741]. Bidiagonalization [694]. Big [877, 1094, 842]. Biographies [682]. Biological [624, 640]. birth [1131]. bit [893, 641, 678, 648, 649]. bit-flip [893]. BlackjackBench [778, 821]. BLACS [165, 177, 204, 239, 402]. BLAS [544, 545, 933, 508, 191, 259, 330, 260, 195, 923, 924, 65, 86, 121, 302, 659, 144]. Blast [549, 550, 509]. Block [125, 747, 184, 248, 259, 260, 195, 385, 428, 429, 430, 834, 71, 114, 137, 386, 387, 840, 693, 439, 472, 795, 892, 894, 282, 330]. Block-Asynchronous [747, 795]. Block-Cyclic [385, 428, 429, 430, 386, 387, 472]. block-Jacobi [892, 894]. Blocking [145, 879]. Blocks [216, 245, 246]. Bombardment [244, 318]. Bonas [964]. Bonn [1030, 1110]. Book [1146, 937, 40, 198]. Boole [615]. Bridging [890]. Bringing [772]. Broadband [710]. Brussels [1012]. Budapest [1093]. Building [216, 246, 684, 617, 245]. built [616]. bulk [670]. Bunch [937]. Butterfly [935]. Bytecode [537, 600].

C [937, 269, 270]. CA

[1044, 1045, 967, 999, 1003, 1005, 956].

Cache [611]. calculating [336]. calculation [670]. Calculations [823, 619, 826, 697]. calculator [937]. Calif [1042]. California [1014, 1068, 953, 1002, 1023, 1049, 1050]. Call [263, 707, 566, 568]. Cambridge [1048]. Canada [1040, 1046, 1089]. Cancun [1000]. capabilities [650]. Capability [785, 800]. capping [900]. Case [389, 660, 886, 532, 847]. Catalyzing [942]. Cathedral [953]. Catherine [987]. CCDSC [808]. CCGrid [1075, 1090]. CCGrid2002 [1056]. CCGrid2003 [1073]. CCGSC [756]. CCI [263]. CD [1094]. CD-ROM [1094]. CELL [673, 647, 663, 690, 692, 712, 662, 710]. Center [1027, 1038, 1104, 1137, 970, 1005, 1041]. CERFACS [122]. Cetraro [1001, 1021, 990]. Challenge [240].

Challenges [873, 154, 189, 779]. Chan [1146]. Changed [865]. Changing [1143, 337, 390, 511]. Characteristics [958]. Characterization [821, 611, 778]. Chateau [964]. Cheaper [721]. Chebyshev [336]. Checkpoint [801]. Checkpoint-on-Failure [801]. Checkpointing [796, 798, 777, 393, 360, 274, 317, 410, 820, 782]. Checksum [359, 361]. Cheju [1004]. chemistry [1009]. Chicago [969, 1055, 1031, 1090]. China [1099, 1117, 1118, 1119, 1120]. choice [583]. Cholesky [860, 254, 329, 382, 487, 391, 491, 688, 689, 733, 760, 811, 886, 663, 692, 857, 717, 847]. Cimmino [184]. City [1124, 1137, 1092, 1041]. Class [653, 702]. clMAGMA [804]. Cloud [940]. Clouds [756, 806, 835, 866, 753, 895, 929]. Cloudy [932]. Cluster [481, 458, 654, 518, 687, 471, 1055, 596, 1056, 1073, 1075, 1090, 1049, 567, 771, 555, 754, 856, 888]. Clusters [577, 609, 467, 512, 520, 630, 756, 806, 835, 866, 797, 803, 578, 554, 613, 706, 753, 807, 895, 929]. CO [1043]. Code [660, 1114, 764]. Collaboration [625]. collaborative [636]. Collection [367, 10, 307, 643, 564, 17]. Collections [373]. Collective [495, 637, 499, 538, 539, 571, 623, 817, 668, 669]. College [987]. Colorado [1098]. Combination [796]. Combining [931]. Communication [504, 819, 506, 175, 238, 310, 590, 592, 593, 771, 543, 800, 459, 300, 905]. Communication-Avoiding [819, 771]. communication/computation [459]. Communications [992, 289, 495, 499, 538, 539, 571, 623, 148, 166, 817, 1103]. Communicators [592, 633]. Community [455, 289, 684, 707, 772]. Compact [937]. Comparative [698, 178, 179]. Comparèes [27]. Comparing [899]. Comparison [449, 450, 451, 805, 41, 51]. Compcon [953]. Compilation [1023]. compiler [345]. Compilers [92, 178, 93, 170, 179].

Compiling [404, 469]. Complete

[437, 443, 362]. complex [908]. Complexity [225, 855]. Component [589, 631, 849, 616]. Component-based [589, 616]. Components [652, 700]. composition [616]. Comput [934]. Computation [987, 359, 941, 990, 984, 1024, 967, 459, 968, 122, 361, 902, 961, 1111, 1013, 40]. Computational [1143, 479, 501, 1085, 1086, 1087, 1088, 912, 327, 328, 378, 379, 1096, 950, 467, 513, 520, 548, 926, 533, 564, 1032, 1064, 1065, 1066, 1080, 1081, 1082, 1083, 772, 154, 189, 512, 514, 613, 687, 706, 913, 855, 890, 921, 930, 565, 1079, 671, 640, 1044, 1045, 1105, 1106, 1107, 1128, 1129, 1121, 1122, 1123, 868, 1141, 1117, 1118, 1119, 1120, 1100, 1101, 1102, 624, 937]. Computations [922, 1143, 641, 678, 628, 71, 114, 770, 275, 934, 699, 314, 1009, 839, 671]. Computed [6, 7, 8, 11, 12, 13]. Computer [909, 1143, 953, 18, 26, 39, 72, 113, 985, 511, 952, 1004, 1062, 101, 962]. Computers [127, 128, 129, 278, 366, 108, 322, 192, 193, 222, 223, 256, 298, 299, 68, 628, 30, 31, 34, 47, 57, 58, 60, 63, 69, 70, 82, 102, 116, 140, 160, 171, 173, 208, 209, 210, 226, 236, 240, 309, 347, 434, 521, 551, 612, 273, 573, 130, 937, 369, 194, 253, 257, 258, 296, 332, 333, 112, 14, 32, 48, 54, 59, 62, 81, 94, 103, 139, 196, 230, 235, 261, 314, 388, 397, 983, 971]. Computing [1014, 1026, 1027, 1038, 1043, 1104, 723, 992, 1143, 1029, 502, 542, 652, 700, 481, 506, 576, 186, 218, 219, 320, 368, 283, 675, 456, 289, 324, 326, 458, 654, 377, 421, 704, 252, 1096, 334, 335, 70, 74, 113, 117, 965, 132, 138, 969, 197, 205, 980, 986, 229, 995, 303, 311, 340, 341, 394, 399, 1018, 467, 489, 511, 513, 518, 520, 522, 548, 553, 1071, 615, 630, 1108, 684, 756, 781, 806, 835, 865, 866, 966, 590, 632, 955, 471, 272, 951, 1055, 940, 1137, 982, 997, 999, 1002, 1004, 1010, 1031, 1048]. Computing [1056, 1057, 1073, 1075, 1090, 1092, 1116, 1049, 691, 711, 869, 715, 1136, 563, 598, 665, 566, 942, 961, 498, 599, 567, 568, 991, 978, 412, 844, 1041, 1053, 602, 604, 772, 846, 848, 606, 1015, 877, 150, 247, 157, 321, 323, 701, 725, 457, 728, 292, 325, 425, 1030, 680, 681, 134, 994, 1001, 1009, 403, 512, 514, 554, 555, 613, 644, 687, 706, 730, 753, 780, 853, 854, 895, 929, 645, 526, 528, 1021, 947, 1003, 1011, 1022, 735, 1012, 695, 1079, 1063, 842, 945, 1052, 1111, 475, 638, 276, 719, 741, 622, 639]. computing [1013, 956, 1103, 775, 338]. Concepts [1143, 343, 395, 589, 396, 433]. Concurrent [185, 186, 192, 193, 222, 223, 256, 298, 45, 56, 149, 194, 257, 258, 296, 276]. Condensed [71, 114, 172, 211, 762, 212, 763]. Condition [627]. Conditioning [652, 700]. Conf [976]. Conference [1006, 1014, 1027, 1042, 1068, 1104, 1044, 1045, 1105, 1106, 1107, 953, 1016, 1046, 1008, 1085, 1086, 1087, 1088, 1121, 1122, 912, 1069, 1096, 658, 969, 980, 1019, 1089, 955, 987, 1020, 975, 981, 1055, 618, 1035, 1137, 959, 982, 997, 998, 1023, 1092, 1005, 1049, 1077, 921, 1141, 1136, 1061, 1036, 1051, 1079, 1032, 1052, 1117, 1118, 1119, 1120, 991, 1064, 1065, 1066, 1080, 1081, 1082, 1083, 1094, 978, 1041, 1053, 1025, 967, 1067, 1095, 1112, 1127, 1133, 1134, 1139, 1140, 1142, 1145, 1128, 1129, 993, 1007, 1123, 965, 1012, 1100, 1101, 1102, 1103]. Conference [597]. Congress [1024]. conjugate [853, 882, 583]. conjugate-gradient [853]. conjunction [1072]. Connecting [513, 514]. Connection [198]. Conquer [381, 787, 444, 476, 793, 884]. Considerations [138, 205]. Considered [334, 335]. Consistent [923, 924]. Constellations [630]. Constructing [252, 264, 388]. Content [372]. Contents [88, 89]. Contributed [932]. Contribution [939]. Control [436]. Convention [1027, 1038, 1104, 1137, 1005]. Convergence [501]. Conversation [794].

Conversion [263]. Cooperative [707]. coordination [803]. Core [698, 343, 395, 646, 443, 803, 382, 487, 391, 396, 433, 754, 807, 825, 788, 916, 817, 770, 836]. cores [887, 915]. Correction [748]. Correlated [803]. Corrigenda [96]. Cosenza [990]. cost [855]. countermeasures [873]. Counters [484, 581, 611, 530, 582]. County [1027]. Couple [524]. Coupled [888, 856]. Coupling [37]. CPU [826, 790, 843, 847, 859]. CPU-GPU [826]. CPUs [734, 846]. Cracow [1017]. CRAY [3, 25, 33, 41, 44, 51, 55, 16]. CRAY-1 [3]. Cross [484, 533, 565, 620]. cross-experiment [620]. Cross-Platform [484, 533, 565]. Crossover [594]. CRPC [253]. CUDA [732, 784, 946]. Current [481, 633, 217]. Cutting [941]. Cyclic [385, 428, 429, 430, 386, 387, 439, 472]. Czestochowa [1095].

D [625, 828]. D.C [1041]. DAG

[725, 775, 776]. DAGuE [725, 726, 775].

Dallas [1038]. Dangers

[320, 368, 334, 335, 321]. DARPX [1125]. Data [1054, 501, 625, 283, 1084, 806, 835, 866, 868, 830, 877, 323, 895, 929, 902, 916, 936, 842]. data-dependent [936]. Database [221, 249]. dataflow [856, 888, 889]. dataflow-based [856, 888]. DCE [358]. Debugging [220, 292]. December [969, 951, 1020, 1109, 1005]. Decomposition [278, 366, 828, 883]. Decompositions [472]. deep [931]. defined [901]. Definite [907]. Demmel [1146]. Denelcor [10, 17, 29]. Denmark [994, 1009, 1108, 1013, 338]. Dense [872, 745, 674, 852, 776, 832, 655, 804, 294, 22, 44, 55, 200, 203, 268, 823, 837, 757, 761, 787, 767, 768, 696, 742, 743, 846, 750, 818, 863, 255, 295, 78, 98, 133, 201, 731, 783, 838, 786, 867, 736, 827, 769, 843, 744, 829]. Denver [1043, 1098]. Dependence [726]. dependent [936]. Deploying [485, 441, 474, 567]. depth [901]. description [170]. Design [872, 322, 415, 800, 424, 254, 256, 294, 299, 329, 333, 382, 39, 70, 143, 226, 232, 268, 391, 557, 632, 123, 976, 869, 619, 242, 275, 828, 191, 255, 295, 332, 487, 94, 213, 231, 558, 838, 885]. Designer [440, 473, 498]. Designing [158, 159, 18, 810]. Detection [796]. detector [878]. Determining [383, 384, 427]. Develop [721]. Developing [502, 416, 417, 64, 79, 84, 605, 110, 112]. Development [502, 185, 186, 507, 949, 580, 68, 965, 665, 149, 119, 389]. Developments [651, 833]. diagonalization [880]. Diagonally [449, 450, 451]. Diego [1005]. Different [18, 39, 70, 650]. differential [996]. Digest [953, 1002]. Digital [283, 323, 288, 976]. Dimensional [238, 166, 175]. Dinos [227]. Diophantine [906]. Direct [60]. direction [217]. Directions [26, 494, 630, 1030]. Discovery [1084]. discretizations [920]. Diskless [360, 274, 317, 410]. Distributed [503, 278, 366, 453, 322, 726, 776, 284, 285, 288, 462, 192, 193, 222, 223, 256, 298, 299, 1069, 224, 172, 207, 211, 228, 232, 234, 1072, 1109, 408, 559, 999, 1004, 1031, 1057, 1059, 1074, 1076, 1091, 1098, 498, 599, 718, 978, 444, 476, 605, 413, 904, 181, 818, 725, 775, 286, 287, 191, 194, 257, 258, 296, 332, 333, 964, 681, 212, 392, 1099, 1052]. Distributed-Memory [726, 718, 904]. Distribution [19, 42, 73, 237, 313, 312]. Distributions [439]. distributive [262]. diverse [933]. Divide [381, 787, 444, 476, 793, 884]. Divide-and-Conquer [381]. Division [113]. do [53]. Documentation [477]. Domain [850, 66]. Dominant [449, 450, 451]. Donato [1146]. Dong [934]. Dongarra [1146, 934, 937, 794, 909, 1143, 939, 942]. Door [156, 153]. dot [670]. double [782]. DPLASMA [726]. Draft [452, 152, 228].

Driven [824]. Dublin [1126]. Dundee [975, 981]. Dynamic [372, 496, 592, 633, 711, 438, 718, 461]. Dynamically [761, 786]. Dynamics [1032].

Early [265]. ECMWF [988]. Edge

[941, 670]. Edinburgh [1057, 1058]. editor [706, 881, 895, 866]. Editorial [199, 705, 671]. Editors [497, 929, 488, 753, 806, 835]. eds [40]. Education [564]. effect [720]. Efficiency [767, 861]. Efficient [872, 225, 920, 770, 911, 818, 262, 616, 885, 900]. effort [642, 555]. eigen [936]. eigen-basis [936]. eigen-spectrum [936]. Eigenproblem [124]. eigensolver [826]. eigensolvers [697]. Eigensystem [2, 1]. Eigenvalue [367, 954, 9, 21, 34, 43, 63, 71, 76, 114, 168, 169, 172, 211, 233, 762, 787, 1111, 444, 476, 480, 20, 52, 77, 206, 212, 763, 936, 829]. Eigenvalues [6, 7, 8, 11, 13, 197, 132, 134]. Eigenvectors [7, 11, 13, 197, 132, 134]. Eighth [1072, 1000]. Eijkhout [1146]. EISPACK [21, 2, 1]. electrical [1037]. Electronic [19, 73, 42, 826, 889, 697]. Electronically [118]. electronics [1037]. Elegant [660]. element [920]. Eleventh [1069]. Elimination [805, 660, 894, 833, 944]. Emerging [481, 914]. Empirical [745, 500, 541]. Enable [940]. Enabled [907, 575, 377, 378, 420, 422, 485, 608, 661, 423, 546]. Enabling [933, 421, 815, 531, 532, 636]. encoding [668]. Encyclopedia [845, 1037]. End [628, 530, 665]. End-user [530]. Energy [872, 798, 594, 767, 831, 861, 885, 900]. Energy-Efficient [872, 885, 900]. engine [725, 775, 710]. Engineering [951, 564, 979, 1009, 931, 1037]. engineers [947]. Enhance [802, 641, 678]. Enhanced [722]. enhancement [1016]. enhancements [281]. Enhancing [709]. Environment [723, 478, 220, 30, 31, 47, 79, 82, 102, 116, 171, 207, 228, 234, 664, 619, 531, 409, 316, 440, 473, 498, 574, 188, 247, 14, 32, 48, 81, 103, 139, 616, 470]. Environments [462, 252, 986, 264, 995, 1018, 511, 593, 1072, 532, 724, 403, 591, 442, 956]. Equation [37]. Equations [10, 31, 47, 82, 102, 116, 140, 171, 208, 209, 210, 236, 309, 521, 551, 612, 663, 692, 846, 14, 17, 30, 32, 48, 81, 103, 139, 235, 996, 662]. erf [598]. Error [748, 796, 759, 813, 809]. Errors [758]. Espoo [1130]. ESSL [263, 266]. Euro [1077, 1008, 1077, 1061]. Euro-Par [1077, 1008, 1077, 1061]. Euro-Par’96 [1008]. Euromicro [1069]. EuroMPI [1135, 1132, 1138]. Europe [961, 1012]. European [1028, 1007, 1017, 1113, 703, 1047, 1135, 954, 1097, 1033, 1039, 1070, 1132, 1060, 1093, 1126, 1110, 1032, 1130, 1138]. EuroPVM [1007, 658, 667]. EuroPVM/MPI [658, 667]. EuroPVMMPI [597]. Evaluating [125, 592, 633, 971]. Evaluation [371, 324, 483, 265, 889, 875]. events [901]. Evolution [133, 411, 276, 78, 98, 925]. Evolves [660]. Example [609, 946, 170]. examples [423]. Exascale [707, 755, 898, 842, 933, 911, 779, 920]. Exception [923, 924]. Exchange [289, 284, 290]. Executing [562]. Execution [625, 726, 731, 856, 888]. exhibition [1012]. Expect [214]. Experience [320, 368, 321]. Experiences [831, 217, 16, 25, 581, 582, 471, 276]. experiment [620]. Experimental [74]. Experiments [124, 535, 574]. explanation [99, 100, 136]. Exploit [938]. Exploiting [802, 144, 648, 649, 971]. Extended [28, 65, 75, 95, 96, 97, 46]. Extending [801]. Extension [129, 2]. Extensions [426, 437, 242, 231, 928]. Extreme [777, 865, 877, 820, 883]. Extreme-Scale [777, 877, 820].

Facility [113, 70]. Factorization [722, 723, 745, 146, 876, 819, 656, 676, 254, 382, 225, 834, 45, 49, 56, 137, 343, 391, 395, 491, 492, 523, 751, 759, 689, 760, 811, 709, 840, 663, 690, 692, 734, 790, 814, 857, 717, 771, 847, 860, 180, 851, 797, 852, 677, 329, 879, 487, 300, 396, 433, 754, 807, 822, 809, 838, 733, 788, 710, 870, 904]. Factorizations [698, 832, 824, 757, 886, 873, 783]. Fail [679]. Fail-Stop [679]. failure [800, 878, 899, 801]. failure-prone [899]. Failures [832, 679]. Farming [460]. Fashion [567]. Fast [860, 673, 482, 195, 43, 144, 497, 885, 887]. Faster [721]. Fault [675, 832, 679, 393, 757, 526, 556, 557, 590, 617, 632, 593, 1002, 359, 360, 361, 274, 317, 410, 441, 474, 724, 701, 803, 783, 527, 558, 496]. Fault-Tolerance [441, 474]. Fault-Tolerant [617, 593, 1002, 359, 317, 410, 724]. Fe [1091, 1062]. Features [845]. February [1026, 1124, 1069]. Fellow [794]. Fermi [765, 789, 739]. FFT [911]. fifteen [78, 98]. Fifth [980, 1004, 988, 1002]. Finding [66]. Fine [893, 762, 763, 826]. Fine-Grained [762, 893, 763, 826]. Finite [583, 920]. Finite-choice [583]. Finland [1130]. First [994, 953, 1103]. FLASH [764]. flip [893]. Floating [648, 649]. Florida [1027, 1104, 1029, 960, 1059]. Fluid [1032]. followed [855]. Forecast [932]. Foreword [703, 729, 752]. Form [248, 294, 172, 197, 211, 737, 282, 255, 295, 132, 134, 212, 38, 719]. Format [689, 688, 733]. forming [890]. Forms [71, 114, 762, 763, 741]. formulated [262]. FORTRAN [90, 91, 109, 5, 8, 14, 32, 33, 46, 48, 79, 96, 97, 197, 446, 319, 27, 28, 30, 31, 47, 60, 64, 75, 81, 82, 84, 102, 103, 116, 134, 139, 171, 308, 344, 345, 350, 404, 469, 537, 600]. Forum [509, 549, 550]. forward [801]. Fourth [969, 1003, 1058, 1032, 984]. FP16 [907, 887]. FP16-Enabled [907]. Framework [834, 229, 785, 840, 562, 572, 605, 603, 640, 624]. France [1008, 1113, 964, 986, 1072, 951, 1076, 1074].

Francis [773]. Francisco [1068, 1044, 1045, 953, 999, 1023, 1050]. Freie [955]. Frontiers [1010, 984]. Ft [1059, 496]. FT-MPI [496]. Fueled [1143]. Fujitsu [41, 51]. Fukuoka [974]. Full [689, 733, 1094]. Fully [745, 76, 711, 52, 77, 688]. function [937]. fundamentals [1030]. Future [575, 339, 516, 630, 1004, 217, 584, 926, 1078].

G [937]. GA [1100, 1101, 1102]. gaps [890]. Gauss [894]. Gaussian [627, 805, 833, 660, 931, 944]. Gdansk [1127, 1145]. GeForce [791]. GEMM [789, 791, 716, 739]. GEMMs [765]. General [449, 450, 451, 197, 946, 132, 134, 358]. General-purpose [946]. Generalized [146, 180, 826]. Generalizing [935]. Generation [453, 224, 552, 588, 529, 38, 635, 109]. Generator [685, 708]. generic [725, 775]. Genetic [594]. Genova [1069]. Geostatistical [922]. Germany [1007, 1030, 1132, 1110, 1061, 1056]. Gers [964]. Gigaflop [55]. Global [707, 645, 1111]. GMRES [927]. Gordon [864, 111, 142, 161, 174]. GPGPU [809]. GPU [933, 722, 748, 830, 852, 834, 732, 784, 785, 826, 840, 886, 887, 788, 813, 734, 789, 790, 717, 916, 946, 770, 843, 696, 719, 741, 742, 743, 744, 772, 793, 846, 859, 870]. GPU-Accelerated [748, 793, 916]. GPU-based [830, 813, 719, 741]. GPU-Resident [834, 840]. GPUs [849, 860, 872, 873, 907, 721, 934, 831, 850, 861, 862, 875, 910, 863, 880, 823, 884, 839, 915, 857, 716, 739, 769, 829, 847]. gradient [853]. gradients [882, 583]. GrADS [507, 626, 610, 535, 640, 624]. GrADSolve [601, 602, 622]. Grained [762, 893, 763, 826]. Grand [240, 154, 189]. Grande [1042]. Grande/ISCOPE [1042]. Granularity [16, 25]. Graphical [149, 185, 186].

Graphics [1143, 747, 110, 795, 894, 109]. Greece [1047, 1135, 959, 1032]. Grid [723, 478, 542, 625, 460, 485, 580, 680, 1056, 1073, 1075, 1090, 1092, 562, 598, 531, 532, 1078, 566, 1063, 569, 604, 606, 610, 681, 555, 899, 603, 622, 640, 1063, 575, 504, 507, 626, 728, 546, 608, 548, 593, 560, 595, 563, 695, 636, 536, 535, 568, 570, 572, 601, 602, 639, 574, 624]. grid-based [622, 601]. Grid-Enabled [608, 546]. Grid-Enabling [531, 532]. grid-shaped [899]. GridRPC [566, 568]. Grids [501, 467, 520, 756, 746, 512, 613, 687, 706, 753]. GridSolve [607, 815, 651]. GridSolve/L [607]. Group [1028, 1113, 703, 1047, 1135, 1097, 1033, 1039, 1132, 1060, 1093, 1126, 1110, 1130, 1138, 1017, 1070]. Growth [1143]. GTX [791]. Guest [706, 881, 895, 929, 488, 753, 806, 835, 866]. Guide [126, 147, 182, 277, 448, 452, 505, 155, 187, 190, 250, 370, 415, 510, 4, 85, 176, 239, 305, 306, 342, 346, 402, 2, 272, 886, 1, 480, 152, 937].

H [67, 950]. Haidar [934]. half [908, 867, 885]. half-complex [908]. half-precision [885]. Handbook [1054, 1115]. Handling [923, 924]. Hardware [698, 479, 776, 484, 821, 525, 581, 660, 836, 938, 812, 530, 934, 778, 582, 611, 825, 839, 841]. Harmful [334, 335]. HARNESS [453, 431, 526, 527, 556, 557, 558, 438]. Harnessing [887]. Hawaii [1094, 1094]. HCW’98 [1029]. Healing [704, 724]. heFFTe [911]. held [948, 1030, 950, 965, 955, 1072, 1034, 990, 1062]. help [555]. HeNCE [152, 186, 187, 188, 218, 247]. Henrici [937]. HEP [10, 17, 29]. Hermitian [367, 9, 20]. Hessenberg [248, 282, 255, 294, 295, 813, 693, 719, 741]. Heterogeneity [802]. Heterogeneous [1029, 150, 151, 156, 218, 219, 220, 157, 320, 368, 776, 456, 334, 335, 229, 240, 303, 431, 683, 837, 405, 406, 436, 1011, 715, 770, 772, 153, 154, 189, 247, 321, 457, 1003, 788]. Heuristics [587]. HI [1011]. Hierarchical [754, 807]. Hierarchically [846]. High [1014, 1026, 1027, 1038, 1043, 1104, 872, 922, 127, 128, 129, 992, 1143, 934, 503, 507, 108, 675, 289, 324, 326, 654, 804, 421, 704, 252, 68, 158, 159, 628, 23, 34, 54, 57, 58, 63, 160, 163, 226, 232, 269, 270, 1001, 306, 340, 341, 346, 394, 399, 432, 434, 489, 494, 513, 514, 548, 615, 630, 707, 751, 780, 781, 853, 758, 590, 632, 1072, 1021, 688, 787, 886, 940, 1137, 982, 997, 999, 1022, 1031, 1057, 1092, 920, 647, 691, 715, 1136, 1012, 717, 766, 767, 816, 665, 858, 1079, 942, 242, 412]. High [844, 1041, 1053, 602, 445, 446, 774, 573, 130, 701, 725, 727, 728, 110, 325, 253, 112, 59, 62, 94, 119, 164, 230, 231, 314, 388, 554, 644, 730, 822, 854, 882, 913, 838, 947, 903, 1062, 475, 638, 775, 1096, 1103]. High-Level [507, 23, 1072]. High-order [920]. High-Performance [1026, 872, 127, 128, 129, 1143, 108, 324, 326, 252, 68, 57, 160, 434, 630, 707, 781, 940, 997, 691, 715, 942, 412, 844, 780, 853, 1012, 816, 858, 774, 130, 110, 112, 94, 119, 164, 230, 388, 644, 882, 913, 838, 903, 1062, 475]. High-Productivity [922]. high-resolution [774]. High-Speed [494]. Highly [248, 704, 911]. Hilaire [986]. Hill [953]. Hilton [970]. History [1125]. Hitachi [41, 51]. Homogeneous [766, 816]. Honolulu [1011]. honor [950]. Honors [1143]. hosts [338]. Hotel [953, 1053, 1041]. Householder [834, 840]. Houston [980]. HP [937]. HPC [878, 265, 337, 390, 825, 836, 938, 899, 1022, 932, 411, 1041, 1053]. HPCC [284, 289, 290, 1103, 264]. HPCN [1012]. HPCS [1125]. HPDC [1057]. HPDC-11 [1057]. HPF [415]. HPL [896]. HPL-AI [896]. Hungary [1039, 1093]. Hybrid [759, 788, 813, 696, 793, 809, 810, 688, 826, 843, 719, 741, 744]. Hybridization [721]. hydrodynamic [336].

IBM [961, 954, 135, 162, 266]. ICA3PP [1020]. ICA3PP/97 [1020]. ICCS [1044, 1045, 1105, 1106, 1107, 1128, 1129, 1121, 1122, 1123, 1141, 1117, 1118, 1119, 1120, 1100, 1101, 1102, 1085, 1086, 1087, 1088, 1064, 1065, 1066, 1080, 1081, 1082, 1083]. ICPP [1089]. Idle [383, 384, 427]. IDR [875]. IEEE [1014, 1104, 909, 1056, 1057, 1073, 1075, 1090, 1027, 953, 1055, 1005, 1049, 1025]. IEEE/ACM [1056, 1073, 1075]. IFIP [967, 1016, 1046]. II [1045, 1105, 1129, 450, 1086, 1122, 965, 550, 1141, 1118, 1065, 1081, 1101, 1134, 1140, 1145]. III [1106, 1087, 1123, 1119, 1066, 1082, 1102]. Illinois [969, 1055, 952, 1031]. IMA [1034]. IMACS [1024]. IML [305, 342]. Impacts [1143]. Implementation [126, 147, 502, 424, 254, 381, 382, 260, 45, 56, 75, 95, 131, 143, 391, 491, 617, 123, 812, 408, 559, 647, 662, 734, 857, 869, 621, 413, 828, 329, 487, 77, 213, 392]. Implementations [805, 49, 859, 833]. implemented [262]. Implementing [797, 22, 44, 55, 646]. Improve [811]. Improved [812, 661, 739]. Improvements [616]. Improving [6, 7, 8, 11, 12, 13, 1144]. in-depth [901]. Including [37]. Incomplete [874, 851, 879]. Increasing [23]. Indefinite [819, 891, 750, 818, 852, 863, 870]. Independent [285, 286, 287]. Indirect [60]. Industrial [937, 1013]. industrial-strength [1013]. Industry [412]. inefficiency [672]. Influence [939]. Informal [950]. Information [284, 470, 409, 1022]. Infrastructure [503, 484]. Initial [105, 107]. Innovations [542]. Inquiry [564]. Inquiry-based [564]. Installation [190, 250, 415, 510, 596]. Installing [105]. Institute [954, 117, 1062, 961]. Instrument [546, 608]. Instrumentation [586]. Int [934]. Integrated [229, 836, 825]. Integration [357]. Intel [137, 825, 836]. Intensive [625, 619, 720]. Interactive [372, 619, 531, 636]. Interconnected [912]. Interface [243, 503, 319, 483, 1113, 1135, 207, 234, 263, 308, 310, 1039, 581, 1132, 1060, 1126, 1110, 1130, 1138, 365, 1028, 415, 1017, 1047, 1097, 680, 350, 1033, 582, 1070, 1093, 597, 634, 695]. Interfaces [418]. Intermediaries [576]. International [1068, 1044, 1045, 1105, 1106, 1107, 957, 973, 974, 953, 1008, 1085, 1086, 1087, 1088, 1121, 1122, 912, 964, 1096, 994, 1019, 1108, 1089, 966, 955, 1072, 951, 1020, 1055, 1109, 1035, 1137, 959, 976, 999, 1002, 1023, 1031, 1048, 1056, 1057, 1058, 1059, 1073, 1075, 1076, 1090, 1091, 1092, 1098, 1116, 1049, 1077, 921, 1141, 1136, 1050, 50, 1061, 273, 1036, 1051, 1079, 1040, 1052, 1117, 1118, 1119, 1120, 1000, 1064, 1065, 1066, 1080, 1081, 1082, 1083, 1094, 1025, 1013, 1067, 1095, 1112, 1127, 1133, 1134, 1139, 1140, 1142, 1145, 1128, 1129, 1123, 1009, 1012]. international [1099, 1063, 1100, 1101, 1102, 1103, 707, 755, 1074]. Internetworking [506]. Interoperation [405, 406, 436]. intersection [902]. Introduction [265, 304, 467, 488, 808, 947, 946, 792, 671]. Invariant [138, 205]. Inverses [874]. Inversion [37, 689, 733]. Investigating [907, 867, 900]. Invited [1096, 1035, 1036, 1051, 1079]. Invocation [601]. IO [528]. IPDPS [1072, 1059, 1076]. iPSC [137]. iPSC/860 [137]. Ireland [1126]. ISC [794]. ISCA [1052]. ISCOPE [1042]. Island [1004, 1094]. ISPA [1109, 1099]. Issue [801, 400, 467, 643, 806, 808, 835, 866, 635, 545, 703, 658, 706, 753, 895, 929, 1062, 671, 792]. Issues [674, 322, 456, 299, 965, 268, 557, 332, 333, 59, 101, 558]. Italy [1069, 1097, 658, 1001, 1070, 1021, 1109, 990, 1103]. iterations [879]. Iterative [748, 216, 244, 245, 246, 655, 462, 305, 310, 342, 648, 664, 850, 875, 892, 318, 515, 885, 887, 915, 649]. IV [1107, 1088, 1120, 1083].

J [934, 937]. Jack

[794, 934, 909, 1143, 939, 942, 943]. Jacobi [892, 894, 879]. James [950, 67, 87]. January [1094]. Japan [974, 1073, 1075, 1092]. Japanese [246]. Java [1026, 1042, 416, 417, 376, 380, 404, 469]. JLAPACK [404, 469]. John [937]. Jordan [894]. Jose [1014]. July [1016, 996, 1031, 1057, 1058, 1092, 1025]. June [1042, 1068, 288, 1085, 1086, 1087, 1088, 1121, 1122, 1123, 1096, 965, 994, 1001, 1108, 1021, 975, 981, 1035, 959, 952, 1002, 1116, 990, 1141, 1036, 1051, 1079, 1080, 1081, 1082, 1083]. JVM [537, 600].

Keeneland [772]. Kepler [791]. Kernel [817, 688, 855]. Kernel-assisted [817]. Kernels [673, 910, 762, 789, 791, 763]. Key [343, 395, 396, 433]. Kingdom [1016, 1058]. Klagenfurt [1077]. Knowledge [1084]. Knoxville [997]. Korea [1004, 1022]. Kraków [1085, 1086, 1087, 1088, 1121, 1122, 1123]. Krylov [830, 862, 875]. Kulisch [40].

L [40, 607]. Laboratory [948, 950, 113]. Lagrangian [774]. Lake [1124, 1137]. Lancaster [996]. Language [665, 214]. Languages [529, 1125]. LAPACK [917, 933, 105, 106, 107, 125, 126, 127, 128, 129, 147, 181, 182, 183, 215, 277, 448, 130, 88, 190, 250, 319, 577, 578, 609, 224, 628, 657, 923, 924, 137, 163, 164, 165, 166, 202, 204, 231, 263, 267, 306, 308, 344, 345, 346, 350, 404, 469, 242]. LAPACK/ESSL [263]. LAPACK90 [432, 435, 445]. LAPACK95 [452, 505]. Large [954, 596, 869, 275, 828, 848, 818, 697, 996]. Large-Scale [275, 848, 996]. Large-scaled [596]. LASL [948]. lattice [904]. Lauderdale [1059]. Learned [581, 582]. learning [931]. Least [652, 700]. least-squares [700]. lecture [925, 615]. lectures [161]. Lens [868]. Lessons [581, 582]. Letters [497]. Level [507, 260, 195, 23, 104, 141, 207, 234, 1072, 688, 760, 811, 144, 191, 121, 622, 80, 86, 115, 131]. Level-3 [260, 195, 688, 760, 811]. Libraries [458, 252, 256, 298, 158, 159, 657, 200, 203, 226, 264, 339, 518, 548, 781, 938, 633, 998, 529, 533, 536, 535, 416, 417, 376, 201, 230, 314, 388, 554, 565]. Library [106, 127, 128, 129, 183, 108, 322, 288, 192, 294, 299, 68, 29, 163, 268, 269, 270, 305, 342, 389, 590, 593, 567, 871, 215, 130, 369, 194, 255, 295, 332, 333, 112, 59, 164, 271]. Lightweight [828]. Limitations [654]. Limited [374, 375]. Linear [848, 872, 917, 721, 106, 127, 128, 129, 183, 449, 450, 451, 652, 674, 799, 279, 280, 216, 244, 245, 246, 108, 322, 508, 544, 545, 776, 653, 655, 804, 16, 577, 192, 193, 259, 294, 297, 298, 299, 331, 381, 547, 68, 628, 10, 15, 18, 22, 24, 25, 28, 29, 31, 44, 46, 47, 55, 57, 58, 75, 82, 95, 96, 97, 102, 104, 116, 131, 135, 140, 141, 162, 163, 171, 173, 175, 200, 203, 208, 209, 210, 226, 230, 232, 236, 238, 268, 306, 309, 310, 339, 346, 432]. Linear [434, 464, 465, 494, 521, 549, 550, 551, 612, 781, 823, 837, 897, 758, 761, 939, 1115, 646, 663, 692, 713, 648, 767, 619, 440, 473, 498, 242, 718, 696, 743, 445, 446, 414, 447, 846, 891, 906, 918, 148, 215, 130, 861, 892, 700, 749, 818, 318, 937, 369, 702, 578, 194, 253, 255, 295, 330, 332, 333, 579, 112, 629, 14, 17, 30, 32, 48, 54, 80, 81, 103, 133, 139, 164, 166, 201, 231, 235, 271, 314, 388, 490, 731, 914, 838, 786, 841, 867, 885, 915, 662, 736, 649, 827, 843]. linear [742, 744, 671, 720, 972, 115]. Linear-Algebra [339]. LINPACK [937, 3, 4, 24, 99, 100, 136, 516, 584, 708, 647, 551, 612, 685]. Linux [525]. Linz [1060]. List [340, 341, 394, 844]. Lists [511]. Liverpool [1028]. Load [910, 547, 579]. Load-balancing [910]. Location [285, 286, 287]. Location-Independent [285, 286, 287]. LOEN [965]. logging [727, 803]. Logistical [454, 506, 607]. long [928]. Look [575, 200, 203, 646, 864, 78, 98, 201]. Looking [827]. Loops [5, 145, 170]. Low [846, 300]. Low-Rank [846]. Luszczek [934]. Lyngby [994, 1009, 1108, 1013]. Lyon [1008].

M7 [494]. M9 [481]. Machine

[453, 1113, 22, 49, 1039, 272, 1060, 1126, 438, 1110, 1130, 1028, 1007, 1017, 1047, 1097, 1033, 1070, 1093, 155, 176, 597, 634, 198]. Madison [1053]. MAGMA [860, 933, 825, 836, 914, 739]. Mail [19, 73, 42]. Malleable [605]. Management [607, 284, 288, 405, 406, 438, 358, 1078]. managers [358]. Managing [436]. Many [1143, 760, 811, 803, 825, 836]. many-core [803, 817]. many-integrated-core [825, 836]. Manycore [696, 744]. March [1026, 992, 1015, 1029, 953, 980]. Market [373, 411]. marketplace [475, 638]. Maryland [1010]. Massachusetts [1048]. Massive [1054]. Massively [625, 1010, 984]. Master [683, 686]. Master-Worker [686]. Math [455, 938]. Mathematical [66, 291, 949, 19, 23, 73, 118, 919, 564, 42, 59, 78, 98, 925]. Mathematics [455, 937, 950, 113, 1024, 1067, 1095, 1112, 1127, 1133, 1134, 1139, 1140, 1142, 1145]. Matrices [907, 723, 278, 366, 627, 9, 22, 71, 114, 886, 768, 873, 908, 880, 20, 858, 903]. Matrix [849, 908, 673, 910, 367, 248, 373, 832, 641, 678, 222, 223, 258, 6, 8, 21, 45, 56, 143, 197, 269, 270, 393, 492, 523, 683, 685, 686, 708, 757, 123, 2, 144, 359, 360, 935, 694, 737, 274, 317, 410, 1, 770, 846, 934, 831, 282, 257, 296, 78, 98, 132, 134, 783, 996, 839, 361, 712, 858, 903, 769, 671, 829, 904]. Matrix-Vector [849, 769]. May [1044, 1045, 1105, 1106, 1107, 1128, 1129, 995, 997, 1022, 1056, 1073, 1075, 1117, 1118, 1119, 1120, 978, 1100, 1101, 1102]. Mcell [608]. McLean [984]. MD [1063]. Measurement [586]. Mechanisms [661]. Meeting [1028, 1113, 703, 1047, 1135, 1097, 1033, 1039, 1070, 966, 1132, 1060, 1093, 1126, 1110, 1130, 1138, 1017]. Melbourne [1020, 1080, 1081, 1082, 1083]. Memory [278, 366, 322, 726, 90, 67, 192, 193, 222, 223, 256, 298, 299, 224, 172, 173, 207, 211, 228, 232, 234, 762, 408, 559, 790, 718, 978, 444, 476, 413, 181, 91, 194, 257, 258, 296, 332, 333, 212, 392, 763, 826, 904, 720]. Memory-Aware [762, 763]. memory-intensive [720]. Message [243, 1113, 1135, 207, 228, 234, 347, 397, 463, 1039, 1132, 1060, 1126, 1110, 316, 1130, 1138, 365, 1028, 369, 727, 803, 1017, 1047, 1097, 680, 348, 1033, 1070, 1093, 597, 634, 695]. Message-Passing [243, 207, 347, 316, 397, 369]. Meta [526, 528]. Meta-computing [526, 528]. Metacomputing [438, 556]. Metascheduler [569, 570]. Meteorology [988]. Method [747, 795, 184]. Methodology [721, 101, 120, 351]. Methods [1143, 245, 246, 626, 305, 310, 342, 966, 951, 124, 664, 439, 472, 918, 893, 937, 963, 610, 1001, 336, 920]. metric [853, 854]. Mexico [948, 977, 1091, 1000]. Middleware [506, 543, 460, 1058, 661, 1050]. Migratable [605]. Migration [441, 474, 572, 603]. MIMD [260, 49]. MIMD-Machine [49]. Mini [53]. Mini-Super [53]. minimal [688]. Minimization [594, 831, 937]. Mining [1084]. Minneapolis [962]. Minnesota [1034]. Miranker [40]. Mississippi [998]. Mixed [907, 922, 748, 641, 655, 678, 896, 915, 647, 927, 847, 918, 699, 887, 662, 936]. Mixed-Precision [907, 922, 896, 647, 847, 915, 918, 887, 936]. Mixing [838]. MN [962]. Model [777, 75, 95, 131, 824, 589, 539, 623, 851, 820, 727, 728]. Model-Driven [824]. Modeling [922, 538, 571]. Modelling [1024, 731, 902]. Models [837, 1072, 1125, 889]. Modern [483, 962]. modes [860]. Modularity [23]. moldable [899]. Moler [937]. Monitoring [525, 563, 598]. Montreal [1089]. Most [33, 34, 63]. moving [779]. MP [16, 41, 44, 51, 55]. MP2 [25]. MPI [1113, 703, 1135, 1097, 1132, 1093, 1126, 1110, 1130, 1138, 746, 1028, 243, 800, 801, 419, 1017, 1047, 658, 304, 1033, 1039, 1070, 357, 405, 406, 436, 496, 526, 527, 528, 556, 557, 558, 617, 592, 633, 437, 1060, 817, 667, 534, 637, 668, 669, 362, 443, 538, 571, 365, 928]. MPI-1 [443]. MPI-2 [592, 437]. MPI Connect [528]. MPIConnect [436]. MPP [293, 348]. MPPs [630]. Multi [698, 732, 646, 790, 770, 849, 754, 807, 784, 788, 720]. multi-component [849]. Multi-Core [698, 646, 770, 754, 807, 788]. Multi-CPU [790]. Multi-GPU [790, 770, 788]. Multi-platform [732, 784]. multi-threaded [720]. Multicore [722, 745, 674, 653, 656, 676, 781, 824, 897, 709, 761, 812, 711, 713, 714, 814, 693, 694, 717, 737, 766, 767, 768, 718, 771, 743, 793, 846, 847, 859, 891, 848, 797, 750, 677, 702, 786, 735, 736, 738, 816, 817, 742]. multicore/many [817]. multicore/many-core [817]. Multiple [698, 722, 832, 393, 758, 734, 847, 829]. Multiplication [849, 673, 223, 258, 144, 908, 831, 712, 769]. multiplications [858, 903]. Multipole [482, 497]. Multiprocessing [16, 25]. Multiprocessor [225]. multiprocessors [300]. multitask [931]. Multitasking [44, 55]. Munich [1007].

NA-NET [167, 684]. NAG [389].

Naleczów [1067]. Naming [285, 286, 287]. Nanjing [1099]. nano [697]. nano-systems [697]. Narrow [449, 450, 451]. Narrow-Banded [449, 450, 451]. Nash [937]. National [950, 113, 284, 289, 290]. native [860]. NATO [1021, 990]. NCA [1048]. Nested [145]. NET [167, 684].

NetBuild [533, 565]. Netherlands [1141, 1064, 1065, 1066]. Netlib [251, 291, 684, 564]. NetSolve [606, 575, 478, 542, 454, 625, 327, 328, 377, 378, 379, 380, 420, 422, 423, 424, 425, 555, 645, 561, 532, 441, 474, 1111, 621, 517]. NetSolve/D [625]. Network [1026, 576, 156, 185, 186, 218, 219, 327, 328, 377, 378, 379, 420, 422, 485, 1069, 334, 335, 229, 240, 303, 1048, 661, 724, 149, 150, 151, 153, 154, 189, 247, 157, 423]. Network-Based [185, 186, 1069, 303, 149]. Network-Enabled [378, 420, 422, 485, 423]. Networked [470, 272, 409, 431]. Networking [1014, 1027, 1038, 1043, 1104, 1137, 1136, 1012]. Networks [334, 335, 386, 387, 393, 360, 410]. Newport [1049]. News [1144]. Next [453, 224, 657, 552, 588, 38]. NHSE [289, 288]. Nice [1072, 1074, 1076]. NM [1062]. No [976]. Node [722]. Non [105, 367, 168, 169, 233, 870]. Non-GPU-resident [870]. Non-Hermitian [367]. Non-Symmetric [168, 169, 233]. Non-Unix [105]. Nonsymmetric [278, 366, 248, 408, 559, 413, 282, 206, 392]. Norfolk [991]. Norway [965]. Note [166, 204, 806, 835, 866, 716, 706, 753, 881, 895, 929, 165]. Notice [934, 65, 121]. novel [826]. November [1006, 1014, 1027, 1038, 1043, 1104, 974, 1017, 988, 1137, 960, 970, 977, 989, 1023, 1136, 1099, 1063, 956]. NT [407, 471]. Numbers [627]. Numerical [183, 479, 368, 458, 577, 609, 252, 158, 159, 805, 138, 163, 205, 264, 389, 434, 464, 465, 490, 494, 518, 548, 552, 588, 587, 642, 751, 823, 913, 631, 975, 981, 939, 711, 936, 440, 473, 498, 536, 535, 567, 962, 944, 871, 918, 215, 937, 1016, 416, 417, 1131, 376, 461, 578, 968, 133, 164, 554, 614, 822, 996, 972, 167]. Numerically [619]. numerics [993]. NVIDIA [791, 857]. NWChem [856, 888].

Oberlech [961]. Object

[993, 232, 242, 231, 271]. Object-Oriented [242, 231]. Obtaining [648, 649]. October [1007, 1046, 1113, 703, 964, 1070, 998, 1010, 1048, 1049, 1060, 1062, 984]. offload [936]. oil [636]. Omiya [1092]. OMPI [746]. One [698, 824, 592, 633, 873, 905]. One-Sided [698, 824, 592, 633, 873, 905]. Open [845]. OpenCL [804, 732, 784]. Opening [153, 156]. OpenMP [897, 936, 891, 871]. Operations [135, 162, 393, 633, 359, 360, 713, 637, 274, 317, 410, 361, 736, 669]. Optimal [798]. Optimisation [547, 579]. Optimization [875, 560, 561, 595, 635, 500, 849, 583, 903, 1013]. optimizations [541]. optimized [905]. Optimizing [504, 712, 769, 599, 883]. Orange [1027]. Orbit [37]. order [851, 920]. Ordinateurs [27]. Oregon [989, 978]. Organizer [618]. Organizers [455]. Orient [589]. Oriented [232, 242, 572, 993, 231, 271, 603]. Orlando [1027, 1029, 960]. orthogonal [936]. Ottawa [1046]. Out-Of-Core [343, 395, 382, 391, 487, 396, 433, 916]. outlook [780]. overhead [300]. Overlap [225, 850, 459]. Overview [316, 242, 568, 363, 364, 540, 573, 231]. Oxford [1016, 987].

PA [1006]. Pacific [1092]. Package [21, 24, 432, 445, 446]. Packages [698]. Packed [426, 689, 733]. PACX [528]. Paderborn [1061]. Palo [1042]. Paper [576]. Papers [801, 1030, 965, 1067, 1133, 1134, 1139, 1140, 1142, 1145, 953, 703, 912, 658, 756, 1035, 1002, 921, 597, 667, 1062, 1036, 1051, 1079, 1094, 1096, 1019, 1108, 1116, 1095, 1112, 1127]. PAPI [524, 525, 555, 901]. Par’96 [1008]. PARA [994, 1108, 1116, 1009, 1013]. PARA95 [311]. PARA96 [338]. Parallel [874, 876, 449, 450, 451, 278, 366, 625, 248, 1007, 419, 458, 653, 656, 676, 677, 948, 1113, 963, 1124, 222, 257, 259, 294, 296, 297, 1069, 964, 547, 382, 1030, 260, 805, 10, 29, 35, 36, 61, 64, 74, 76, 79, 83, 84, 117, 965, 969, 143, 168, 169, 170, 176, 980, 986, 233, 240, 995, 311, 338, 343, 391, 395, 1018, 1039, 518, 553, 1071, 1108, 837, 897, 122, 1089, 966, 528, 955, 987, 123, 471, 272, 1072, 1020, 762, 763, 812, 408, 559, 594, 988, 976, 998, 1010, 1023, 1059, 1074, 1076, 1091, 1098]. Parallel [958, 1116, 990, 1060, 664, 1126, 693, 694, 737, 619, 50, 1110, 534, 273, 214, 961, 498, 567, 945, 1130, 1040, 1052, 984, 1000, 991, 476, 601, 605, 845, 413, 967, 1067, 1142, 906, 972, 1028, 1015, 750, 188, 282, 1008, 109, 292, 1017, 702, 255, 295, 330, 1047, 579, 487, 1097, 833, 17, 52, 77, 120, 206, 213, 262, 1009, 396, 1019, 403, 433, 1033, 1070, 731, 1034, 1035, 1077, 1093, 1061, 1036, 1051, 983, 962, 622, 1013, 672, 956, 155, 223, 258, 331, 969, 994, 1109, 597, 634, 1099, 961, 1095, 1112, 1127]. Parallel [1133, 1134, 1139, 1140, 1145]. Parallelism [709]. Parallelizing [444, 170]. Paravirtualization [720]. ParCo97 [1030]. ParILUT [876]. Paris [1113, 703]. PARKBENCH [351, 352, 307]. PaRSEC [922, 797, 802]. Part [1044, 1045, 1105, 1106, 1107, 1128, 1129, 1121, 1122, 1123, 760, 1141, 1117, 1118, 1119, 1120, 1100, 1101, 1102, 1133, 1134, 1139, 1140, 1142, 1145, 1085, 1086, 1087, 1088, 1064, 1065, 1066, 1080, 1081, 1082, 1083]. Partial [790, 814, 731, 822, 996]. Pasadena [1002]. Passing [243, 1113, 1135, 207, 228, 234, 347, 463, 1039, 1132, 1060, 597, 634, 1126, 695, 1110, 316, 1130, 1138, 365, 1028, 369, 1017, 1047, 1097, 680, 348, 397, 1033, 1070, 1093]. Past [575, 516, 78, 98, 584]. Paths [72]. Patterns [90, 664, 91, 672]. PB [259, 330]. PB-BLAS [259, 330]. PC [596]. PDE [849]. PDE-based [849]. PDS [221, 249]. Perform [934]. Performance [1014, 1026, 1027, 1038, 1043, 1104, 849, 872, 127, 128, 129, 183, 215, 992, 1143, 503, 221, 108, 322, 371, 675, 775, 289, 324, 326, 419, 483, 484, 641, 654, 678, 804, 293, 421, 462, 704, 252, 299, 1096, 68, 158, 159, 225, 300, 805, 14, 23, 30, 31, 32, 34, 47, 48, 57, 58, 59, 60, 63, 81, 82, 102, 103, 116, 139, 140, 160, 163, 171, 208, 209, 210, 226, 232, 235, 236, 266, 267, 269, 270, 306, 309, 340, 341, 346, 347, 394, 399, 432, 434, 489, 513, 521, 525, 548]. Performance [551, 581, 586, 612, 615, 630, 707, 751, 781, 732, 758, 557, 590, 632, 592, 688, 811, 787, 886, 940, 1137, 982, 997, 999, 1031, 1057, 1092, 647, 691, 648, 715, 1136, 530, 766, 767, 665, 534, 942, 637, 669, 599, 242, 412, 844, 1041, 1053, 538, 571, 572, 602, 445, 446, 859, 905, 573, 933, 130, 861, 875, 249, 701, 725, 727, 728, 110, 325, 253, 332, 333, 112, 54, 62, 94, 101, 119, 164, 230, 231, 1001, 314, 388, 397, 514, 554, 582, 644, 730, 731, 780, 822, 853, 854, 882, 913]. performance [784, 558, 810, 838, 899, 1021, 947, 855, 1022, 764, 901, 649, 1012, 816, 931, 858, 903, 1062, 1079, 1144, 620, 475, 638, 603, 774, 1103]. Performance-portable [732, 784]. Performances [27]. Performant [717]. Period [798]. personal [230]. Perspective [50, 440, 473, 498, 41, 230, 919]. Petascale [522, 764]. Peter [937]. Petersburg [1080, 1081, 1082, 1083]. Phase [404]. Phi [825, 836]. Philadelphia [937]. physics [1009]. Pioneer [909]. Pioneering [1143]. Piotr [934]. Pipeline [22]. Pipelining [225, 646]. Pitfalls [72]. Pittsburgh [1006]. Pivoting [790, 814, 822]. planet [930]. PLASMA [897, 871]. Platform [484, 471, 533, 635, 732, 784, 565]. Platforms [456, 683, 686, 878, 457, 262, 899, 841, 788, 817]. PlayStation [654, 691]. PLW [665]. pocket [937]. Point [648, 649]. Poland [1017, 1085, 1086, 1087, 1088, 1121, 1122, 1123, 1067, 1095, 1112, 1127, 1133, 1134, 1139, 1140, 1142, 1145]. Poly [244, 318]. Poly-Iterative [244, 318]. Port [836, 825]. Portable [106, 127, 128, 129, 183, 322, 483, 299, 821, 158, 159, 79, 163, 581, 825, 215, 130, 332, 333, 778, 164, 582, 732, 784, 688]. Porting [871]. Portland [989, 978]. Porto [1019, 1035, 1036, 1051, 1079]. Portugal [1019, 1035, 1036, 1051, 1079]. Position [576]. Positive [907]. Post [800]. Post-failure [800]. potential [885]. Power [767, 900]. Poznan [1112]. Pozo [1146]. pp [937]. pp. [937]. PPAM [1095, 1112, 1127, 1133, 1134, 1139, 1140, 1142, 1145, 1067]. PPoPP [1124]. Practical [320, 321, 368, 480]. Practice [1124]. Precision [907, 922, 748, 641, 655, 678, 896, 647, 927, 847, 918, 892, 699, 867, 885, 887, 915, 662, 936]. precisions [908]. Preconditioned [862]. preconditioners [851]. Preconditioning [874, 892, 894, 879]. Prediction [922, 371, 462, 661, 621]. Preface [545, 486, 666, 658, 349, 398, 400, 466, 519, 520, 549, 550, 585, 613, 740]. Preliminary [791]. Preparing [562]. Present [575, 516, 584]. Principles [1124]. Prize [864, 111, 142, 161, 174]. Problem [9, 43, 76, 168, 169, 172, 211, 233, 685, 708, 591, 787, 595, 529, 532, 440, 473, 498, 444, 476, 20, 52, 77, 206, 212]. Problem-Solving [529]. Problems [279, 280, 367, 327, 328, 378, 379, 954, 21, 762, 480, 78, 98, 336, 996, 763, 843, 829]. Procedure [566, 568]. Proceedings [1006, 1014, 1027, 1042, 1105, 1106, 1107, 1128, 1129, 974, 1029, 1085, 1086, 1087, 1088, 1121, 1122, 1123, 948, 954, 969, 995, 1018, 1039, 1070, 966, 955, 1072, 951, 1055, 1034, 1137, 952, 960, 970, 977, 989, 997, 998, 999, 1004, 1023, 1031, 1058, 1059, 1075, 1076, 1005, 990, 1060, 1141, 1136, 38, 1032, 1117, 1118, 1119, 1120, 1000, 1064, 1065, 1066, 1080, 1081, 1082, 1083, 1094, 978, 1041, 1053, 1068, 1104, 1044, 1045, 1007, 1016, 1008, 1017, 1124, 1069, 964, 950, 1009, 1089, 975, 981, 988, 1022, 1048, 1056, 1057, 1073, 1074, 1092, 1098, 1049, 1050, 1012, 1061, 1063]. proceedings [961, 1040, 1052, 1100, 1101, 1102, 1024, 1025, 1013, 1028, 1113, 1047, 1135, 980, 1033, 1109, 1132, 1126, 1110, 1130, 1138, 967, 1097, 994, 959, 1091, 1077, 1093, 1099, 1103]. Process [405, 406, 436, 632, 919, 358, 1035]. processes [931]. Processing [747, 1069, 969, 980, 1089, 470, 1072, 1020, 1109, 1059, 1074, 1076, 1091, 1098, 50, 409, 1040, 1000, 991, 1067, 1095, 1112, 1127, 1133, 1134, 1139, 1140, 1142, 1145, 795, 1008, 1019, 1034, 1077, 1061, 1036, 1051, 1099, 945]. Processing-Systems [1019]. Processor [673, 61, 83, 663, 690, 692, 662, 712, 647]. Processors [483, 948, 260, 143, 123, 988, 976, 646, 711, 713, 1025, 967, 894, 120, 213, 736, 650]. Product [910, 683, 686]. Productive [940]. Productivity [922]. Profiling [767]. Program [635, 1125]. Programmes [27]. Programming [483, 1124, 101, 836, 837, 732, 589, 1072, 869, 38, 442, 946, 292, 120, 825, 784, 616, 889]. Programs [90, 419, 64, 75, 79, 84, 95, 131, 562, 534, 91, 109]. Progress [844, 181]. Progressing [478]. Project [380, 422, 580, 610, 112, 541, 507, 626, 726, 609, 707, 755, 500]. Projects [50]. prone [899]. Properties [805]. Proposal [297, 28, 207, 234, 308, 331, 46, 80, 115, 350]. Proposed [923, 924]. Prospectus [129, 68, 628, 657]. protection [893]. Protein [594, 621]. protocol [801]. Protocols [777, 820, 803]. Providing [503, 325, 461, 399, 785]. Provisional [88, 89]. Publ [976]. Public [66, 273]. PULSAR [869]. PUMMA [223, 258]. Purpose [674, 946]. Put [455]. PVM [1113, 703, 1097, 1033, 1093, 1126, 1110, 1130, 1028, 155, 188, 217, 218, 281, 477, 1007, 1017, 293, 1047, 176, 229, 240, 389, 1039, 1070, 357, 358, 407, 241, 272, 1060, 276]. PVM/MPI [1113, 703, 1097, 1093, 1126, 1110, 1130, 1033, 1028, 1047, 1039, 1060, 1017, 1070].

PVMPI [357, 405, 406]. Python [665].

QCG [746]. QCG-OMPI [746]. QL [124]. QR [677, 807, 124, 828]. QR/QL [124]. quadtree [668]. Quality [454, 1016]. quantum [670]. Quebec [1089]. Quest [522]. Quick [250]. QZ [336].

R [937]. Race [898]. Random [627, 935]. Randomization [799, 749]. randomized [818, 916]. Rank [846]. ranking [853, 854]. rationale [800]. Reading [1105, 1106, 1107, 988]. Receive [909]. Reconfigurable [431]. Reconfiguration [438]. Recovery [664, 800, 801]. Rectangular [689, 733]. Recursive [491, 492, 523, 751, 822]. Redesigning [727, 15]. Redistribution [385, 428, 429, 430, 386, 387, 439, 472, 855]. Reducing [731, 916]. Reduction [248, 294, 71, 114, 172, 197, 211, 212, 762, 813, 693, 737, 766, 768, 830, 851, 282, 255, 295, 132, 134, 884, 763, 816, 719, 741]. reductions [928]. Reference [305, 342, 437, 443, 362]. Refinement [748, 655, 648, 885, 887, 915, 649, 936]. Region [1092]. relations [351]. Relaxation [747, 795, 893]. Release [105, 107]. Reliability [599]. Reliable [968, 628]. Remote [479, 566, 568, 601]. Renaissance [1053]. Report [142, 174, 311, 355, 468, 551, 612, 273, 181]. Repositories [283, 372, 418, 285, 323, 286, 287]. Repository [291]. Request [504, 815]. Requirements [513, 514]. Rescheduling [626, 610]. Research [70, 113, 965, 494, 1021, 253]. reservoir [636]. Resident [834, 840, 870]. Resilience [758]. Resilient [759, 813, 809]. resolution [774]. Resource [373, 358, 1078]. Resources [251, 374, 375, 513, 514]. Restarted [927]. Result [821]. Resulted [1143]. Results [107, 92, 383, 384, 791, 93, 427, 170, 351, 352, 855]. Retraction [934]. Reuse [326]. Reversal [9, 37, 20]. Reverse [310, 359, 361]. Review [937, 419, 40, 534]. Reviews [1146, 198]. Revised [1096, 1108, 1116, 1067, 1095, 1112, 1127, 1133, 1134, 1139, 1140, 1142, 1145]. Revisited [693]. Revisiting [686, 782, 648, 649]. Rhinos [227]. Richardson [1052]. rigid [899]. RISC [135, 162]. roadmap [755]. Rockefeller [970]. ROM [1094]. Romine [1146]. Rouge [1128, 1129]. Routines [917, 183, 254, 547, 382, 10, 266, 391, 2, 760, 811, 646, 567, 1, 215, 191, 329, 579, 487, 17, 731]. RPC [728, 560, 595, 601, 602, 622]. RS [266]. RS/6000 [266]. RS/6000-550 [266]. Rules [865]. Run [240]. Runtime [661, 828, 724]. Russia [1080, 1081, 1082, 1083].

S [41, 51]. S-810 [41, 51]. S-810/20 [41, 51]. Saint [986]. Salt [1124, 1137]. San [1014, 1068, 1044, 1045, 953, 999, 1023, 1005, 1050]. SANS [555, 642]. SANS-Effort [555]. Santa [1003, 1091, 1062]. Santorini [1047, 1135]. Santorini/Thera [1047]. SC’06 [1104]. SC’11 [1136]. SC2000 [1038]. SC2001 [1043]. SC97 [1014, 1014]. SC98 [1027, 1027]. Scalability [802, 268]. Scalable [576, 625, 458, 484, 704, 192, 193, 256, 298, 628, 200, 203, 518, 982, 997, 998, 717, 409, 771, 724, 194, 201, 271, 914, 843, 470]. ScaLAPACK [322, 369, 370, 415, 192, 193, 194, 254, 299, 329, 332, 333, 510, 381, 382, 426, 487, 657, 271, 353, 354, 389, 391, 785, 535]. Scale [777, 954, 869, 275, 828, 848, 877, 820, 883, 697, 996]. scaled [596]. Scanning [635]. SCHEDULE [64, 85, 84]. Scheduled [761, 786]. Scheduler [711]. Scheduling [626, 460, 580, 385, 428, 429, 430, 713, 714, 736, 738, 650, 718, 574, 610, 622]. school [996]. Schwarz [905]. Science [1105, 1106, 1107, 1143, 1085, 1086, 1087, 1088, 1121, 1122, 912, 327, 328, 378, 379, 1096, 113, 902, 941, 1141, 564, 1062, 1117, 1118, 1119, 1120, 1064, 1065, 1066, 772, 1044, 1045, 1128, 1129, 1123, 1009, 913, 988, 890, 921, 926, 1079, 1080, 1081, 1082, 1083, 1100, 1101, 1102, 912, 921]. Sciences [455, 951, 1094]. Scientific [957, 502, 283, 948, 377, 485, 546, 704, 40, 969, 980, 986, 995, 1018, 467, 511, 520, 1108, 684, 756, 781, 806, 835, 866, 1116, 529, 691, 735, 498, 991, 1024, 699, 323, 1046, 425, 994, 403, 512, 613, 687, 706, 753, 895, 929, 900, 945]. scientists [947]. SCIRUN [532]. scope [801]. Scotland [1057]. Seamless [478]. Seattle [1136, 1053]. Second [995, 1019, 955, 976, 1008, 1009, 311]. Section [50, 680, 681, 687, 695]. Secular [37]. Secure [479]. Select [597]. Selected [909, 912, 1096, 266, 352, 1019, 614, 1108, 756, 1116, 921, 667, 1112, 1127, 1133, 1134, 1139, 1140, 1142, 1145, 703, 658, 1035, 1036, 1051, 1079]. selection [668]. Self [724, 577, 578, 609, 704, 629, 552, 588, 587, 631, 567, 538, 571, 604, 639, 642]. Self-Adapting [588, 587, 631, 578, 629, 552, 642]. Self-Healing [704, 724]. Semantic [618]. Semantics [632]. semi [774]. semi-Lagrangian [774]. semiconductor [670]. Seminar [957, 976, 961]. Semiseparable [846]. Seoul [1022]. Sept [1030]. September [1028, 948, 1113, 703, 1047, 1135, 1097, 658, 950, 117, 986, 1019, 1033, 1039, 1070, 955, 987, 1055, 1034, 1132, 1060, 1093, 1126, 1110, 1032, 1130, 1138, 1067, 1095, 1112, 1127, 1133, 1134, 1139, 1140, 1142, 1145, 1103]. Sequence [624, 640]. Sequencing [504, 815]. Server [575, 327, 328, 378, 379, 420, 485, 249, 221]. Servers [661]. Service [454, 940, 815]. Services [251, 1058, 1050]. Session [265]. Set [917, 508, 259, 297, 28, 75, 95, 96, 97, 104, 131, 141, 544, 803, 330, 331, 46, 80, 115].

Sets [1054]. Setting [503]. Seventh [1029, 1031, 1092]. shaped [899]. Shared [173, 790]. Shopping [118]. Short [673, 712]. short-vector [712]. Should [214]. SHPCC [982]. SHPCC-92 [982]. SIAM [969, 980, 991]. SICEDR [8]. Sided [698, 824, 592, 633, 694, 737, 873, 714, 738, 905]. Significant [1143]. SIGPLAN [1124]. Silent [796]. SIMD [673, 712]. Simple [596]. Simplified [606]. Simulated [594, 574]. Simulation [531, 1041, 1053, 764]. Simulation/Visualization [531]. Simulations [485, 546, 608, 636]. single [720]. single-and [720]. Singular [12, 812, 859, 883]. Sites [315, 401, 356]. Sixth [951, 1010, 978]. size [894]. Sizes [935]. Skinny [723]. SLHPF [415]. slicing [936]. Small [673, 16, 25, 886, 908, 880, 858, 903]. SmartGridRPC [728]. SNIPE [470, 409]. Society [909, 953, 937, 1004]. Soft [758, 759, 809]. Software [698, 721, 1143, 479, 66, 477, 507, 283, 372, 418, 284, 285, 289, 290, 291, 324, 326, 577, 609, 252, 256, 294, 298, 949, 580, 628, 19, 23, 26, 31, 47, 73, 82, 102, 116, 118, 140, 171, 210, 237, 264, 309, 312, 313, 314, 435, 463, 464, 521, 551, 552, 588, 587, 612, 707, 755, 897, 631, 990, 50, 533, 564, 983, 601, 414, 500, 848, 323, 1016, 1046, 286, 287, 425, 461, 578, 253, 255, 295, 629, 14, 30, 32, 42, 48, 59, 78, 81, 98, 103, 133, 139, 235, 388, 490]. software [779, 919, 925, 901, 827, 565, 1062, 541, 720, 642, 447]. software-defined [901]. Solution [652, 216, 244, 245, 246, 655, 732, 689, 700, 480, 318, 784, 996, 733, 915]. Solutions [799, 156, 749]. Solve [857]. Solver [758, 891, 875, 750, 423, 515]. Solvers [872, 907, 449, 450, 451, 377, 422, 34, 35, 36, 63, 743, 830, 862, 892, 810, 838, 841, 867, 885, 887, 742]. solves [850]. Solving [863, 154, 189, 327, 328, 378, 379, 21, 24, 37, 61, 83, 173, 208, 209, 236, 595, 529, 663, 692, 532, 440, 473, 498, 846, 906, 818, 879, 262, 591, 662, 843]. Some [674, 3, 35, 36, 45, 49, 56]. Sonic [1092]. Sons [937]. Sorrento [1097, 658, 1109, 1103]. Sourcebook [553, 1071]. Sources [949]. Spain [1096, 1033]. Sparse [849, 874, 910, 641, 678, 208, 209, 236, 269, 270, 491, 492, 523, 850, 861, 892, 879, 671]. Special [674, 801, 703, 67, 658, 680, 681, 142, 174, 400, 467, 643, 687, 806, 808, 835, 866, 895, 929, 50, 695, 635, 1062, 545, 706, 753, 671, 792]. Specialist [976]. Specific [1025]. spectra [336]. Spectral [278, 366]. Spectrum [599, 936]. Speed [494, 887]. sphere [774]. Spin [37]. spreadsheets [606]. Spring [953]. Squares [652, 700]. Squeezing [33, 34, 63]. SRS [605]. St [987, 1080, 1081, 1082, 1083]. Stability [195, 859, 336, 810, 944]. Stack [53]. Stage [768, 884]. Standard [243, 509, 30, 31, 47, 82, 102, 116, 140, 171, 208, 209, 210, 228, 236, 309, 521, 549, 550, 551, 612, 871, 801, 14, 32, 48, 81, 103, 139, 235, 348, 365, 304]. Standards [214]. Stanford [956, 967]. Stanimire [934]. State [607, 950, 1108, 998, 1116, 697, 1078, 670]. State-of-the-art [697]. states [670]. Static [457]. Statistical [1084]. Status [780, 217]. Stewart [937]. Stochastic [462]. Stop [679]. Storage [506, 426, 1137, 1136, 599, 543, 688]. Strategies [971]. Strategy [529]. strength [1013]. Structural [561]. Structure [631, 594, 826, 889, 621, 697]. Studies [847, 389]. Study [698, 886, 178, 532, 300, 179]. Stuttgart [1132]. Subprograms [917, 544, 545, 297, 46, 238, 549, 550, 148, 330, 80, 166, 508, 259, 331, 28, 75, 95, 96, 97, 104, 115, 131, 141, 175]. Subroutine [8, 688]. Subroutines [197, 134]. Subspaces [138, 205]. Suite [92, 93, 170].

Summer [996, 117]. Sunriver [993]. Super [53]. Supercomputer [315, 401, 161, 356]. Supercomputers [106, 957, 38, 363, 364, 540, 154, 189, 971]. Supercomputing [1006, 1068, 1104, 979, 973, 974, 156, 185, 960, 970, 977, 989, 1005, 149, 151, 153, 963, 988, 959]. superhighway [1022]. Support [502, 507, 283, 546, 608, 528, 770, 323]. Supporting [496, 358]. Supportive [1072]. Supports [229]. survey [918, 833, 62]. SVD [880, 884, 916]. Sweden [1116]. Switzerland [1025]. Symbolic [776]. Symmetric [907, 819, 43, 76, 168, 169, 233, 124, 762, 787, 768, 444, 476, 891, 750, 818, 852, 863, 52, 77, 763, 769, 829, 870]. Symmetric-Indefinite [819]. Symmetry [9, 37, 20]. Symposium [973, 974, 992, 1124, 950, 1072, 951, 1109, 952, 999, 1002, 1010, 1031, 1048, 1056, 1057, 1059, 1073, 1074, 1075, 1076, 1090, 1091, 1098, 984, 1000, 1041, 1053, 1062, 1099]. System [799, 625, 477, 377, 35, 36, 758, 759, 631, 526, 556, 560, 561, 595, 596, 790, 814, 869, 438, 1094, 846, 892, 749, 425, 809, 645, 867, 1111, 276, 621, 601, 622, 542, 135, 162]. System/6000 [135, 162]. Systems [449, 450, 451, 216, 244, 245, 246, 250, 655, 293, 485, 381, 547, 24, 37, 61, 83, 173, 303, 1019, 431, 463, 525, 586, 824, 837, 357, 1004, 529, 663, 692, 648, 599, 718, 771, 1025, 696, 605, 793, 906, 750, 818, 863, 318, 459, 879, 579, 271, 754, 780, 853, 854, 915, 662, 649, 1062, 961, 1052, 843, 744, 697]. Systolic [828, 797, 956].

takes [241]. talks [1035, 1051, 1079, 1036]. Tall [723]. Tampa [1104]. Task [726, 460, 441, 474, 718, 891]. Taskers [358]. tasks [826, 936]. tau [336]. tau-QZ [336]. TC2 [1046, 1016]. TC2/WG2.5 [1046, 1016]. Technical [302, 549, 550]. Techniques [872, 184, 799, 655, 208, 209, 236, 1023, 749, 885, 903]. Technologies [421, 337, 390, 399, 1041, 1053, 325]. technology [1001, 1021]. Telescoping [529]. Templates [279, 280, 480, 245, 246, 310, 914]. Tennessee [997]. tensor [887, 915]. Terascale [586]. Terms [805]. Tertiary [594, 621]. Test [367, 373, 92, 75, 95, 131, 93, 170]. Testing [105, 435]. Texas [1052]. theory [889]. Thera [1047]. Third [1018, 1035, 999, 1036, 1013, 1007, 1099, 1063, 1050]. Thirty [953]. Thirty-first [953]. threaded [720]. Three [555]. Threshold [876]. Tile [751, 709, 761, 787, 734, 766, 768, 771, 822, 786, 738, 816]. Tiled [653, 656, 676, 750, 677, 702]. Tiles [714, 693]. Tiling [374, 375, 459, 383, 384, 457, 427]. Time [798, 383, 384, 9, 37, 661, 427, 20, 731]. Time-Reversal [37]. Timings [3]. Tingxing [934]. tiny [873]. TLB [611]. TN [995]. Today [489]. Tokyo [1073, 1075, 1092]. Tolerance [675, 832, 679, 757, 632, 441, 474, 701, 783]. Tolerant [393, 496, 526, 556, 557, 590, 617, 593, 1002, 359, 360, 274, 317, 410, 724, 803, 527, 558, 361]. Tomov [934]. Tool [84, 143, 123, 596, 563, 598, 109]. Tools [185, 186, 218, 219, 90, 91, 419, 458, 64, 119, 120, 213, 986, 995, 1018, 518, 524, 643, 530, 534, 149, 110, 403, 555, 1125, 442]. Top [488, 497, 493, 511]. Top-500 [511]. TOP500 [315, 340, 341, 356, 394, 401, 844, 355, 468]. topology [817]. topology-aware [817]. Toronto [1040]. Torun [1139, 1140]. Touvet [986]. Townsend [995]. Trace [764]. Trace-based [764]. tracer [774]. Track [511]. Tracking [726]. trading [815]. Traditional [938]. Transfer [661, 830, 931]. Transformation [942]. Transformations [834, 840, 714, 738]. Transforms [935]. Transient [813]. Translation [537, 600]. Translational [919]. Transparent [533, 565]. transport [774]. Transpose [222, 257, 296]. traversing [188]. Tree [828]. Tree-Based [828]. Trends [481, 513, 554, 615, 644, 1004, 514, 730, 1078, 638, 276]. triangular [850, 879]. Tridiagonal [294, 381, 197, 124, 768, 444, 255, 295, 132, 134, 741]. Tridiagonalization [829]. Tromsø [965]. Truss [561]. Trustworthy [576]. tune [731]. Tuned [499, 414, 447]. Tuning [484, 587, 857, 716, 495]. Turing [1143, 925]. Tutorial [481, 354, 494, 272, 353]. Tweaked [938]. Twenty [1002]. Twenty-fifth [1002]. Two [166, 175, 238, 714, 694, 737, 768, 884, 738]. Two-Sided [694, 737, 714, 738]. Two-Stage [768]. TX [1038, 980, 288].

U.S. [340, 341, 394]. UK

[1028, 1105, 1106, 1107, 987, 996, 988].

Ulrich [40]. Umeå [1116]. Uncertain [932]. Understanding [143, 123, 213]. Unified [777, 820]. uniform [461]. United [1016, 1058]. Uniting [289]. Units [747, 795]. Universal [223, 258]. Universität [955]. University [1034, 952, 956]. Unix [105, 250]. Unrolling [5]. Unstable [664]. Untrusted [576]. Update [65, 524, 121]. Updated [508, 544]. Updating [851]. Upper [248, 282, 719, 741]. Upper-Hessenberg [248, 282]. Urbana [952]. USA [1038, 1044, 1045, 1128, 1129, 1124, 1137, 1052, 1100, 1101, 1102, 967, 1014, 1027, 1068, 1104, 992, 1029, 995, 999, 1048, 1059, 1005, 1049, 1062, 1063, 991]. Usage [557, 558]. Use [506, 380, 988, 543, 670]. User [1028, 182, 277, 448, 452, 505, 155, 187, 937, 370, 418, 1113, 1047, 1135, 1097, 4, 176, 207, 234, 239, 306, 346, 402, 1033, 1039, 1070, 272, 1132, 1060, 1093, 1126, 1110, 1130, 1138, 152, 1017, 530]. User-Level [207, 234]. Users [703, 85, 423]. Using [907, 799, 576, 484, 641, 678, 377, 425, 879, 30, 31, 44, 47, 55, 82, 102, 116, 140, 171, 208, 209, 210, 236, 237, 240, 309, 521, 525, 551, 611, 612, 897, 617, 560, 561, 359, 360, 663, 692, 530, 410, 828, 891, 574, 624, 928, 849, 873, 749, 863, 154, 189, 880, 14, 32, 48, 81, 103, 139, 235, 312, 313, 393, 554, 751, 822, 616, 884, 762, 763, 915, 594, 595, 361, 714, 693, 738, 766, 816, 768, 936, 640]. UT [1137]. Utah [1124]. utilizing [918].

V [306, 346, 305, 342]. v1.1 [402]. VA [992, 991]. Valencia [1096]. Value [812, 859, 883]. Values [12]. Variable [894, 547, 579]. Variable-size [894]. Variants [125, 805]. Various [30, 31, 47, 82, 102, 116, 140, 171, 208, 209, 210, 236, 309, 347, 521, 551, 612, 859, 14, 32, 48, 81, 103, 139, 235, 397]. VECPAR [1096, 1019, 1036, 1035, 1051, 1079]. Vector [849, 673, 1015, 910, 948, 260, 22, 60, 965, 173, 1019, 1035, 1051, 712, 769, 956, 928, 1019, 1036]. Vectorizing [92, 93, 178, 179]. Vendor [265]. Venice [1070]. Versailles [951]. Version [187, 424]. Versions [105]. Very [886, 858, 903]. VI [951]. via [830, 19, 42, 73]. Vienna [1138]. Virginia [982, 984]. Virtual [453, 155, 285, 288, 1113, 176, 1039, 272, 1060, 597, 634, 1126, 438, 1110, 1130, 828, 1028, 1007, 286, 287, 1017, 1047, 1097, 1033, 1070, 1093, 546, 608]. visP [598]. visPerf [563]. Visual [292, 316]. Visualization [220, 531]. Volume [67, 437, 443]. Vorst [1146]. VP [41, 51]. VP-200 [41, 51]. VPE [316]. vs [798].

W [937, 40]. WA [1136]. Washington [1041, 1053]. Web [373]. WG [967]. WG2.5 [1046, 1016]. Which [1143]. while [641, 678]. Wide [599]. Wiley [937, 1037]. Wilkinson [67, 950, 87]. Willard [40]. Williamsburg [982]. Windows [407, 471]. Within [144, 726, 785]. Work [455]. Worker [683, 686]. workflow [650]. workflows [815]. Working [1016, 1046, 165, 166, 204, 967]. Workshop [1026, 1015, 1029, 948, 964, 954, 86, 177, 204, 994, 995, 302, 311, 1018, 403, 613, 1108, 1072, 1021, 1034, 988, 1004, 1058, 1116, 990, 1050, 38, 1013, 1001, 1009, 1003, 1011, 1063, 165, 338, 756]. Workshops [1040]. Workstation [156]. Workstations [386, 387, 393, 360, 410, 348]. World [1143, 912, 496, 1024, 241]. World-Changing [1143]. Wrapper [665]. Wroclaw [1133, 1134]. Wyndham [1041].

x [937, 16, 25, 41, 44, 51, 55]. X-MP-2 [16]. X-MP-4 [41, 44, 51, 55]. X-MP2 [25]. XaaS [940]. Xeon [825, 836]. XNETLIB [237, 313, 312].

years [864, 78, 98, 921]. York [937, 970].

Zurich [1025].

References

Smith:1976:MER

[1] B. T. Smith, J. M. Boyle, J. J. Dongarra, B. S. Garbow, Y. Ikebe, V. C. Klema, and C. B. Moler. Matrix Eigensystem Routines: EISPACK Guide. Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1976. CODEN LNCSD9. ISBN 0-387-07546-1. ISSN 0302-9743 (print), 1611-3349 (electronic). vii + 551 pp. LCCN QA193 .M371 1976. See also [2].

Garbow:1977:MER

[2] B. S. Garbow, J. M. Boyle, J. J. Dongarra, and C. B. Moler. Matrix Eigensystem Routines: EISPACK Guide Extension, volume 51 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1977. ISBN 0-387-08254-9, 3-540-08254-9. viii + 343 pp. LCCN QA193 .M381, QA267.A1,L43 no. 51.

Dongarra:1978:SLT

[3] J. Dongarra. Some LINPACK timings on the CRAY-1. In Buzbee and Morrison [948], pages 58–75. U.S. Scientific Laboratory, Los Alamos, NM Conference proceedings LA-7491-C.

Dongarra:1979:LUG

[4] J. J. Dongarra, J. R. Bunch, C. B. Moler, and G. W. Stewart. LINPACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1979. ISBN 0-89871-172-X (paperback). 320 pp. LCCN QA76.73 .L22 L5 1979; QA184 .L56 1982; QA214 .L56 1979.

Dongarra:1979:ULF

[5] J. J. Dongarra and A. R. Hinds. Unrolling loops in FORTRAN. Software— Practice and Experience, 9(3):219–226, March 1979. CODEN SPEXBL. ISSN 0038-0644 (print), 1097-024X (elec-

tronic). URL http://www.netlib.
org/utk/people/JackDongarra/PAPERS/
Unrolling-Loops-in-Fortran.pdf.

Dongarra:1980:IAC

[6] J. J. Dongarra. Improving the Accuracy of Computed Matrix Eigenvalues. Ph.D. thesis, Department of Computer Science, University of New Mexico, Albuquerque, NM, USA,

1980. URL http://proquest.umi.
com/pqdweb?did=748354481&sid=4&
Fmt=1&clientId=9456&RQT=309&VName=
PQD.

Dongarra:1981:IAC

[7] J. J. Dongarra, C. B. Moler, and J. H. Wilkinson. Improving the accuracy of computed eigenvalues and eigenvectors. Technical Report ANL-81-43, Applied Mathematics Division, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1981. Published as [11, 12, 13].

Dongarra:1982:ASF

[8] Jack J. Dongarra. Algorithm 589: SICEDR: A FORTRAN subroutine for improving the accuracy of computed matrix eigenvalues. ACM Transactions on Mathematical Software, 8(4):371–375, December 1982. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/ALGORITHM-
589-SICEDR.pdf.

Dongarra:1982:EPH

[9] J. J. Dongarra, J. R. Gabriel, and J. H. Wilkinson. The eigenvalue problem for Hermitian matrices with time reversal symmetry. Technical Memorandum 3, Mathematics and Computer Science Division, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1982. Published as [20] where D. D. Koelling is also an author.

Dongarra:1983:CPL

[10] J. Dongarra and R. Hiromoto. A collection of parallel linear equations routines for the Denelcor HEP. ?? ANL/MCS-TM-15, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, September 1983.

Dongarra:1983:IACa

[11] J. J. Dongarra, C. B. Moler, and J. H. Wilkinson. Improving the accuracy of computed eigenvalues and eigenvectors. SIAM Journal on Numerical Analysis, 20(1):23–45, February 1983. CODEN SJNAEQ. ISSN 0036-1429 (print), 1095-7170 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/Improving-
the-Accuracy-of-Computed-Eigenvalues-
and-Eigenvectors.pdf.

Dongarra:1983:IACb

[12] J. J. Dongarra. Improving the accuracy of computed singular values. SIAM Journal on Scientific and Statistical Computing, 4(4):712–719, December 1983. CODEN SIJCD4. ISSN 0196-

5204. URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
Improving-the-Accuracy-of-Computed-
Singular-Values.pdf.

Dongarra:1983:IACc

[13] J. J. Dongarra, C. B. Moler, and J. H. Wilkinson. Improving the accuracy of computed eigenvalues and eigenvectors. ACM SIGNUM Newsletter, 20:23–45, 1983. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dongarra:1983:PVC

[14] J. J. Dongarra. Performance of various computers using standard linear equations software in a FORTRAN environment. ACM SIGARCH Computer Architecture News, 11(5):22–27, December 1983. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Dongarra:1983:RLA

[15] J. J. Dongarra. Redesigning linear algebra algorithms. Bull. Dir. Etud. Rech. C (France), Bulletin de la Direction des Etudes et Recherches, Serie C, C(1):

51–60 (or 51–59??), ???? 1983. CO-DEN EDBCAA. ISSN 0013-4511.

Chen:1984:MLA

[16] Steven S. Chen, Jack J. Dongarra, and Christopher C. Hsiung. Multiprocessing linear algebra algorithms on the Cray X-MP-2: Experiences with small granularity. Journal of Parallel and Distributed Computing, 1(1):22–31, August 1984. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/Multiprocessing-
Linear-Algebra-Algorithms-on-the-
CRAY.pdf.

Dongarra:1984:CPL

[17] Jack J. Dongarra and Robert E. Hiromoto. A collection of parallel linear equations routines for the Denelcor HEP. Parallel Computing, 1(2):133–142, December 1984. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/A-Collection-
of-Parallel-Linear-Equations-Routines-
for-the-Denelcor-HEP.pdf.

Dongarra:1984:DAL

[18] Jack J. Dongarra. Designing algorithms in linear algebra for different computer architectures. In Proceedings of the 23rd IEEE Conference on Decision & Control, December 12–14, 1984, Las Vegas Hilton, Las Vegas, Nevada, page 661. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984. CODEN PCDCDZ. ISSN 0191-2216. Three volumes. IEEE catalog number 84CH2093-3.

Dongarra:1984:DMS

[19] J. Dongarra and E. Grosse. Distribution of mathematical software via electronic mail. Technical Report MCS-TM-48, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1984.

Dongarra:1984:EPH

[20] J. J. Dongarra, J. R. Gabriel, D. D. Koelling, and J. H. Wilkinson. The eigenvalue problem for Hermitian matrices with time reversal symmetry. Linear Algebra and its Applications, 60(????):27–42, August 1984. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/The-
Eigenvalue-Problem-for-Hermitian-
Matrices.pdf.

Dongarra:1984:EPS

[21] J. J. Dongarra and C. B. Moler. EIS-PACK — A package for solving matrix eigenvalue problems. In Cowell [949], pages 68–87. ISBN 0-13-823501-

5. LCCN QA76.95 .S68 1984.

Dongarra:1984:ILA

[22] J. J. Dongarra, F. G. Gustavson, and A. Karp. Implementing linear algebra algorithms for dense matrices on a vector pipeline machine. SIAM Review, 26(1):91–112, January 1984. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/Algorithms-
for-Dense-Matrices-on-Vector-Pipeline-
Machine.pdf.

Dongarra:1984:IPM

[23] J. Dongarra. Increasing the performance of mathematical software through high-level modularity. In Glowinski and Lions [951], pages 239– 248. ISBN 0-444-87597-2. LCCN QA297 .I57 1983.

Dongarra:1984:LPS

[24] J. J. Dongarra and G. W. Stewart. LINPACK — A package for solving linear systems. In Cowell [949], pages 20–48. ISBN 0-13-823501-5. LCCN QA76.95 .S68 1984.

Dongarra:1984:MLA

[25] J. Dongarra, S. Chen, and S. Hsiung. Multiprocessing linear algebra algorithms on the CRAY X-MP2: Experiences with small granularity. Technical Report MCS-TM-24, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, February 1984. ???? pp.

Dongarra:1984:NDS

[26] J. Dongarra, E. Lusk, R. Overbeek, B. Smith, and D. Sorensen. New directions in software for advanced computer architectures. Technical Report MCS-TM-32, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, August 1984. ???? pp.

Dongarra:1984:PCO

[27] J. J. Dongarra. Performances comparèes de 80 ordinateurs sur des programmes Fortran. Technique et Science Informatiques, 3(5):355–360, 1984. CODEN TTSIDJ. ISSN 0752-4072,

0264-7419. URL http://www.netlib.
org/utk/people/JackDongarra/PAPERS/
Performances-Comparees-de-80-Ordinateurs-
sur-des-Programmes-Fortran.pdf.

Dongarra:1984:PES

[28] J. Dongarra, J. DuCroz, S. Hammarling, and R. Hanson. A proposal for an extended set of Fortran Basic Linear Algebra Subprograms. Mathematics and Computer Science Division Technical Memo MCS-TM-41, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, December 1984.

Dongarra:1984:PLA

[29] J. Dongarra and D. Sorensen. A parallel linear algebra library for the Denelcor HEP. Technical Report ANL/MCS/TM-33, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, ?? 1984.

Dongarra:1984:PVCa

[30] J. J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. Technical Memorandum 23, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1984. 8 pp.

Dongarra:1984:PVCb

[31] J. J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. ACM SIGNUM Newsletter, 19(1):23–26, January 1984. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/Performance-

                                        of-Various-Computers-in-Fortran-
                                        (Signum-News).pdf.

Dongarra:1984:PVCc

[32] J. J. Dongarra. Performance of various computers using standard linear equations software in a FOR-TRAN environment. Technique et Science Informatiques, 3(5):317– 320, 1984. CODEN TTSIDJ, TTSIEK. ISSN 0752-4072, 0264-

                                        7419. URL http://www.netlib.org/
                                        utk/people/JackDongarra/PAPERS/
                                        Performance-of-Various-Computers-
                                        Using-Linear-Equations-Software.
                                        pdf.

Dongarra:1984:SMA

[33] Jack J. Dongarra and Stanley C. Eisenstat. Squeezing the most out of an algorithm in CRAY FORTRAN. ACM Transactions on Mathematical Software, 10(3):219–230, September 1984. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

                                        URL http://www.netlib.org/utk/
                                        people/JackDongarra/PAPERS/Algorithm-
                                        in-CRAY-FORTRAN.pdf.

Dongarra:1984:SME

[34] J. Dongarra, L. Kaufman, and S. Hammarling. Squeezing the most out of eigenvalue solvers on high performance computers. Technical Report MCS-TM-46, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1984.

Dongarra:1984:SPBa

[35] J. Dongarra and A. Sameh. On some parallel banded system solvers. Technical Report ANL/MCS-TM-27, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, ?? 1984.

Dongarra:1984:SPBb

[36] Jack J. Dongarra and Ahmed H. Sameh. On some parallel banded system solvers. Parallel Computing, 1(3–4):223–235, December 1984. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/On-Parallel-
Banded-System-Solvers.pdf.

Dongarra:1984:SSE

[37] J. J. Dongarra, J. R. Gabriel, D. D. Koelling, and J. H. Wilkinson. Solving the secular equation including spin orbit coupling for systems with inversion and time-reversal symmetry. Journal of Computational Physics, 54(2):278–288, May 1984. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/Inversion-
and-Time-Reversal-Symmetry.pdf;
http://www.sciencedirect.com/science/
article/pii/0021999184901190.

Messina:1984:PFA

[38] P. Messina, B. Smith, and J. Dongarra. Proceedings form the Argonne workshop on programming the next generation of supercomputers. Technical Report MCS-TM-34, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1984.

Dongarra:1985:ADD

[39] J. J. Dongarra, B. T. Smith, and D. Sorensen. Algorithm design for different computer architectures.

IEEE Software, 2(4):79–80, July 1985. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/Algorithm-
Design-for-Different-Computer-Architectures.
pdf.

Dongarra:1985:BRB

[40] Jack Dongarra. Book review: A New Approach to Scientific Computation (Ulrich W. Kulisch and Willard L. Miranker, eds.). SIAM Review, 27(2):267–268, ???? 1985. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Dongarra:1985:CCX

[41] J. Dongarra and A. Hinds. Comparison of the CRAY X-MP-4, the Fujitsu VP-200, and the Hitachi S-810/20: An Argonne perspective. Technical Report ANL-8579, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, October 1985.

Dongarra:1985:DMS

[42] J. J. Dongarra and E. Grosse. Distribution of mathematical software via electronic mail. ACM SIGNUM Newsletter, 20(3):45–47, July 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dongarra:1985:FAS

[43] J. J. Dongarra and D. C. Sorensen. A fast algorithm for the symmetric eigenvalue problem. In Hwang [952], pages 338–342. ISBN 0-8186-0632-0 (paperback), 0-8186-8632-4 (hard), 0-8186-4632-2 (microfiche). LCCN QA76.9.C62 S95 1985. IEEE catalog number 85CH2146-9.

Dongarra:1985:IDL

[44] J. Dongarra and T. Hewitt. Implementing dense linear algebra algorithms using multitasking on the CRAY X-MP-

4. Technical Report MCS-TM-55, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, August 1985.

Dongarra:1985:ISC

[45] J. J. Dongarra, A. H. Sameh, and D. C. Sorensen. Implementation of some concurrent algorithms for matrix factorization. In Proceedings of the Eighteenth Hawaii International Conference on System Science, Honolulu, HI, USA, 2–4 January 1985, pages 39–46. Western Periodicals Co.,, North Hollywood, CA, 1985. CODEN PHISD7. ISBN ???? ISSN 0073-1129. LCCN ????

Dongarra:1985:PES

[46] J. J. Dongarra, J. Du Croz, S. Hammarling, and R. J. Hanson. A proposal for an extended set of FORTRAN basic linear algebra subprograms. ACM SIGNUM Newsletter, 20(1):2–18, January 1985. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (elec-

tronic). URL http://www.netlib.
org/utk/people/JackDongarra/PAPERS/
Proposal-for-Extended-Set-of-Fortran-
BLAS.pdf.

Dongarra:1985:PVCa

[47] J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. ?? MCA-TM-23, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, ??

Dongarra:1985:PVCc

[48] J. J. Dongarra. Performance of various computers using standard linear equations software in a FORTRAN environment. ACM SIGARCH Computer Architecture News, 13(1):3–11, March 1985. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Dongarra:1985:SIF

[49] J. J. Dongarra, A. H. Sameh, and D. C. Sorensen. Some implementations of the QR-factorization on an MIMD-machine. Parallel Computing, ??(??):??, October 1985. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Martin:1985:SSI

[50] J. L. Martin, R. Baron, R. Rashid, E. Siegel, A. Tevanian, M. Young, P. C. Patton, J. R. Gurd, J. Sargeant, K. Arvind, A. Gottlieb, G. C. Fox, P. A. Rigsbee, W. K. Giloi, T. Hoag, J. C. Browne, J. J. Dongarra, B. T. Smith, and D. C. Sorensen. Special section — international parallel processing projects — a software perspective. IEEE Software, 2(4):65–80, ???? 1985. CODEN IESOEG. ISSN 0740-7459 (print), 1937-4194 (electronic).

Dongarra:1986:CCX

[51] Jack J. Dongarra and Alan Hinds. Comparison of the CRAY X-MP-4, Fujitsu VP-200, and Hitachi S-810/20. Simulation, 47(3):93–107, September 1986. CODEN SIMUA2. ISSN 0037-5497 (print), 1741-3133 (electronic).

Dongarra:1986:FPA

[52] J. J. Dongarra and D. C. (Danny C.) Sorensen. A fully parallel algorithm for the symmetric eigenvalue problem. Technical Report CSRD-542, University of Illinois at Urbana-Champaign, Center for Supercomputing Research and Development, Urbana, IL 61801, USA, 1986. 22 pp. Also available as Argonne Report MCS-TM-62.

Dongarra:1986:HDM

[53] J. Dongarra. How do the ‘mini-supers’ stack up? Computer, 19(3):93, 100, March 1986. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (elec-

tronic). URL http://www.netlib.
org/utk/people/JackDongarra/PAPERS/
How-Do-the-Minisupers-Stack-Up.
pdf.

Dongarra:1986:HPC

[54] J. J. Dongarra and D. C. Sorensen. High performance computers and algorithms from linear algebra. In Cullum and Willoughby [954], pages 15– 36. ISBN 0-444-70074-9. LCCN QA193 .I261 1985.

Dongarra:1986:IDL

[55] Jack J. Dongarra and Tom Hewitt. Implementing dense linear algebra algorithms using multitasking on the CRAY X-MP-4 (or, Approaching the gigaflop). SIAM Journal on Scientific and Statistical Computing, 7(1):347–350, January 1986. CODEN SIJCD4. ISSN 0196-

5204. URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
Implementing-Dense-Linear-Algebra-
Algorithms-Using-Multitasking.pdf.

Dongarra:1986:ISC

[56] J. J. Dongarra, A. H. Sameh, and D. C. Sorensen. Implementation of some concurrent algorithms for matrix factorization. Parallel Computing, 3(1):25–34, March 1986. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

                                        URL http://www.netlib.org/utk/
                                        people/JackDongarra/PAPERS/Implementation-
                                        of-Some-Concurrent-Algorithms-for-
                                        Matrix-Factorization.pdf.

Dongarra:1986:LAHa

[57] J. J. Dongarra and D. C. Sorensen. Linear algebra on high-performance computers. In Feilmeier et al. [955], pages 3–32. ISBN 0-444-70009-9. LCCN QA76.6 .I5471 1985.

Dongarra:1986:LAHb

[58] J. J. Dongarra and D. C. Sorensen. Linear algebra on high performance computers. Applied Mathematics and Computation, 20(1–2):57–88, September 1986. CODEN AMHCBQ. ISSN 0096-3003 (print), 1873-5649 (electronic).

                                        URL http://www.netlib.org/utk/
                                        people/JackDongarra/PAPERS/Linear-
                                        Algebra-on-High-Performance-Computers.
                                        pdf.

Dongarra:1986:PLI

[59] J. J. Dongarra and D. C. Sorensen. Performance and library issues for mathematical software on high performance computers. In Wouk [956], pages 112–

135. ISBN 0-89871-201-7. LCCN

QA76.9.A73 N49 1986.

Dongarra:1986:PVC

[60] J. Dongarra and I. Duff. Performance of vector computers for direct and indirect addressing in Fortran. Harwell report, AERE Harwell Laboratory, Chilton, Oxon, England, ?? 1986.

Dongarra:1986:SBS

[61] J. J. Dongarra and L. Johnsson. Solving banded systems on a parallel processor. Technical Report MCS-TM-85, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, November 1986.

Dongarra:1986:SHP

[62] Jack J. Dongarra. A survey of high performance computers. In Bell [953], pages 8–11. CODEN PCICDQ. ISBN 0-8186-0692-4 (paperback), 0-8186-4692-6 (microfiche). LCCN QA75.5.C58 1986. IEEE catalog number 86CH2285-.

Dongarra:1986:SME

[63] Jack J. Dongarra, Linda Kaufman, and Sven Hammarling. Squeezing the most out of eigenvalue solvers on high performance computers. Linear Algebra and its Applications, 77(??):113–136, ?? 1986. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/Squeezing-
the-Most-out-of-Eigenvalue-Solvers.

pdf. Special volume on parallel computing.

Dongarra:1986:STD

[64] J. Dongarra and D. Sorensen. SCHED-ULE: Tools for developing and analyzing parallel Fortran programs. Technical Report ANL-MCS-TM-86, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, November 1986.

Dongarra:1986:UNE

[65] J. Dongarra, J. DuCroz, S. Hammarling, and R. Hanson. An update notice on the extended BLAS. ACM SIGNUM Newsletter, 21(4):2–4, 1986. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Astfalk:1987:FPD

[66] Greg Astfalk, Jack Dongarra, and Eric Grosse. Finding public domain mathematical software. Numerical Analysis Manuscript 87-5, AT&T Bell Laboratories, Murray Hill, NJ, USA, 1987.

Chatelin:1987:SVM

[67] F. Chatelin, J. Dongarra, and I. Duff. Special volume in memory of James H. Wilkinson. Linear Algebra and its Applications, 88/89:??, 1987. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Demmel:1987:PDL

[68] James Demmel, Jack Dongarra, Jeremy Du Croz, Anne Greenbaum, Sven Hammarling, and Danny Sorensen. Prospectus for the development of a linear algebra library for high-performance computers. LA-PACK Working Note and Mathematics and Computer Science Division Report 01 and ANL/MCS-TM-97, Mathematics and Computer Science Division, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, September

1987. URL http://www.netlib.org/
lapack/lawns/lawn01.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn01.pdf. ANL, MCS-TM-97, September 1987, and LAPACK Working Note #1.

Dongarra:1987:AAC

[69] J. Dongarra and I. Duff. Advanced architecture computers. Technical Report ANL-MCS-TM-57 (Revision 1), Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, October 1987.

Dongarra:1987:ACR

[70] J. J. Dongarra and E. L. Lusk. Advanced Computing Research Facility and algorithm design for different computers. In Ahmed K. Noor, editor, Parallel computations and their impact on mechanics / presented at the Winter Annual Meeting of the American Society of Mechanical Engineers, Boston, Massachusetts, December 13-18, 1987, volume 86 of AMD (Series), pages 49– 53. American Society of Mechanical Engineers, 345 E. 47th St., New York, NY 10017, USA, 1987. CODEN AMD-VAS. ISSN 0160-8835. LCCN QA76.6 .A43 1987.

Dongarra:1987:BRM

[71] Jack Dongarra, Sven Hammarling, and Danny Sorensen. Block reduction of matrices to condensed forms for eigenvalue computations. LA-PACK Working Note 02, Mathematics and Computer Science Division, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, September

1987. URL http://www.netlib.org/
lapack/lawns/lawn02.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn02.pdf. ANL, MCS-TM-99, September 1987.

Dongarra:1987:CBP

[72] Jack Dongarra, Joanne L. Martin, and Jack Worlton. Computer benchmarking: Paths and pitfalls. IEEE Spectrum, 24(7):38–43, July 1987. CODEN IEESAM. ISSN 0018-9235 (print), 1939-9340 (electronic).

                                        URL http://www.netlib.org/utk/
                                        people/JackDongarra/PAPERS/Computer-
                                        Benchmarking-Paths-and-Pitfalls.
                                        pdf.

Dongarra:1987:DMS

[73] Jack J. Dongarra and Eric Grosse. Distribution of mathematical software via electronic mail. Communications of the ACM, 30(5):403–407, May 1987. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

                                        URL http://www.netlib.org/utk/
                                        people/JackDongarra/PAPERS/Distribution-
                                        of-Mathematical-Software-via-Electronic-
                                        Mail.pdf.

Dongarra:1987:EPC

[74] J. J. Dongarra, editor. Experimental Parallel Computing Architectures. North-Holland, Amsterdam, The Netherlands, 1987. ISBN 0-444-70234-

2. xiii + 303 pp. LCCN QA76.5 .E985

1987.

Dongarra:1987:ESF

[75] J. J. Dongarra, J. Du Croz, S. Hammarling, and R. J. Hanson. An extended set of Fortran Basic Linear Algebra Subprograms: Model implementation and test programs. Technical Report MCS-TM-81, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, January 1987.

Dongarra:1987:FPA

[76] J. J. Dongarra and D. C. Sorensen. A fully parallel algorithm for the symmetric eigenvalue problem. SIAM Journal on Scientific and Statistical Computing, 8(2):S139–S154, March 1987. CODEN SIJCD4. ISSN 0196-

5204. URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
A-Fully-Parallel-Algorithm-for-
the-Symmetric-Eigenvalue-Problem.
pdf;   http://www.netlib.org/utk/
people/JackDongarra/PAPERS/siamdc.
ps.

Dongarra:1987:IFP

[77] J. J. Dongarra and D. C. Sorensen. On the implementation of a fully parallel algorithm for the symmetric eigenvalue problem. Proceedings of the SPIE — The International Society for Optical Engineering, 696:45–53, 1987. CODEN PSISDG. ISSN 0277-786X (print), 1996-756X (electronic).

Dongarra:1987:LAE

[78] J. J. Dongarra and D. C. Sorensen. A look at the evolution of mathematical software for dense matrix problems over the past fifteen years. In Anonymous [957], pages 203–216.

Dongarra:1987:PED

[79] J. J. Dongarra and D. C. Sorensen. A portable environment for developing parallel FORTRAN programs. Parallel Computing, 5(1–2):175–186, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/Portable-
Environment-for-Developing-Parallel-
FORTRAN-Programs.pdf.

Dongarra:1987:PSLa

[80] Jack Dongarra, Jeremy Du Croz, Iain Duff, and Sven Hammarling. A proposal for a set of Level 3 basic linear algebra subprograms. ACM SIGNUM Newsletter, 22(3):2–14, July 1987. CO-DEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dongarra:1987:PVCa

[81] J. J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. In Walter J. Karplus, editor, Multiprocessors and array processors: proceedings of the Third Conference on Multiprocessors and Array Processors: 14–16 January 1987, San Diego, California, pages 15–32. Society for Computer Simulation, San Diego, CA, USA, 1987. ISSN 0735-9276. LCCN QA76.5 .C61923.

Dongarra:1987:PVCb

[82] Jack J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. Simulation, 49(2):51–62, August 1987. CODEN SIMUA2. ISSN 0037-5497 (print), 1741-3133 (electronic).

Dongarra:1987:SBS

[83] Jack J. Dongarra and Lennart Johnsson. Solving banded systems on a parallel processor. Parallel Computing, 5(1–2):219–246, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/Solving-
Banded-Systems-on-a-Parallel-Processor.

pdf. Proceedings of the international conference on vector and parallel computing—issues in applied research and development (Loen, 1986).

Dongarra:1987:STD

[84] J. J. Dongarra and D. C. Sorensen. Schedule: A tool for developing and analyzing parallel Fortran programs. In Jamieson et al. [958], pages 363– 394 (??). ISBN 0-262-10036-3. LCCN QA76.6 .C42981 1987.

Dongarra:1987:SUG

[85] J. J. Dongarra and D. C. Sorensen. SCHEDULE Users Guide. Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, June 1987. ???? pp.

Dongarra:1987:WLB

[86] J. J. Dongarra. Workshop on the Level 3 BLAS. Technical Report MCS-TM-89, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, March 1987. ???? pp.

Golub:1987:JW

[87] Gene H. Golub, Miki Neumann, James W. Demmel, Paul Saylor, James M. Boyle, Iain Duff, and Jack Dongarra. James Wilkinson (1919– 1986). Annals of the History of Computing, 9(2):205–210, April/June 1987. CODEN AHCOE5. ISSN 0164-1239. From the introduction: “A series of lightly edited extracts from messages that were sent over various computer networks during the period October 5, 1986–February 13, 1987”.

Bischof:1988:LPC

[88] C. Bischof, J. Demmel, J. Dongarra, J. Du Croz, A. Greenbaum, S. Hammarling, and D. Sorensen. LAPACK provisional contents. Mathematics and Computer Science Division Report ANL-88-38, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, September 1988. (LAPACK Working Note #5).

Bischof:1988:PC

[89] C. Bischof, J. Demmel, J. Dongarra, J. Du Croz, A. Greenbaum, S. Hammarling, and D. Sorensen. Provisional contents. LAPACK Working Note 05, Mathematics and Computer Science Division, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, September

1988. URL http://www.netlib.org/
lapack/lawns/lawn05.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn05.pdf. ANL, MCS-TM-38, September 1988.

Brewer:1988:TAAa

[90] O. Brewer, J. Dongarra, and D. Sorensen. Tools to aid in the analysis of memory access patterns for FORTRAN programs. LAPACK Working Note 06, Mathematics and Computer Science Division, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, June

1988. URL http://www.netlib.org/
lapack/lawns/lawn06.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn06.pdf. ANL, MCS-TM-120 (or ANL-MCS-TM-119??), June 1988.

Brewer:1988:TAAb

[91] Orlie Brewer, Jack Dongarra, and Danny Sorensen. Tools to aid in the analysis of memory access patterns for FORTRAN programs. Parallel Computing, 9(1):25–35, December 1988. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/Tools-
to-Aid-Analysis-of-Memory-Access-
Patterns-for-FORTRAN-Programs.pdf.

Callahan:1988:VCTa

[92] D. Callahan, J. Dongarra, and D. Levine. Vectorizing compilers: a test suite and results. Technical Memorandum 109, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, March 1988. To appear: Proceedings of the Supercomputing ’88 Conference.

Callahan:1988:VCTb

[93] David Callahan, Jack Dongarra, and David Levine. Vectorizing compilers: a test suite and results. In IEEE [960], pages 98–105. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Dongarra:1988:ADH

[94] J. J. Dongarra and D. C. Sorensen. Algorithm design for high-performance computers. In Paul and Almasi [961], pages 83–97. ISBN 0-444-70371-3. LCCN QA76.5 .I147 1986.

Dongarra:1988:AES

[95] Jack J. Dongarra, Jeremy Du Croz, Sven Hammarling, and Richard J. Hanson. Algorithm 656: An extended set of Basic Linear Algebra Subprograms: Model implementation and test programs. ACM Transactions on Mathematical Software, 14(1): 18–32, March 1988. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.

acm.org/pubs/citations/journals/
toms/1988-14-1/p18-dongarra/;
http://www.netlib.org/utk/people/
JackDongarra/PAPERS/An-Extended-
Set-of-BLAS-Model-Implementation-
and-Test-Programs.pdf.

Dongarra:1988:CES

[96] Jack J. Dongarra, Jeremy Du Croz, Sven Hammarling, and Richard J. Hanson. Corrigenda: “An extended set of FORTRAN Basic Linear Algebra Subprograms”. ACM Transactions on Mathematical Software, 14(4): 399, December 1988. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [97].

Dongarra:1988:ESF

[97] Jack J. Dongarra, Jeremy Du Croz, Sven Hammarling, and Richard J. Hanson. An extended set of FOR-TRAN Basic Linear Algebra Subprograms. ACM Transactions on Mathematical Software, 14(1):1–17, March 1988. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.

acm.org/pubs/citations/journals/
toms/1988-14-1/p1-dongarra/;   http:
//www.netlib.org/utk/people/JackDongarra/
PAPERS/An-Extended-Set-of-FORTRAN-

Basic-Linear-Algebra-Subprograms.

pdf. See also [96].

Dongarra:1988:LAE

[98] J. J. Dongarra and D. C. Sorensen. A look at the evolution of mathematical software for dense matrix problems over the past fifteen years. In Schultz [962], pages 29–36. ISBN 0-387-96733-8. LCCN QA76.5.N79 1988.

Dongarra:1988:LBEa

[99] J. J. Dongarra. The LINPACK benchmark: an explanation. In Evaluating Supercomputers: Strategies for Exploiting, Evaluating and Benchmarking Computers with Advanced Architecture, pages 150–167. Unicom Seminars, Uxbridge, UK, ???? 1988. ISBN ???? LCCN ????

Dongarra:1988:LBEb

[100] J. J. Dongarra. The LINPACK benchmark: an explanation. In Houstis et al. [959], pages 456– 474. CODEN LNCSD9. ISBN 0-387-18991-2, 3-540-18991-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.297.

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/The-
LINPACK-Benchmark-An-Explanation.

pdf. The conference was organized and sponsored by the Computer Technology Institute (C.T.I.) of Greece.

Dongarra:1988:PMP

[101] Jack J. Dongarra, Danny C. Sorensen, Kathryn Connolly, and Jim Patterson. Programming methodology and performance issues for advanced computer architectures. Parallel Computing, 8(1–3):41–58, October 1988.

CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

                                        URL http://www.netlib.org/utk/
                                        people/JackDongarra/PAPERS/Programming-
                                        Methodology-and-Performance-Issues-
                                        for-Advanced-Computer.pdf.

Dongarra:1988:PVCa

[102] J. J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. Technical Report MCS-TM-23, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, January 1988. ???? pp.

Dongarra:1988:PVCb

[103] J. J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. ACM SIGARCH Computer Architecture News, 16(1):47–69, March 1988. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Dongarra:1988:SLB

[104] J. J. Dongarra, J. Du Croz, I. S. Duff, and S. Hammarling. A set of level 3 Basic Linear Algebra Subprograms. ?? ANL-MCS-TM-88, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, ?? 1988.

Anderson:1989:ITI

[105] E. Anderson and J. Dongarra. Installing and testing the initial release of LAPACK — Unix and non-Unix versions. LAPACK Working Note 10, Mathematics and Computer Science Division, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, May

1989. URL http://www.netlib.org/
lapack/lawns/lawn10.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn10.pdf. ANL, MCS-TM-130, May 1989.

Anderson:1989:LPL

[106] E. Anderson, Z. Bai, C. Bischof, J. Demmel, J. Dongarra, J. DuCroz, A. Greenbaum, S. Hammarling, A. McKenney, and D. Sorensen. LA-PACK: a portable linear algebra library for supercomputers. ??, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, ?? 1989.

Anderson:1989:RIR

[107] E. Anderson and J. Dongarra. Results from the initial release of LA-PACK. LAPACK Working Note and Computer Science Dept. Technical Report 16 and CS-89-89, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, November

1989. URL http://www.netlib.org/
lapack/lawns/lawn16.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn16.pdf. UT-CS-89-89, November 1989. (Replaced by LAWN 41 or 81!!). LAPACK Working Note #16.

Bischof:1989:LAL

[108] C. H. Bischof and J. J. Dongarra. A linear algebra library for highperformance computers. In Carey [963], pages 45–56. ISBN 0-471-92436-9. LCCN M89.E02452; QA76.6. Preprint MCS-P105-0989, Mathematics and Computer, September 1989.

Brewer:1989:GTA

[109] O. Brewer, J. Dongarra, and D. Sorensen. A graphics tool to aid in the generation of parallel FORTRAN programs. In Stanley Y. W. Su and George J. Knafl, editors, Proceedings: the thirteenth annual International Computer Software & Applications Conference, pages 89– 93. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN 0-8186-1964-3. LCCN QA76.6.C6295 1989. Computer Society order number 1964. IEEE catalog number 89CH2743-3.

Brewer:1989:GTD

[110] O. Brewer, J. Dongarra, D. Levine, and D. Sorensen. Graphics tools for developing high-performance algorithms. In Evans and Sutti [966], pages 39– 50. ISBN 0-85274-224-x. LCCN QA76.5.I5775 1988.

Browne:1989:GBP

[111] J. Browne, J. Dongarra, A. Karp, K. Kennedy, and D. Kuck. 1988 Gordon Bell Prize. IEEE Software, 6(3): 78–85, May 1989. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Demmel:1989:PDL

[112] J. Demmel, J. J. Dongarra, J. DuCroz, and A. Greenbaum. A project for developing a linear algebra library for highperformance computers. In Wright [967], pages 87–92. ISBN 0-444-87310-4. LCCN QA76.5 .I2775 1988.

Dongarra:1989:ACR

[113] Jack Dongarra. Advanced Computing Research Facility, Mathematics and Computer Science Division, Argonne National Laboratory. International Journal of Supercomputer Applications, 3(4):6–8, December 1989. CO-DEN IJSAE9. ISSN 0890-2720. URL

http://journals.sagepub.com/doi/
pdf/10.1177/109434208900300402.

Dongarra:1989:BRM

[114] Jack J. Dongarra, Danny C. Sorensen, and Sven J. Hammarling. Block reduction of matrices to condensed forms for eigenvalue computations. Journal of Computational and Applied Mathematics, 27(1–2):215–227, September 1989. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/Block-
Reduction-of-Matrices-to-Condensed-
Forms-for-Eigenvalue-Computations.

pdf. (LAPACK Working Note #2).

Dongarra:1989:PSL

[115] Jack J. Dongarra, Jeremy Du Croz, Iain S. Duff, and Sven Hammarling. A proposal for a set of Level 3 Basic Linear Algebra Subprograms. In Rodrigue [945], pages 40–44. ISBN 0-89871-228-9. LCCN QA76.5 .S515 1987.

Dongarra:1989:PVC

[116] J. J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. Technical Memorandum 23, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, June 4, 1989.

Dongarra:1989:SIP

[117] J. Dongarra and E. Lusk. Summer Institute in Parallel Computing: September 5–15, 1989. Technical Report MCS-TM-136, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, December 1989.

Dongarra:1989:SMS

[118] Jack Dongarra and Eric Grosse. Shopping for mathematical software electronically. IEEE Potentials, 8(1):37–38, February 1989. CODEN IEPTDF. ISSN 0278-6648 (print), 1558-1772 (electronic).

                                        URL http://www.netlib.org/utk/
                                        people/JackDongarra/PAPERS/Shopping-
                                        for-Mathematical-Software-Electronically.

pdf. condensed version of CACM pa-

per.

Dongarra:1989:TAD

[119] J. Dongarra, D. Sorensen, and O. Brewer. Tools to aid in the development of high-performance algorithms. In Cosnard et al. [964], pages 89–100. ISBN 0-444-87367-8. LCCN QA76.5 .I6191 1988.

Dongarra:1989:TMP

[120] J. J. Dongarra, D. C. Sorensen, and O. Brewer. Tools and methodology for programming parallel processors. In Wright [967], pages 125–137. ISBN 0-444-87310-4. LCCN QA76.5 .I2775 1988.

Dongarra:1989:UNL

[121] J. Dongarra, I. Duff, J. DuCroz, and S. Hammarling. An update notice on the level 3 BLAS. ACM SIGNUM Newsletter, 24(1):9–10, January 1989. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Duff:1989:PCA

[122] Iain S. Duff. Parallel computation at CERFACS. In Dongarra et al. [969], pages 66–67. ISBN 0-89871-262-9. LCCN QA76.58.S55 1989.

Fineberg:1989:TAD

[123] J. Dongarra, O. Brewer, S. Fineberg, and J. A. Kohl. A tool to aid in the design, implementation, and understanding of matrix algorithms for parallel processors. Technical Report MCS-P115-1189, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, 1989.

Greenbaum:1989:EQQ

[124] A. Greenbaum and J. Dongarra. Experiments with QR/QL methods for the symmetric tridiagonal eigenproblem. LAPACK Working Note and Computer Science Dept. Technical Report 17 and CS-89-92, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, November

1989. URL http://www.netlib.org/
lapack/lawns/lawn17.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn17.pdf. UT-CS-89-92, November 1989. LAPACK Working Note #17.

Anderson:1990:EBA

[125] E. Anderson and J. Dongarra. Evaluating block algorithm variants in LAPACK. LAPACK Working Note and Computer Science Dept. Technical Report 19 and CS-90-103, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, April

1990. URL http://www.netlib.org/

lapack/lawns/lawn19.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn19.pdf. LAPACK Working Note #19. UT-CS-90-103, April 1990.

Anderson:1990:IGL

[126] E. Anderson and J. Dongarra. Implementation guide for LAPACK. LA-PACK Working Note 18, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, April

1990. URL http://www.netlib.org/
lapack/lawns/lawn18.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn18.pdf. UT-CS-90-101, April 1990.

Anderson:1990:LPLa

[127] E. Anderson, Z. Bai, C. Bischof, J. Demmel, J. Dongarra, J. Du Croz, A. Greenbaum, S. Hammarling, A. McKenney, and D. Sorensen. LA-PACK: a portable linear algebra library for high-performance computers. LAPACK Working Note 20, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, May

1990. URL http://www.netlib.org/
lapack/lawns/lawn20.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn20.pdf. UT-CS-90-105, May 1990.

Anderson:1990:LPLb

[128] E. Anderson, Z. Bai, C. Bischof, J. Demmel, J. Dongarra, J. DuCroz, A. Greenbaum, S. Hammarling, A. McKenney, and D. Sorensen. LA-PACK: a portable linear algebra library for high-performance computers. In IEEE [970], pages 2–11. ISBN 0-8186-2056-0 (paperback) (IEEE Computer Society), 0-89791-412-0 (paperback) (ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Anderson:1990:PEL

[129] E. Anderson, C. Bischof, J. Demmel, J. Dongarra, J. Du Croz, S. Hammarling, and W. Kahan. Prospectus for an extension to LAPACK: a portable linear algebra library for highperformance computers. LAPACK Working Note and Computer Science Dept. Technical Report 26 and CS-90-118, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, November

1990. URL http://www.netlib.org/
lapack/lawns/lawn26.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn26.pdf. LAPACK Working Note #26. UT-CS-90-118, November 1990.

Angerson:1990:LPL

[130] E. Angerson, Z. Bai, J. Dongarra, A. Greenbaum, A. McKenney, J. Du Croz, S. Hammarling, J. Demmel, C. Bischof, and D. Sorensen. LA-PACK: a portable linear algebra library for high-performance computers. In IEEE [970], pages 2–11. ISBN 0-8186-2056-0 (paperback) (IEEE Computer Society), 0-89791-412-0 (paperback) (ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

Dongarra:1990:ASL

[131] Jack J. Dongarra, Jeremy Du Croz, Sven Hammarling, and Iain Duff.

Algorithm 679: a set of Level 3 Basic Linear Algebra Subprograms: Model implementation and test programs. ACM Transactions on Mathematical Software, 16(1):18– 28, March 1990. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        toms/1990-16-1/p18-dongarra/;
                                        http://www.netlib.org/utk/people/
                                        JackDongarra/PAPERS/ALGORITHM-679-
                                        A-Set-of-Level-3-BLAS.pdf.  See

also [144, 195, 260].

Dongarra:1990:CEE

[132] J. Dongarra, G. A. Geist, and C. Romine. Computing the eigenvalues and eigenvectors of a general matrix by reduction to tridiagonal form. Technical Report ORNL/TM-11669, Oak Ridge National Laboratory, Knoxville, TN, USA, ?? 1990. Published in [197].

Dongarra:1990:ENS

[133] Jack Dongarra and Sven Hammarling. Evolution of numerical software for dense linear algebra. In Cox and Hammarling [968], pages 297–327. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Dongarra:1990:FSC

[134] J. J. Dongarra, George A. Geist, and C. H. Romine. Fortran subroutines for computing the eigenvalues and eigenvectors of a general matrix by reduction to general tridiagonal form. Technical report CS-90-116, University of Tennessee, Computer Science Dept., Knoxville, TN, USA, October 1990. 13 pp.

Dongarra:1990:IRS

[135] J. Dongarra, P. Mayes, and G. Radicati. The IBM RISC System/ 6000 and linear algebra operations. LAPACK Working Note 28, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, December

1990. URL http://www.netlib.org/
lapack/lawns/lawn28.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn28.pdf. UT-CS-90-122, December 1990.

Dongarra:1990:LBE

[136] J. J. Dongarra. The LINPACK benchmark: an explanation. In Van der Steen [971], pages 1–21. ISBN 0-412-37860-4, 0-442-31198-2 (U.S.). LCCN QA76.9.E94 E93 1990; QA76.88.E93 1990.

Dongarra:1990:LBF

[137] J. Dongarra and S. Ostrouchov. LA-PACK block factorization algorithms on the Intel iPSC/860. LAPACK Working Note and Computer Science Dept. Technical Report 24 and CS-90-115, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, October

1990. URL http://www.netlib.org/
lapack/lawns/lawn24.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn24.pdf. LAPACK Working Note #24. UT-CS-90-115, October, 1990.

Dongarra:1990:NCC

[138] J. Dongarra, S. Hammarling, and J. H. Wilkinson. Numerical considerations in computing invariant subspaces. LAPACK Working Note and Computer Science Dept. Technical Report 25 and CS-90-117, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, October

                                        1990. URL http://www.netlib.org/
                                        lapack/lawns/lawn25.ps;   http:/
                                        /www.netlib.org/lapack/lawnspdf/

lawn25.pdf. LAPACK Working Note #25. UT-CS-90-117, October, 1990.

Dongarra:1990:PVCa

[139] J. J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. ?? CS-89-85, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, ?? 1990.

Dongarra:1990:PVCb

[140] Jack J. Dongarra. Performance of various computers using standard linear equations software. ACM SIGARCH Computer Architecture News, 18(1):17– 31, March 1990. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Dongarra:1990:SLB

[141] Jack J. Dongarra, Jeremy Du Croz, Sven Hammarling, and Iain Duff. A set of level 3 Basic Linear Algebra Subprograms. ACM Transactions on Mathematical Software, 16(1): 1–17, March 1990. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.

acm.org/pubs/citations/journals/
toms/1990-16-1/p1-dongarra/;   http:
//www.netlib.org/utk/people/JackDongarra/
PAPERS/Set-of-Level-3-Basic-Linear-
Algebra-Subprograms.pdf.

Dongarra:1990:SRG

[142] Jack Dongarra, Alan H. Karp, Ken Kennedy, and David Kuck. Special report: 1989 Gordon Bell Prize. IEEE Software, 7(3):100–104, May 1990. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

Dongarra:1990:TAD

[143] Jack Dongarra, Orlie Brewer, James Arthur Kohl, and Samuel Fineberg. A tool to aid in the design, implementation, and understanding of matrix algorithms for parallel processors. Journal of Parallel and Distributed Computing, 9(2):185–202, June 1990. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/Tool-
for-Design-Implementation-and-Understanding-
of-Matrix-Algorithms.pdf.

Higham:1990:EFM

[144] Nicholas J. Higham. Exploiting fast matrix multiplication within the level 3 BLAS. ACM Transactions on Mathematical Software, 16(4):352–368, December 1990. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.

acm.org/pubs/citations/journals/
toms/1990-16-4/p352-higham/. De-

scribes algorithms based on Strassen’s method which are asymptotically faster than the standard N^{3} algorithm, and in practice, faster for N ≈ 100, and examines their numerical stability. See [131, 195, 260].

Schreiber:1990:ABN

[145] Robert Schreiber and Jack Dongarra. Automatic blocking of nested loops. Technical report CS-90-108, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, May 1990. URL http://www.

                                        netlib.org/utk/papers/autoblock/
                                        paper.html;   http://www.netlib.
                                        org/utk/papers/autoblock/paper.
                                        ps;   http://www.netlib.org/utk/
                                        people/JackDongarra/pdf/autoblock.
                                        pdf.

Anderson:1991:GFA

[146] E. Anderson, Z. Bai, and J. Dongarra. Generalized QR factorization and its applications. LA-PACK Working Note 31, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, April

                                        1991. URL http://www.netlib.org/
                                        lapack/lawns/lawn31.ps;   http:/
                                        /www.netlib.org/lapack/lawnspdf/

lawn31.pdf. UT-CS-91-131, April 1991.

Anderson:1991:IGL

[147] E. Anderson, J. Dongarra, and S. Ostrouchov. Implementation guide for LAPACK. LAPACK Working Note and Computer Science Dept. Technical Report 35 and CS-91-138, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, August

                                        1991. URL http://www.netlib.org/
                                        lapack/lawns/lawn35.ps;   http:/
                                        /www.netlib.org/lapack/lawnspdf/

lawn35.pdf. LAPACK Working Note #35. UT-CS-91-138, August 1991.

Anderson:1991:SDM

[148] E. Anderson, A. Benzoni, J. Dongarra, S. Moulton, S. Ostrouchov, B. Tourancheau, and R. van de Geijn. Basic linear algebra communications subprograms. In Stout and Wolfe [978], pages 287– 290. ISBN 0-8186-2290-3 (paperback), 0-8186-2291-1 (fiche). LCCN QA76.5 .D58 1991.

Beguelin:1991:GDT

[149] A. Beguelin, J. J. Dongarra, G. A. Geist, R. Manchek, and V. S. Sunderam. Graphical development tools for network-based concurrent supercomputing. In IEEE [977], pages 435– 444. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Beguelin:1991:HNC

[150] A. Beguelin, J. J. Dongarra, G. A. Geist, R. Manchek, and V. S. Sunderam. Heterogeneous network computing. In IEE [976], pages 94–99. ISBN 0-85296-519-2. ISSN 0537-9989. LCCN QA76.58.I567 1991.

Beguelin:1991:HNS

[151] A. Beguelin, J. Dongarra, A. Geist, R. Manchek, and V. S. Sunderam. Heterogeneous network supercomputing. Supercomputing Review, ??(??):??, August 1991. CODEN SURVEG. ISSN 1048-6836.

Beguelin:1991:HUG

[152] A. Beguelin, J. Dongarra, A. Geist, R. Manchek, K. Moore, and R. Wade. HeNCE: a user’s guide (Draft). Technical report, Oak Ridge National Laboratory, Knoxville, TN, USA, November 1991.

Beguelin:1991:ODH

[153] A. Beguelin, J. Dongarra, A. Geist, R. Manchek, and V. S. Sunderam. Opening the door to heterogeneous network supercomputing. Supercomputing Review, ??(??):44–45, September 1991. CODEN SURVEG. ISSN 1048-6836.

Beguelin:1991:SCG

[154] A. Beguelin, J. Dongarra, A. Geist, R. Manchek, and V. S. Sunderam. Solving computational grand challenges using a network of heterogeneous supercomputers. In Dongarra et al. [980], page ?? ISBN 0-89871-303-X. LCCN QA76.58 .P76 1992.

Beguelin:1991:UGP

[155] A. Beguelin, J. J. Dongarra, A. Geist, R. Mancheck, and V. Sunderam. A users’ guide to PVM (Parallel Virtual Machine). Technical Report ORNL/TM-11826, Mathematical Sciences Section, Oak Ridge National Laboratory, Knoxville, TN, USA, July 1991.

Beguelin:1991:WSO

[156] Adam Beguelin, Jack Dongarra, Al Geist, Robert Manchek, and Vaidy Sunderam. Workstation solutions: Opening the door to heterogeneous network supercomputing. Supercomputing Review, 4(9):44–45, September 1991. CODEN SURVEG. ISSN 1048-

6836. URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
Opening-the-Door-to-Heterogeneous-
Network-Supercomputing.pdf.

Benguelin:1991:HNC

[157] A. Benguelin, J. J. Dongarra, G. A. Geist, R. Manchek, and V. S. Sunderam. Heterogeneous network computing. In Second International Specialist Seminar on the Design and Application of Parallel Digital Processors: 15–19 April 1991, venue, the Gulbenkian Foundation, Lisbon, Portugal, number 334 in IEE conference publication, page ?? IEE, London, UK, 1991. ISBN 0-85296-519-2. ISSN 0537-9989. LCCN QA76.58 .I567 1991.

Demmel:1991:DPHa

[158] James Demmel, Jack Dongarra, and W. Kahan. On designing portable high performance numerical libraries. LAPACK Working Note 39, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, July

1991. URL http://www.netlib.org/
lapack/lawns/lawn39.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn39.pdf. UT-CS-91-141, July, 1991.

Demmel:1991:DPHb

[159] J. Demmel, J. Dongarra, and W. Kahan. On designing portable high performance numerical libraries. In Griffiths and Watson [975], page ?? ISBN 0-582-08908-5. LCCN QA297 .D85 1991.

Dongarra:1991:BHP

[160] J. Dongarra and W. Gentzsch. Benchmarking of high-performance computers. Parallel Computing, 17(10–11): 1067–1069, ???? 1991. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dongarra:1991:GBP

[161] J. J. Dongarra, A. Karp, K. Miura, and H. D. Simon. Gordon Bell Prize lectures (supercomputer applications). In IEEE [977], pages 328– 337. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894

                                        1991. URL http://www.netlib.org/
                                        utk/people/JackDongarra/PAPERS/
                                        Gordon-Bell-Prize-Lectures.pdf.

ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Dongarra:1991:IRS

[162] J. J. Dongarra, P. Mayes, and G. Radicati di Brozolo. The IBM RISC System/6000 and linear algebra operations. Supercomputer, 8(4):15–30, July 1991. CO-DEN SPCOEL. ISSN 0168-7875.

                                        URL http://www.netlib.org/utk/
                                        people/JackDongarra/PAPERS/The-
                                        IBM-RISC-System-6000-and-Linear-
                                        Algebra-Operations.pdf.

Dongarra:1991:LPHa

[163] J. Dongarra. LAPACK: a portable high performance numerical library for linear algebra. In Anonymous [974], pages 73–76. ISBN 4-87378-284-8. LCCN QA76.88 .I587 1991.

Dongarra:1991:LPHb

[164] J. Dongarra and J. Demmel. LAPACK: a portable high-performance numerical library for linear algebra. Supercomputer, 8(6):33–38, November 1991. CO-DEN SPCOEL. ISSN 0168-7875.

Dongarra:1991:LWNa

[165] J. Dongarra. LAPACK Working Note 34: Workshop on the BLACS. Computer Science Dept. Technical Report CS-91-134, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, May 1991.

Dongarra:1991:LWNb

[166] Jack J. Dongarra and Robert A. van de Geijn. LAPACK working note 37: Two dimensional basic linear algebra communications subprograms. Technical report, Oak Ridge National Laboratory, Knoxville, TN, USA, October 28, 1991. Proposed Standard.

Dongarra:1991:NNA

[167] Jack J. Dongarra and Bill Rosener. NA-NET: Numerical analysis NET. Technical Report CS-91-146, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, September 1991. 21 pp.

Dongarra:1991:PANa

[168] J. J. Dongarra and M. Sidani. A parallel algorithm for the nonsymmetric eigenvalue problem. Report ORNL/TM-12003, Mathematical Sciences Section, Oak Ridge National Laboratory, Knoxville, TN, USA, ?? 1991.

Dongarra:1991:PANb

[169] J. Dongarra and M. Sidani. A parallel algorithm for the non-symmetric eigenvalue problem. Computer Science Dept. Technical Report CS-91-137, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, 1991.

Dongarra:1991:PLT

[170] J. Dongarra, M. Furtney, S. Reinhardt, and J. Russell. Parallel loops — a test suite for parallelizing compilers: description and example results. Parallel Computing, 17(10–11):1247–1255, December 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/Parallel-
Loops-A-Test-Suite-for-Parallelizing-
Compilers.pdf.

Dongarra:1991:PVC

[171] J. Dongarra. Performance of various computers using standard linear equations software in a Fortran environment. ?? CS-89-85, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, ?? 1991.

Dongarra:1991:RCF

[172] J. J. Dongarra and R. A. van de Geijn. Reduction to condensed form for the eigenvalue problem on distributed memory architectures. LA-PACK Working Note and Computer Science Dept. Technical Report 30 and CS-91-130 and ORNL/TM-12006, Department of Computer Science, University of Tennessee, Knoxville and Mathematical Sciences Section, Oak Ridge National Laboratory, Knoxville, TN 37996, USA and Knoxville, TN, USA, April

1991. URL http://www.netlib.org/
lapack/lawns/lawn30.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn30.pdf. LAPACK Working Note #30, to appear in Parallel Computing. UT-CS-91-130, April 1991.

Dongarra:1991:SLS

[173] Jack J. Dongarra, Iain S. Duff, Danny C. Sorensen, and Henk A. van der Vorst. Solving Linear Systems on Vector and Shared Memory Computers. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1991. ISBN 0-89871-270-X. x + 256 pp. LCCN QA184 .S65 1991.

Dongarra:1991:SRG

[174] Jack Dongarra, Alan Karp, Ken Miura, and Horst Simon. Special report: 1990 Gordon Bell Prize. IEEE Software, 8(3):92–98, May 1991. CODEN IESOEG. ISSN 0740-7459 (print), 0740-7459 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/1990-
Gordon-Bell-Prize-Winners.pdf.

Dongarra:1991:TDB

[175] Jack J. Dongarra and Robert A. van de Geijn. Two dimensional Basic Linear Algebra Communication Subprograms. LAPACK Working Note and Computer Science Dept. Technical Report 37 and CS-91-138, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, October

1991. URL http://www.netlib.org/
lapack/lawns/lawn37.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn37.pdf. LAPACK Working Note #37. UT-CS-91-138, October, 1991.

Dongarra:1991:UGP

[176] Jack Dongarra et al. A Users’ Guide to PVM Parallel Virtual Machine. Oak Ridge National Laboratory, Knoxville, TN, USA, July 1991. ?? pp.

Dongarra:1991:WB

[177] J. J. Dongarra. Workshop on the BLACS. LAPACK Working Note and Computer Science Dept. Technical Report 34 and CS-91-134, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, May

                                        1991. URL http://www.netlib.org/
                                        lapack/lawns/lawn34.ps;   http:/
                                        /www.netlib.org/lapack/lawnspdf/

lawn34.pdf. LAPACK Working Note #34. UT-CS-91-134, May 1991.

Levine:1991:CSAa

[178] D. Levine, D. Callahan, and J. Dongarra. A comparative study of automatic vectorizing compilers. ?? MCS-P218-0391, Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, ?? 1991.

Levine:1991:CSAb

[179] D. Levine, D. Callahan, and J. Dongarra. A comparative study of automatic vectorizing compilers. Parallel Computing, 17(10–11):1223–1244, December 1991. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.netlib.
                                        org/utk/people/JackDongarra/PAPERS/
                                        Comparative-Study-of-Automatic-
                                        Vectorizing-Compilers.pdf.

Anderson:1992:GFA

[180] E. Anderson, Z. Bai, and J. Dongarra. Generalized QR factorization and its applications. Linear Algebra and its Applications, 162/164:243–271, 1992. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/Generalized-
QR-Factorization-and-Its-Applications.

pdf. Directions in matrix theory (Auburn, AL, 1990).

Anderson:1992:LDM

[181] E. Anderson, A. Benzoni, J. Dongarra, S. Moulton, S. Ostrouchov, B. Tourancheau, and R. van de Geijn. LAPACK for distributed memory architectures: progress report. In Dongarra et al. [980], pages 625–630. ISBN 0-89871-303-X. LCCN QA76.58 .P76 1992.

Anderson:1992:LUG

[182] E. Anderson, Z. Bai, C. Bischof, J. Demmel, J. Dongarra, J. Du Croz, A. Greenbaum, S. Hammarling, A. McKenney, S. Ostrouchov, and D. Sorensen. LAPACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1992. ISBN 0-89871-294-7. xv + 235 pp. LCCN QA76.73.F25 L36 1992.

Anderson:1992:PLP

[183] Edward Anderson and Jack Dongarra. Performance of LAPACK: a portable library of numerical linear algebra routines. LAPACK Working Note 44, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, May

1992. URL http://www.netlib.org/
lapack/lawns/lawn44.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn44.pdf. UT-CS-92-156, May 1992.

Arioli:1992:TAB

[184] Mario Arioli, Iain S. Duff, Daniel Ruiz, and Miloud Sadkane. Techniques for accelerating the block Cimmino method. In J. Dongarra, K. Kennedy, P. Messina, D. C. Sorensen, and R. G. Voigt, editors, Proceedings of Fifth SIAM Conference on Parallel Processing for Scientific Computing, pages 98– 104. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1992. Also reprinted in Proceedings of One-day Workshop on Parallel Numerical Analysis, 21 June 1991. Editors D B Duncan, K I M Mc Kinnon, and F Plab. Report EPCC-TR92-05, Edinburgh Parallel Computing Centre, 1-7.

Beguelin:1992:GDT

[185] Adam Beguelin, Jack J. Dongarra, et al. Graphical development tools for network-based concurrent supercomputing. HeNCE document, 1992.

Beguelin:1992:HGD

[186] Adam Beguelin, Jack J. Dongarra, Al Geist, Robert Manchek, Keith Moore, Reed Wade, and Vaidy Sunderam. HeNCE: Graphical development tools for network-based concurrent computing. In IEEE [982], pages 129–136. ISBN 0-8186-2775-1. LCCN QA76.76.A65 S33 1992. IEEE catalog number 92TH0432-5.

Beguelin:1992:HUG

[187] Adam Beguelin, Jack J. Dongarra, et al. HeNCE: a Users’ Guide Version 1.2. Oak Ridge National Laboratory, Knoxville, TN, USA, February 1992. ?? pp.

Beguelin:1992:PHT

[188] A. Beguelin, J. Dongarra, A. Geist, R. Manchek, and V. Sunderam. PVM and HeNCE: traversing the parallel environment. CRAY Channels, 14(4):22– 25, Fall 1992. CODEN CRCHE8.

Beguelin:1992:SCG

[189] A. Beguelin, J. Dongarra, A. Geist, R. Manchek, and V. Sunderam. Solving computational grand challenges using a network of heterogeneous supercomputers. In Dongarra et al. [980], pages 596–601. ISBN 0-89871-303-X. LCCN QA76.58 .P76 1992.

Blackford:1992:IGL

[190] Susan Blackford and Jack Dongarra. Installation guide for LA-PACK. LAPACK Working Note 41, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, March

1992. URL http://www.netlib.org/
lapack/lawns/lawn41.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn41.pdf. UT-CS-92-151, March, 1992.

Choi:1992:DDL

[191] J. Choi, J. J. Dongarra, and D. W. Walker. The design of distributed level 3 BLAS routines. ??, 1992. in preparation.

Choi:1992:SAS

[192] J. Choi, J. J. Dongarra, R. Pozo, and D. W. Walker. ScaLAPACK: a scalable linear algebra library for distributed memory concurrent computers. In Siegel [984], pages 120–127. ISBN 0-8186-2772-7 (hardback), 0-8186-2771-9 (microfiche). LCCN QA76.58 .S95 1992.

Choi:1992:SSLa

[193] Jaeyoung Choi, Jack J. Dongarra, Roldan Pozo, and David W. Walker. ScaLAPACK: a scalable linear algebra for distributed memory concurrent computers. LAPACK Working Note 55, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, November 1992. URL http://www.

netlib.org/lapack/lawns/lawn55.
ps;   http://www.netlib.org/lapack/
lawnspdf/lawn55.pdf. UT-CS-92-

181, November 1992.

Choi:1992:SSLb

[194] J. Choi, J. J. Dongarra, R. Pozo, and D. W. Walker. ScaLAPACK: a scalable linear algebra library for distributed memory concurrent computers. In Siegel [984], pages 120–127. ISBN 0-8186-2772-7 (hardback), 0-8186-2771-9 (microfiche). LCCN QA76.58 .S95 1992. IEEE catalog number 92CH3185-6.

Demmel:1992:SBA

[195] James W. Demmel and Nicholas J. Higham. Stability of block algorithms with fast level-3 BLAS. ACM Transactions on Mathematical Software, 18(3):274–291, September 1992. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.

acm.org/pubs/citations/journals/
toms/1992-18-3/p274-demmel/. See

[131, 144, 260].

Dongarra:1992:AAC

[196] Jack J. Dongarra and Iain S. Duff. Advanced architecture computers. In H. Adeli, editor, Supercomputing in Engineering Analysis, pages 19–62. Marcel Dekker, New York, NY, USA, 1992.

Dongarra:1992:AFS

[197] J. J. Dongarra, G. A. Geist, and C. H. Romine. Algorithm 710: FOR-TRAN subroutines for computing the eigenvalues and eigenvectors of a general matrix by reduction to general tridiagonal form. ACM Transactions on Mathematical Software, 18(4):392– 400, December 1992. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.

acm.org/pubs/citations/journals/
toms/1992-18-4/p392-dongarra/;
http://www.netlib.org/utk/people/
JackDongarra/PAPERS/ALGORITHM-710-
FORTRAN-Subroutines-for-Computing-
Eigenvalues.pdf.

Dongarra:1987:BRC

[198] Jack Dongarra. Book reviews: The Connection Machine. International Journal of Supercomputer Applications, 1(1):112, March 1987. CO-DEN IJSAE9. ISSN 0890-2720. URL

http://journals.sagepub.com/doi/
pdf/10.1177/109434208700100110.

Dongarra:1992:E

[199] Jack Dongarra. Editorial. International Journal of Supercomputer Applications, 6(4):313, Winter 1992. CO-DEN IJSAE9. ISSN 0890-2720.

Dongarra:1992:LASa

[200] J. J. Dongarra, R. van de Geijn, and D. W. Walker. A look at scalable dense linear algebra libraries. Technical Report ORNL/TM-12126, Mathematical Sciences Section, Oak Ridge National Laboratory, Knoxville, TN, USA, ?? 1992.

Dongarra:1992:LASb

[201] J. Dongarra, R. van de Geijn, and D. Walker. A look at scalable dense linear algebra libraries. In IEEE [982], pages 372–379. ISBN 0-8186-2775-1. LCCN QA76.76.A65 S33 1992. IEEE catalog number 92TH0432-5.

Dongarra:1992:LNA

[202] Jack Dongarra. LAPACK is now available. ACM SIGNUM Newsletter, 27(1):3–4, January 1992. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dongarra:1992:LSD

[203] Jack Dongarra, Robert van de Geijn, and David Walker. A look at scalable dense linear algebra libraries. LAPACK Working Note 43, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, April

                                        1992. URL http://www.netlib.org/
                                        lapack/lawns/lawn43.ps;   http:/
                                        /www.netlib.org/lapack/lawnspdf/

lawn43.pdf. UT-CS-92-155, April, 1992.

Dongarra:1992:LWN

[204] Jack J. Dongarra. LAPACK working note 34: Workshop on the BLACS. Oak Ridge National Laboratory, February 6, 1992.

Dongarra:1992:NCC

[205] Jack J. Dongarra, Sven Hammarling, and James H. Wilkinson. Numerical considerations in computing invariant subspaces. SIAM Journal on Matrix Analysis and Applications, 13(1):145–161, January 1992. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/Numerical-
Considerations-in-Computing-Invariant-
Subspaces.pdf.

Dongarra:1992:PAN

[206] J. J. Dongarra and M. Sidani. A parallel algorithm for the nonsymmetric eigenvalue problem. In Griffiths and Watson [981], pages 85–102. ISBN 0-582-08908-5. LCCN QA297.D85 1991.

Dongarra:1992:PUL

[207] Jack J. Dongarra, Rolf Hempel, Anthony J. G. Hey, and David W. Walker. A proposal for a user-level messagepassing interface in a distributed memory environment. Technical Report TM-12231, Oak Ridge National Laboratory, Knoxville, TN, USA, October 1992.

Dongarra:1992:PVCa

[208] Jack J. Dongarra and Henk A. van der Vorst. Performance of various computers using standard techniques for solving sparse linear equations. Technical Report CS-92-168, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, 1992. ???? pp.

Dongarra:1992:PVCb

[209] Jack J. Dongarra and Henk A. van der Vorst. Performance of various computers using standard sparse linear equations solving techniques. Supercomputer, 9(5):17–29, September 1992. CODEN SPCOEL. ISSN 0168-

7875. URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
Performance-of-Various-Computers-
Using-Standard-Techniques.pdf.

Dongarra:1992:PVCc

[210] Jack J. Dongarra. Performance of various computers using standard linear equations software. ACM SIGARCH Computer Architecture News, 20(3):22– 44, June 1992. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Dongarra:1992:RCFa

[211] J. J. Dongarra and M. Sidani. Reduction to condensed form for the eigenvalue problem on distributed memory architectures. Report ORNL/TM-12006, Mathematical Sciences Section, Oak Ridge National Laboratory, Knoxville, TN, USA, ?? 1992.

Dongarra:1992:RCFb

[212] Jack J. Dongarra and Robert A. van de Geijn. Reduction to condensed form for the eigenvalue problem on distributed memory architectures. Parallel Computing, 18(9):973–982, September 1992. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/Reduction-
to-Condensed-Form-for-the-Eigenvalue-
Problem-on-Distributed-Memory.pdf.

Dongarra:1992:TAD

[213] J. Dongarra, D. Sorensen, and O. Brewer. Tools to aid in the design, implementation, and understanding of algorithms for parallel processors. In Perrott [983], pages 195–219. ISBN 0-412-39960-1. LCCN QA76.58.S63 1992.

Pancake:1992:WSW

[214] C. M. Pancake, R. Bailey, D. Barkai, B. Smith, J. Brandenburg, J. Dongarra, M. Kalos, M. Snir, and D. Gelernter. What should we expect from parallel language standards? The International Journal of Supercomputer Applications and High Performance Computing, 6(1):112–117, ???? 1992. CODEN IJSCFG. ISSN 1078-3482.

Anderson:1993:PLP

[215] E. C. Anderson and J. Dongarra. Performance of LAPACK: a portable library of numerical linear algebra routines. Proceedings of the IEEE, 81(8):1094–1102, August 1993. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/Performance-
of-LAPACK-A-Portable-Library.pdf.

Barrett:1993:BBI

[216] R. Barrett, T. Chan, J. Demmel, J. Donato, J. Dongarra, V. Eijkhout, V. Pozo, Romime C., and H. van der Vorst. Building blocks for iterative solution of linear systems. in preparation, 1993.

Beguelin:1993:PEC

[217] A. Beguelin, J. Dongarra, A. Geist, R. Manchek, Otto, S., and J. Walpole. PVM: Experiences, current status and future direction. In IEEE [989], pages 765–766. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Beguelin:1993:PHT

[218] A. Beguelin, J. Dongarra, A. Geist, R. Manchek, K. Moore, and V. Sunderam. PVM and HeNCE: Tools for heterogeneous network computing. In Kowalik and Grandinetti [990], page ?? ISBN 3-540-56451-9 (Berlin), 0-387-56451-9 (New York). LCCN QA76.58 .S629 1993.

Beguelin:1993:THN

[219] A. Beguelin, J. Dongarra, A. Geist, and R. Manchek. Tools for heterogeneous network computing. In Sincovec [991], pages 854–861. ISBN 0-89871-315-3. LCCN QA 76.58 S55 1993.

URL http://www.netlib.org/utk/
papers/pvm4/pvm4.html;   http:/
/www.netlib.org/utk/papers/pvm4/
pvm4.ps;   http://www.netlib.org/
utk/people/JackDongarra/pdf/pvm4.

pdf. Two volumes.

Beguelin:1993:VDH

[220] Adam Beguelin, Jack Dongarra, Al Geist, and Vaidy Sunderam. Visualization and debugging in a heterogeneous environment. Computer, 26 (6):88–95, June 1993. CODEN CP-TRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://www.

netlib.org/utk/people/JackDongarra/
PAPERS/Visualization-and-Debugging-
in-a-Heterogeneous-Environment.
pdf.

Berry:1993:PPD

[221] Michael W. Berry, Jack J. Dongarra, and Brian H. Larose. PDS: A Performance Database Server. Scientific Computing, ??(??):??, ?? 1993. CODEN SCHRCU. ISSN 1930-5753 (print), 1930-6156 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/PDS-
A-Performance-Database-Server.pdf.

(to appear).

Choi:1993:PMT

[222] Jaeyoung Choi, Jack J. Dongarra, and David W. Walker. Parallel matrix transpose algorithms on distributed memory concurrent computers. LAPACK Working Note 65, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, November

1993. URL http://www.netlib.org/
lapack/lawns/lawn65.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn65.pdf. UT-CS-93-215, November, 1993.

Choi:1993:PPU

[223] Jaeyoung Choi, Jack J. Dongarra, and David W. Walker. PUMMA: Parallel Universal Matrix Multiplication Algorithms on distributed memory concurrent computers. LA-PACK Working Note 57, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, May

1993. URL http://www.netlib.org/
lapack/lawns/lawn57.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn57.pdf. UT-CS-93-187, May

Demmel:1993:LDM

[224] J. Demmel, J. Dongarra, R. Van de Geijn, and D. Walker. LAPACK for distributed memory architectures: The next generation. In Sincovec [991], pages 323–329. ISBN 0-89871-315-3. LCCN QA 76.58 S55 1993. Two volumes.

Desprez:1993:PCF

[225] F. Desprez, J. Dongarra, and B. Tourancheau. Performance complexity of LU factorization with efficient pipelining and overlap on a multiprocessor. LAPACK Working Note 67, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, December

                                        1993. URL http://www.netlib.org/
                                        lapack/lawns/lawn67.ps;   http:/
                                        /www.netlib.org/lapack/lawnspdf/

lawn67.pdf. UT-CS-93-218, December, 1993.

Dongarra:1993:DLA

[226] Jack J. Dongarra and David W. Walker. The design of linear algebra libraries for high performance computers. LAPACK Working Note and Computer Science Dept. Technical Report 58 and ORNL/TM-12404, Department of Computer Science, University of Tennessee, Knoxville and Mathematical Sciences Section, Oak Ridge National Laboratory, Knoxville, TN 37996, USA and Knoxville, TN, USA, June

                                        1993. URL http://www.netlib.org/
                                        lapack/lawns/lawn58.ps;   http:/
                                        /www.netlib.org/lapack/lawnspdf/

lawn58.pdf. UT-CS-93-188, June 1993.

Dongarra:1993:DR

[227] J. J. Dongarra. From dinos to rhinos. In Fincham and Ford [987], pages 1–10. ISBN 0-19-853680-1. ISSN 0960-2526. LCCN QA76.58.P3755 1993.

Dongarra:1993:DSM

[228] J. J. Dongarra, R. Hempel, A. J. G. Hey, and D. W. Walker. A draft standard for message passing in a distributed memory environment. In Hoffmann and Kauranne [988], pages 465–

481. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

Dongarra:1993:IPF

[229] Jack Dongarra, G. A. Geist, Robert Manchek, and V. S. Sunderam. Integrated PVM framework supports heterogeneous network computing. Computers in Physics, 7(2):166–174, March–April 1993. CODEN CPHYE2. ISSN 0894-1866 (print), 1558-4208 (electronic).

URL http://www.netlib.org/utk/
papers/comp-phy7/comp-phy7.html;
http://www.netlib.org/utk/papers/
comp-phy7/comp-phy7.ps;   http:/
/www.netlib.org/utk/people/JackDongarra/
pdf/comp-phy7.pdf;   https://aip.
scitation.org/doi/10.1063/1.4823162.

Dongarra:1993:LAL

[230] J. Dongarra. Linear algebra libraries for high-performance computers: a personal perspective. IEEE parallel and distributed technology: systems and applications, 1(1):17–24, February 1993. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/Linear-

Algebra-Libraries-for-High-Performance-
Computers.pdf.

Dongarra:1993:LDO

[231] Jack J. Dongarra, Roldan Pozo, and David W. Walker. LAPACK++: a design overview of object-oriented extensions for high performance linear algebra. In IEEE [989], pages 162–

171. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Dongarra:1993:OOD

[232] J. Dongarra, R. Pozo, and D. Walker. An object oriented design for high performance linear algebra on distributed memory architectures. LA-PACK Working Note 61, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, August

1993. URL http://www.netlib.org/
lapack/lawns/lawn61.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn61.pdf. UT-CS-93-200, August 1993.

Dongarra:1993:PANb

[233] Jack J. Dongarra and Majed Sidani. A parallel algorithm for the nonsymmetric eigenvalue problem. SIAM Journal on Scientific Computing, 14 (3):542–569, May 1993. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic). URL http:/

/www.netlib.org/tennessee/ut-cs-
91-137.ps;   http://www.netlib.
org/utk/people/JackDongarra/pdf/
ut-cs-91-137.pdf.

Dongarra:1993:PUM

[234] J. Dongarra, R. Hempel, A. Hay, and D. Walker. A proposal for a userlevel message passing interface in a distributed memory environment. Technical Report ORNL/TM-12231, Oak Ridge National Laboratory, Knoxville, TN, USA, February 1993.

Dongarra:1993:PVCa

[235] J. Dongarra. Performance of various computers using standard linear equations software. Computer science dept. technical report, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, January 1993.

Dongarra:1993:PVCb

[236] Jack J. Dongarra and Henk A. van der Vorst. Performance of various computers using standard sparse linear equations solving techniques. In Dongarra and Gentzsch [985], pages 177– 188. ISBN 0-444-81518-X. LCCN QA76.9.E94 C63 1993.

Dongarra:1993:SDU

[237] J. Dongarra, T. Rowan, and R. Wade. Software distribution using XNETLIB. Computer Science Dept. Technical Report CS-93-214, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, ?? 1993.

Dongarra:1993:TDB

[238] J. J. Dongarra, R. A. Van de Geijn, and R. Clint Whaley. Two dimensional basic linear algebra communication subprograms. In Sincovec [991], pages 347– 352. ISBN 0-89871-315-3. LCCN QA 76.58 S55 1993. Two volumes.

Dongarra:1993:UGB

[239] J. J. Dongarra, R. A. van de Geijn, and R. C. Whaley. A users’ guide to the BLACS. Manuscript. Department of Computer Science, University of Tennessee, Knoxville, TN 37996., 1993.

Dongarra:1993:UPR

[240] J. J. Dongarra, A. Geist, R. Manchek, and W. Jiang. Using PVM 3.0 to run grand challenge applications on a heterogeneous network of parallel computers. In Sincovec [991], pages 873–877. ISBN 0-89871-315-3. LCCN QA 76.58 S55 1993. Two volumes.

Geist:1993:PTW

[241] A. Geist, J. Dongarra, A. Beguelin, B. Manchek, and Weicheng Jiang. PVM takes over the world. In IEEE [989], page 618. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Pozo:1993:LDO

[242] R. Pozo, J. J. Dongarra, and D. W. Walker. LAPACK++: a design overview of object-oriented extensions for high performance linear algebra. In IEEE [989], pages 162–171. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/sc93-
++.ps.

Anonymous:1994:MMI

[243] Anonymous. MPI: a message-passing interface standard. The International Journal of Supercomputer Applications and High Performance Computing, 8(3/4):159–416, Fall/Winter 1994. CODEN IJSAE9. ISSN 0890-2720.

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/.pdf.

Edited by J. Dongarra.

Barrett:1994:ABI

[244] Richard Barrett, Michael Berry, Jack Dongarra, Victor Eijkhout, and Charles Romine. Algorithmic bombardment for the iterative solution of linear systems: a poly-iterative approach. LAPACK Working Note 76, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, August

1994. URL http://www.netlib.org/
lapack/lawns/lawn76.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn76.pdf. UT-CS-94-239, August, 1994.

Barrett:1994:TSLa

[245] Richard Barrett, Michael Berry, Tony F. Chan, James Demmel, June Donato, Jack Dongarra, Victor Eijkhout, Roldan Pozo, Charles Romine, and Henk van der Vorst. Templates for the Solution of Linear Systems: Building Blocks for Iterative Methods. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-328-5. xiii + 112 pp. LCCN QA297.8 .T45 1994. URL ftp:/

/cs.utk.edu/linalg/templates.ps.

Barrett:1994:TSLb

[246] Richard Barrett, Michael Berry, Tony F. Chan, James W. Demmel, June Donato, Jack Dongarra, Victor Eijkhout, Roldan Pozo, Charles Romine, and Henk van der Vorst. Templates for the Solution of Linear Systems: Building Blocks for Iterative Methods (Japanese). Asakura Shoten, Tokyo, Japan, 1994. ISBN 4-254-11401-X. ???? pp. LCCN ???? See book review [1146].

Beguelin:1994:HHN

[247] Adam Beguelin, Jack J. Dongarra, George Al Geist, Robert Manchek, and Keith Moore. HeNCE: a heterogeneous network computing environment. Scientific Programming, 3(1):49–60, Spring 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

                                        URL http://www.netlib.org/utk/
                                        people/JackDongarra/PAPERS/HeNCE-
                                        A-Heterogeneous-Network-Computing-
                                        Environment.pdf.

Berry:1994:HPA

[248] Michael W. Berry, Jack J. Dongarra, and Youngbae Kim. A highly parallel algorithm for the reduction of a nonsymmetric matrix to block upper-Hessenberg form. LA-PACK Working Note 68, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, February

                                        1994. URL http://www.netlib.org/
                                        lapack/lawns/lawn68.ps;   http:/
                                        /www.netlib.org/lapack/lawnspdf/

lawn68.pdf. UT-CS-94-221, February 1994.

Berry:1994:PPD

[249] M. W. Berry, J. J. Dongarra, B. H. Larosei, and T. A. Letsche. PDS: a performance database server. Scientific Programming, 3(2):147–156, Summer 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Blackford:1994:QIG

[250] S. Blackford and J. Dongarra. Quick installation guide for LAPACK on Unix systems. LAPACK Working Note 81, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, September

1994. URL http://www.netlib.org/
lapack/lawns/lawn81.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn81.pdf. UT-CS-94-249, September, 1994.

Browne:1994:NSR

[251] Shirley V. Browne, Jack J. Dongarra, Stan C. Green, Keith Moore, Thomas H. Rowan, and Reed C. Wade. Netlib services and resources. Report ORNL/TM-12680, Oak Ridge National Laboratory, Knoxville, TN, USA, April 1994. 42 pp.

Choi:1994:CNS

[252] J. Choi, J. J. Dongarra, R. Pozo, and D. W. Walker. Constructing numerical software libraries for highperformance computing environments. Lecture Notes in Computer Science, 879:147–168, 1994. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Choi:1994:CRL

[253] Jaeyoung Choi, Jack J. Dongarra, Roldan Pozo, Danny C. Sorensen, and David W. Walker. CRPC research into linear algebra software for high performance computers. The International Journal of Supercomputer Applications and High Performance Computing, 8(2):99–118, Summer 1994. CODEN IJSAE9. ISSN 0890-2720.

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/CRPC-
Research-into-Linear-Algebra-Software-
for-High-Performance.pdf.

Choi:1994:DIS

[254] J. Choi, J. J. Dongarra, S. Ostrouchov, A. P. Petitet, D. W. Walker, and R. C. Whaley. The design and implementation of the ScaLAPACK LU, QR, and Cholesky factorization routines. LAPACK Working Note 80, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, September

1994. URL http://www.netlib.org/
lapack/lawns/lawn80.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn80.pdf. UT-CS-94-246, September, 1994.

Choi:1994:DPD

[255] J. Choi, J. J. Dongarra, and D. W. Walker. The design of a parallel, dense linear algebra software library: reduction to Hessenberg, tridiagonal, and bidiagonal form. In Dongarra and Tourancheau [995], pages 98–111. ISBN 0-89871-343-9. LCCN QA76.58.I568 1994.

Choi:1994:DSS

[256] Jaeyoung Choi, J. J. Dongarra, and D. W. Walker. The design of scalable software libraries for distributed memory concurrent computers. In Siegel [1000], pages 792–799. ISBN 0-8186-5602-6, 0-8186-5601-8. ISSN 1063-7133. LCCN QA 76.58 I56 1994. IEEE catalog number 94TH0652-8.

Choi:1994:PMT

[257] Jaeyoung Choi, J. J. Dongarra, and D. W. Walker. Parallel matrix transpose algorithms on distributed memory concurrent computers. In IEEE [998], pages 245–252. ISBN 0-8186-4980-1 (paper), 0-8186-4981-X (microfiche). LCCN QA76.58 .S34 1993.

Choi:1994:PPU

[258] Jaeyoung Choi, Jack J. Dongarra, and David W. Walker. PUMMA: Parallel Universal Matrix Multiplication Algorithms on distributed memory concurrent computers. Concurrency: practice and experience, 6(7):543–570, October 1994. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/PUMMA-
Parallel-Universal-Matrix-Multiplication-
Algorithms.pdf.

Choi:1994:PSP

[259] Jaeyoung Choi, J. J. Dongarra, and D. W. Walker. PB-BLAS: a set of parallel block Basic Linear Algebra Subprograms. In IEEE [997], pages 534–

541. ISBN 0-8186-5680-8, 0-8186-5681-

6. LCCN QA76.58.S32 1994. IEEE catalog number 94TH0637-9.

Dayde:1994:PBI

[260] Michael J. Daydé, Iain S. Duff, and Antoine Petitet. A parallel block implementation of level-3 BLAS for MIMD vector processors. ACM Transactions on Mathematical Software, 20(2): 178–193, June 1994. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.

acm.org/pubs/citations/journals/

toms/1994-20-2/p178-dayde/. See

[131, 144, 195].

Dongarra:1994:AAC

[261] J. J. Dongarra and Iain S. Duff. Advanced architecture computers. Technical Report CS-89-90, University of Tennessee, Knoxville, TN, USA, 1994. Revision of 1987 Report AERE R12415, HMSO, London. An earlier version appeared in Federal Supercomputer Programs and Policies, US Government Printing Office, Washington DC, 710– 799.

Dongarra:1994:AEP

[262] Jack Dongarra, Andrew Lumsdaine, Xinhui Niu, Roldan Pozo, and Karin Remington. Accurate and efficient (parallel) algorithms have been formulated and implemented on distributive platforms for solving . . .. In ????, editor, OONSKI’94: Proceedings of the second annual objectoriented numerics conference, Sunriver, Oregon, April 24–27, 1994, page ?? Rogue Wave Software, Corvallis, OR, USA, 1994. URL http://www.

netlib.org/utk/papers/oonski94.
ps;   http://www.netlib.org/utk/
people/JackDongarra/pdf/oonski94.
pdf.

Dongarra:1994:CCI

[263] J. Dongarra and M. Kolatis. Call conversion interface (CCI) for LAPACK/ ESSL. LAPACK Working Note 82, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, August

1994. URL http://www.netlib.org/
lapack/lawns/lawn82.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn82.pdf. UT-CS-94-250, August, 1994.

Dongarra:1994:CNS

[264] Jack Dongarra. Constructing numerical software libraries for HPCC environments. In IEEE [999], pages 4– ?? ISBN 0-8186-6395-2, 0-8186-6396-0. LCCN QA76.9.D5I328 1994. IEEE catalog number 94TH0667-6.

Dongarra:1994:IHE

[265] J. Dongarra and D. Reed. Introduction to the HPC early evaluation vendor session. In Anonymous [992], pages 131–134.

Dongarra:1994:IRP

[266] Jack Dongarra and Michael Kolatis. IBM RS/6000-550 & -590 performance for selected routines in ESSL. LAPACK Working Note 71, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, April

1994. URL http://www.netlib.org/
lapack/lawns/lawn71.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn71.pdf. UT-CS-94-231, April 1994.

Dongarra:1994:PL

[267] Jack Dongarra. Performance of LA-PACK. In Gilbert and Kershaw [996], pages 55–68 (or 55–67??). ISBN 0-19-853463-9. LCCN QA374.L335 1994.

Dongarra:1994:SIA

[268] J. J. Dongarra, R. A. Van de Geijn, and D. W. Walker. Scalability issues affecting the design of a dense linear algebra library. Journal of Parallel and Distributed Computing, 22(3): 523–537, September 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.idealibrary.com/links/doi/
10.1006/jpdc.1994.1108/production;
 http://www.idealibrary.com/links/
doi/10.1006/jpdc.1994.1108/production/
pdf;   http://www.netlib.org/utk/
people/JackDongarra/PAPERS/Scalability-
Issues-Affecting-the-Design-of-
a-Dense-Linear-Algebra-Library.
pdf;   https://www.math.utah.edu/
pub/bibnet/authors/d/dongarra-jack-
j.bib.

Dongarra:1994:SMLa

[269] J. Dongarra, A. Lumsdaine, X. Niu, R. Pozo, and K. Remington. A sparse matrix library in C++ for high performance architectures. LA-PACK Working Note 74, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, July

1994. URL http://www.netlib.org/
lapack/lawns/lawn74.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn74.pdf. UT-CS-94-236, July 1994.

Dongarra:1994:SMLb

[270] J. Dongarra, A. Lumsdaine, X. Niu, R. Pozo, and K. Remington. A sparse matrix library in C++ for high performance architectures. In Anonymous [993], pages 214–218. URL http:

//www.netlib.org/netlib/lapack/
lawns/lawn74.ps;   http://www.netlib.
org/netlib/lapack/lawnspdf/lawn74.
pdf.

Dongarra:1994:SOO

[271] J. J. Dongarra, R. Pozo, and D. W. Walker. ScaLAPACK++: an object oriented linear algebra library for scalable systems. In IEEE [998], pages 216–223. ISBN 0-8186-4980-1 (paper), 0-8186-4981-X (microfiche). LCCN QA76.58 .S34 1993.

Geist:1994:PPV

[272] Al Geist, Adam Beguelin, Jack Dongarra, Weicheng Jiang, Robert Manchek, and Vaidy Sunderam. PVM: Parallel Virtual Machine: a Users’ Guide and Tutorial for Networked Parallel Computing. Scientific and engineering computation. MIT Press, Cambridge, MA, USA, 1994. ISBN 0-262-57108-0 (paperback). xvii + 279 pp. LCCN QA76.58 .P85 1994. US$19.95.

PARKBENCH:1994:PRP

[273] PARKBENCH Committee/Assembled by R.Hockney (Chairman) and M. Berry (Secretary). PARKBENCH report: Public international benchmarks for parallel computers. Scientific Programming, 3(2):101–146, Summer 1994. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Plank:1994:ABD

[274] J. S. Plank, Y. Kim, and J. J. Dongarra. Algorithm-based diskless checkpointing for fault tolerant matrix operations. LAPACK Working Note 90, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, December

1994. URL http://www.netlib.org/
lapack/lawns/lawn90.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn90.pdf. UT-CS-94-268, December 1994.

Sullivan:1987:ADL

[275] Francis Sullivan and Jack Dongarra. Algorithm design for large-scale computations. International Journal of Supercomputer Applications, 1(1):99– 105, March 1987. CODEN IJSAE9. ISSN 0890-2720. URL http://

journals.sagepub.com/doi/pdf/10.
1177/109434208700100107.

Sunderam:1994:PCC

[276] V. S. Sunderam, G. A. Geist, J. Dongarra, and R. Manchek. The PVM concurrent computing system: Evolution, experiences, and trends. Parallel Computing, 20(4):531–545, March 31, 1994. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/The-
PVM-Concurrent-Computing-System-
Evolution-Experiences-Trends.pdf.

Anderson:1995:LUG

[277] E. Anderson, Z. Bai, C. Bischof, J. Demmel, J. Dongarra, J. Du Croz, A. Greenbaum, S. Hammarling, A. McKenney, S. Ostrouchov, and D. Sorensen. LAPACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, second edition, 1995. ISBN 0-89871-345-5 (paperback). xix + 325 pp. LCCN QA76.73.F25 L36 1995.

Bai:1995:SDN

[278] Z. Bai, J. Demmel, J. Dongarra, A. Petitet, H. Robinson, and K. Stanley. The spectral decomposition of nonsymmetric matrices on distributed memory parallel computers. LAPACK Working Note 91, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, January

1995. URL http://www.netlib.org/

lapack/lawns/lawn91.ps;   http:/
/www.netlib.org/lapack/lawnspdf/
lawn91.pdf;   http://www.netlib.
org/utk/papers/sign/sign.html. UT-

CS-95-273, January 1995.

Bai:1995:TLAa

[279] Z. Bai, D. Day, J. Demmel, J. Dongarra, M. Gu, A. Ruhe, and H. van der Vorst. Templates for linear algebra problems. LAPACK Working Note 106, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, October

1995. URL http://www.netlib.org/
lapack/lawns/lawn106.ps;   http:
//www.netlib.org/lapack/lawnspdf/

lawn106.pdf. UT-CS-95-311, October 1995. Published in [280].

Bai:1995:TLAb

[280] Z. Bai, D. Day, J. Demmel, and J. Dongarra. Templates for linear algebra problems. Lecture Notes in Computer Science, 1000:115–??, 1995. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://www.netlib.org/utk/
papers/etemplates.ps;   http://
www.netlib.org/utk/papers/etemplates/
paper.html.

Beguelin:1995:REP

[281] Adam Beguelin, Jack Dongarra, Al Geist, Robert Manchek, and Vaidy Sunderam. Recent enhancements to PVM. The International Journal of Supercomputer Applications and High Performance Computing, 9(2):108–127, Summer 1995. CODEN IJSCFG. ISSN 1078-3482. URL http://www.

netlib.org/utk/papers/pvm-ijsa/
ijsa.html.

Berry:1995:PAR

[282] Michael W. Berry, Jack J. Dongarra, and Youngbae Kim. A parallel algorithm for the reduction of a nonsymmetric matrix to block upper-Hessenberg form. Parallel Computing, 21(8):1189–1211, August 10, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Boisvert:1995:DSD

[283] Ronald Boisvert, Shirley Browne, and Jack Dongarra. Digital software and data repositories for support of scientific computing. In A Forum on Research and Technology Advances in Digital Libraries (DL’95), May 15– 19, 1995, McLean VA. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., May

                                        1995.  URL ftp://netlib.bell-
                                        labs.com/netlib/srwn/srwn09.ps.
                                        gz;   http://www.netlib.org/srwn/
                                        srwn09.ps;   http://www.netlib.
                                        org/utk/papers/dig-lib/main.html;
                                         http://www.netlib.org/utk/people/
                                        JackDongarra/pdf/srwn09.pdf. McLean,

Virginia.

Browne:1995:DIM

[284] Shirley Browne, Jack Dongarra, Geoffrey C. Fox, Ken Hawick, Ken Kennedy, Rick Stevens, Robert Olson, and Tom Rowan. Distributed information management in the National HPCC Software Exchange. In Karin [1005], pages 463–477. ISBN 0-7803-3604-6, 0-89791-862-2, 0-7803-3605-4. ISSN 1063-9535. LCCN QA76.88 .S856

                                        1995. URL http://www.netlib.org/
                                        srwn/srwn10.html;   http://www.
                                        supercomp.org/sc95/proceedings/
                                        588_BROW/SC95.HTM. ACM order num-

ber 415932. IEEE Order Plan catalog number 95CB35990. These proceedings are not available in printed form. However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

Browne:1995:LINa

[285] Shirley Browne, Jack Dongarra, Stan Green, Keith Moore, Theresa Pepin, Tom Rowan, Reed Wade, and Eric Grosse. Location-independent naming for virtual distributed software repositories. In Symposium on Software Reusability, pages 179–185. ACM Press, New York, NY 10036, USA,

April 1995. URL ftp://netlib.bell-
labs.com/netlib/srwn/srwn07.ps.

gz. Seattle, Washington.

Browne:1995:LINb

[286] Shirley Browne, Jack Dongarra, Stan Green, Keith Moore, Theresa Pepin, Tom Rowan, and Reed Wade. Location-independent naming for virtual distributed software repositories. ACM SIGSOFT Software Engineering Notes, 20(SI):179–185, 1995. CODEN SFENDP. ISSN 0163-5948 (print), 1943-5843 (electronic).

Browne:1995:LNV

[287] Shirley Browne, Jack Dongarra, Stan Green, Keith Moore, Theresa Pepin, Tom Rowan, Reed Wade, and Eric Grosse. Location-independent naming for virtual distributed software repositories. In Mansur H. Samadzadeh and Mansour K. Zand, editors, Proceedings of the ACM SIGSOFT Symposium on Software Reusability (SSR ’95), April 28–30, 1995, Seattle, Washington, USA: co-located with the 17th International Conference on Software Engineering (ICSE-17), April 23–30, 1995, pages 179–185. ACM Press, New York, NY 10036, USA, August 1995. CODEN SFENDP. ISBN 0-89791-739-1. ISSN 0163-5948 (print), 1943-5843 (electronic). LCCN QA76.6 .S62

v.20 1995. URL http://www.acm.org/
pubs/contents/proceedings/ssr/211782/
;   http://www.netlib.org/srwn/
srwn07.ps;   http://www.netlib.
org/utk/papers/lifn/main.html;
http://www.netlib.org/utk/people/
JackDongarra/pdf/srwn07.pdf. ACM

order number 595950.

Browne:1995:MNV

[288] Shirley Browne, Jack Dongarra, Geoffrey C. Fox, Ken Hawick, Ken Kennedy, Rick Stevens, Robert Olson, and Tom Rowan. Management of the NHSE — A Virtual Distributed Digital Library: June 11–13, 1995, Austin, TX. In Frank M. Shipman, Richard Furuta, and David M. Levy, editors, Proceedings of Digital Libraries ’95: the Second Annual Conference on the Theory and Practice of Digital Libraries, June 11–13, 1995, Austin, Texas, page ?? Hypermedia Research Laboratory, Texas A&M University, College Station, TX, 1995. LCCN Z699.A1

D54 1995. URL http://www.netlib.
org/srwn/srwn10.html;   http://
www.netlib.org/srwn/srwn11.ps;
http://www.netlib.org/utk/people/
JackDongarra/pdf/srwn11.pdf.

Browne:1995:NHSa

[289] Shirley Browne, Jack Dongarra, Jeff Horner, Paul McMahan, and Scott Wells. National HPCC Software Exchange (NHSE): Uniting the high performance computing and communications community. D-Lib magazine: the magazine of the Digital Library Forum, May 15, 1995. ISSN 1082-9873. URL

http://www.dlib.org/dlib/may98/
browne/05browne.html;   http://
www.ncstrl.org:8900/ncstrl/servlet/
search?formname=detail&id=oai%3Ancstrlh%
3Acnri_dlib%3Acnri.dlib%2F%2Fmay98-
browne.

Browne:1995:NHSb

[290] Shirley Browne, Jack Dongarra, Stan Green, Keith Moore, Tom Rowan, Reed Wade, Geoffrey Fox, Ken Hawick, Ken Kennedy, Jim Pool, Rick Stevens, Bob Olson, and Terry Disz. The National HPCC Software Exchange. IEEE Computational Science & Engineering, 2(2):62–69, Summer 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (elec-

tronic). URL http://www.computer.
org/cse/cs1998/c2062abs.htm;   http:
//www.netlib.org/srwn/srwn08.ps;
 http://www.netlib.org/utk/papers/
nse-cse/nse-cse.html;   http://
www.netlib.org/utk/people/JackDongarra/
pdf/srwn08.pdf.

Browne:1995:NMS

[291] Shirley Browne, Jack Dongarra, Eric Grosse, and Tom Rowan. The Netlib Mathematical Software Repository. D-Lib magazine: the magazine of the Digital Library Forum, September 1995. ISSN 1082-9873. URL http://

www.cnri.reston.va.us/home/dlib.
html;   http://WWW.CNRI.Reston.
VA.US/home/dlib/september95/netlib/
09browne.html;   http://www.ncstrl.
org:8900/ncstrl/servlet/search?
formname=detail&id=oai%3Ancstrlh%

                                        3Acnri_dlib%3Acnri.dlib%2F%2Fmay98-
                                        browne.

Browne:1995:VPD

[292] James C. Browne, Syed I. Hyder, Jack Dongarra, Keith Moore, and Peter Newton. Visual programming and debugging for parallel computing. IEEE parallel and distributed technology: systems and applications, 3(1):75–83, Spring 1995. CODEN IPDTEX. ISSN 1063-6552 (print), 1558-1861 (electronic).

Casanova:1995:PPM

[293] Henri Casanova, Jack Dongarra, and Weicheng Jiang. The performance of PVM on MPP systems. Technical report, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, August 1995.

                                        URL http://www.netlib.org/utk/
                                        papers/pvmmpp.ps;   http://www.
                                        netlib.org/utk/papers/pvmmpp/pvmmpp.
                                        html;   http://www.netlib.org/utk/
                                        people/JackDongarra/pdf/pvmmpp.
                                        pdf.

Choi:1995:DPDa

[294] J. Choi, J. Dongarra, and D. Walker. The design of a parallel dense linear algebra software library: Reduction to Hessenberg, tridiagonal, and bidiagonal form. LAPACK Working Note 92, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, February

                                        1995. URL http://www.netlib.org/
                                        lapack/lawns/lawn92.ps;   http:/
                                        /www.netlib.org/lapack/lawnspdf/

lawn92.pdf. UT-CS-95-275, February 1995.

Choi:1995:DPDb

[295] Jaeyoung Choi, Jack J. Dongarra, and David W. Walker. The design of a parallel dense linear algebra software library: reduction to Hessenberg, tridiagonal, and bidiagonal form. Numerical Algorithms, 10(3–4):379–399, October 1995. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (electronic).

Choi:1995:PMT

[296] Jaeyoung Choi, Jack J. Dongarra, and David W. Walker. Parallel matrix transpose algorithms on distributed memory concurrent computers. Parallel Computing, 21(9):1387– 1405, September 12, 1995. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Choi:1995:PSP

[297] J. Choi, J. Dongarra, S. Ostrouchov, A. Petitet, D. Walker, and R. C. Whaley. A proposal for a set of parallel basic linear algebra subprograms. LAPACK Working Note 100, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, May

1995. URL http://www.netlib.org/
lapack/lawns/lawn100.ps;   http:
//www.netlib.org/lapack/lawnspdf/
lawn100.pdf;   http://www.netlib.
org/utk/papers/pblas.ps;   http:
//www.netlib.org/utk/papers/pblas/
pblas.html;   http://www.netlib.
org/utk/people/JackDongarra/pdf/

pblas.pdf. LAPACK Working Note #100. UT-CS-95-292, May 1995.

Choi:1995:SLA

[298] Jaeyoung Choi and J. J. Dongarra. Scalable linear algebra software libraries for distributed memory concurrent computers. In IEEE [1004], pages 170–177. ISBN 0-8186-7125-4. LCCN QA76.9.D5 I24 1995. IEEE catalog number 95TB8106.

Choi:1995:SPL

[299] J. Choi, J. Demmel, I. Dhillon, J. Dongarra, S. Ostrouchov, A. Petitet, K. Stanley, D. Walker, and R. C. Whaley. ScaLAPACK: a portable linear algebra library for distributed memory computers — design issues and performance. LAPACK Working Note 95, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, March

                                        1995. URL http://www.netlib.org/
                                        lapack/lawns/lawn95.ps;   http:/
                                        /www.netlib.org/lapack/lawnspdf/
                                        lawn95.pdf;   http://www.netlib.
                                        org/utk/papers/scalapack/paper.

html. LAPACK Working Note #95. UT-CS-95-283, March 1995.

Desprez:1995:PSF

[300] F. Desprez, J. J. Dongarra, and B. Tourancheau. Performance study of LU factorization with low communication overhead on multiprocessors. Parallel Processing Letters, 5(2):157–169, June 1995. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Dongarra:1995:A

[301] J. J. Dongarra, H. W. Meuer, and E. Strohmaier. The aims. Supercomputer, 11(2–3):4–5, ???? 1995. CODEN SPCOEL. ISSN 0168-7875.

Dongarra:1995:BTW

[302] J. Dongarra, S. Hammarling, and S. Ostrouchov. BLAS technical workshop. LAPACK Working Note 109, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, November

1995. URL http://www.netlib.org/
lapack/lawns/lawn109.ps;   http:
//www.netlib.org/lapack/lawnspdf/

lawn109.pdf. UT-CS-95-317, November 1995.

Dongarra:1995:HNC

[303] J. Dongarra. Heterogeneous networkbased computing systems. In Dongarra et al. [1001], pages 5–16. ISBN 0-444-82163-5. ISSN 0927-5452. LCCN QA76.88.H55 1995.

Dongarra:1995:IMS

[304] Jack Dongarra, Steve W. Otto, Marc Snir, and David Walker. An introduction to the MPI Standard. Technical report CS-95-274, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, January

1995.  URL http://www.netlib.
org/tennessee/ut-cs-95-274.ps;
http://www.netlib.org/utk/papers/
intro-mpi/intro-mpi.html;   http:
//www.netlib.org/utk/people/JackDongarra/
pdf/ut-cs-95-274.pdf. Appears in

CACM [348].

Dongarra:1995:IVI

[305] J. Dongarra, A. Lumsdaine, R. Pozo, and K. Remington. IML++ v. 1.2: Iterative methods library reference guide. LAPACK Working Note 102, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, August

1995. URL http://www.netlib.org/
lapack/lawns/lawn102.ps;   http:
//www.netlib.org/lapack/lawnspdf/

lawn102.pdf. UT-CS-95-303, August 1995.

Dongarra:1995:LVH

[306] J. Dongarra, R. Pozo, and D. Walker. LAPACK++ V. 1.0: High performance linear algebra users’ guide. LAPACK Working Note 98, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, May

                                        1995. URL http://www.netlib.org/
                                        lapack/lawns/lawn98.ps;   http:/
                                        /www.netlib.org/lapack/lawnspdf/

lawn98.pdf. UT-CS-95-290, May 1995.

Dongarra:1995:PBC

[307] J. J. Dongarra and T. Hey. The Park-Bench benchmark collection. Supercomputer, 11(2–3):94–114, June 1995. CODEN SPCOEL. ISSN 0168-7875.

Dongarra:1995:PFI

[308] J. J. Dongarra, J. Du Croz, S. Hammarling, J. Wasniewski, and A. Zemla. A proposal for a Fortran 90 interface for LAPACK. LAPACK Working Note 101, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, July

                                        1995. URL http://www.netlib.org/
                                        lapack/lawns/lawn101.ps;   http:
                                        //www.netlib.org/lapack/lawnspdf/

lawn101.pdf. UT-CS-95-295, July 1995.

Dongarra:1995:PVC

[309] Jack J. Dongarra. Performance of various computers using standard linear equations software. Technical Report CS-89-85, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, February 1995. 34 pp.

Dongarra:1995:RCI

[310] J. Dongarra, V. Eijkhout, and A. Kalhan. Reverse communication interface for linear algebra templates for iterative methods. LA-PACK Working Note 99, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, May

1995. URL http://www.netlib.org/
lapack/lawns/lawn99.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn99.pdf. UT-CS-95-291, May 1995.

Dongarra:1995:RSW

[311] Jack Dongarra and Jerzy Waśniewski. Report on the Second Workshop on Applied Parallel Computing, PARA95. ACM SIGNUM Newsletter, 30(4):28– ??, October 1995. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dongarra:1995:SDU

[312] J. Dongarra, T. Rowan, and R. Wade. Software distribution using Xnetlib. ACM Transactions on Mathematical Software, 21(1):79–88, March 1995. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

URL http://www.netlib.org/srwn/
srwn04.ps;   http://www.netlib.
org/utk/papers/xnetlib/xnetlib.
html;   http://www.netlib.org/utk/
people/JackDongarra/pdf/srwn04.
pdf.

Dongarra:1995:SDX

[313] Jack Dongarra, Tom Rowan, and Reed Wade. Software distribution using XNETLIB. ACM Transactions on Mathematical Software, 21(1):79– 88, March 1995. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.

acm.org/pubs/citations/journals/
toms/1995-21-1/p79-dongarra/.

Dongarra:1995:SLL

[314] Jack J. Dongarra and David W. Walker. Software libraries for linear algebra computations on high performance computers. SIAM Review, 37(2):151–180, June 1995. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

URL http://www.netlib.org/utk/
papers/siam-review93/paper.html;
 http://www.netlib.org/utk/papers/
siam-review93/paper.ps;   http:/
/www.netlib.org/utk/papers/siam_
cover.ps;   http://www.netlib.org/
utk/papers/siam_cover/siam_cover.
html;   http://www.netlib.org/utk/
people/JackDongarra/pdf/siam93.
pdf;   http://www.netlib.org/utk/
people/JackDongarra/pdf/siam_cover.
pdf.

Dongarra:1995:TSS

[315] J. J. Dongarra, H. W. Meuer, and E. Strohmaier. TOP500 supercomputer sites. Supercomputer, 11(2-3): 133–163 (or 164–194??), June 1995. CODEN SPCOEL. ISSN 0168-7875.

Newton:1995:OVV

[316] P. Newton and J. Dongarra. Overview of VPE: a visual environment for message-passing. In IEEE [1003], pages 85–92. ISBN 0-8186-7121-1. LCCN QA76.9.D5 H48 1995. URL http://

www.netlib.org/utk/papers/hcw95.

ps;   http://www.netlib.org/utk/
people/JackDongarra/pdf/hcw95.pdf.

Plank:1995:ADC

[317] James S. Plank, Youngbae Kim, and Jack J. Dongarra. Algorithm-based diskless checkpointing for fault-tolerant matrix operations. In IEEE [1002], pages 351–360. CODEN DPFTDL. ISBN 0-8186-7079-7, 0-8186-7145-9. ISSN 0731-3071. LCCN QA 76.9 F38

I57 1995. URL http://www.cs.utk.
edu/~plank/plank/papers/FTCS25.
1995.html;   http://www.netlib.
org/utk/papers/fault.ps;   http:
//www.netlib.org/utk/people/JackDongarra/

pdf/fault.pdf. IEEE catalog number 95CH35823.

Barrett:1996:ABI

[318] R. Barrett, M. Berry, J. Dongarra, V. Eijkhout, Romine, and C. Algorithmic bombardment for the iterative solution of linear systems: a poly-iterative approach. Journal of Computational and Applied Mathematics, 74(1–2):91–109, ???? 5, 1996. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/bombard.
ps;   http://www.netlib.org/utk/
people/JackDongarra/pdf/bombard.
pdf.

Blackford:1996:FIL

[319] L. Susan Blackford, Jack J. Dongarra, Jeremy Du Croz, Sven Hammarling, and Jerzy Wasniewski. A Fortran 90 interface for LAPACK. LAPACK Working Note 117, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, December

1996. URL http://www.netlib.org/
lapack/lawns/lawn117.ps;   http:
//www.netlib.org/lapack/lawnspdf/

lawn117.pdf. UT-CS-96-341, December 1996.

Blackford:1996:PEDa

[320] L. S. Blackford, A. Cleary, J. Demmel, I. Dhillon, J. Dongarra, S. Hammarling, A. Petitet, H. Ren, K. Stanley, and R. C. Whaley. Practical experience in the dangers of heterogeneous computing. LAPACK Working Note 112, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, July

1996. URL http://www.netlib.org/
lapack/lawns/lawn112.ps;   http:
//www.netlib.org/lapack/lawnspdf/

lawn112.pdf. UT-CS-96-330, July 1996.

Blackford:1996:PEDb

[321] L. S. Blackford, A. Cleary, J. Demmel, I. Dhillon, J. Dongarra, S. Hammarling, A. Petitet, H. Ren, K. Stanley, and R. C. Whaley. Practical experience in the dangers of heterogeneous computing. In Waśniewski et al. [1013], pages 57–64. ISBN 3-540-62095-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58

.P35 1996. URL http://www.netlib.
org/utk/papers/practical-hetro/
paper.html;   http://www.netlib.
org/utk/papers/practical-hetro/
paper.ps;   http://www.netlib.org/
utk/people/JackDongarra/pdf/prac-
het.pdf.

Blackford:1996:SPL

[322] Laura Susan Blackford, J. Choi, A. Cleary, A. Petitet, R. C. Whaley, J. Demmel, I. Dhillon, K. Stanley, J. Dongarra, S. Hammarling, G. Henry, and D. Walker. ScaLAPACK: a portable linear algebra library for distributed memory computers — design issues and performance. In ACM [1006], page 5. ISBN 0-89791-854-1. LCCN A76.88 .S8573 1996. URL http:/

/www.netlib.org/utk/papers/sc96-
scalapack/paper.html;   http://
www.netlib.org/utk/papers/sc96-
scalapack/paper.ps;   http://www.
netlib.org/utk/people/JackDongarra/
pdf/scala96.pdf;   http://www.supercomp.
org/sc96/proceedings/SC96PROC/DONGARRA/

INDEX.HTM. ACM Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Boisvert:1996:DSD

[323] R. Boisvert, S. Browne, J. Dongarra, and E. Grosse. Digital software and data repositories for support of scientific computing. In Nabil R. Adam, B. K. Bhargava, M. Halem, and Y. Yesha, editors, Digital libraries: research and technology advances: ADL ’95 Forum, McLean, Virginia, USA, May 15–17, 1995: selected papers, pages 103–114. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61410-9. LCCN QA267.A1 L43 no.1082; Z699 .D532 1996.

Browne:1996:EHP

[324] Shirley Browne, Jack Dongarra, and Tom Rowan. Evaluation of highperformance computing software. In ????, editor, Proceedings of the Scalable Parallel Libraries Conference III SPLC96, Mississippi State University, October 1996, page ?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. URL http:/

                                        /www.netlib.org/utk/papers/eval_
                                        hpc_sw/eval_hpc_sw.html;   http:
                                        //www.netlib.org/utk/papers/eval_
                                        hpc_sw/eval_hpc_sw.ps;   http://
                                        www.netlib.org/utk/people/JackDongarra/
                                        pdf/eval_hpc_sw.pdf.

Browne:1996:PAH

[325] S. Browne, H. Casanova, and J. Dongarra. Providing access to high performance computing technologies. In Waśniewski et al. [1013], pages 123–

133. ISBN 3-540-62095-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35

                                        1996. URL http://www.netlib.org/
                                        utk/papers/nhse-netsolve/paper.
                                        html;   http://www.netlib.org/utk/
                                        papers/nhse-netsolve/paper.ps;
                                        http://www.netlib.org/utk/people/
                                        JackDongarra/pdf/nhsenet.pdf.

Browne:1996:SRH

[326] S. Browne, J. Dongarra, G. Fox, K. Hawick, and T. Rowan. Software reuse in high-performance computing. Technical report, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, 1996. URL http://www.

                                        netlib.org/netlib/utk/people/JackDongarra/
                                        PAPERS/wisr.ps;   http://www.netlib.
                                        org/netlib/utk/people/JackDongarra/
                                        pdf/wisr.pdf.

Casanova:1996:NNSa

[327] Henri Casanova and Jack Dongarra. NetSolve: a network server for solving computational science problems. Technical report CS-96-328, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, May 1996. URL http://www.

netlib.org/utk/papers/NETSOLVE/
INDEX.HTM;   http://www.netlib.
org/utk/papers/NETSOLVE/paper.ps;
http://www.netlib.org/utk/people/
JackDongarra/pdf/netsolve.pdf.

Casanova:1996:NNSb

[328] Henri Casanova and Jack Dongarra. NetSolve: a network server for solving computational science problems. In ACM [1006], page 40. ISBN 0-89791-854-1. LCCN A76.88 .S8573 1996. URL http://www.

supercomp.org/sc96/proceedings/
SC96PROC/CASANOVA/INDEX.HTM. ACM

Order Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Choi:1996:DIS

[329] Jaeyoung Choi, J. J. Dongarra, L. S. Ostrouchov, Petitet, A. P., D. W. Walker, and R. C. Whaley. Design and implementation of the ScaLAPACK LU, QR, and Cholesky factorization routines. Scientific Programming, 5(3): 173–184, Fall 1996. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://www.

netlib.org/netlib/lapack/lawns/
lawn80.ps;   http://www.netlib.
org/netlib/lapack/lawnspdf/lawn80.
pdf;   http://www.netlib.org/utk/
papers/factor/ftcover.html.

Choi:1996:PBS

[330] Jaeyoung Choi, Jack J. Dongarra, and David W. Walker. PB-BLAS: a set of parallel block basic linear algebra subprograms. Concurrency: practice and experience, 8(7):

517–535, September 1996. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http://

www3.interscience.wiley.com/cgi-
bin/abstract?ID=23295.

Choi:1996:PSP

[331] Jaeyoung Choi, J. Dongarra, S. Ostrouchov, A. Petitet, D. Walker, and R. C. Whaley. A proposal for a set of Parallel Basic Linear Algebra Subprograms. In Dongarra et al. [1009], pages 107– 114. CODEN LNCSD9. ISBN 3-540-60902-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1995.

Choi:1996:SPLa

[332] J. Choi, J. Demmel, I. Dhillon, J. Dongarra, S. Ostrouchov, A. Petitet, K. Stanley, D. Walker, and R. C. Whaley. ScaLAPACK: a portable linear algebra library for distributed memory computers — design issues and performance. Computer Physics Communications, 97(1– 2):1–15, August 2, 1996. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/0010465596000173.

Choi:1996:SPLb

[333] J. Choi, J. Demmel, I. Dhillon, J. Dongarra, S. Ostrouchov, A. Petitet, K. Stanley, D. Walker, and R. C. Whaley. ScaLAPACK: a portable linear algebra library for distributed memory computers. design issues and performance. In Dongarra et al. [1009], pages 95–106. CODEN LNCSD9. ISBN 3-540-60902-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1995.

Demmel:1996:DHNa

[334] Jim Demmel, Jack Dongarra, Sven Hammarling, Susan Ostrouchov, and Ken Stanley. The dangers of heterogeneous network computing: Heterogeneous networks considered harmful. Technical report, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, March 1996. URL

http://www.netlib.org/tennessee/
hetero.ps;   http://www.netlib.
org/utk/people/JackDongarra/pdf/
hetero.pdf.

Demmel:1996:DHNb

[335] J. Demmel, J. Dongarra, S. Hammarling, and S. Ostrouchov. The dangers of heterogeneous network computing: Heterogeneous networks considered harmful. In IEEE [1011], pages 64–71.

Dongarra:1996:CTA

[336] J. J. Dongarra, B. Straughan, and D. W. Walker. Chebyshev tau-QZ algorithm methods for calculating spectra of hydrodynamic stability problems. Applied Numerical Mathematics: Transactions of IMACS, 22(4): 399–434, December 1996. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http:/

/www.netlib.org/utk/papers/ctqz.
ps;   http://www.netlib.org/utk/
people/JackDongarra/pdf/ctqz.pdf.

Dongarra:1996:CTH

[337] J. J. Dongarra, H. W. Meuer, H. D. Simon, and E. Strohmaier. Changing technologies of HPC. Lecture Notes in Computer Science, 1067:875–879, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongarra:1996:DHW

[338] Jack Dongarra and Jerzy Waśniewski. Denmark hosts Workshop on Applied Parallel Computing (PARA96). ACM SIGNUM Newsletter, 31(3):28–??, ???? 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Dongarra:1996:FLA

[339] Jack Dongarra. Future linear-algebra libraries. IEEE Computational Science & Engineering, 3(2):38–40, Summer 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (elec-

                                        tronic). URL http://www.computer.
                                        org/cse/cs1998/c2038abs.htm;   http:
                                        //www.netlib.org/utk/papers/future-
                                        la/future-la.ps;   http://www.netlib.
                                        org/utk/papers/future-la/index.
                                        html;   http://www.netlib.org/utk/
                                        people/JackDongarra/pdf/future-
                                        la.pdf.

Dongarra:1996:HPCa

[340] Jack Dongarra and Horst D. Simon. High performance computing in the U.S. in 1995 — an analysis on the basis of the TOP500 list. Technical report CS-96-318, University of Tennessee, Knoxville, Knoxville, TN 37996, USA,

                                        1996.  URL http://www.netlib.
                                        org/utk/papers/top500-us/top500-
                                        us.ps;   http://www.netlib.org/
                                        utk/papers/top500-us/us.html;
                                        http://www.netlib.org/utk/people/
                                        JackDongarra/pdf/top500-us.pdf.

Dongarra:1996:HPCb

[341] J. J. Dongarra and H. D. Simon. High performance computing in the U.S. in 1995 — an analysis on the basis of the TOP500 list. Supercomputer, 12(1):16– 22, January 1996. CODEN SPCOEL. ISSN 0168-7875.

Dongarra:1996:IVI

[342] J. Dongarra, A. Lumsdaine, R. Pozo, and K. Remington. IML++ v. 1.2: Iterative methods library reference guide. Technical report, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, 1996. URL http://www.

netlib.org/netlib/lapack/lawns/
lawn102.ps;   http://www.netlib.
org/netlib/lapack/lawnspdf/lawn102.
pdf.

Dongarra:1996:KCP

[343] Jack J. Dongarra, Sven Hammarling, and David W. Walker. Key concepts for parallel out-of-core LU factorization. LAPACK Working Note 110, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, April

1996. URL http://www.netlib.org/
lapack/lawns/lawn110.ps;   http:
//www.netlib.org/lapack/lawnspdf/
lawn110.pdf;   http://www.netlib.
org/utk/papers/outofcore/outofcore.
ps;   http://www.netlib.org/utk/
papers/outofcore/paper.html;   http:
//www.netlib.org/utk/people/JackDongarra/
pdf/outofcore.pdf. UT-CS-96-324,

April 1996.

Dongarra:1996:LF

[344] J. J. Dongarra, J. Du Croz, S. Hammarling, J. Waśniewski, and A. Zemla. LAPACK for Fortran 90. Applied Mathematics and Computer Science, 6(2):375–382, ???? 1996. CODEN AMCPE9. ISSN 0867-

857X. URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
lapack90.ps.

Dongarra:1996:LFC

[345] J. J. Dongarra, J. Du Croz, S. Hammarling, J. Waśniewski, and A. Zemla. LAPACK for Fortran 90 compiler. In Liddell et al. [1012], pages 826–833. ISBN 3-540-61142-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .H52 1996.

Dongarra:1996:LVH

[346] J. Dongarra, R. Pozo, and D. Walker. LAPACK++ V. 1.0: High performance linear algebra users’ guide. Technical report, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, 1996. URL http://www.

                                        netlib.org/netlib/lapack/lawns/
                                        lawn98.ps;   http://www.netlib.
                                        org/netlib/lapack/lawnspdf/lawn98.
                                        pdf.

Dongarra:1996:MPP

[347] Jack Dongarra and Tom Dunigan. Message-passing performance of various computers. Technical report CS-95-299, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, 1996. URL http://www.

                                        netlib.org/utk/papers/commperf.
                                        ps;   http://www.netlib.org/utk/
                                        papers/latbw.ps;   http://www.netlib.
                                        org/utk/papers/latbw/commperf.html;
                                         http://www.netlib.org/utk/people/
                                        JackDongarra/pdf/commperf.pdf;
                                        http://www.netlib.org/utk/people/
                                        JackDongarra/pdf/latbw.pdf.

Dongarra:1996:MPS

[348] Jack J. Dongarra, Steve W. Otto, Marc Snir, and David Walker. A message passing standard for MPP and workstations. Communications of the ACM, 39(7):84–90, July 1996. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Dongarra:1996:P

[349] J. J. Dongarra, H. W. Meuer, and E. Strohmaier. Preface. Supercomputer, 12(1):4–5, ???? 1996. CODEN SPCOEL. ISSN 0168-7875.

Dongarra:1996:PFI

[350] J. J. Dongarra, J. Du Croz, S. Hammarling, J. Waśniewski, and A. Zemla. A proposal for a Fortran 90 interface for LAPACK. In Dongarra et al. [1009], pages 158–165. CODEN LNCSD9. ISBN 3-540-60902-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1995.

Dongarra:1996:PMR

[351] J. J. Dongarra, T. Hey, and E. Strohmaier. PARKBENCH: methodology, relations and results. In Liddell et al. [1012], pages 770–777. ISBN 3-540-61142-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .H52 1996.

Dongarra:1996:SRP

[352] J. J. Dongarra, T. Hey, and E. Strohmaier. Selected results from the PARK-BENCH benchmark. In Bouge [1008], pages 251–254. ISBN 3-540-61626-8 (vol. 1), 3-540-61627-6 (vol. 2). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I554 1996.

Dongarra:1996:STa

[353] J. Dongarra and A. Petitet. ScaLA-PACK tutorial. In Dongarra et al. [1009], pages 166–176. CODEN LNCSD9. ISBN 3-540-60902-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1995.

Dongarra:1996:STb

[354] J. Dongarra and L. S. Blackford. ScaLAPACK tutorial. Lecture Notes in Computer Science, 1184:204–215, 1996. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://www.netlib.org/utk/
papers/scalapack-tutorial.ps;
http://www.netlib.org/utk/people/
JackDongarra/pdf/scalapack-tutorial.
pdf.

Dongarra:1996:TR

[355] Jack Dongarra, Hans Meuer, and Erich Strohmaier. Top500 report. Technical report, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, November 1996. URL http://

www.netlib.org/benchmark/top500/
lists/top500_9611.ps;   http://
www.netlib.org/benchmark/top500/
top500.list.html;   http://www.
netlib.org/utk/people/JackDongarra/
pdf/top500_9611.pdf.

Dongarra:1996:TSS

[356] J. J. Dongarra, H. W. Meuer, and E. Strohmaier. TOP500 supercomputer sites. Supercomputer, 12(1):91– 120, January 1996. CODEN SPCOEL. ISSN 0168-7875.

Fagg:1996:PIP

[357] Graham Fagg and Jack Dongarra. PVMPI: An integration of PVM and MPI systems. Calculateurs Parallèles, 8(2):151–166, 1996. CODEN ???? ISSN 1260-3198. URL http://

www.netlib.org/utk/papers/pvmpi/
paper.html;   http://www.netlib.
org/utk/papers/pvmpi/pvmpi.ps;

http://www.netlib.org/utk/people/
JackDongarra/pdf/pvmpi.pdf.

Fagg:1996:TGR

[358] G. E. Fagg, K. S. London, and J. J. Dongarra. Taskers and general resource managers: PVM supporting DCE process management. In Bode et al. [1007], pages 180– 187. ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E975 1996

Bar. URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
europvm96.ps;   http://www.netlib.
org/utk/people/JackDongarra/pdf/
europvm96.pdf.

Kim:1996:FMO

[359] Youngbae Kim, James S. Plank, and Jack J. Dongarra. Fault-tolerant matrix operations using checksum and reverse computation. In IEEE [1010], page ?? ISBN 0-8186-7551-9 (paper), 0-8186-7553-5 (microfiche). ISSN 1088-4955. LCCN QA76.58 .S95 1996. IEEE Computer Society Press Order Number PR07551. IEEE order plan catalog number 96TB100062.

Kim:1996:FTMa

[360] James S. Plank, Youngbae Kim, and Jack Dongarra. Fault tolerant matrix operations for networks of workstations using diskless checkpointing. Technical report, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, February 1996. URL http://www.

cs.utk.edu/~plank/plank/papers/
ADCKP.html.

Kim:1996:FTMb

[361] Youngbae Kim, J. S. Plank, and J. J. Dongarra. Fault tolerant matrix operations using checksum and reverse computation. In Frontiers’96: proceedings / the Sixth Symposium on the Frontiers of Massively Parallel Computing, October 27–31, 1996, Annapolis, Maryland, pages 70–77. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7551-9. LCCN QA76.58 .S95 1996. IEEE catalog number 96TB100062.

Snir:1996:MCR

[362] Marc Snir, Steve W. Otto, Steven Huss-Lederman, David W. Walker, and Jack Dongarra. MPI: the complete reference. MIT Press, Cambridge, MA, USA, 1996. ISBN 0-262-69184-1. xii + 336 pp. LCCN QA76.642.M65 1996. US$27.50.

vanderSteen:1996:ORSa

[363] Aad J. van der Steen and Jack Dongarra. Overview of recent supercomputers. Technical report, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, 1996. URL http://www.

                                        netlib.org/utk/papers/advanced-
                                        computers/;   http://www.netlib.
                                        org/utk/papers/advanced-computers/
                                        advanced-computers.ps;   http://
                                        www.netlib.org/utk/people/JackDongarra/
                                        pdf/advanced-computers.pdf.

vanderSteen:1996:ORSb

[364] Aad J. van der Steen and Jack Dongarra. Overview of recent supercomputers. NHSE Review, 1(1):??, February 10, 1996.

Walker:1996:MSM

[365] D. W. Walker and J. J. Dongarra. MPI: a standard message passing interface.

Supercomputer, 12(1):56–68, January 1996. CODEN SPCOEL. ISSN 0168-7875.

Bai:1997:SDN

[366] Z. Bai, J. Demmel, J. Dongarra, A. Petitet, H. Robinson, and K. Stanley. The spectral decomposition of nonsymmetric matrices on distributed memory parallel computers. SIAM Journal on Scientific Computing, 18(5):1446–1461, September 1997. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/28136;   http:/
/www.netlib.org/utk/papers/sign/
sign.html;   http://www.netlib.
org/utk/papers/sign/sign.ps;   http:
//www.netlib.org/utk/people/JackDongarra/
pdf/sign.pdf.

Bai:1997:TMC

[367] Z. Bai, D. Day, J. Demmel, and J. Dongarra. A test matrix collection for non-Hermitian eigenvalue problems. LA-PACK Working Note 123 and CS-97-355, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, March

1997. URL http://www.netlib.org/
lapack/lawns/lawn123.ps;   http:
//www.netlib.org/lapack/lawnspdf/

lawn123.pdf. LAPACK Working Note

123. UT-CS-97-355, March 1997.

Blackford:1997:PEN

[368] L. S. Blackford, A. Cleary, A. Petitet, R. C. Whaley, J. Demmel, I. Dhillon, H. Ren, K. Stanley, J. Dongarra, and S. Hammarling. Practical experience in the numerical dangers of heterogeneous computing. ACM Transactions on Mathematical Software, 23(2):

133–147, June 1997. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.

                                        acm.org/pubs/citations/journals/
                                        toms/1997-23-2/p133-blackford/.

Blackford:1997:SLA

[369] L. S. Blackford, J. Choi, A. Cleary, E. D’Azevedo, J. Demmel, I. Dhillon, J. Dongarra, S. Hammarling, G. Henry, A. Petitet, K. Stanley, D. Walker, and R. C. Whaley. ScaLAPACK: a linear algebra library for messagepassing computers. In Proceedings of the Eighth SIAM Conference on Parallel Processing for Scientific Computing (Minneapolis, MN, 1997), page 15 (electronic). Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1997.

                                        URL http://www.netlib.org/utk/
                                        papers/siam397-scalapack/index.
                                        html;   http://www.netlib.org/utk/
                                        papers/siam397-scalapack/siam397-
                                        scalapack.ps;   http://www.netlib.
                                        org/utk/people/JackDongarra/pdf/
                                        siam397-scalapack.pdf.

Blackford:1997:SUG

[370] L. S. Blackford, J. Choi, A. Cleary, E. D’Azevedo, J. Demmel, I. Dhillon, J. Dongarra, S. Hammarling, G. Henry, A. Petitet, K. Stanley, D. Walker, and R. C. Whaley. ScaLAPACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1997. ISBN 0-89871-400-1 (paperback), 0-89871-401-X (CD-ROM), 0-89871-397-8 (set). xxvi + 325 pp. LCCN QA185.D37 S33 1997.

Bode:1997:PEP

[371] A. Bode and J. Dongarra. Performance evaluation and prediction. Lecture Notes in Computer Science, 1300:969– 970, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Boisvert:1997:IDC

[372] Ronald Boisvert, Shirley V. Browne, Jack J. Dongarra, Eric Grosse, and Bruce Miller. Interactive and dynamic content in software repositories. Technical report UT-CS-97-351, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, February 1997.

URL http://www.netlib.org/utk/
papers/dl97/main.html;   http:/
/www.netlib.org/utk/papers/dl97/
main.ps;   http://www.netlib.org/
utk/people/JackDongarra/pdf/main.
pdf.

Boisvert:1997:MMW

[373] Ronald F. Boisvert, Roldan Pozo, Karin Remington, Richard F. Barrett, and Jack J. Dongarra. Matrix market: A. Web resource for test matrix collections. In Boisvert [1016], pages 125–137. ISBN 0-412-80530-8. LCCN QA297 .I35 1996.

URL http://www.netlib.org/utk/
papers/matrixmarket/matrixmarket.
html;   http://www.netlib.org/utk/
papers/matrixmarket/matrixmarket.
ps;   http://www.netlib.org/utk/
people/JackDongarra/pdf/matrixmarket.
pdf.

Calland:1997:TLRa

[374] Pierre-Yves Calland, Jack Dongarra, and Yves Robert. Tiling with limited resources. Technical report CS-97-350, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, February 1997.

                                        URL http://www.netlib.org/utk/
                                        papers/tiling/main.html;   http:
                                        //www.netlib.org/utk/papers/tiling/
                                        tiling.ps;   http://www.netlib.
                                        org/utk/people/JackDongarra/pdf/
                                        tiling.pdf.

Calland:1997:TLRb

[375] P. Calland, J. Dongarra, and Y. Robert. Tiling with limited resources. In Thiele [1025], pages 229–238. ISBN 0-8186-7959-X (casebound), 0-8186-7958-1, 0-8186-7960-3 (microfiche). ISSN 1063-6862. LCCN TK7874.6 .I58 1997.

Casanova:1997:JAN

[376] Henri Casanova, Jack Dongarra, and David M. Doolin. Java access to numerical libraries. Concurrency: practice and experience, 9(11):1279–1291, November 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

                                        URL http://www.netlib.org/utk/
                                        people/JackDongarra/PAPERS/jaccess.
                                        ps;   http://www.netlib.org/utk/
                                        people/JackDongarra/pdf/jaccess.
                                        pdf;   http://www3.interscience.
                                        wiley.com/cgi-bin/abstract?ID=13823;
                                         http://www3.interscience.wiley.
                                        com/cgi-bin/fulltext?ID=13823&PLACEBO=

IE.pdf. Special Issue: Java for computational science and engineering — simulation and modeling II.

Casanova:1997:NES

[377] H. Casanova and J. Dongarra. Network enabled solvers for scientific computing using the NetSolve system. In Goscinski et al. [1020], pages 17–33. ISBN 0-7803-4229-1 (softbound), 0-7803-4230-5 (microfiche). LCCN QA76.58 .I528 1997.

Casanova:1997:NNE

[378] Henri Casanova and Jack Dongarra. NetSolve: a network-enabled server for solving computational science problems. The International Journal of Supercomputer Applications and High Performance Computing, 11(3):212–223, Fall 1997. CODEN IJSCFG. ISSN 1078-

3482. URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
netsolve.ps;   http://www.netlib.
org/utk/people/JackDongarra/pdf/
netsolve.pdf.

Casanova:1997:NNS

[379] H. Casanova and J. Dongarra. Net-Solve: a network server for solving computational science problems. In Anonymous [1015], pages 24–31. ISSN 1421-6337.

Casanova:1997:UJN

[380] H. Casanova and J. Dongarra. The use of Java in the NetSolve project. In Sydow [1024], pages 791–796. ISBN 3-89685-550-6 (set), 3-89685-551-4 (vol. 1), 3-89685-552-2 (vol. 2), 3-89685-553-0 (vol. 3),3-89685-554-9 (vol. 4), 3-89685-555-7 (vol. 5), 3-89685-556-5 (vol. 6). LCCN Q183.9 .I46

1997. URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
netsolve-imacs897.ps;   http://
www.netlib.org/utk/people/JackDongarra/
pdf/netsolve-imacs897.pdf. In co-

operation with R.-P. Schafer, W. Rufeger, and H. Lehmann.

Cleary:1997:ISD

[381] A. Cleary and J. Dongarra. Implementation in ScaLAPACK of divideand-conquer algorithms for banded and tridiagonal linear systems. LA-PACK Working Note 125 and CS-97-358, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, April

1997. URL http://www.netlib.org/
lapack/lawns/lawn125.ps;   http:
//www.netlib.org/lapack/lawnspdf/

lawn125.pdf. UT-CS-97-358, April 1997.

DAzevedo:1997:DIP

[382] Ed F. D’Azevedo and J. Dongarra. The design and implementation of the parallel out-of-core ScaLAPACK LU, QR and Cholesky factorization routines. Technical report, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, January

1997. URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
ooc-scalapack.ps;   http://www.
netlib.org/utk/people/JackDongarra/
pdf/ooc-scalapack.pdf.

Desprez:1997:DITa

[383] F. Desprez, J. Dongarra, F. Rastello, and Y. Robert. Determining the idle time of a tiling: New results. In IEEE [1023], pages 307–317. ISBN 0-8186-8090-3. LCCN QA76.58 .I5445 1997.

Desprez:1997:DITb

[384] F. Desprez, J. Dongarra, F. Rastello, and Y. Robert. Determining the idle time of a tiling: New results. Journal of Information Science and Engineering, 14(1):167–190, March 1997. CODEN JINEEY. ISSN 1016-2364.

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/idle-
tiling.ps.

Desprez:1997:SBC

[385] Frederic Desprez, Jack Dongarra, Antoine Petitet, Cyril Randriamaro, and Yves Robert. Scheduling blockcyclic array redistribution. LAPACK Working Note 120 and UT-CS-97-349, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, February

1997. URL http://www.netlib.org/
lapack/lawns/lawn120.ps;   http:
//www.netlib.org/lapack/lawnspdf/
lawn120.pdf;   http://www.netlib.
org/utk/papers/redist/redist.ps;
http://www.netlib.org/utk/people/
JackDongarra/pdf/redist.pdf. LA-

PACK Working Note 120. UT-CS-97-349, February 1997.

Dongarra:1997:BAR

[386] J. J. Dongarra, F. Desprez, A. Petitet, and C. Randriamaro. Block-cyclic array redistribution on networks of workstations. Lecture Notes in Computer Science, 1332:343–350, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongarra:1997:BCA

[387] J. J. Dongarra, F. Desprez, A. Petitet, and C. Randriamaro. Block-cyclic array redistribution on networks of workstations. Lecture Notes in Computer Science, 1332:343–350, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongarra:1997:CLA

[388] Jack J. Dongarra and David W. Walker. Constructing linear algebra software libraries for high-performance computers. In Iterative methods in scientific computing (Hong Kong, 1995), pages 111–167. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997.

Dongarra:1997:CSD

[389] J. J. Dongarra, S. Hammarling, and A. Petitet. Case studies on the development of ScaLAPACK and the NAG numerical PVM library. In Boisvert [1016], pages 236–248. ISBN 0-412-80530-8. LCCN QA297 .I35

1996. URL http://www.netlib.org/
utk/papers/woco96/woco96.html;
http://www.netlib.org/utk/papers/
woco96/woco96.ps;   http://www.
netlib.org/utk/people/JackDongarra/
pdf/woco96.pdf.

Dongarra:1997:CTH

[390] J. Dongarra, H. W. Meuer, H. D. Simon, and E. Strohmaier. Changing technologies of HPC. Future Generation Computer Systems, 12(5):461–474, April 1997. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (elec-

tronic). URL http://www.netlib.
org/utk/people/JackDongarra/PAPERS/
ct-hpc.ps;   http://www.netlib.
org/utk/people/JackDongarra/pdf/
ct-hpc.pdf.

Dongarra:1997:DIP

[391] J. J. Dongarra and E. F. D’Azevedo. The design and implementation of the parallel out-of-core ScaLAPACK LU, QR, and Cholesky factorization routines. LAPACK Working Note 118, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, January

1997. URL http://www.netlib.org/
lapack/lawns/lawn118.ps;   http:
//www.netlib.org/lapack/lawnspdf/

lawn118.pdf. UT-CS-97-347, January 1997.

Dongarra:1997:DMI

[392] Jack Dongarra, Greg Henry, and David Watkins. A distributed memory implementation of the nonsymmetric QR algorithm. In Proceedings of the Eighth SIAM Conference on Parallel Processing for Scientific Computing (Minneapolis, MN, 1997), page 8 (electronic). Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1997.

Dongarra:1997:FTM

[393] J. Dongarra, Y. Kim, and J. Plank. Fault tolerant matrix operations for networks of workstations using multiple checkpointing. In IEEE [1022], pages 460–465. ISBN 0-8186-7901-8, 0-8186-7902-6 (casebound) , 0-8186-7903-4 (microfiche). LCCN QA76.88 .H653 1997. URL http://

www.netlib.org/utk/papers/hpc97/
hpc97.ps;   http://www.netlib.org/
utk/papers/hpc97/index.html;   http:
//www.netlib.org/utk/people/JackDongarra/

pdf/hpc97.pdf. IEEE Order Plan number 97TB100110.

Dongarra:1997:HPC

[394] J. J. Dongarra and H. D. Simon. High performance computing in the U.S. in 1996 — an analysis on the basis of the TOP500 list. Supercomputer, 13(1):19– 28, ???? 1997. CODEN SPCOEL. ISSN 0168-7875.

Dongarra:1997:KCPa

[395] Jack Dongarra, Sven Hammarling, and David W. Walker. Key concepts for parallel out-of-core LU factorization. Technical report, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, 1997.

                                        URL http://www.netlib.org/utk/
                                        people/JackDongarra/PAPERS/ooc.
                                        ps;   http://www.netlib.org/utk/
                                        people/JackDongarra/pdf/ooc.pdf.

Dongarra:1997:KCPb

[396] Jack J. Dongarra, Sven Hammarling, and David W. Walker. Key concepts for parallel out-of-core LU factorization. Parallel Computing, 23(1–2): 49–70, April 16, 1997. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Dongarra:1997:MPP

[397] Jack J. Dongarra and Tom Dunigan. Message-passing performance of various computers. Concurrency: practice and experience, 9(10): 915–926, October 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http://

                                        www3.interscience.wiley.com/cgi-
                                        bin/abstract?ID=13808;   http://
                                        www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=13808&PLACEBO=IE.
                                        pdf.

Dongarra:1997:P

[398] J. J. Dongarra, H. W. Meuer, and E. Strohmaier. Preface. Supercomputer, 13(1):4–5, ???? 1997. CODEN SPCOEL. ISSN 0168-7875.

Dongarra:1997:PAH

[399] J. Dongarra, S. Browne, and H. Casanova. Providing access to high performance computing technologies. Lecture Notes in Computer Science, 1215:24–34, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongarra:1997:PSI

[400] Jack Dongarra and Bernard Tourancheau. Preface to the special issue. The International Journal of Supercomputer Applications and High Performance Computing, 11(2):83, Summer 1997. CODEN IJSCFG. ISSN 1078-3482.

Dongarra:1997:TSS

[401] J. J. Dongarra, H. W. Meuer, and E. Strohmaier. TOP500 supercomputer sites. Supercomputer, 13(1):89– 120, ???? 1997. CODEN SPCOEL. ISSN 0168-7875.

Dongarra:1997:UGB

[402] J. Dongarra and R. C. Whaley. A user’s guide to the BLACS v1.1. LAPACK Working Note 94, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, May 5,

1997. URL http://www.netlib.org/
lapack/lawns/lawn94.ps;   http:/
/www.netlib.org/lapack/lawnspdf/

lawn94.pdf. Updated May 5, 1997 (Version 1.1).

Dongarra:1997:WET

[403] J. Dongarra and B. Tourancheau. Workshop on environments and tools for parallel scientific computing. Parallel Computing, 23(1–2):1–4, April 16, 1997. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Doolin:1997:JCL

[404] David M. Doolin and Jack Dongarra. JLAPACK — compiling LAPACK Fortran to Java, phase 1. Technical report CS-97-367, University of Tennessee, Knoxville, Knoxville, TN 37996, USA,

1997. URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
f2jpaper.ps;   http://www.netlib.
org/utk/people/JackDongarra/pdf/
f2jpaper.pdf.

Fagg:1997:HMAa

[405] G. Fagg, J. Dongarra, and A. Geist. Heterogeneous MPI application interoperation and process management under PVMPI. Technical report CS-97-???, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, June 1997. URL http://www.

netlib.org/utk/papers/pvmmpi97.
ps;   http://www.netlib.org/utk/
people/JackDongarra/pdf/pvmmpi97.
pdf.

Fagg:1997:HMAb

[406] G. E. Fagg, J. J. Dongarra, and A. Geist. Heterogeneous MPI application interoperation and process management under PVMPI. Lecture Notes in Computer Science, 1332:91– 98, 1997. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Fischer:1997:AAP

[407] Markus Fischer and Jack Dongarra. Another architecture: PVM on Windows 95/NT. In ????, editor, Concurrent Computing Conference, Atlanta, GA, March 10– 11, 1994, page ?? ????, ????,

1997. URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
nt-paper.ps;   http://www.netlib.
org/utk/people/JackDongarra/pdf/
nt-paper.pdf.

Henry:1997:PIN

[408] G. Henry, D. Watkins, and J. Dongarra. A parallel implementation of the nonsymmetric QR algorithm for distributed memory architectures. LAPACK Working Note 121, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, March

1997. URL http://www.netlib.org/
lapack/lawns/lawn121.ps;   http:
//www.netlib.org/lapack/lawnspdf/

lawn121.pdf. UT-CS-97-352, March 1997.

Moore:1997:SNI

[409] Keith Moore, Graham E. Fagg, Al Geist, and Jack Dongarra. Scalable networked information processing environment (SNIPE). In ACM [1014], page ?? ISBN 0-89791-985-8. LCCN QA76.9.A25 A265

1997. URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
snipe.html;   http://www.supercomp.
org/sc97/proceedings/TECH/MOORE/

INDEX.HTM. ACM SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Plank:1997:FTM

[410] James S. Plank, Youngbae Kim, and Jack J. Dongarra. Faulttolerant matrix operations for networks of workstations using diskless checkpointing. Journal of Parallel and Distributed Computing, 43(2): 125–138, June 15, 1997. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.idealibrary.com/links/doi/
10.1006/jpdc.1997.1336/production;
 http://www.idealibrary.com/links/

                                        doi/10.1006/jpdc.1997.1336/production/
                                        pdf;   http://www.idealibrary.com/
                                        links/doi/10.1006/jpdc.1997.1336/
                                        production/ref.

Strohmaier:1997:EHM

[411] E. Strohmaier, J. J. Dongarra, H. W. Meuer, and H. D. Simon. Evolution of the HPC market. In Grandinetti et al. [1021], pages 27–44. ISBN 0-7923-4550-

9. LCCN QA76.88 .A38 1997.

Strohmaier:1997:HPC

[412] E. Strohmaier, J. J. Dongarra, H. W. Meuer, and H. D. Simon. High-performance computing in industry. Supercomputer, 13(1):74–88, ???? 1997. CO-DEN SPCOEL. ISSN 0168-7875.

                                        URL http://www.netlib.org/utk/
                                        papers/top500-indust/paper.html;
                                         http://www.netlib.org/utk/papers/
                                        top500-indust/paper.ps;   http:/
                                        /www.netlib.org/utk/people/JackDongarra/
                                        pdf/top500ind.pdf.

Watkins:1997:PIN

[413] Greg Henry, David Watkins, and Jack Dongarra. A parallel implementation of the nonsymmetric QR algorithm for distributed memory architectures. Technical report CS-97-352, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, March 1997. URL http://www.

                                        netlib.org/netlib/lapack/lawns/
                                        lawn121.ps;   http://www.netlib.
                                        org/netlib/lapack/lawnspdf/lawn121.

pdf. LAPACK Working Note 121.

Whaley:1997:ATL

[414] R. Whaley and J. Dongarra. Automatically tuned linear algebra software. LAPACK Working Note 131, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, December

1997. URL http://www.netlib.org/
lapack/lawns/lawn131.ps;   http:
//www.netlib.org/lapack/lawnspdf/

lawn131.pdf. UT-CS-97-366, December 1997.

Blackford:1998:IGD

[415] L. S. Blackford, J. J. Dongarra, C. A. Papadopoulos, and R. C. Whaley. Installation guide and design of the HPF 1.1 interface to ScaLAPACK, SLHPF. LAPACK Working Note 137, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, August

1998. URL http://www.netlib.org/
lapack/lawns/lawn137.ps;   http:
//www.netlib.org/lapack/lawnspdf/

lawn137.pdf. UT-CS-98-396, August 1998.

Boisvert:1998:DNLa

[416] Ronald F. Boisvert, Jack J. Dongarra, Roldan Pozo, Karin A. Remington, and G. W. Stewart. Developing numerical libraries in Java. In ACM [1026], page ?? ISBN ????

LCCN ????  URL http://www.
cs.ucsb.edu/conferences/java98/
papers/jnt.pdf;   http://www.cs.
ucsb.edu/conferences/java98/papers/
jnt.ps;   http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
jnt.ps;   http://www.netlib.org/
utk/people/JackDongarra/pdf/jnt.

pdf. Possibly unpublished, except electronically.

Boisvert:1998:DNLb

[417] Ronald F. Boisvert, Jack J. Dongarra, Roldan Pozo, Karin A. Remington, and G. W. Stewart. Developing numerical libraries in Java. Concurrency: practice and experience, 10(11–13):1117–1129, September 1998. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http://www3.

                                        interscience.wiley.com/cgi-bin/
                                        abstract?ID=10050395;   http://
                                        www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=10050395&PLACEBO=

IE.pdf. Special Issue: Java for Highperformance Network Computing.

Boisvert:1998:UIS

[418] R. Boisvert, Shirley Browne, Jack Dongarra, E. Grosse, and B. Miller. User interfaces for software repositories. Technical report, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, 1998. URL http:

                                        //www.cs.utk.edu/~swells/papers/

ir_draft.doc. Submitted to Information Retrieval Journal.

Browne:1998:RPA

[419] Shirley Browne, Jack Dongarra, and Kevin London. Review of performance analysis tools for MPI parallel programs. NHSE Review, 3,

1998.  CODEN ????  ISSN

                                        ???? URL http://www.cs.utk.edu/
                                        ~browne/perftools-review/.  Ac-

cepted, to appear.

Casanova:1998:ANN

[420] Henri Casanova and Jack Dongarra. Applying NetSolve’s network-enabled server. IEEE Computational Science & Engineering, 5(3):57–67, July/ September 1998. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://dlib.

                                        computer.org/cs/books/cs1998/pdf/

c3057.pdf;   http://www.computer.
org/cse/cs1998/c3057abs.htm;   http:
//www.netlib.org/utk/people/JackDongarra/
PAPERS/netsolve-nes.ps;   http:/
/www.netlib.org/utk/people/JackDongarra/
pdf/netsolve-nes.pdf.

Casanova:1998:ETH

[421] H. Casanova and J. Dongarra. Enabling technologies for high performance computing. In Papailiou [1032], pages 71–76. ISBN 0-471-98579-1 (vol. 1), 0-471-98580-5 (vol. 2). LCCN QA911 .E95 1998. Three volumes in two books.

Casanova:1998:NES

[422] H. Casanova, J. J. Dongarra, and K. Moore. Network-enabled solvers and the NetSolve project. SIAM News, 31(1):??, 1998. ISSN 0036-

1437. URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
netsolve-siam.ps;   http://www.
netlib.org/utk/people/JackDongarra/
pdf/netsolve-siam.pdf.

Casanova:1998:NNE

[423] H. Casanova and J. J. Dongarra. Net-Solve: a network-enabled solver; examples and users. In Antonio [1029], pages 19–28. ISBN 0-8186-8365-1, 0-8186-8367-8 (microfiche). LCCN QA76.88 .H48 1998. IEEE catalog number 98EX126.

Casanova:1998:NVD

[424] H. Casanova and J. Dongarra. Net-Solve version 1.2: Design and implementation. LAPACK Working Note 140, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, November

1998. URL http://www.netlib.org/
lapack/lawns/lawn140.ps;   http:
//www.netlib.org/lapack/lawnspdf/

lawn140.pdf. UT-CS-98-406, Nov 1998.

Casanova:1998:UAB

[425] Henri Casanova and Jack Dongarra. Using agent-based software for scientific computing in the NetSolve system. Parallel Computing, 24(12–13): 1777–1790, November 1, 1998. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/cas/tree/store/
parco/sub/1998/24/12-13/1354.pdf;
 http://www.netlib.org/utk/people/
JackDongarra/PAPERS/netsolve-agent.
ps;   http://www.netlib.org/utk/
people/JackDongarra/pdf/netsolve-
agent.pdf.

DAzevedo:1998:PSE

[426] E. D’Azevedo and J. Dongarra. Packed storage extensions for ScaLA-PACK. LAPACK Working Note 135, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, April

1998. URL http://www.netlib.org/
lapack/lawns/lawn135.ps;   http:
//www.netlib.org/lapack/lawnspdf/

lawn135.pdf. UT-CS-98-385, April 1998.

Desprez:1998:DIT

[427] F. Desprez, J. Dongarra, F. Rastello, and Y. Robert. Determining the idle time of a tiling: New results. Journal of Information Science and Engineering, 14(1):167–190, ???? 1998. CODEN JI-NEEY. ISSN 1016-2364.

Desprez:1998:MSB

[428] F. Desprez, S. Domas, J. Dongarra, and A. Petitet. More on scheduling block-cyclic array redistribution. Lecture Notes in Computer Science, 1511:275–287, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/lcr98.
ps.

Desprez:1998:SBA

[429] F. Desprez, J. Dongarra, A. Petitet, and C. Randriamaro. Scheduling block-cyclic array redistribution. In D’Hollander et al. [1030], pages 227– 234. ISBN 0-444-82882-6. LCCN QA76.58.P3795 1997.

Desprez:1998:SBC

[430] F. Desprez, J. Dongarra, A. Petitet, C. Randriamaro, and Y. Robert. Scheduling block-cyclic array redistribution. IEEE Transactions on Parallel and Distributed Systems, 9(2):192–205, February 1998. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

URL http://dlib.computer.org/
td/books/td1998/pdf/l0192.pdf;
http://www.computer.org/tpds/td1998/
l0192abs.htm.

Dongarra:1998:HHA

[431] J. Dongarra, G. Fagg, A. Geist, and J. A. Kohl. HARNESS: Heterogeneous adaptable reconfigurable NEtworked systems. In IEEE [1031], pages 358–359. ISBN 0-8186-8579-4, 0-8186-8581-6. ISSN 1082-8907. LCCN QA76.9.D5 I157 1998. IEEE Order Plan Catalog Number 98TB100244.

IEEE Computer Society Press order number PR08579.

Dongarra:1998:HPL

[432] J. Dongarra and J. Waśniewski. High performance linear algebra package LAPACK90. Lecture Notes in Computer Science, 1388:387–391, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Dongarra:1998:KCP

[433] J. J. Dongarra, S. Hammarling, and D. W. Walker. Key concepts for parallel out-of-core LU factorization. Computers and Mathematics with Applications, 35(7):13–31, ???? 1998. CODEN CMAPDK. ISSN 0898-1221 (print), 1873-7668 (electronic).

Dongarra:1998:NLA

[434] Jack J. Dongarra, Iain S. Duff, Danny C. Sorensen, and Henk A. van der Vorst. Numerical Linear Algebra for High-Performance Computers. Software, Environments, and Tools. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1998. ISBN 0-89871-428-1. xviii + 342 pp. LCCN QA76.88 .N86 1998. URL http://

                                        www.loc.gov/catdir/enhancements/
                                        fy0726/98044444-d.html;   http:/
                                        /www.loc.gov/catdir/enhancements/
                                        fy0726/98044444-t.html.

Dongarra:1998:TSL

[435] J. Dongarra, W. Owczarz, J. Wasniewski, and P. Yalamov. Testing software for LAPACK90. LA-PACK Working Note 138, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, September

1998. URL http://www.netlib.org/
lapack/lawns/lawn138.ps;   http:
//www.netlib.org/lapack/lawnspdf/

lawn138.pdf. UT-CS-98-401, Sept 1998.

Fagg:1998:MMH

[436] G. E. Fagg, K. S. London, and J. J. Dongarra. MPIConnect: Managing heterogeneous MPI applications interoperation and process control. Lecture Notes in Computer Science, 1497:93– 96, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Gropp:1998:MCR

[437] William Gropp, Steven Huss-Lederman, Andrew Lumsdaine, Ewing Lusk, Bill Nitzberg, William Saphir, and Marc Snir. MPI: The Complete Reference. Volume 2, The MPI-2 Extensions. Scientific and Engineering Computation. MIT Press, Cambridge, MA, USA, second edition, 1998. ISBN 0-262-57123-4 (vol. 2), 0-262-69216-3 (set). 350 pp. LCCN QA76.642 .M65 1998. US$30 (paperback). URL http:/

/mitpress.mit.edu/book-home.tcl?
isbn=0262571234. See also volume 1

[443].

Migliardi:1998:DRV

[438] M. Migliardi, J. Dongarra, A. Geist, and V. Sunderam. Dynamic reconfiguration and virtual machine management in the Harness metacomputing system. Lecture Notes in Computer Science, 1505:127–134, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

URL http://www.netlib.org/utk/

                                        people/JackDongarra/PAPERS/harness1.
                                        ps.

Petitet:1998:ARM

[439] A. Petitet and J. Dongarra. Algorithmic redistribution methods for block cyclic distributions. LA-PACK Working Note 133, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, March

                                        1998. URL http://www.netlib.org/
                                        lapack/lawns/lawn133.ps;   http:
                                        //www.netlib.org/lapack/lawnspdf/

lawn133.pdf. UT-CS-98-383, March 1998.

Petitet:1998:NLA

[440] A. Petitet, H. Casanova, J. Dongarra, Y. Robert, and R. C. Whaley. A numerical linear algebra problem solving environment designer’s perspective. LAPACK Working Note 139, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, October

                                        1998. URL http://www.netlib.org/
                                        lapack/lawns/lawn139.ps;   http:
                                        //www.netlib.org/lapack/lawnspdf/

lawn139.pdf. UT-CS-98-405, Oct 1998.

Plank:1998:DFT

[441] J. S. Plank, H. Casanova, M. Beck, and J. Dongarra. Deploying faulttolerance and task migration with Net-Solve. Lecture Notes in Computer Science, 1541:418–432, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Saltz:1998:PTE

[442] Joel H. Saltz, Alan Sussman, Susan Graham, James Demmel, Scott Baden, and Jack Dongarra. Programming tools and environments. Communications of the ACM, 41(11):64–73, November 1998. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

URL http://www.acm.org:80/pubs/
citations/journals/cacm/1998-41-
11/p64-sussman/.

Snir:1998:MCR

[443] Marc Snir, Steve W. Otto, Steven Huss-Lederman, David W. Walker, and Jack Dongarra. MPI: The Complete Reference. Volume 1, The MPI-1 Core. MIT Press, Cambridge, MA, USA, second edition, September 1998. ISBN 0-262-69215-5. 450 pp. LCCN QA76.642 .M65 1998. US$35 (paperback). URL http:/

/mitpress.mit.edu/book-home.tcl?
isbn=0262692155. See also volume 2

[437].

Tisseur:1998:PDC

[444] F. Tisseur and J. Dongarra. Parallelizing the divide and conquer algorithm for the symmetric tridiagonal eigenvalue problem on distributed memory architectures. LAPACK Working Note 132, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, March

1998. URL http://www.netlib.org/
lapack/lawns/lawn132.ps;   http:
//www.netlib.org/lapack/lawnspdf/

lawn132.pdf. UT-CS-98-382, March 1998.

Wasniewski:1998:HPLa

[445] J. Wasniewski and J. Dongarra. High performance linear algebra package — LAPACK90. LAPACK Working Note 134, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, April

                                        1998. URL http://www.netlib.org/
                                        lapack/lawns/lawn134.ps;   http:
                                        //www.netlib.org/lapack/lawnspdf/

lawn134.pdf. UT-CS-98-384, April

1998.

Wasniewski:1998:HPLb

[446] J. Waśniewski and J. Dongarra. High performance linear algebra package for FORTRAN 90. Lecture Notes in Computer Science, 1541:579–581, 1998. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic).

Whaley:1998:ATL

[447] R. Clint Whaley and Jack J. Dongarra. Automatically Tuned Linear Algebra Software (ATLAS). In ACM [1027], page ???? ISBN ???? LCCN

                                        ???? URL http://www.netlib.org/
                                        utk/people/JackDongarra/PAPERS/
                                        atlas-sc98.ps;   http://www.supercomp.
                                        org/sc98/TechPapers/sc98_FullAbstracts/
                                        Whaley814/INDEX.HTM. Best Paper

Award for Systems.

Anderson:1999:LUG

[448] E. Anderson, Z. Bai, C. Bischof, S. Blackford, J. Demmel, J. Dongarra, J. Du Croz, A. Greenbaum, S. Hammarling, A. McKenney, and D. Sorensen. LAPACK Users’ Guide. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, third edition, 1999. ISBN 0-89871-447-

8. xxi + 407 pp. LCCN QA76.73.F25

L36 1999.

Arbenz:1999:CPSa

[449] P. Arbenz, A. Cleary, J. Dongarra, and M. Hegland. A comparison of parallel solvers for diagonally dominant and general narrow-banded linear systems. LAPACK Working Note 142, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, February 1999. 18

pp. URL http://www.netlib.org/
lapack/lawns/lawn142.ps;   http:
//www.netlib.org/lapack/lawnspdf/
lawn142.pdf;   http://www.netlib.
org/utk/people/JackDongarra/PAPERS/

band1.ps. UT-CS-99-414, Feb 1999.

Arbenz:1999:CPSb

[450] P. Arbenz, A. Cleary, J. Dongarra, and M. Hegland. A comparison of parallel solvers for diagonally dominant and general narrow-banded linear systems II. LAPACK Working Note 143, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, May 1999.

10 pp. URL http://www.netlib.org/
lapack/lawns/lawn143.ps;   http:
//www.netlib.org/lapack/lawnspdf/
lawn143.pdf;   http://www.netlib.
org/utk/people/JackDongarra/PAPERS/

band2.ps. UT-CS-99-415, May 1999.

Arbenz:1999:CPSc

[451] P. Arbenz, A. Cleary, J. Dongarra, and M. Hegland. A comparison of parallel solvers for diagonally dominant and general narrow-banded linear systems. Parallel and Distributed Computing Practices, 2(4):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL http:

//www.cs.okstate.edu/~pdcp/vols/
vol02/vol02no4abs.html#arbenz.

Barker:1999:LUG

[452] V. Barker, S. Blackford, J. Dongarra, J. DuCroz, S. Hammarling, J. Waśniewski, and P. Yalamov. LA-PACK95 users’ guide (draft). Technical report, University of Tennessee, Knoxville, Knoxville, TN 37996, USA,

                                        1999. URL http://www.netlib.org/
                                        utk/people/JackDongarra/PAPERS/
                                        lapack95ug.ps.

Beck:1999:HNG

[453] Micah Beck, Jack J. Dongarra, Graham E. Fagg, G. Al Geist, Paul Gray, James Kohl, Mauro Migliardi, Keith Moore, Terry Moore, Philip Papadopoulous, Stephen L. Scott, and Vaidy Sunderam. HARNESS: a next generation distributed virtual machine. Future Generation Computer Systems, 15(5–6):571–582, October 1, 1999. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/19/19/30/21/20/abstract.html;
                                         http://www.netlib.org/utk/people/
                                        JackDongarra/PAPERS/harness2.ps.

Beck:1999:LQS

[454] Micah Beck, Henri Casanova, Jack Dongarra, Terry Moore, Jim Plank, Francine Berman, and Rich Wolski. Logistical quality of service in NetSolve. Computer Communications, 22(11):1034–1044, 1999. CODEN COCOD7. ISSN 0140-3664 (print), 1873-703X (electronic).

                                        URL http://www.netlib.org/utk/
                                        people/JackDongarra/PAPERS/final-
                                        lqos-in-netsolve.pdf.

Berry:1999:AOP

[455] Mike Berry and Jack Dongarra. Atlanta organizers put mathematics to work for the math sciences community. SIAM News, 32(6), 1999. ISSN 0036-

                                        1437. URL http://www.netlib.org/

utk/people/JackDongarra/PAPERS/
atlanta.pdf.

Boulet:1999:AIH

[456] Pierre Boulet, Jack Dongarra, Fabrice Rastello, Yves Robert, and Frederic Vivien. Algorithmic issues on heterogeneous computing platforms. Parallel Processing Letters, 9(2):197–213, 1999. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/blackberry-
99.ps.

Boulet:1999:STH

[457] Pierre Boulet, Jack Dongarra, Yves Robert, and Frédéric Vivien. Static tiling for heterogeneous computing platforms. Parallel Computing, 25 (5):547–568, May 1, 1999. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.com/cas/tree/store/
parco/sub/1999/25/5/1404.pdf;
http://www.netlib.org/utk/people/
JackDongarra/PAPERS/static-tiling-
hetero.ps.

Browne:1999:NLT

[458] Shirley Browne, Jack Dongarra, and Anne Trefethen. Numerical libraries and tools for scalable parallel cluster computing. In ????, editor, IEEE Cluster Computing BOF at SC99, November, 1999, Portland OR, page ?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA,

1999. URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
libraries-tools1.pdf.

Calland:1999:TSC

[459] Pierre-Yves Calland, Jack Dongarra, and Yves Robert. Tiling on systems with communication/ computation overlap. Concurrency: practice and experience, 11(3):139–153, March 1999. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (elec-

                                        tronic). URL http://www.netlib.
                                        org/utk/people/JackDongarra/PAPERS/
                                        tiling-concur.ps;   http://www3.
                                        interscience.wiley.com/cgi-bin/
                                        abstract?ID=61003665;   http://
                                        www3.interscience.wiley.com/cgi-
                                        bin/fulltext?ID=61003665&PLACEBO=
                                        IE.pdf.

Casanova:1999:AST

[460] Henri Casanova, MyungHo Kim, James S. Plank, and Jack J. Dongarra. Adaptive scheduling for task farming with grid middleware. The International Journal of High Performance Computing Applications, 13 (3):231–240, Fall 1999. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http://

                                        journals.sagepub.com/doi/pdf/10.
                                        1177/109434209901300306;   http:
                                        //www.netlib.org/utk/people/JackDongarra/
                                        PAPERS/netsolve-blackberry98.ps.

Casanova:1999:PUD

[461] H. Casanova and J. Dongarra. Providing uniform dynamic access to numerical software. In Heath et al. [1034], pages 345–366. ISBN 0-387-98680-4. LCCN QA76.58 .A543 1999.

                                        URL http://www.loc.gov/catdir/
                                        enhancements/fy0817/98033425-t.
                                        html.

Casanova:1999:SPP

[462] Henri Casanova, Michael G. Thomason, and Jack J. Dongarra. Stochastic performance prediction for iterative algorithms in distributed environments. Journal of Parallel and Distributed Computing, 58(1): 68–91, July 1999. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.idealibrary.com/links/doi/
10.1006/jpdc.1999.1541/production;
 http://www.idealibrary.com/links/
doi/10.1006/jpdc.1999.1541/production/
pdf;   http://www.idealibrary.com/
links/doi/10.1006/jpdc.1999.1543/
production;   http://www.idealibrary.
com/links/doi/10.1006/jpdc.1999.
1543/production/pdf;   http://www.
netlib.org/utk/people/JackDongarra/
PAPERS/henri-thesis-jpdc.ps.

Dongarra:1999:MPS

[463] Jack Dongarra, Graham Fagg, Rolf Hempel, and David Walker. Message passing software systems. In Webster [1037], page ?? ISBN 0-471-13946-7. LCCN TK9 .W55 1999.

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/mpss-

ency.ps. 24 volumes.

Dongarra:1999:NLAa

[464] Jack Dongarra and Victor Eijkhout. Numerical linear algebra algorithms and software. Journal CAM (Numerical) Linear Algebra, 31(4):??, October 28, 1999. CODEN ???? ISSN

???? URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
la-na20th-jcam.ps.

Dongarra:1999:NLAb

[465] Jack Dongarra and Victor Eijkhout. Numerical linear algebra. In Allen Kent and James Williams, editors, Encyclopedia of Computer Science and Technology, pages 207–233. Marcel Dekker, New York, NY, USA, August

                                        1999. URL http://www.netlib.org/
                                        utk/people/JackDongarra/PAPERS/
                                        la-ency.ps.

Dongarra:1999:P

[466] J. Dongarra and E. J. Kontoghiorghes. Preface. Parallel and Distributed Computing Practices, 2(4):??, ???? 1999. CODEN ???? ISSN 1097-2803. URL

                                        http://www.cs.okstate.edu/~pdcp/
                                        vols/vol02/vol02no4preface.html.

Dongarra:1999:SII

[467] Jack J. Dongarra and Bernard Tourancheau. Special issue introduction: Clusters and computational grids for scientific computing. The International Journal of High Performance Computing Applications, 13(3):179, Fall 1999. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL

                                        http://journals.sagepub.com/doi/
                                        pdf/10.1177/109434209901300301.

Dongarra:1999:TR

[468] Jack Dongarra, Hans-Werner Meuer, and Erich Strohmaier. Top500 report. Technical report, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, 1999. URL http://

                                        www.netlib.org/benchmark/top500/
                                        lists/top500_9911.ps.

Doolin:1999:JCL

[469] D. Doolin, J. Dongarra, and K. Seymour. JLAPACK — compiling LAPACK Fortran to Java. Scientific Programming, 7(2):111–138, 1999. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/f2jrep~

1.pdf. The software is available on the World-Wide Web at http://www.cs.

utk.edu/f2j/.

Fagg:1999:SNI

[470] Graham E. Fagg, Keith Moore, and Jack J. Dongarra. Scalable Networked Information Processing Environment (SNIPE). Future Generation Computer Systems, 15(5–6): 595–605, October 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

http://www.elsevier.com/gej-ng/
10/19/19/30/21/22/abstract.html;
 http://www.netlib.org/utk/people/
JackDongarra/PAPERS/snipe-fgcs.
ps.

Fischer:1999:EWN

[471] Markus Fischer and Jack Dongarra. Experiences with Windows 95/NT as a cluster computing platform for parallel computing. Parallel and Distributed Computing Practices, 2(2):??, February 1999. CODEN ???? ISSN 1097-2803.

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/pdcp.
pdf.

Petitet:1999:ARM

[472] A. P. Petitet and J. J. Dongarra. Algorithmic redistribution methods for block-cyclic decompositions. IEEE Transactions on Parallel and Distributed Systems, 10(12):1201–1216, December 1999.

CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

                                        URL http://dlib.computer.org/
                                        td/books/td1999/pdf/l1201.pdf;
                                        http://www.computer.org/tpds/td1999/
                                        l1201abs.htm;   http://www.netlib.
                                        org/utk/people/JackDongarra/PAPERS/
                                        alg-dist.ps;   http://www.netlib.
                                        org/utk/people/JackDongarra/pdf/
                                        alg-dist.pdf.

Petitet:1999:NLA

[473] A. Petitet, H. Casanova, R. Whaley, J. Dongarra, and Y. Robert. A numerical linear algebra problem solving environment designer’s perspective. In SIAM Annual Meeting, Atlanta, GA, May 13, 1999. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA,

                                        1999. URL http://www.netlib.org/
                                        utk/people/JackDongarra/PAPERS/
                                        la-handbook-chp10.ps.

Plank:1999:DFT

[474] James S. Plank, Henri Casanova, Micah Beck, and Jack J. Dongarra. Deploying fault-tolerance and task migration with NetSolve. Future Generation Computer Systems, 15(5–6): 745–755, October 1, 1999. CO-DEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/19/19/30/21/34/abstract.html;
                                         http://www.netlib.org/utk/people/
                                        JackDongarra/PAPERS/netsolve-ft-
                                        tm.pdf.

Strohmaier:1999:MHP

[475] Erich Strohmaier, Jack J. Dongarra, Hans W. Meuer, and Horst D. Simon. The marketplace of highperformance computing. Parallel Computing, 25(13–14):1517–1544, December 1999. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.nl/gej-ng/10/35/21/
32/36/24/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/35/
21/32/36/24/article.pdf;   http:
//www.netlib.org/utk/people/JackDongarra/
PAPERS/pc-benchmarking.pdf.

Tisseur:1999:PDC

[476] Françoise Tisseur and Jack Dongarra. A parallel divide and conquer algorithm for the symmetric eigenvalue problem on distributed memory architectures. SIAM Journal on Scientific Computing, 20(6):2223–2236, November 1999. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

URL http://epubs.siam.org/sam-
bin/dbq/article/33695;   http://
www.netlib.org/utk/people/JackDongarra/
PAPERS/lawn132.ps;   http://www.
netlib.org/utk/people/JackDongarra/
pdf/lawn132.pdf.

Beguelin:19xx:PSS

[477] A. Beguelin, J. J. Dongarra, G. A. Geist, R. Manchek, and V. S. Sunderam. PVM software system and documentation. Email to

netlib@ornl.gov, 19xx.

Arnold:2000:NEP

[478] Dorian C. Arnold and Jack Dongarra. The NetSolve environment: Progressing towards the seamless grid. In Sadayappan [1040], pages 199–

206. ISBN 0-7695-0771-9, 0-7695-0773-5 (microfiche). LCCN QA76.58

.I575 2000. URL http://ieeexplore.
ieee.org/lpdocs/epic03/;   http:
//www.netlib.org/utk/people/JackDongarra/
PAPERS/netsolve-icpp2000.pdf. IEEE

Computer Society Order Number PR00771.

Arnold:2000:SRA

[479] Dorian C. Arnold, Shirley Browne, Jack Dongarra, Graham Fagg, and Keith Moore. Secure remote access to numerical software and computational hardware. In ????, editor, Proceedings of the DoD HPC Users Group Conference (HPCUG 2000), Albuquerque, New Mexico, Albuquerque Hilton Hotel, June 5–8, 2000, page ?? ????, ????, 2000. URL http:

//www.hpcm.dren.net/Htdocs/UGC/
UGC00/index.html;   http://www.
hpcmo.hpc.mil/Htdocs/UGC/UGC00/;
 http://www.hpcmo.hpc.mil/Htdocs/
UGC/UGC00/paper/dorian_arnold_paper.
pdf;   http://www.hpcmo.hpc.mil/
Htdocs/UGC/UGC00/present/dorian_
arnold_pr.pdf;   http://www.netlib.
org/utk/people/JackDongarra/PAPERS/
netsolve-dod.pdf.

Bai:2000:TSA

[480] Zhaojun Bai, James Demmel, Jack Dongarra, and Axel Ruhe, editors. Templates for the solution of algebraic eigenvalue problems: a practical guide. Software, Environments, and Tools. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2000. ISBN 0-89871-471-0 (paperback), 0-89871-958-5 (ebook). xxx + 410 pp. LCCN QA193 .T46 2000. URL http://

www.loc.gov/catdir/enhancements/
fy0726/00059507-d.html;   http:/
/www.loc.gov/catdir/enhancements/

fy0726/00059507-t.html. A practical

guide.

Baker:2000:TMC

[481] Mark Baker, Rajkumar Buyya, and Jack Dongarra. Tutorial M9: Current and emerging trends in cluster computing. In ACM [1038], pages 23–24. ISBN ???? LCCN

QA76.88. URL http://www.sc2000.
org/proceedings/info/fp.pdf.

Board:2000:FMA

[482] John Board and Klaus Schulten. The fast multipole algorithm. Computing in Science and Engineering, 2 (1):76–79, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

URL http://dlib.computer.org/
cs/books/cs2000/pdf/c1076.pdf;
http://www.computer.org/cse/cs1999/

c1076abs.htm. See correspondence [497].

Browne:2000:PPI

[483] S. Browne, J. Dongarra, N. Garner, G. Ho, and P. Mucci. A portable programming interface for performance evaluation on modern processors. The International Journal of High Performance Computing Applications, 14(3):189–204, Fall 2000. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/papi-
journal-final.pdf;   https://www.
math.utah.edu/pub/tex/bib/ijsa.
bib.

Browne:2000:SCP

[484] S. Browne, J. Dongarra, N. Garner, K. London, and P. Mucci. A scalable cross-platform infrastructure for application performance tuning using hardware counters. In ACM [1038], page 65. ISBN ???? LCCN QA76.88.

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/papi-
sc2000.pdf;   http://www.sc2000.
org/proceedings/techpapr/papers/
pap256.pdf.

Casanova:2000:NES

[485] Henri Casanova, Satoshi Matsuoka, and J. Dongarra. Network-enabled server systems: Deploying scientific simulations on the grid. In Tentner [1041], page ?? ISBN 1-56555-197-4. LCCN QA76.88 .H53 2000.

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/nes-
astc2000.pdf.

Darema:2000:P

[486] Frederica Darema, Jack Dongarra, and Subhash Saini. Preface. The International Journal of High Performance Computing Applications, 14(3): 179, Fall 2000. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

DAzevedo:2000:DIP

[487] Eduardo D’Azevedo and Jack Dongarra. The design and implementation of the parallel out-of-core ScaLA-PACK LU, QR, and Cholesky factorization routines. Concurrency: practice and experience, 12(15):1481– 1493, ???? 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic). URL http://www3.

interscience.wiley.com/cgi-bin/
abstract/76505648/START;   http:
//www3.interscience.wiley.com/cgi-

bin/fulltext?ID=76505648&PLACEBO=
IE.pdf.

Dongarra:2000:GEI

[488] Jack Dongarra and Francis Sullivan. Guest Editors’ introduction: The top 10 algorithms. Computing in Science and Engineering, 2 (1):22–23, January/February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

URL http://dlib.computer.org/
cs/books/cs2000/pdf/c1022.pdf;
http://www.computer.org/cse/cs1999/

c1022abs.htm. See correspondence [497].

Dongarra:2000:HPC

[489] Jack Dongarra, Hans Meuer, Horst Simon, and Erich Strohmaier. High performance computing today. In Peter T. Cummings, Phillip R. Westmoreland, and Brice Carnahan, editors, Foundations of molecular modeling and simulation: proceedings of the First International Conference on Molecular Modeling and Simulation, Keystone, Colorado, July 23–28, 2000, volume 97(325) of AIChE symposium series, page ?? American Institute of Chemical Engineers, New York, NY,

2000. URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
hpc-today.pdf.

Dongarra:2000:NLA

[490] J. J. Dongarra and V. Eijkhout. Numerical linear algebra algorithms and software. Journal of Computational and Applied Mathematics, 123(1–2): 489–514, ???? 2000. CODEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

Dongarra:2000:NRI

[491] Jack Dongarra and Padma Raghavan. A new recursive implementation of sparse Cholesky factorization. In ????, editor, IMACS 2000, Lausanne, Switzerland, page ?? ????, ????,

                                        2000. URL http://www.netlib.org/
                                        utk/people/JackDongarra/PAPERS/
                                        imacs-sparse.pdf.

Dongarra:2000:RAS

[492] Jack Dongarra, Victor Eijkhout, and Piotr Luszczek. Recursive approach in sparse matrix LU factorization. In Proceedings of the 1st SGI Users Conference, pages 409–418. ACC Cyfronet UMM, Cracow, Poland, October 2000.

Dongarra:2000:TA

[493] J. Dongarra and F. Sullivan. The Top 10 algorithms. Computing in Science and Engineering, 2(1):22–23, January/ February 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Dongarra:2000:TMH

[494] Jack Dongarra, Iain Duff, and Danny Sorensen. Tutorial M7: High-speed numerical linear algebra: Algorithms and research directions. In ACM [1038], pages 21–22. ISBN ???? LCCN

                                        QA76.88. URL http://www.sc2000.
                                        org/proceedings/info/fp.pdf.

Fagg:2000:AAC

[495] Graham E. Fagg, Sathish S. Vadhiyar, and Jack J. Dongarra. ACCT: Automatic Collective Communications Tuning. Lecture Notes in Computer Science, 1908:354–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http://

link.springer-ny.com/link/service/
series/0558/bibs/1908/19080354.
htm;   http://link.springer-ny.
com/link/service/series/0558/papers/
1908/19080354.pdf.

Fagg:2000:FMF

[496] Graham E. Fagg and Jack J. Dongarra. FT-MPI: Fault Tolerant MPI, supporting dynamic applications in a dynamic world. Lecture Notes in Computer Science, 1908:346–??, 2000. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1908/19080346.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1908/19080346.
pdf;   http://www.netlib.org/utk/
people/JackDongarra/PAPERS/ft-mpi.
pdf.

Makino:2000:LEF

[497] Jun Makino, John Board, Klaus Schulten, Peter Borcherds, and Rubin D. Orduz Z. Letters to the editors: “The Fast Multipole Algorithm” and “The Top 10 Algorithms”. Computing in Science and Engineering, 2(3):4–5, May/June 2000. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

URL http://dlib.computer.org/
cs/books/cs2000/pdf/c3004.pdf. See

[488, 482].

Petitet:2000:PDS

[498] A. Petitet, H. Casanova, J. Dongarra, Y. Robert, and R. Whaley. Parallel and distributed scientific computing: A numerical linear algebra problem solving environment designer’s perspective. In Jacek Blazewicz et al., editors, Handbook on Parallel and Distributed Processing, page ?? Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. ISBN 3-540-66441-6. LCCN QA76.58 .H36

                                        2000. URL http://www.netlib.org/
                                        utk/people/JackDongarra/PAPERS/
                                        la-handbook.ps.

Vadhiyar:2000:ATC

[499] Sathish S. Vadhiyar, Graham E. Fagg, and Jack Dongarra. Automatically tuned collective communications. In ACM [1038], page 46. ISBN ???? LCCN QA76.88.

                                        URL http://www.netlib.org/utk/
                                        people/JackDongarra/PAPERS/atcc.
                                        pdf;   http://www.sc2000.org/proceedings/
                                        techpapr/papers/pap270.pdf.

Whaley:2000:AEO

[500] R. C. Whaley, A. Petitet, and J. Dongarra. Automated empirical optimization of software and the ATLAS Project. LAPACK Working Note 147, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, September

                                        2000. URL http://www.netlib.org/
                                        lapack/lawns/lawn147.ps;   http:
                                        //www.netlib.org/lapack/lawnspdf/

lawn147.pdf. UT-CS-00-448, Septem-

ber 2000.

Arnold:2001:CCD

[501] D. C. Arnold, S. S. Vahdiyar, and J. J. Dongarra. On the convergence of computational and data grids. Parallel Processing Letters, 11(2–3):187–202, 2001. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic). URL http:

//ejournals.wspc.com.sg/ppl/11/
sample/S012962640100052X.html;
http://www.netlib.org/utk/people/
JackDongarra/PAPERS/convergence-
data-grids.pdf.

Arnold:2001:DAS

[502] Dorian C. Arnold and Jack Dongarra. Developing an architecture to support the implementation and development of scientific computing applications. In Boisvert and Tang [1046], pages 39–56. ISBN 0-7923-7339-1. LCCN QA76.758 .I345

2000. URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
ifips-wg2-5.pdf.

Arnold:2001:PII

[503] D. Arnold, W. Lee, J. Dongarra, and M. Wheeler. Providing infrastructure and interface to high performance applications in a distributed setting. In Tentner [1041], page ?? ISBN 1-56555-197-4. LCCN QA76.88 .H53

2000. URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
Netsolve-ipars.ps.

Arnold:2001:RSO

[504] Dorian C. Arnold, Dieter Bachmann, and Jack Dongarra. Request sequencing: Optimizing communication for the Grid. Lecture Notes in Computer Science, 1900:1213–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/1900/19001213.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/1900/19001213.
pdf;   http://www.netlib.org/utk/

people/JackDongarra/PAPERS/sequencing.
pdf.

Barker:2001:LUG

[505] V. A. Barker, L. S. Blackford, J. J. Dongarra, J. J. Du Croz, S. J. Hammarling, M. Marinova, J. Waśniewski, and P. Yalamov. LAPACK95 Users’ Guide. Software, environments, and tools. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 2001. ISBN 0-89871-504-0 (paperback), 0-89871-820-1 (e-book). xviii + 258 pp. LCCN QA76.73.F25

L36 2001. URL http://www.loc.
gov/catdir/enhancements/fy0726/
2001042995-d.html;   http://www.
loc.gov/catdir/enhancements/fy0726/
2001042995-t.html.

Beck:2001:LCI

[506] Micah Beck, Dorian Arnold, Alessandro Bassi, Fran Berman, Henri Casanova, Jack Dongarra, Terry Moore, Graziano Obertelli, James Plank, Martin Swany, Sathish Vadhiyar, and Rich Wolski. Logistical computing and internetworking: Middleware for the use of storage in communication. In Lee [1050], pages 12–21. ISBN 0-7695-1528-

2. LCCN QA76.76.M54 I58 2001.

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/loci-

sc2001.pdf. IEEE Computer Society Order Number PR01528.

Berman:2001:GPS

[507] Francine Berman, Andrew Chien, Keith Cooper, Jack Dongarra, Ian Foster, Dennis Gannon, Lennart Johnsson, Ken Kennedy, Carl Kesselman, John Mellor-Crummey, Dan Reed, Linda Torczon, and Rich Wolski. The GrADS Project: Software support for highlevel Grid application development. The International Journal of High Performance Computing Applications, 15 (4):327–344, Winter 2001. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http://

journals.sagepub.com/doi/pdf/10.
1177/109434200101500401;   http:
//www.netlib.org/utk/people/JackDongarra/
PAPERS/gradspaper.pdf.

Blackford:2001:USB

[508] L. S. Blackford, J. Demmel, J. Dongarra, I. Duff, S. Hammarling, G. Henry, M. Heroux, L. Kaufman, A. Lumsdaine, A. Petitet, R. Pozo, K. Remington, and R. C. Whaley. An updated set of Basic Linear Algebra Subprograms (BLAS), February 2001. Submitted to ACM Transactions on Mathematical Software.

BLAST:2001:BLA

[509] Basic Linear Algebra Subprograms Technical (BLAST) Forum. Basic linear algebra subprograms technical (BLAST) forum standard. Technical report, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, January 23, 2001. viii + 209 pp.

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/blas-
report.pdf.

Choi:2001:IGS

[510] J. Choi, J. Demmel, I. Dhillon, J. Dongarra, S. Ostrouchov, A. Petitet, K. Stanley, D. Walker, and R. C. Whaley. Installation guide for ScaLA-PACK. LAPACK Working Note 93, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, August 31,

                                        2001. URL http://www.netlib.org/
                                        lapack/lawns/lawn93.ps;   http:/
                                        /www.netlib.org/lapack/lawnspdf/

lawn93.pdf. Updated August 31, 2001 (Version 1.7).

Dongarra:2001:BTC

[511] Jack Dongarra, Hans Meuer, Horst Simon, and Erich Strohmaier. Biannual Top-500 computer lists track changing environments for scientific computing. SIAM News, 34(9): ??, November 2001. ISSN 0036-

                                        1437. URL http://www.siam.org/
                                        siamnews/11-01/top500.pdf.

Dongarra:2001:CCG

[512] Jack Dongarra, Masaaki Shimasaki, and Bernard Tourancheau. Clusters and computational grids for scientific computing. Parallel Computing, 27(11):1401–1402, October 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/35/21/47/41/27/abstract.html;
                                         http://www.elsevier.nl/gej-ng/
                                        10/35/21/47/41/27/article.pdf.

Dongarra:2001:HPCa

[513] Jack Dongarra. High performance computing and trends: Connecting computational requirements with computing resources. Lecture Notes in Computer Science, 2150:33–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2150/21500033.htm;   http:/
                                        /link.springer-ny.com/link/service/

series/0558/papers/2150/21500033.
pdf.

Dongarra:2001:HPCb

[514] J. Dongarra. High performance computing and trends: connecting computational requirements with computing resources. In Katz [1049], page

135. ISBN 0-7695-1116-3; 0-7695-1117-1 (bookbroker); 0-7695-1118-X (microfiche). ISSN 0272-5428. LCCN QA76.58 .I38 2001.

Dongarra:2001:ISB

[515] Jack Dongarra, Victor Eijkhout, and Henk van der Vorst. An iterative solver benchmark. Scientific Programming, 9(4):223–231, 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic). URL http://

iospress.metapress.com/app/home/
contribution.asp%3Fwasp=64t4wprhwk589ellmv56%
26referrer=parent%26backto=issue%
2C3%2C4%3Bjournal%2C3%2C12%3Blinkingpublicationresults%
2C1%2C1;   http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
sparse-bench.pdf.

Dongarra:2001:LBP

[516] Jack Dongarra, Piotr Luszczek, and Antoine Petitet. The LINPACK benchmark: Past, present, and future. Technical report, University of Tennessee, Knoxville, Knoxville, TN 37996, USA,

2001. URL http://www.netlib.org/
utk/people/JackDongarra/PAPERS/
hpl.pdf.

Dongarra:2001:NA

[517] J. Dongarra. Netsolve and its application. In IEEE [1048], page 21. ISBN 0-7695-1432-4; 0-7695-1433-2 (case); 0-7695-1434-0 (microfiche). LCCN TK5105.5 .I323 2001.

Dongarra:2001:NLT

[518] Jack Dongarra, Shirley Moore, and Anne Trefethen. Numerical libraries and tools for scalable parallel cluster computing. The International Journal of High Performance Computing Applications, 15(2):175–180, Summer 2001. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Dongarra:2001:P

[519] Jack Dongarra and Bernard Tourancheau. Preface. The International Journal of High Performance Computing Applications, 15(3):199, Fall 2001. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Dongarra:2001:PCC

[520] J. Dongarra and B. Tourancheau. Preface: Clusters and computational grids for scientific computing. Parallel Processing Letters, 11(2–3):185–??, 2001. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Dongarra:2001:PVC

[521] Jack J. Dongarra. Performance of various computers using standard linear equations software. Technical Report CS-89-85, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, August 2001. 58 pp.

Dongarra:2001:QPC

[522] Jack J. Dongarra and David W. Walker. The quest for petascale computing. Computing in Science and Engineering, 3(3):32–39, May/June 2001. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

URL http://dlib.computer.org/
cs/books/cs2001/pdf/c3032.pdf;
http://ieeexplore.ieee.org/iel5/
5992/19880/00919263.pdf;   http:
//www.computer.org/cse/cs1999c3032abs.
htm;   http://www.netlib.org/utk/
people/JackDongarra/PAPERS/petascale.
pdf.

Dongarra:2001:RAS

[523] Jack Dongarra, Victor Eijkhout, and Piotr Luszczek. Recursive approach in sparse matrix LU factorization. Scientific Programming, 9(1):51–60, 2001. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL http://www.netlib.
org/netlib/utk/people/JackDongarra/
PAPERS/recur-sparse-sciprog.pdf;
 http://www.netlib.org/utk/people/
JackDongarra/PAPERS/rlu03.pdf.

Dongarra:2001:UCT

[524] Jack Dongarra. An update of a couple of tools: ATLAS and PAPI. Slide show., University of Tennessee, Knoxville, Knoxville, TN 37996, USA, 2001. 11 pp.

URL http://www.netlib.org/utk/
people/JackDongarra/SLIDES/salishan.
ps.

Dongarra:2001:UPH

[525] Jack Dongarra, Kevin London, Shirley Moore, Phil Mucci, and Dan Terpstra. Using PAPI for hardware performance monitoring on Linux systems. In ????, editor, Linux Clusters: The HPC Revolution, June 25– 27, 2001, National Center for Supercomputing Applications (NCSA), University of Illinois, Urbana, IL, page ?? ????, ????, 2001. URL http:/

                                        /www.linuxclustersinstitute.org/
                                        Linux-HPC-Revolution/Archive/PDF01/
                                        smoore_Utk.pdf;   http://www.netlib.
                                        org/utk/people/JackDongarra/PAPERS/
                                        papi-linux.pdf. Submitted.

Fagg:2001:FTM

[526] Graham E. Fagg, Antonin Bukovsky, and Jack J. Dongarra. Fault tolerant MPI for the HARNESS metacomputing system. Lecture Notes in Computer Science, 2073:355–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2073/20730355.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2073/20730355.
                                        pdf.

Fagg:2001:HFT

[527] Graham E. Fagg, Antonin Bukovsky, and Jack J. Dongarra. HARNESS and fault tolerant MPI. Parallel Computing, 27(11):1479–1495, October 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL

                                        http://www.elsevier.com/gej-ng/
                                        10/35/21/47/41/32/abstract.html;
                                         http://www.elsevier.nl/gej-ng/
                                        10/35/21/47/41/32/article.pdf;
                                        http://www.netlib.org/utk/people/
                                        JackDongarra/PAPERS/harness-ftmpi-
                                        pc.pdf.

Fagg:2001:PIS

[528] Graham E. Fagg, Edgar Gabriel, Michael Resch, and Jack J. Dongarra. Parallel IO support for metacomputing applications: MPI Connect IO applied to PACX–MPI. Lecture Notes in Computer Science, 2131:135– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310135.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310135.
pdf;   http://www.netlib.org/utk/
people/JackDongarra/PAPERS/epvm2001-
pio.pdf.

Kennedy:2001:TLS

[529] Ken Kennedy, Bradley Broom, Keith Cooper, Jack Dongarra, Rob Fowler, Dennis Gannon, Lennart Johnsson, John Mellor-Crummey, and Linda Torczon. Telescoping languages: a strategy for automatic generation of scientific problem-solving systems from annotated libraries. Journal of Parallel and Distributed Computing, 61(12): 1803–1826, December 1, 2001. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.idealibrary.com/links/doi/
10.1006/jpdc.2001.1724;   http:
//www.idealibrary.com/links/doi/
10.1006/jpdc.2001.1724/pdf;   http:
//www.idealibrary.com/links/doi/
10.1006/jpdc.2001.1724/ref;   http:
//www.netlib.org/netlib/utk/people/
JackDongarra/PAPERS/Telescope.pdf.

London:2001:EUT

[530] K. London, J. Dongarra, S. Moore, P. Mucci, K. Seymour, and T. Spencer. End-user tools for application performance analysis using hardware counters. In Sha [1052], page ?? ISBN 1-880843-39-0. LCCN QA76.58 .I5443

2001. URL http://www.netlib.org/

                                        utk/people/JackDongarra/PAPERS/
                                        papi_end_user.pdf.

Miller:2001:GEI

[531] Michelle Miller, Christopher Moulding, Jack Dongarra, and Christopher Johnson. Grid-enabling an interactive simulation/visualization environment. In Tentner [1053], page ?? ISBN 1-56555-237-7. LCCN

                                        ???? URL http://www.netlib.org/
                                        utk/people/JackDongarra/PAPERS/
                                        scirun_netsolve_hpc.pdf.

Miller:2001:GEP

[532] Michelle Miller, Christopher Moulding, Jack Dongarra, and Christopher Johnson. Grid-enabling problem solving environments: a case study of SCIRUN and NetSolve. In Tentner [1053], page ?? ISBN 1-56555-237-7. LCCN

                                        ???? URL http://www.netlib.org/
                                        utk/people/JackDongarra/PAPERS/
                                        scirun_netsolve.pdf.

Moore:2001:NTC

[533] Keith Moore and Jack Dongarra. Net-Build: Transparent cross-platform access to computational software libraries. Concurrency and Computation: Practice and Experience, ??(??): ??, 2001. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

                                        tronic). URL http://www.netlib.
                                        org/utk/people/JackDongarra/PAPERS/
                                        netbuild-C-PE.pdf. Submitted.

Moore:2001:RPA

[534] Shirley Moore, David Cronk, Kevin London, and Jack Dongarra. Review of performance analysis tools for MPI parallel programs. Lecture Notes in Computer Science, 2131:241– ??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer-
ny.com/link/service/series/0558/
bibs/2131/21310241.htm;   http:/
/link.springer-ny.com/link/service/
series/0558/papers/2131/21310241.
pdf;   http://www.netlib.org/utk/
people/JackDongarra/PAPERS/perftools-
review2.pdf.

Petitet:2001:NLGa

[535] Antoine Petitet, Susan Blackford, Jack Dongarra, Brett Ellis, Graham Fagg, Kenneth Roche, and Sathish Vadhiyar. Numerical libraries and the Grid: The GrADS experiments with ScaLA-PACK. In ACM [1043], page ?? ISBN 1-58113-293-X. LCCN QA76.88 .S85

2001. URL http://www.sc2001.org/
papers/pap.pap177.pdf.

Petitet:2001:NLG

[536] Antoine Petitet, Susan Blackford, Jack Dongarra, Brett Ellis, Graham Fagg, Kenneth Roche, and Sathish Vadhiyar. Numerical libraries and the Grid. The International Journal of High Performance Computing Applications, 15 (4):359–374, Winter 2001. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http://

journals.sagepub.com/doi/pdf/10.
1177/109434200101500403;   http:
//www.netlib.org/utk/people/JackDongarra/
PAPERS/version8-ijsa.pdf.

Seymour:2001:ATF

[537] Keith Seymour and Jack Dongarra. Automatic translation of Fortran to JVM bytecode. In ACM [1042], pages 126–133. ISBN 1-58113-359-6. LCCN QA76.9.O35 A26

                                        2001. URL http://www.netlib.org/
                                        utk/people/JackDongarra/PAPERS/
                                        f2jreport.pdf;   http://www.philippsen.
                                        com/JGI2001/camerareadyabstracts/
                                        51.html;   http://www.philippsen.
                                        com/JGI2001/finalpapers/18500126.
                                        ps.

Vadhiyar:2001:PMS

[538] Sathish S. Vadhiyar, Graham E. Fagg, and Jack J. Dongarra. Performance modeling for self adapting collective communications for MPI. In Oldehoeft [1062], page ?? CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

                                        URL http://www.netlib.org/utk/
                                        people/JackDongarra/PAPERS/coll-
                                        lacsi-2001.pdf.

Vadhiyar:2001:TAM

[539] Sathish S. Vadhiyar, Graham E. Fagg, and Jack J. Dongarra. Towards an accurate model for collective communications. Lecture Notes in Computer Science, 2073:41–??, 2001. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

                                        tronic). URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        bibs/2073/20730041.htm;   http:/
                                        /link.springer-ny.com/link/service/
                                        series/0558/papers/2073/20730041.
                                        pdf.

vanderSteen:2001:ORS

[540] Aad J. van der Steen and Jack J. Dongarra. Overview of recent supercomputers. Technical report, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, 2001. URL http://www.

                                        nhse.org/NHSEreview/ORS/.

Whaley:2001:AEO

[541] R. Clint Whaley, Antoine Petitet, and Jack J. Dongarra. Automated empirical optimizations of software and the ATLAS project. Parallel Computing, 27(1–2):3–35, January 2001. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http://

www.elsevier.nl/gej-ng/10/35/21/
47/25/23/abstract.html;   http:
//www.elsevier.nl/gej-ng/10/35/
21/47/25/23/article.pdf;   http:
//www.netlib.org/utk/people/JackDongarra/
PAPERS/atlas_pub.pdf.

Arnold:2002:ING

[542] Dorian C. Arnold, Henri Casanova, and Jack Dongarra. Innovations of the NetSolve Grid Computing System. Concurrency and Computation: Practice and Experience, 14(13– 15):1457–1479, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www.netlib.
org/netlib/utk/people/JackDongarra/
PAPERS/cpe678.pdf.

Beck:2002:MUS

[543] Micah Beck, Dorian Arnold, Alessandro Bassi, Fran Berman, Henri Casanova, Jack Dongarra, Terry Moore, Graziano Obertelli, James Plank, and Martin Swany. Middleware for the use of storage in communication. Parallel Computing, 28(12):1773–1787, December 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

tronic). URL http://www.netlib.
org/netlib/utk/people/JackDongarra/
PAPERS/middleware-storage-pc.pdf.

Blackford:2002:USB

[544] L. Susan Blackford, James Demmel, Jack Dongarra, Iain Duff, Sven Hammarling, Greg Henry, Michael Heroux, Linda Kaufman, Andrew Lumsdaine, Antoine Petitet, Roldan Pozo, Karin Remington, and R. Clint Whaley. An updated set of Basic Linear Algebra Subprograms (BLAS). ACM Transactions on Mathematical Software, 28(2):135–151, June 2002. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

                                        URL http://www.netlib.org/utk/
                                        people/JackDongarra/PAPERS/blast-
                                        toms.pdf.

Boisvert:2002:PSI

[545] Ronald F. Boisvert and Jack J. Dongarra. Preface to the special issue on the Basic Linear Algebra Subprograms (BLAS). ACM Transactions on Mathematical Software, 28(2):133–134, June 2002. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Casanova:2002:VIS

[546] Henri Casanova, Thomas Bartol, Francine Berman, Adam Birnbaum, Jack Dongarra, Mark Ellisman, Marcio Faerman, Erhan Gockay, Michelle Miller, Graziano Obertelli, Stuart Pomerantz, Terry Sejnowski, Joel Stiles, and Rich Wolski. The Virtual Instrument: Support for Grid-enabled scientific simulations. Parallel and Distributed Computing Practices, ??(??):??, ???? 2002. CODEN ???? ISSN 1097-

                                        2803. URL http://www.netlib.org/
                                        netlib/utk/people/JackDongarra/
                                        PAPERS/vi-itr-jpdc.pdf. Submitted.

Cuencal:2002:AOP

[547] Javier Cuencal, Domingo Giminez, Josi Gonzalez, Jack Dongarra, and Kenneth Roche. Automatic optimisation of parallel linear algebra routines in systems with variable load. In Monien and Feldmann [1061], page ?? ISBN 3-540-44049-6 (softcover). LCCN QA76.58 .I553

2002. URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/la-spain.pdf. Submitted.

Dongarra:2002:HPC

[548] Jack Dongarra. High performance computing, computational Grid, and numerical libraries. Lecture Notes in Computer Science, 2474:1–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (elec-

tronic). URL http://link.springer.
de/link/service/series/0558/bibs/
2474/24740001.htm;   http://link.
springer.de/link/service/series/
0558/papers/2474/24740001.pdf.

Dongarra:2002:PBLa

[549] Jack Dongarra. Preface: Basic Linear Algebra Subprograms Technical (Blast) Forum Standard I. The International Journal of High Performance Computing Applications, 16 (1):1–111, Spring 2002. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http://

journals.sagepub.com/doi/pdf/10.
1177/10943420020160010101.

Dongarra:2002:PBLb

[550] Jack Dongarra. Preface: Basic Linear Algebra Subprograms Technical (Blast) Forum Standard II. The International Journal of High Performance Computing Applications, 16(2): 115, Summer 2002. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic).

Dongarra:2002:PVC

[551] Jack J. Dongarra. Performance of various computers using standard linear equations software (Linpack benchmark report). Technical report CS-89-85, University of Tennessee, Knoxville, Knoxville, TN 37996, USA,

                                        2002. URL http://www.netlib.org/
                                        benchmark/performance.ps.

Dongarra:2002:SAN

[552] Jack Dongarra and Victor Eijkhout. Self-adapting numerical software for next generation applications. LA-PACK Working Note 157, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, August

                                        2002. URL http://www.netlib.org/
                                        lapack/lawns/lawn157.ps;   http:
                                        //www.netlib.org/lapack/lawnspdf/

lawn157.pdf. UT-CS-02-484, August 2002.

Dongarra:2002:SPC

[553] Jack Dongarra, Ian Foster, Geoffrey Fox, William Gropp, Ken Kennedy, Linda Torczon, and Andy White, editors. The Sourcebook of Parallel Computing. Morgan Kaufmann Publishers, San Francisco, CA, USA, 2002. ISBN 1-55860-871-0. xvi + 842 + 8 pp. LCCN QA76.58 S638 2003. US$59.95.

Dongarra:2002:THP

[554] J. Dongarra. Trends in high performance computing and using numerical libraries on clusters. In Gropp [1055], page 172. ISBN 0-7695-1745-5. LCCN QA76.58 .I38 2002.

Dongarra:2002:TTH

[555] J. Dongarra. Three tools to help with cluster and grid computing: SANS-Effort, PAPI, and NetSolve. In IEEE [1056], page 2. ISBN 0-7695-1582-7, 0-7695-1583-5 (bookbroker), 0-7695-1584-3 (microfiche). LCCN QA76.9.C58 I33 2002a. IEEE Computer Society order number PR01582.

Fagg:2002:FTM

[556] Graham E. Fagg, Antonin Bukovsky, Sathish Vadhiyar, and Jack J. Dongarra. Fault tolerant MPI for the HARNESS MetaComputing system. Technical report ????, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, 2002. 14

pp. URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/ft-mpi-iccs-gef.pdf.

Fagg:2002:HFTa

[557] Graham E. Fagg and Jack J. Dongarra. HARNESS fault tolerant MPI design, usage and performance issues. Technical report ????, University of Tennessee, Knoxville, Knoxville, TN 37996, USA,

2002. URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/ft-mpi-fgcs-grid-se.pdf.

Fagg:2002:HFTb

[558] Graham E. Fagg and Jack J. Dongarra. HARNESS fault tolerant MPI design, usage and performance issues. Future Generation Computer Systems, 18(8): 1127–1142, October 2002. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Henry:2002:PIN

[559] Greg Henry, David Watkins, and Jack Dongarra. A parallel implementation of the nonsymmetric QR algorithm for distributed memory architectures. SIAM Journal on Scientific Computing, 24(1):284–311, January 2002. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (elec-

tronic). URL http://epubs.siam.
org/sam-bin/dbq/article/32516.

Hiroyasu:2002:OSU

[560] T. Hiroyasu, M. Miki, H. Shimosaka, Y. Tanimura, and J. Dongarra. Optimization system using Grid RPC. In ????, editor, The Japan Society of Mechanical Engineers, Kyoto University, Kyoto, Japan, October 12–13, 2002, page ?? ????, ????, 2002. ISBN ???? LCCN

???? URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/Japanese-GridRPC.pdf.

Hiroyasu:2002:TSO

[561] T. Hiroyasu, M. Miki, H. Shimosaka, M. Sano, Y. Tanimura, Y. Mimura, S. Yoshimura, and J. Dongarra. Truss structural optimization using NetSolve system. In ????, editor, The Japan Society of Mechanical Engineers, Kyoto University, Kyoto, Japan, October 12–13, 2002, page ?? ????, ????, 2002. ISBN ???? LCCN

???? URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/Japanese-NetSolve.pdf.

Kennedy:2002:TFP

[562] Ken Kennedy, Mark Mazina, John Mellor-Crummey, Keith Cooper, Linda Torczon, Fran Berman, Andrew Chien, Holly Dail, Otto Sievert, Dave Angulo, Ian Foster, Dennis Gannon, Lennart Johnsson, Carl Kesselman, Ruth Aydt, Daniel Reed, Jack Dongarra, Sathish Vadhiyar, and Rich Wolski. Toward a framework for preparing and executing adaptive grid programs. In IEEE [1059], page ?? ISBN 0-7695-1573-8, 0-7695-1574-6, 0-7695-1575-4. LCCN QA76.58.I583

2002. URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/GrADSIPDPS02.pdf. To ap-

pear.

Lee:2002:VMT

[563] DongWoo Lee, Jack J. Dongarra, and R. S. Ramakrishna. vis-Perf: Monitoring tool for Grid computing. In Parashar [1063], page ?? ISBN 3-540-00133-6 (softcover). LCCN QA76.9.C58 G74

2002. URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/visperf.pdf. Also available

via the World Wide Web.

Moore:2002:ANA

[564] Shirley Moore, A. J. Baker, Jack Dongarra, Christian Halloy, and Chung Ng. Active Netlib: An active mathematical software collection for inquiry-based computational science and engineering education. Journal of Digital Information, 2(4):??, May 2002. CODEN ???? ISSN 1368-

7506. URL http://jodi.ecs.soton.
ac.uk/Articles/v02/i04/Moore;
http://journals.tdl.org/jodi/article/

                                        view/56;   http://journals.tdl.
                                        org/jodi/article/view/56/59.

Moore:2002:NTC

[565] Keith Moore and Jack Dongarra. Net-Build: transparent cross-platform access to computational software libraries. Concurrency and Computation: Practice and Experience, 14(13– 15):1445–1456, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

                                        URL http://www.netlib.org/utk/
                                        people/JackDongarra/PAPERS/netbuild-
                                        C-PE.pdf.

Nakada:2002:GRP

[566] Hidemoto Nakada, Satoshi Matsuoka, Keith Seymour, Jack Dongarra, Craig Lee, and Henri Casanova. GridRPC: a remote procedure call API for grid computing. In Parashar [1063], pages xi + 318. ISBN 3-540-00133-6 (softcover). LCCN QA76.9.C58 G74

                                        2002. URL http://www.netlib.org/
                                        netlib/utk/people/JackDongarra/
                                        PAPERS/grpc.pdf. Submitted.

Roche:2002:DPN

[567] Kenneth J. Roche and Jack J. Dongarra. Deploying parallel numerical library routines to cluster computing in a self adapting fashion. Parallel Computing, ??(??):??, 2002. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (elec-

                                        tronic). URL http://www.netlib.
                                        org/netlib/utk/people/JackDongarra/
                                        PAPERS/dyn_pnumlibs.pdf. Submit-

ted.

Seymour:2002:OGR

[568] Keith Seymour, Hidemoto Nakada, Satoshi Matsuoka, Jack Dongarra, Craig Lee, and Henri Casanova. Overview of GridRPC: a remote procedure call API for Grid computing. Lecture Notes in Computer Science, 2536:274–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

//link.springer.de/link/service/
series/0558/bibs/2536/25360274.
htm;   http://link.springer.de/
link/service/series/0558/papers/
2536/25360274.pdf.

Vadhiyar:2002:MGa

[569] Sathish S. Vadhiyar and Jack J. Dongarra. A metascheduler for the grid. In IEEE [1058], page ?? ISBN 0-7695-1721-8, 0-7695-1723-4 (microfiche). LCCN QA76.76.M54 I58

2002. URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/hpdc_meta.pdf.

Vadhiyar:2002:MGb

[570] S. S. Vadhiyar and J. J. Dongarra. A metascheduler for the Grid. In IEEE [1057], pages 343–351. ISBN 0-7695-1686-6; 0-7695-1688-2 (microfiche). LCCN QA76.9.D5I593 2002c.

Vadhiyar:2002:PMS

[571] Sathish S. Vadhiyar, Graham E. Fagg, and Jack J. Dongarra. Performance modeling for self adapting collective communications for MPI. In Oldehoeft [1062], page ?? CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

URL http://www.netlib.org/utk/
people/JackDongarra/PAPERS/coll-
lacsi-2001.pdf.

Vadhiyar:2002:POM

[572] Sathish S. Vadhiyar and J. Dongarra. A performance oriented migration framework for the Grid. Technical report ????, University of Tennessee, Knoxville, Knoxville, TN 37996, USA,

                                        2002. 9 pp. URL http://www.netlib.
                                        org/netlib/utk/people/JackDongarra/
                                        PAPERS/vadhiyar-migration.pdf.

vanderSteen:2002:OHP

[573] Aad van der Steen and Jack Dongarra. Overview of high performance computers. In Handbook of Massive Data Sets [1054], page ?? ISBN 1-4020-0489-3. LCCN QA76.9.D3 H3474

                                        2002. URL http://www.netlib.org/
                                        utk/people/JackDongarra/PAPERS/
                                        overview-hpc.pdf.

YarKhan:2002:ESU

[574] Asim YarKhan and Jack J. Dongarra. Experiments with scheduling using simulated annealing in a Grid environment. Lecture Notes in Computer Science, 2536:232–??, 2002. CODEN LNCSD9. ISSN 0302-9743 (print), 1611-3349 (electronic). URL http:

                                        //link.springer.de/link/service/
                                        series/0558/bibs/2536/25360232.
                                        htm;   http://link.springer.de/
                                        link/service/series/0558/papers/
                                        2536/25360232.pdf;   http://www.
                                        netlib.org/netlib/utk/people/JackDongarra/
                                        PAPERS/annealing-scheduler.pdf.

Agrawal:2003:NPP

[575] Sudesh Agrawal, Jack Dongarra, Keith Seymour, and Sathish Vadhiyar. Net-Solve: Past, present, and future — a look at a Grid enabled server. In Fran Berman, Geoffrey Fox, and Anthony J. G. Hey, editors, Grid Computing: Making the Global Infrastructure a Reality, Wiley series in communications networking and distributed systems, page ?? Wiley, New York, NY, USA, 2003. ISBN 0-470-85319-0. LCCN QA76.9.C58 G755 2003.

URL http://www.loc.gov/catdir/
description/wiley039/2002192438.
html;   http://www.loc.gov/catdir/
toc/wiley031/2002192438.html;
http://www.netlib.org/netlib/utk/
people/JackDongarra/PAPERS/netsolve-
ppf.pdf.

Beck:2003:STN

[576] Micah Beck, Jack Dongarra, Victor Eijkhout, Mike Langston, Terry Moore, and Jim Plank. Scalable, trustworthy network computing using untrusted intermediaries: a position paper. In ????, editor, DOE/NSF Workshop on New Directions in Cyber-Security in Large-Scale Networks: Deployment Obstacles, National Conference Center, Lansdowne, Virginia, March, 2003, page ?? ????, ????, 2003. ISBN ???? LCCN

???? URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/position-trustworthy.pdf.

Chen:2003:SASa

[577] Zizhong Chen, Jack Dongarra, Piotr Luszczek, and Kenneth Roche. Self adapting software for numerical linear algebra and LAPACK for clusters. LAPACK Working Note 160, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, January

2003. URL http://www.netlib.org/
lapack/lawns/lawn160.ps;   http:

                                        //www.netlib.org/lapack/lawnspdf/
                                        lawn160.pdf;   http://www.netlib.
                                        org/netlib/utk/people/JackDongarra/
                                        PAPERS/lfc-pc.pdf. UT-CS-03-499,

January 2003. Submitted to Parallel Computing.

Chen:2003:SASb

[578] Z. Z. Chen, J. Dongarra, P. Luszczek, and K. Roche. Self-adapting software for numerical linear algebra and LA-PACK for clusters. Parallel Computing, 29(11–12):1723–1743, ???? 2003. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cuenca:2003:AOP

[579] J. Cuenca, D. Gimenez, J. Gonzalez, J. Dongarra, and K. Roche. Automatic optimisation of parallel linear algebra routines in systems with variable load. In Clematis [1069], pages 409– 416. ISBN 0-7695-1875-3. ISSN 1066-6192. LCCN QA76.58. IEEE Computer Society Order Number PR01875.

Dail:2003:SGA

[580] Holly Dail, Otto Sievert, Fran Berman, Henri Casanova, Asim YarKhan, Sathish Vadhiyar, Jack Dongarra, Chuang Liu, Lingyun Yang, Dave Angulo, and Ian Foster. Scheduling in the grid application development software project. In Nabrzyski et al. [1078], page ?? ISBN 1-4020-7575-8. LCCN QA76.9.C58 G78

                                        2004. URL http://www.netlib.org/
                                        netlib/utk/people/JackDongarra/
                                        PAPERS/grads-kluwer2003.pdf.

Dongarra:2003:ELLa

[581] Jack Dongarra, Kevin London, Shirley Moore, Philip Mucci, Daniel Terpstra, Haihang You, and Min Zhou. Experiences and lessons learned with a portable interface to hardware performance counters. In IEEE [1076], page ?? ISBN 0-7695-1926-1. LCCN QA76.58 .M47

2003. URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/papi-lessons.pdf.

Dongarra:2003:ELLb

[582] J. Dongarra, K. London, S. Moore, P. Mucci, D. Terpstra, Haihang You, and Min Zhou. Experiences and lessons learned with a portable interface to hardware performance counters. In IEEE [1074], pages lxv + 304. ISBN 0-7695-1926-1. ISSN 1530-2075. LCCN QA76.58.

Dongarra:2003:FCA

[583] Jack Dongarra and Victor Eijkhout. Finite-choice algorithm optimization in Conjugate Gradients. LA-PACK Working Note 159, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, January

2003. URL http://www.netlib.org/
lapack/lawns/lawn159.ps;   http:
//www.netlib.org/lapack/lawnspdf/

lawn159.pdf. UT-CS-03-502, January 2003.

Dongarra:2003:LBP

[584] Jack J. Dongarra, Piotr Luszczek, and Antoine Petitet. The LIN-PACK Benchmark: past, present and future. Concurrency and Computation: Practice and Experience, 15(9):803–820, August 10, 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

                                        URL http://www.netlib.org/utk/
                                        people/JackDongarra/PAPERS/hpl.
                                        pdf.

Dongarra:2003:P

[585] Jack Dongarra and Bernard Tourancheau. Preface. Parallel Processing Letters, 13(2):93–??, June 2003. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Dongarra:2003:PIM

[586] Jack Dongarra, Allen Malony, Shirley Moore, Philip Mucci, and Sameer Shende. Performance instrumentation and measurement for terascale systems. In Sloot et al. [1081], pages 53–62. CODEN LNCSD9. ISBN 3-540-40195-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2003 pt.1-4

                                        2003. URL http://www.netlib.org/
                                        netlib/utk/people/JackDongarra/
                                        PAPERS/instr-meas.pdf.

Dongarra:2003:SANa

[587] Jack Dongarra and Victor Eijkhout. Self-adapting numerical software and automatic tuning of heuristics. In Sloot et al. [1080], page ?? CO-DEN LNCSD9. ISBN 3-540-40194-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2003 pt.1-4

                                        2003. URL http://www.netlib.org/
                                        netlib/utk/people/JackDongarra/
                                        PAPERS/sans-iccs2003.pdf.

Dongarra:2003:SAN

[588] Jack Dongarra and Victor Eijkhout. Self-adapting numerical software for next generation applications. The International Journal of High Performance Computing Applications, 17 (2):125–131, Summer 2003. CO-DEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL

http://journals.sagepub.com/doi/
pdf/10.1177/1094342003017002002;
 http://www.netlib.org/netlib/utk/
people/JackDongarra/PAPERS/sans-
position.pdf;   http://www.netlib.
org/utk/people/JackDongarra/PAPERS/
sans-ijhpca.pdf.

Eidson:2003:AAO

[589] Thomas Eidson, Jack Dongarra, and Victor Eijkhout. Applying aspectorient programming concepts to a component-based programming model. In ????, editor, Proceedings. International Parallel and Distributed Processing Symposium, 2003. 22– 26 April 2003, page ?? ????, ????, 2003. ISBN ???? LCCN

???? URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/ipdps-cca-2003.pdf.

Fagg:2003:FTC

[590] Graham E. Fagg, Edgar Gabriel, Zizhon Chen, Thara Angskun, George Bosilca, Antonin Bukovsky, and Jack J. Dongarra. Fault tolerant communication library and applications for high performance computing. In ????, editor, Los Alamos Computer Science Institute Symposium, Santa Fe, NM, October 27–29, 2003, page ?? ????, ????, 2003. ISBN ???? LCCN

???? URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/lacsi2003-ftmpi-fagg.pdf.

Fox:2003:PSE

[591] G. Fox, J. J. Dongarra, D. Arnold, H. Casanova, Ann C. Catlin, T. Haupt, E. N. Houstis, and J. R. Rice. Problem solving environments. In Dongarra et al. [1071], pages 409–442. ISBN 1-55860-871-0. LCCN QA76.58 S638 2003. US$59.95.

Gabriel:2003:EPM

[592] Edgar Gabriel, Graham Fagg, and Jack Dongarra. Evaluating the performance of MPI-2 dynamic communicators and one-sided communication. In Dongarra et al. [1070], page ?? CO-DEN LNCSD9. ISBN 3-540-20149-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E973

                                        2003. URL http://www.netlib.org/
                                        netlib/utk/people/JackDongarra/
                                        PAPERS/europvm-mpi-2003-mpi2.pdf.

Gabriel:2003:FTC

[593] Edgar Gabriel, Graham E. Fagg, Antonin Bukovsky, Thara Angskun, and Jack J. Dongarra. A faulttolerant communication library for Grid environments. In ????, editor, 17th Annual ACM International Conference on Supercomputing (ICS’03) International Workshop on Grid Computing and e-Science, June 21, 2003, San Francisco, page ?? ????, ????, 2003. ISBN ???? LCCN

                                        ???? URL http://www.netlib.org/
                                        netlib/utk/people/JackDongarra/
                                        PAPERS/FTMPI-SF-gabriel.pdf.

Hiroyasu:2003:EMP

[594] Tomoyuki Hiroyasu, Mitsunori Miki, Shinya Ogura, Keiko Aoi, Takeshi Yoshida, Yuko Okamoto, and Jack Dongarra. Energy minimization of protein tertiary structure by parallel simulated annealing using genetic crossover. Journal of Genetic Programming and Evolvable Machines, ??(??):??, ???? 2003. CODEN ???? ISSN 1389-

2576. URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/02_gemp_hiroyasu.pdf. Spe-

cial Issue on Biological Applications of Genetic and Evolutionary Computation.

Hiroyasu:2003:OPS

[595] Tomoyuki Hiroyasu, Mitsunori Miki, Hisashi Shimosaka, and Jack Dongarra. Optimization problem solving system using Grid RPC. In IEEE [1075], page ?? ISBN 0-7695-1919-9. LCCN

???? URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/hiroyasu_ccgrid_poster.pdf.

Hiroyasu:2003:SIA

[596] Tomoyuki Hiroyasu, Mitsunori Miki, Kenzo Kodama, Junichi Uekawa, and Jack Dongarra. A simple installation and administration tool for the large-scaled PC cluster system. In ????, editor, ClusterWorld Conference and Expo, San Jose, California, June 24–26, 2003, page ?? ????, ????, 2003. ISBN ???? LCCN

???? URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/05_ccwc_hiroyasu_abst.pdf.

Kranzlmuller:2003:RAP

[597] Dieter Kranzlmüller, Peter Kacsuk, Jack Dongarra, and Jens Volkert. Recent advances in Parallel Virtual Machine and Message Passing Interface (select papers from the EuroPVMMPI 2002 Conference). The International Journal of High Performance Computing Applications, 17 (1):3–5, Spring 2003. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http://

                                        journals.sagepub.com/doi/pdf/10.
                                        1177/1094342003017001001.

Lee:2003:VMT

[598] DongWoo Lee, Jack J. Dongarra, and R. S. Ramakrishna. visP erf: Monitoring tool for grid computing. In Sloot et al. [1082], pages 233– 243. CODEN LNCSD9. ISBN 3-540-40196-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2003 pt.1-4

                                        2003. URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        tocs/t2659.htm.

Plank:2003:OPR

[599] James S. Plank, Micah Beck, Jack Dongarra, Rich Wolski, and Heni Casanova. Optimizing performance and reliability in distributed computing systems through wide spectrum storage. In Gerndt [1072], page ?? ISBN 0-7695-1880-X. LCCN QA76.642 .I586

                                        2003. URL http://www.netlib.org/
                                        netlib/utk/people/JackDongarra/
                                        PAPERS/ibp-ipdps2003.pdf.

Seymour:2003:ATF

[600] Keith Seymour and Jack Dongarra. Automatic translation of Fortran to JVM bytecode. Concurrency and Computation: Practice and Experience, 15(3–5):207–222, March/April 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

                                        tronic). URL http://www.netlib.
                                        org/netlib/utk/people/JackDongarra/
                                        PAPERS/f2jreport.pdf.

Vadhiyar:2003:GGB

[601] Sathish Vadhiyar and Jack Dongarra. GrADSolve — A Grid-based RPC system for remote invocation of parallel software. Parallel and Distributed Computing Practices, ??(??): ??, 2003. CODEN ???? ISSN 1097-

2803. URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/gradsolve-jpdc-mod-2003.

pdf. Submitted.

Vadhiyar:2003:GRH

[602] Sathish Vadhiyar and Jack Dongarra. GrADSolve — RPC for high performance computing on the Grid. In Kosch et al. [1077], page ?? ISBN 3-540-40788-X. LCCN QA76.58 .I5443

2003. URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/gradsolve-europar-2003.pdf.

Vadhiyar:2003:POM

[603] S. S. Vadhiyar and J. J. Dongarra. A performance oriented migration framework for the grid. In IEEE [1073], pages 130–137. ISBN 0-7695-1919-9. LCCN QA76.9.C58.

Vadhiyar:2003:SAG

[604] Sathish Vadhiyar and Jack Dongarra. Self adaptability in grid computing. Concurrency and Computation: Practice and Experience, ??(??):??, ???? 2003. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (elec-

tronic). URL http://www.netlib.
org/netlib/utk/people/JackDongarra/
PAPERS/sans-grid-edinburgh.pdf.

Submitted.

Vadhiyar:2003:SFD

[605] Sathish S. Vadhiyar and Jack J. Dongarra. SRS — A framework for developing malleable and migratable parallel applications for distributed systems. Parallel Processing Letters, 13(2):291– ??, June 2003. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Abramson:2004:SGC

[606] D. Abramson, J. Dongarra, E. Meek, P. Roe, and Zhiao Shi. Simplified grid computing through spreadsheets and NetSolve. In IEEE [1092], pages 19–24. ISBN 0-7695-2138-X. LCCN QA76.88.

Beck:2004:ALS

[607] Micah Beck, Jack Dongarra, Jian Huang, Terry Moore, and James S. Plank. Active logistical state management in GridSolve/L. In IEEE [1090], page ?? ISBN 0-7803-8430-X. LCCN QA76.9.C58

I42 2004.  URL http://www.
mcs.anl.gov/ccgrid2004;   http:/
/www.netlib.org/netlib/utk/people/
JackDongarra/PAPERS/GridSolveL.

pdf. IEEE catalog number 04EX836.

Casanova:2004:VIS

[608] Henri Casanova, Francine Berman, Thomas Bartol, Erhan Gokcay, Terry Sejnowski, Adam Birnbaum, Jack Dongarra, Michelle Miller, Mark Ellisman, Marcio Faerman, Graziano Obertelli, Rich Wolski, Stuart Pomerantz, and Joel Stiles. The Virtual Instrument: Support for Grid-enabled Mcell simulations. The International Journal of High Performance Computing Applications, 18(1):3–17, February 2004.

CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL

http://hpc.sagepub.com/content/
18/1/3.full.pdf+html.

Chen:2004:LCP

[609] Zizhong Chen, Jack Dongarra, Piotr Luszczek, and Kenneth Roche. The LAPACK for Clusters Project: An example of self adapting numerical software. In Sprague [1094], pages 282– 291. ISBN 0-7695-2056-1. LCCN

Q350. URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/lfc-hicss.pdf. HICSS-37,

Hilton Waikoloa Village, Big Island, Hawaii.

Cooper:2004:NGS

[610] K. Cooper, A. Dasgupta, K. Kennedy, C. Koelbel, A. Mandal, G. Marin, M. Mazina, J. Mellor-Crummey, F. Berman, H. Casanova, A. Chien, H. Dail, X. Liu, A. Olugbile, O. Sievert, H. Xia, L. Johnsson, B. Liu, M. Patel, D. Reed, W. Deng, C. Mendes, Z. Shi, A. YarKhan, and J. Dongarra. New grid scheduling and rescheduling methods in the GrADS project. In IEEE [1091], pages 199–?? ISBN 0-7695-2132-0. LCCN QA76.58.

Dongarra:2004:ACT

[611] Jack Dongarra, Shirley Moore, Phil Mucci, Keith Seymour, and Haihang You. Accurate cache and TLB characterization using hardware counters. In Bubak et al. [1085], page ?? CO-DEN LNCSD9. ISBN 3-540-22114-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13

2004. URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/cache-tlb-iccs2004.pdf.

Dongarra:2004:PVC

[612] Jack J. Dongarra. Performance of various computers using standard linear equations software, (Linpack benchmark report). Technical report CS-89-85, University of Tennessee, Knoxville, Knoxville, TN 37996, USA,

                                        2004. URL http://www.netlib.org/
                                        benchmark/performance.ps.

Dongarra:2004:PWC

[613] J. Dongarra and B. Tourancheau. Preface: Workshop on clusters and computational grids for scientific computing. The International Journal of High Performance Computing Applications, 18(3):283, Fall 2004. ISSN 1094-3420 (print), 1741-2846 (electronic).

Dongarra:2004:SNA

[614] Jack Dongarra, Kaj Madsen, and Jerzy Waśniewski. Selected numerical algorithms. Future Generation Computer Systems, 20(3):349–351, April 1, 2004. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dongarra:2004:THPa

[615] Jack Dongarra. The Boole Lecture: Trends in high performance computing. The Computer Journal, 47(4):399–403, July 2004. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (elec-

                                        tronic). URL http://www3.oup.co.
                                        uk/computer_journal/hdb/Volume_
                                        47/Issue_04/470399.sgm.abs.html;
                                         http://www3.oup.co.uk/computer_
                                        journal/hdb/Volume_47/Issue_04/
                                        pdf/470399.pdf.

Eidson:2004:IEC

[616] T. Eidson, V. Eijkhout, and J. Dongarra. Improvements in the efficient composition of applications built using a component-based programming environment. In IEEE [1091], page 198. ISBN 0-7695-2132-0. LCCN QA76.58.

Fagg:2004:BUF

[617] Graham E. Fagg and Jack J. Dongarra. Building and using a faulttolerant MPI implementation. The International Journal of High Performance Computing Applications, 18 (3):353–361, Fall 2004. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

//hpc.sagepub.com/content/18/3/
353.full.pdf+html.

Heinrich:2004:SCO

[618] Kevin Heinrich, Michael W. Berry, Jack J. Dongarra, and Sathish Vadhiyar. The semantic conference organizer. In Bozdogan [1084], page 588. ISBN 1-58488-344-8. LCCN QA76.9.D343 S685

2004. URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/sco.pdf.

Luszczek:2004:DIE

[619] Piotr Luszczek and Jack Dongarra. Design of interactive environment for numerically intensive parallel linear algebra calculations. In Bubak et al. [1085], page ?? CO-DEN LNCSD9. ISBN 3-540-22114-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13

2004. URL http://www.netlib.org/
netlib/utk/people/JackDongarra/
PAPERS/conie-iccs2004.pdf.

Song:2004:ACE

[620] F. Song, F. Wolf, N. Bhatia, J. Dongarra, and S. Moore. An algebra for cross-experiment performance analysis. In Eigenmann [1089], pages 63–72. ISBN 0-7695-2197-5. LCCN QA76.6 .I548 2004.

Tanimura:2004:IPT

[621] Y. Tanimura, K. Aoi, T. Hiroyasu, M. Miki, Y. Okamoto, and J. Dongarra. Implementation of protein tertiary structure prediction system with NetSolve. In IEEE [1092], pages 320– 327. ISBN 0-7695-2138-X. LCCN QA76.88.

Vadhiyar:2004:GGB

[622] Sathish S. Vadhiyar and Jack J. Dongarra. GrADSolve — a grid-based RPC system for parallel computing with application-level scheduling. Journal of Parallel and Distributed Computing, 64 (6):774–783, June 2004. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Vadhiyar:2004:TAM

[623] Sathish S. Vadhiyar, Graham E. Fagg, and Jack J. Dongarra. Towards an accurate model for collective communications. The International Journal of High Performance Computing Applications, 18(1):159–167, February 2004. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL

                                        http://hpc.sagepub.com/content/
                                        18/1/159.full.pdf+html.

YarKhan:2004:BSA

[624] Asim YarKhan and Jack Dongarra. Biological sequence alignment on the Computational Grid using the Grads Framework. Journal of Grid Computing, 1(1):??, ???? 2004. ISSN 1570-7873 (print), 1572-9184 (elec-

tronic). URL http://www.netlib.
org/netlib/utk/people/JackDongarra/
PAPERS/fasta_grads.pdf.

Beck:2005:NDM

[625] M. Beck, J. Dongarra, and J. S. Plank. NetSolve/D: a massively parallel grid execution system for scalable data intensive collaboration. In IEEE [1098], pages 223a–223a. ISBN 0-7695-2312-9. LCCN QA76.58 .I583 2005. IEEE Computer Society Order Number P2312.

Berman:2005:NGS

[626] F. Berman, H. Casanova, A. Chien, K. Cooper, H. Dail, A. Dasgupta, W. Deng, J. Dongarra, L. Johnsson, K. Kennedy, C. Koelbel, B. Liu, X. Liu, A. Mandal, G. Marin, M. Mazina, J. Mellor-Crummey, C. Mendes, A. Olugbile, M. Patel, D. Reed, Z. Shi, O. Sievert, H. Xia, and A. YarKhan. New Grid scheduling and rescheduling methods in the GrADS Project. International Journal of Parallel Programming, 33(2–3):209–229, June 2005. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0885-7458&
volume=33&issue=2&spage=209.

Chen:2005:CNG

[627] Zizhong Chen and Jack J. Dongarra. Condition numbers of Gaussian random matrices. SIAM Journal on Matrix Analysis and Applications, 27(3):603–620, July 2005.

CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

                                        URL http://epubs.siam.org/sam-
                                        bin/dbq/article/61641.

Demmel:2005:LPR

[628] Jim Demmel and Jack Dongarra. LA-PACK 2005 prospectus: Reliable and scalable software for linear algebra computations on high end computers. LAPACK Working Note 164, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, February

                                        2005. URL http://www.netlib.org/
                                        lapack/lawns/lawn164.ps;   http:
                                        //www.netlib.org/lapack/lawnspdf/

lawn164.pdf. UT-CS-05-546, February 2005.

Demmel:2005:SAL

[629] J. Demmel, J. Dongarra, V. Eijkhout, E. Fuentes, A. Petitet, R. Vuduc, R. C. Whaley, and K. Yelick. Selfadapting linear algebra algorithms and software. Proceedings of the IEEE, 93 (2):293–312, February 2005. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Dongarra:2005:HPC

[630] Jack Dongarra, Thomas Sterling, Horst Simon, and Erich Strohmaier. Highperformance computing: Clusters, constellations, MPPs, and future directions. Computing in Science and Engineering, 7(2):51–59, March/April 2005. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

                                        URL http://csdl.computer.org/
                                        dl/mags/cs/2005/02/c2051.htm;
                                        http://csdl.computer.org/dl/mags/
                                        cs/2005/02/c2051.pdf.

Eijkhout:2005:CSS

[631] Victor Eijkhout, Erika Fuentes, Thomas Eidson, and Jack Dongarra. The component structure of a self-adapting numerical software system. International Journal of Parallel Programming, 33(2– 3):137–143, June 2005. CODEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic). URL http:

//www.springerlink.com/openurl.
asp?genre=article&issn=0885-7458&
volume=33&issue=2&spage=137.

Fagg:2005:PFT

[632] Graham E. Fagg, Edgar Gabriel, Zizhong Chen, Thara Angskun, George Bosilca, Jelena Pjesivac-Grbovic, and Jack J. Dongarra. Process fault tolerance: Semantics, design and applications for high performance computing. The International Journal of High Performance Computing Applications, 19(4):465–477, Winter 2005. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

//hpc.sagepub.com/content/19/4/
465.full.pdf+html.

Gabriel:2005:EDC

[633] Edgar Gabriel, Graham E. Fagg, and Jack J. Dongarra. Evaluating dynamic communicators and one-sided operations for current MPI libraries. The International Journal of High Performance Computing Applications, 19(1):67–79, Spring 2005. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http://

hpc.sagepub.com/content/19/1/67.
full.pdf+html.

Kranzlmuller:2005:RAP

[634] Dieter Kranzlmüller, Peter Kacsuk, and Jack Dongarra. Recent advances in Parallel Virtual Machine and Message Passing Interface. The International Journal of High Performance Computing Applications, 19 (2):99–101, Summer 2005. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http://

                                        hpc.sagepub.com/content/19/2/99.
                                        full.pdf+html.

Moura:2005:SIS

[635] J. M. F. Moura, M. Puschel, D. Padua, and J. Dongarra. Scanning the issue: Special issue on program generation, optimization, and platform adaptation. Proceedings of the IEEE, 93 (2):211–215, February 2005. CODEN IEEPAD. ISSN 0018-9219 (print), 1558-2256 (electronic).

Parashar:2005:EIC

[636] Manish Parashar, Rajeev Muralidhar, Wonsuck Lee, Dorian Arnold, Jack Dongarra, and Mary Wheeler. Enabling interactive and collaborative oil reservoir simulations on the Grid. Concurrency and Computation: Practice and Experience, 17(11):1387–1414, September 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Pjesivac-Grbovic:2005:PAM

[637] J. Pjesivac-Grbovic, T. Angskun, G. Bosilca, G. E. Fagg, E. Gabriel, and J. J. Dongarra. Performance analysis of MPI collective operations. In IEEE [1098], pages 272a–272a. ISBN 0-7695-2312-9. LCCN QA76.58 .I583 2005.

IEEE Computer Society Order Number P2312.

Strohmaier:2005:RTM

[638] Erich Strohmaier, Jack J. Dongarra, Hans W. Meuer, and Horst D. Simon. Recent trends in the marketplace of high performance computing. Parallel Computing, 31(3–4):261– 273, March/April 2005. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Vadhiyar:2005:SAG

[639] Sathish S. Vadhiyar and Jack J. Dongarra. Self adaptivity in Grid computing. Concurrency and Computation: Practice and Experience, 17(2–4):235– 257, February/April 2005. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

YarKhan:2005:BSA

[640] Asim YarKhan and Jack J. Dongarra. Biological sequence alignment on the computational grid using the GrADS framework. Future Generation Computer Systems, 21(6):980–986, June 2005. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Buttari:2006:UMP

[641] Alfredo Buttari, Jack J. Dongarra, Jakub Kurzak, Piotr Luszczek, and Stanimire Tomov. Using mixed precision for sparse matrix computations to enhance the performance while achieving 64-bit accuracy. LAPACK Working Note 180, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, October 22, 2006. URL http:/

/www.netlib.org/lapack/lawnspdf/
lawn180.pdf.

Dongarra:2006:SAN

[642] J. Dongarra, G. Bosilca, Z. Chen, V. Eijkhout, G. E. Fagg, E. Fuentes, J. Langou, P. Luszczek, J. Pjesivac-Grbovic, K. Seymour, H. You, and S. S. Vadhiyar. Self-Adapting Numerical Software (SANS) effort. IBM Journal of Research and Development, 50(2/3): 223–238, March /May 2006. CODEN IBMJAE. ISSN 0018-8646 (print), 2151-8556 (electronic). URL http://

                                        www.research.ibm.com/journal/rd/
                                        502/dongarra.html.

Dongarra:2006:SIT

[643] Jack Dongarra and Bernard Tourancheau. Special issue on tools in the ACTS Collection 2004. The International Journal of High Performance Computing Applications, 20 (3):317, Fall 2006. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

                                        //hpc.sagepub.com/content/20/3/
                                        317.full.pdf+html.

Dongarra:2006:THPb

[644] J. Dongarra. Trends in highperformance computing. IEEE Circuits & Devices, 22(1):22–27, ???? 2006. ISSN 8755-3996.

Emad:2006:AAN

[645] Nahid Emad, S.-A. Shahzadeh-Fazeli, and Jack Dongarra. An asynchronous algorithm on the NetSolve global computing system. Future Generation Computer Systems, 22(3):279– 290, February 2006. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Kurzak:2006:ILA

[646] Jakub Kurzak and Jack Dongarra. Implementing linear algebra routines on multi-core processors with pipelining and a look ahead. LA-PACK Working Note 178, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, September 2006.

11 pp. URL http://www.netlib.
org/lapack/lawnspdf/lawn178.pdf;
 http://www.netlib.org/lapack/lawnspdf/

lawn178.ps. Also available as UT-CS-06-581.

Kurzak:2006:IMP

[647] Jakub Kurzak and Jack Dongarra. Implementation of the mixed-precision high performance LINPACK benchmark on the CELL Processor. LA-PACK Working Note 177, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, September 2006.

12 pp. URL http://www.netlib.
org/lapack/lawnspdf/lawn177.pdf;
 http://www.netlib.org/lapack/lawnspdf/

lawn177.ps. Also available as UT-CS-06-580.

Langou:2006:EPBa

[648] Julie Langou, Julien Langou, Piotr Luszczek, Jakub Kurzak, Alfredo Buttari, and Jack Dongarra. Exploiting the performance of 32 bit floating point arithmetic in obtaining 64 bit accuracy (revisiting iterative refinement for linear systems). LAPACK Working Note 175, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, June 2006.

17 pp. URL http://www.netlib.
org/lapack/lawnspdf/lawn175.pdf;

                                         http://www.netlib.org/lapack/lawnspdf/
                                        lawn175.ps.

Langou:2006:EPBb

[649] Julie Langou, Julien Langou, Piotr Luszczek, Jakub Kurzak, Alfredo Buttari, and Jack Dongarra. Exploiting the performance of 32 bit floating point arithmetic in obtaining 64 bit accuracy (revisiting iterative refinement for linear systems). In ACM [1104], page ?? ISBN 0-7695-2700-0. LCCN QA76.5 .P742 2006. Contains one CD-ROM.

Shi:2006:SWA

[650] Zhiao Shi and Jack J. Dongarra. Scheduling workflow applications on processors with different capabilities. Future Generation Computer Systems, 22(6):665–675, May 2006. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

YarKhan:2006:RDG

[651] Asim YarKhan, Keith Seymour, Kiran Sagi, Zhiao Shi, and Jack Dongarra. Recent developments in Grid-Solve. The International Journal of High Performance Computing Applications, 20(1):131–141, Spring 2006. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL

                                        http://hpc.sagepub.com/content/
                                        20/1/131.full.pdf+html.

Baboulin:2007:CCC

[652] Marc Baboulin, Jack J. Dongarra, Serge Gratton, and Julien Langou. Computing the conditioning of the components of a linear least squares solution. LAPACK Working Note 193, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, September

2007. URL http://www.netlib.org/
lapack/lawnspdf/lawn193.pdf.

Buttari:2007:CPT

[653] Alfredo Buttari, Julien Langou, Jakub Kurzak, and Jack J. Dongarra. A class of parallel tiled linear algebra algorithms for multicore architectures. LAPACK Working Note 191, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, September

2007. URL http://www.netlib.org/
lapack/lawnspdf/lawn191.pdf.

Buttari:2007:LPH

[654] Alfredo Buttari, Jack J. Dongarra, and Jakub Kurzak. Limitations of the PlayStation 3 for high performance cluster computing. LAPACK Working Note 185, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA,

May 2007. URL http://www.netlib.
org/lapack/lawnspdf/lawn185.pdf.

Buttari:2007:MPI

[655] Alfredo Buttari, Jack Dongarra, Julie Langou, Julien Langou, Piotr Luszczek, and Jakub Kurzak. Mixed precision iterative refinement techniques for the solution of dense linear systems. The International Journal of High Performance Computing Applications, 21(4):457–466, November 2007. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL

http://hpc.sagepub.com/content/
21/4/457.full.pdf+html.

Buttari:2007:PTF

[656] Alfredo Buttari, Julien Langou, Jakub Kurzak, and Jack J. Dongarra. Parallel tiled QR factorization for multicore architectures. LAPACK Working Note 190, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, July

2007. URL http://www.netlib.org/
lapack/lawnspdf/lawn190.pdf. Pub-

lished in [676].

Demmel:2007:PNL

[657] James W. Demmel, Jack J. Dongarra, Beresford N. Parlett, William Kahan, Ming Gu, David S. Bindel, Yozo Hida, Xiaoye S. Li, Osni A. Marques, E. Jason Riedy, Christof Vömel, Julien Langou, Piotr Luszczek, Jakub Kurzak, Alfredo Buttari, Julie Langou, and Stanimire Tomov. Prospectus for the next LAPACK and ScaLAPACK libraries. LAPACK Working Note 181, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, March 11,

2007. URL http://www.netlib.org/
lapack/lawnspdf/lawn181.pdf.

DiMartino:2007:SIS

[658] Beniamino Di Martino, Dieter Kranzlmüller, and Jack Dongarra. Special issue on selected papers from the EuroPVM/MPI 2005 Conference , Sorrento, Italy, 18-21 September 2005 — preface. The International Journal of High Performance Computing Applications, 21(2):129–131, Summer 2007. ISSN 1094-3420 (print), 1741-2846 (electronic).

Dongarra:2007:B

[659] Jack Dongarra, Victor Eijkhout, and Julien Langou. BLAS. In Hogben [1115], page ?? ISBN 1-58488-510-6 (hardcover), 1-4200-1057-3 (e-book). LCCN QA184.2 .H36

2007. URL http://www.crcnetbase.
com/isbn/9781420010572;   http:/
/www.crcnetbase.com/isbn/9781584885108;
http://www.loc.gov/catdir/enhancements/
fy0647/2006045491-d.html. Asso-

ciate editors: Richard Brualdi, Anne Greenbaum and Roy Mathias.

Dongarra:2007:HEC

[660] Jack Dongarra and Piotr Luszczek. How elegant code evolves with hardware: The case of Gaussian elimination. In Oram and Wilson [1114], chapter 14, pages 229–252. ISBN 0-596-51004-7 (paperback). LCCN QA76.758 .B428 2007; QA76.758 .B43 2007; QA76.758 .B48 2007. URL http:/

/proquest.safaribooksonline.com/
9780596510046;   http://www.oreilly.
com/catalog/9780596510046.

Jeannot:2007:IRT

[661] Emmanuel Jeannot, Keith Seymour, Asym Yarkhan, and Jack J. Dongarra. Improved runtime and transfer time prediction mechanisms in a network enabled servers middleware. Parallel Processing Letters, 17(1):47–59, March 2007. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Kurzak:2007:IMP

[662] Jakub Kurzak and Jack Dongarra. Implementation of mixed precision in solving systems of linear equations on the Cell processor. Concurrency and Computation: Practice and Experience, 19(10):1371–1385, July 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kurzak:2007:SSL

[663] Jakub Kurzak, Alfredo Buttari, and Jack J. Dongarra. Solving systems of linear equations on the CELL processor using Cholesky factorization. LA-PACK Working Note 184, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, May 2007. URL http:/

/www.netlib.org/lapack/lawnspdf/
lawn184.pdf.

Langou:2007:RPI

[664] J. Langou, Z. Chen, G. Bosilca, and J. Dongarra. Recovery patterns for iterative methods in a parallel unstable environment. SIAM Journal on Scientific Computing, 30(1):102–116, 2007. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Luszczek:2007:HPD

[665] Piotr Luszczek and Jack Dongarra. High performance development for high end computing with Python Language Wrapper (PLW). The International Journal of High Performance Computing Applications, 21 (3):360–369, August 2007. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

//hpc.sagepub.com/content/21/3/
360.full.pdf+html.

DiMartino:2007:P

[666] Beniamino Di Martino, Dieter Kranzlmüller, and Jack Dongarra. Preface. The International Journal of High Performance Computing Applications, 21(2):129–131, May 2007. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

//hpc.sagepub.com/content/21/2/
129.full.pdf+html.

Mohr:2007:SPE

[667] Bernd Mohr, Jesper Larsson Traff, Joachim Worringen, and Jack Dongarra. Selected papers from EuroPVM/ MPI 2006. Parallel Computing, 33(9): 593–594, ???? 2007. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pjesivac-Grbovic:2007:MCA

[668] Jelena Pješivac-Grbović, George Bosilca, Graham E. Fagg, Thara Angskun, and Jack J. Dongarra. MPI collective algorithm selection and quadtree encoding. Parallel Computing, 33(9):613– 623, September 2007. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Pjesivac-Grbovic:2007:PAM

[669] Jelena Pjesivac-Grbovic, Thara Angskun, George Bosilca, Graham E. Fagg, Edgar Gabriel, and Jack J. Dongarra. Performance analysis of MPI collective operations. The Journal of Networks, Software Tools, and Cluster Computing, 10(2):127–143, ???? 2007. ISSN 1386-7857.

Vomel:2007:UBS

[670] Christof Vömel, Stanimire Z. Tomov, Lin-Wang Wang, Osni A. Marques, and Jack J. Dongarra. The use of bulk states to accelerate the band edge state calculation of a semiconductor quantum dot. Journal of Computational Physics, 223 (2):774–782, May 1, 2007. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0021999106004736.

Wasniewski:2007:EIS

[671] Jerzy Waśniewski, Jack Dongarra, Kaj Madsen, Sivan Toledo, and Zahari Zlatev. Editorial introduction to the special issue on computational linear algebra and sparse matrix computations. Applicable algebra in engineering, communication and computing, 18 (3):205–207, 2007. CODEN AAE-CEW. ISSN 0938-1279 (print), 1432-0622 (electronic).

Wolf:2007:AAI

[672] Felix Wolf, Bernd Mohr, Jack Dongarra, and Shirley Moore. Automatic analysis of inefficiency patterns in parallel applications. Concurrency and Computation: Practice and Experience, 19(11):1481–1496, August 10, 2007. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Alvaro:2008:FSS

[673] Wesley Alvaro, Jakub Kurzak, and Jack J. Dongarra. Fast and small short vector SIMD matrix multiplication kernels for the CELL processor. LAPACK Working Note 189, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, January

                                        2008. URL http://www.netlib.org/
                                        lapack/lawnspdf/lawn189.pdf.

Baboulin:2008:SID

[674] Marc Baboulin, Jack J. Dongarra, and Stanimire Tomov. Some issues in dense linear algebra for multicore and special purpose architectures. LAPACK Working Note 200, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA,

May 2008. URL http://www.netlib.
org/lapack/lawnspdf/lawn200.pdf.

Bosilca:2008:ABF

[675] George Bosilca, Remi Delmas, Jack J. Dongarra, and Julien Langou. Algorithmic based fault tolerance applied to high performance computing. LAPACK Working Note 205, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, May 23,

2008. URL http://www.netlib.org/
lapack/lawnspdf/lawn205.pdf.

Buttari:2008:PTF

[676] Alfredo Buttari, Julien Langou, Jakub Kurzak, and Jack Dongarra. Parallel tiled QR factorization for multicore architectures. Concurrency and Computation: Practice and Experience, 20 (13):1573–1590, September 2008. CO-DEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Buttari:2008:PTQ

[677] Alfredo Buttari, Julien Langou, Jakub Kurzak, and Jack Dongarra. Parallel tiled QR factorization for multicore architectures. Concurrency and Computation: Practice and Experience, 20 (13):1573–1590, September 10, 2008. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Buttari:2008:UMP

[678] Alfredo Buttari, Jack Dongarra, Jakub Kurzak, Piotr Luszczek, and Stanimir Tomov. Using mixed precision for sparse matrix computations to enhance the performance while achieving 64-bit accuracy. ACM Transactions on Mathematical Software, 34(4): 17:1–17:22, July 2008. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Chen:2008:ABF

[679] Zizhong Chen and Jack Dongarra. Algorithm-based fault tolerance for fail-stop failures. IEEE Transactions on Parallel and Distributed Systems, 19 (12):1628–1641, December 2008. CO-DEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

DiMartino:2008:SSG

[680] Beniamino Di Martino, Dieter Kranzlmueller, and Jack Dongarra. Special section: Grid computing and the message passing interface. Future Generation Computer Systems, 24(2): 119–120, ???? 2008. CODEN FG-SEVI. ISSN 0167-739x (print), 1872-7115 (electronic).

Dimov:2008:SSA

[681] Ivan Dimov, Jack Dongarra, Kaj Madsen, Jerzy Wasniewski, and Zahari Zlatev. Special section: Applications of distributed and grid computing. Future Generation Computer Systems, 24 (6):582–584, June 2008. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Dongarra:2008:B

[682] Jack Dongarra and Thomas Haigh. Biographies. IEEE Annals of the History of Computing, 30(2):74–81, April/June 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Dongarra:2008:MPH

[683] Jack Dongarra, Jean-François Pineau, Yves Robert, and Frédéric Vivien. Matrix product on heterogeneous master– worker platforms. In ACM [1124], pages 53–62. ISBN 1-59593-795-1. LCCN QA76.642 .A27 2008.

Dongarra:2008:NNB

[684] Jack Dongarra, Gene H. Golub, Eric Grosse, Cleve Moler, and Keith Moore. Netlib and NA-Net: Building a scientific computing community. IEEE Annals of the History of Computing, 30 (2):30–41, April/June 2008. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=4544554.

Dongarra:2008:PLB

[685] Jack J. Dongarra and Julien Langou. The problem with the Linpack benchmark matrix generator. LAPACK Working Note 206, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, June 12,

2008. URL http://www.netlib.org/
lapack/lawnspdf/lawn206.pdf. Ver-

sion 1; version 2 is dated 18 September 2008.

Dongarra:2008:RMP

[686] Jack Dongarra, Jean-François Pineau, Yves Robert, Zhiao Shi, and Frédéric Vivien. Revisiting matrix product on master-worker platforms. International Journal of Foundations of Computer Science, 19(6):1317–1336, December 2008. CODEN IFCSEN. ISSN 0129-0541 (print), 1793-6373 (electronic).

Dongarra:2008:SSC

[687] Jack Dongarra and Bernard Tourancheau. Special section: Cluster and computational grids for scientific computing. Future Generation Computer Systems, 24(1):30, January 2008. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Gustavson:2008:LCK

[688] Fred G. Gustavson, Jerzy Wasniewski, and Jack Dongarra. Level-3 Cholesky kernel subroutine of a fully portable high performance minimal storage hybrid format Cholesky algorithm. LAPACK Working Note 211, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, December

                                        2008. URL http://www.netlib.org/
                                        lapack/lawnspdf/lawn211.pdf.

Gustavson:2008:RFP

[689] Fred G. Gustavson, Jerzy Wasniewski, Jack J. Dongarra, and Julien Langou. Rectangular full packed format for Cholesky’s algorithm: Factorization, solution and inversion. LA-PACK Working Note 199, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, April 2008. URL http:/

                                        /www.netlib.org/lapack/lawnspdf/
                                        lawn199.pdf.

Kurzak:2008:FCP

[690] Jakub Kurzak and Jack J. Dongarra. QR factorization for the CELL processor. LAPACK Working Note 201, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA,

May 2008. URL http://www.netlib.
org/lapack/lawnspdf/lawn201.pdf.

Kurzak:2008:PHP

[691] Jakub Kurzak, Alfredo Buttari, Piotr Luszczek, and Jack Dongarra. The PlayStation 3 for high-performance scientific computing. Computing in Science and Engineering, 10(3):84–87, May/June 2008. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kurzak:2008:SSL

[692] Jakub Kurzak, Alfredo Buttari, and Jack Dongarra. Solving systems of linear equations on the CELL processor using Cholesky factorization. IEEE Transactions on Parallel and Distributed Systems, 19(9):1175–1186, September 2008. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ltaief:2008:PBH

[693] Hatem Ltaief, Jakub Kurzak, and Jack Dongarra. Parallel block Hessenberg reduction using algorithmsby-tiles for multicore architectures revisited. LAPACK Working Note 208, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, August

2008. URL http://www.netlib.org/
lapack/lawnspdf/lawn208.pdf.

Ltaief:2008:PBT

[694] Hatem Ltaief, Jakub Kurzak, and Jack Dongarra. Parallel band two-sided matrix bidiagonalization for multicore architectures. LAPACK Working Note 209, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, October

                                        2008. URL http://www.netlib.org/
                                        lapack/lawnspdf/lawn209.pdf.

Martino:2008:SSG

[695] Beniamino Di Martino, Dieter Kranzlmüller, and Jack Dongarra. Special section: Grid computing and the Message Passing Interface. Future Generation Computer Systems, 24(2):119– 120, February 2008. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Tomov:2008:TDL

[696] Stanimire Tomov, Jack Dongarra, and Marc Baboulin. Towards dense linear algebra for hybrid GPU accelerated manycore systems. LAPACK Working Note 210, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, October 2008. URL http:/

                                        /www.netlib.org/lapack/lawnspdf/
                                        lawn210.pdf.

Vomel:2008:SAE

[697] Christof Vömel, Stanimire Z. Tomov, Osni A. Marques, A. Canning, Lin-Wang Wang, and Jack J. Dongarra. State-of-the-art eigensolvers for electronic structure calculations of large scale nano-systems. Journal of Computational Physics, 227(15):7113–7124, 2008. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Agullo:2009:CSO

[698] Emmanuel Agullo, Bilel Hadri, Hatem Ltaief, and Jack Dongarra. Comparative study of one-sided factorizations with multiple software packages on multi-core hardware. LAPACK Working Note 217, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, April 28, 2009. URL http://

www.netlib.org/lapack/lawnspdf/

lawn217.pdf. UT-CS-09-640.

Baboulin:2009:ASC

[699] Marc Baboulin, Alfredo Buttari, Jack Dongarra, Jakub Kurzak, Julie Langou, Julien Langou, Piotr Luszczek, and Stanimire Tomov. Accelerating scientific computations with mixed precision algorithms. Computer Physics Communications, 180(12): 2526–2533, December 2009. CODEN CPHCBZ. ISSN 0010-4655 (print), 1879-2944 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0010465508003846.

Baboulin:2009:CCC

[700] Marc Baboulin, Jack Dongarra, Serge Gratton, and Julien Langou. Computing the conditioning of the components of a linear least-squares solution. Numerical Linear Algebra with Applications, 16(7):517–533, 2009. CODEN NLAAEM. ISSN 1070-5325 (print), 1099-1506 (electronic).

Bosilca:2009:ABF

[701] George Bosilca, Rémi Delmas, Jack Dongarra, and Julien Langou. Algorithmbased fault tolerance applied to high performance computing. Journal of Parallel and Distributed Computing, 69 (4):410–416, April 2009. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Buttari:2009:CPT

[702] Alfredo Buttari, Julien Langou, Jakub Kurzak, and Jack Dongarra. A class of parallel tiled linear algebra algorithms for multicore architectures. Parallel Computing, 35(1):38–53, January 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Cappello:2009:FSI

[703] Franck Cappello, Thomas Herault, and Jack Dongarra. Foreword [special issue: selected papers from the 14th European PVM/MPI Users Group Meeting, Paris, September 30–October 3, 2007]. Parallel Computing, 35(12): 571, December 2009. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Chen:2009:HSS

[704] Zizhong Chen and J. Dongarra. Highly scalable self-healing algorithms for high performance scientific computing. IEEE Transactions on Computers, 58(11):1512– 1524, November 2009. CODEN IT-COB4. ISSN 0018-9340 (print), 1557-9956 (electronic). URL http://

                                        ieeexplore.ieee.org/stamp/stamp.
                                        jsp?tp=&arnumber=4799775.

Dongarra:2009:E

[705] Jack Dongarra and Bernard Tourancheau. Editorial. The International Journal of High Performance Computing Applications, 23 (3):195, August 2009. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

                                        //hpc.sagepub.com/content/23/3/
                                        195.full.pdf+html.

Dongarra:2009:GEN

[706] Jack Dongarra and Bernard Tourancheau. Guest editor’s note: special issue on clusters and computational grids for scientific computing. Parallel Processing Letters, 19(3):379, 2009. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Dongarra:2009:IES

[707] Jack Dongarra, Pete Beckman, Patrick Aerts, Frank Cappello, Thomas Lippert, Satoshi Matsuoka, Paul Messina, Terry Moore, Rick Stevens, Anne Trefethen, and Mateo Valero. The International Exascale Software Project: a call to cooperative action by the global high-performance community. The International Journal of High Performance Computing Applications, 23(4): 309–322, November 2009. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

//hpc.sagepub.com/content/23/4/
309.full.pdf+html.

Dongarra:2009:PLB

[708] Jack J. Dongarra and Julien Langou. The problem with the LINPACK Benchmark 1.0 Matrix Generator. The International Journal of High Performance Computing Applications, 23(1): 5–13, February 2009. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http://hpc.

sagepub.com/content/23/1/5.full.
pdf+html.

Hadri:2009:EPT

[709] Bilel Hadri, Hatem Ltaief, Emmanuel Agullo, and Jack Dongarra. Enhancing parallelism of tile QR factorization for multicore architectures. LAPACK Working Note 222, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, September

                                        4, 2009. URL http://www.netlib.
                                        org/lapack/lawnspdf/lawn222.pdf.

UT-CS-09-645.

Kurzak:2009:FCB

[710] Jakub Kurzak and Jack Dongarra. QR factorization for the Cell Broadband Engine. Scientific Programming, 17 (1–2):31–42, ???? 2009. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Kurzak:2009:FDS

[711] Jakub Kurzak and Jack Dongarra. Fully dynamic scheduler for numerical computing on multicore processors. LAPACK Working Note 220, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, June 4,

                                        2009. URL http://www.netlib.org/
                                        lapack/lawnspdf/lawn220.pdf. UT-

CS-09-643.

Kurzak:2009:OMM

[712] Jakub Kurzak, Wesley Alvaro, and Jack Dongarra. Optimizing matrix multiplication for a short-vector SIMD architecture — CELL processor. Parallel Computing, 35(3):138–150, March 2009. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Kurzak:2009:SLA

[713] Jakub Kurzak, Hatem Ltaief, Jack Dongarra, and Rosa M. Badia. Scheduling linear algebra operations on multicore processors. LAPACK Working Note 213, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, February 2009. URL http:/

/www.netlib.org/lapack/lawnspdf/
lawn213.pdf.

Kurzak:2009:STS

[714] Jakub Kurzak, Hatem Ltaief, and Jack Dongarra. Scheduling two-sided transformations using algorithms-by-tiles on multicore architectures. LA-PACK Working Note 214, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, February

2009. URL http://www.netlib.org/
lapack/lawnspdf/lawn214.pdf.

Lastovetsky:2009:HPH

[715] Alexey Lastovetsky and J. J. Dongarra. High-Performance Heterogeneous Computing. Wiley series on parallel and distributed computing. Wiley, New York, NY, USA, 2009. ISBN 0-470-04039-4 (cloth). xii + 267 + 2 pp. LCCN QA76.88 .L38 2009.

Li:2009:NAT

[716] Yinan Li, Jack Dongarra, and Stanimire Tomov. A note on auto-tuning GEMM for GPUs. LAPACK Working Note 212, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, January 2009. URL http:/

/www.netlib.org/lapack/lawnspdf/
lawn212.pdf.

Ltaief:2009:SHP

[717] Hatem Ltaief, Stanimire Tomov, Rajib Nath, Peng Du, and Jack Dongarra. A scalable high performant Cholesky factorization for multicore with GPU accelerators. LAPACK Working Note 223, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, November

25, 2009. URL http://www.netlib.
org/lapack/lawnspdf/lawn223.pdf.

UT-CS-09-646.

Song:2009:DTS

[718] Fengguang Song, Asim YarKhan, and Jack Dongarra. Dynamic task scheduling for linear algebra algorithms on distributed-memory multicore systems. LAPACK Working Note 221, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, April 13,

2009. URL http://www.netlib.org/
lapack/lawnspdf/lawn221.pdf. UT-

CS-09-638.

Tomov:2009:ARU

[719] Stanimire Tomov and Jack Dongarra. Accelerating the reduction to upper Hessenberg form through hybrid GPUbased computing. LAPACK Working Note 219, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, May 24, 2009. URL http://

www.netlib.org/lapack/lawnspdf/

lawn219.pdf. UT-CS-09-642.

Youseff:2009:PES

[720] Lamia Youseff, Keith Seymour, Haihang You, Dmitrii Zagorodnov, Jack Dongarra, and Rich Wolski. Paravirtualization effect on single-and multithreaded memory-intensive linear algebra software. The Journal of Networks, Software Tools, and Cluster Computing, 12(2):101–122, ???? 2009. ISSN 1386-7857.

Agullo:2010:FCB

[721] Emmanuel Agullo, Cedric Augonnet, Jack Dongarra, Hatem Ltaief, Raymond Namyst, Samuel Thibault, and Stanimire Tomov. Faster, cheaper, better — a hybridization methodology to develop linear algebra software for GPUs. LAPACK Working Note 230, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, September 15, 2010. URL http:/

/www.netlib.org/lapack/lawnspdf/

lawn230.pdf. UT-CS-10-658. To appear in GPU Computing GEMs, vol. 2.

Agullo:2010:FMN

[722] E. Agullo, C. Augonnet, J. Dongarra, M. Faverge, H. Ltaief, S. Thibault, and S. Tomov. QR factorization on a multicore node enhanced with multiple GPU accelerators. LAPACK Working Note 233, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, October

2010. URL http://www.netlib.org/
lapack/lawnspdf/lawn233.pdf. UT-

CS-10-XXX, published in Proceedings of IPDPS 2011.

Agullo:2010:QFT

[723] Emmanuel Agullo, Camille Coti, Jack Dongarra, Thomas Herault, and Julien Langou. QR factorization of tall and skinny matrices in a grid computing environment. LAPACK Working Note 224, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, April 6,

2010. URL http://www.netlib.org/
lapack/lawnspdf/lawn224.pdf. UT-

CS-10-651. Published in the Proceedings of IPDPS 2010: 24th IEEE International Parallel and Distributed Processing Symposium Atlanta GA April 2010.

Angskun:2010:SHN

[724] Thara Angskun, Graham Fagg, George Bosilca, Jelena Pješivac-Grbović, and Jack Dongarra. Self-healing network for scalable fault-tolerant runtime environments. Future Generation Computer Systems, 26(3):479–485, March 2010. CODEN FGSEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Bosilca:2010:DGD

[725] G. Bosilca, A. Bouteiller, A. Danalis, T. Herault, P. Lemarinier, and J. Dongarra. DAGuE: a generic distributed DAG engine for high performance computing. LAPACK Working Note 231, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, September

                                        15, 2010. URL http://www.netlib.
                                        org/lapack/lawnspdf/lawn231.pdf.

UT-CS-10-659.

Bosilca:2010:DMT

[726] G. Bosilca, A. Bouteiller, A. Danalis, M. Faverge, H. Haidar, T. Herault, J. Kurzak, J. Langou, P. Lemarinier, H. Ltaief, P. Luszczekl, A. YarKhan, and J. Dongarra. Distributed-memory task execution and dependence tracking within DAGuE and the DPLASMA Project. LAPACK Working Note 232, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, September

                                        15, 2010. URL http://www.netlib.
                                        org/lapack/lawnspdf/lawn232.pdf.

UT-CS-10-660.

Bouteiller:2010:RML

[727] Aurelien Bouteiller, George Bosilca, and Jack Dongarra. Redesigning the message logging model for high performance. Concurrency and Computation: Practice and Experience, 22(16): 2196–2211, November 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Brady:2010:SNR

[728] Thomas Brady, Jack Dongarra, Michele Guidolin, Alexey Lastovetsky, and Keith Seymour. SmartGridRPC: The new RPC model for high performance Grid computing. Concurrency and Computation: Practice and Experience, 22(18):2467–2487, December 25, 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dongarra:2010:F

[729] Jack Dongarra. Foreword. In CUDA [946], pages xiii–xiv. ISBN 0-13-138768-5. LCCN QA76.76.A65.

Dongarra:2010:RTH

[730] Jack J. Dongarra, Hans W. Meuer, Horst D. Simon, and Erich Strohmaier. Recent trends in high performance computing. In Bultheel and Cools [1131], pages 93–107. ISBN 981-283-625-X. LCCN QA297 .B54 2010.

Dongarra:2010:RTT

[731] Jack Dongarra and Piotr Luszczek. Reducing the time to tune parallel dense linear algebra routines with partial execution and performance modelling. LAPACK Working Note 235, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, October 8,

2010. URL http://www.netlib.org/
lapack/lawnspdf/lawn235.pdf. UT-

CS-10-661.

Du:2010:COT

[732] Peng Du, Rick Weber, Piotr Luszczek, Stanimire Tomov, Gregory Peterson, and Jack Dongarra. From CUDA to OpenCL: Towards a performanceportable solution for multi-platform GPU programming. LAPACK Working Note 228, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, September 6, 2010. URL http:/

/www.netlib.org/lapack/lawnspdf/

lawn228.pdf. UT-CS-10-656.

Gustavson:2010:RFP

[733] Fred G. Gustavson, Jerzy Waśniewski, Jack J. Dongarra, and Julien Langou. Rectangular full packed format for Cholesky’s algorithm: factorization, solution, and inversion. ACM Transactions on Mathematical Software, 37(2): 18:1–18:21, April 2010. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Kurzak:2010:ITF

[734] Jakub Kurzak, Rajib Nath, Peng Du, and Jack Dongarra. An implementation of the tile QR factorization for a GPU and multiple CPUs. LA-PACK Working Note 229, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, September 15,

2010. URL http://www.netlib.org/
lapack/lawnspdf/lawn229.pdf. UT-

CS-10-657. Submitted to PARA’10.

Kurzak:2010:SCM

[735] Jakub Kurzak, David A. Bader, and J. J. Dongarra, editors. Scientific computing with multicore and accelerators, volume 10 of Chapman and Hall/CRC computational science. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2010. ISBN 1-4398-2536-X (hardback). xxxiii + 480 pp. LCCN Q183.9 .S325 2010.

Kurzak:2010:SDL

[736] Jakub Kurzak, Hatem Ltaief, Jack Dongarra, and Rosa M. Badia. Scheduling dense linear algebra operations on multicore processors. Concurrency and Computation: Practice and Experience, 22(1):15–44, January 2010. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Ltaief:2010:PTS

[737] Hatem Ltaief, Jakub Kurzak, and Jack Dongarra. Parallel two-sided matrix reduction to band bidiagonal form on multicore architectures. IEEE Transactions on Parallel and Distributed Systems, 21(4):417–423, April 2010. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Ltaief:2010:STS

[738] Hatem Ltaief, Jakub Kurzak, Jack Dongarra, and Rosa M. Badia. Scheduling two-sided transformations using tile algorithms on multicore architectures. Scientific Programming, 18(1):35–50, ???? 2010. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (electronic).

Nath:2010:IMG

[739] Rajib Nath, Stanimire Tomov, and Jack Dongarra. An improved MAGMA GEMM for Fermi GPUs. LA-PACK Working Note 227, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, July 29,

2010. URL http://www.netlib.org/
lapack/lawnspdf/lawn227.pdf. UT-

CS-10-655.

Sloot:2010:P

[740] Peter Sloot, Peter Coveney, and Jack Dongarra. Preface. Journal of Computational Science, 1(1):3–4, May 2010. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750310000177.

Tomov:2010:ARU

[741] Stanimire Tomov, Rajib Nath, and Jack Dongarra. Accelerating the reduction to upper Hessenberg, tridiagonal, and bidiagonal forms through hybrid GPU-based computing. Parallel Computing, 36(12):645–654, December 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Tomov:2010:DLA

[742] Stanimire Tomov, Rajib Nath, Hatem Ltaief, and Jack Dongarra. Dense linear algebra solvers for multicore with GPU accelerators. In 2010 IEEE International Symposium on Parallel & Distributed Processing, Workshops and Phd Forum (IPDPSW). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, April 2010.

Tomov:2010:DLAa

[743] Stanimire Tomov, Rajib Nath, Hatem Ltaief, and Jack Dongarra. Dense linear algebra solvers for multicore with GPU accelerators. LAPACK Working Note 225, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, April 18, 2010. URL http://

www.netlib.org/lapack/lawnspdf/

lawn225.pdf. UT-CS-09-649. Published in the Proceedings of IPDPS 2010: 24th IEEE International Parallel and Distributed Processing Symposium Atlanta GA April 2010.

Tomov:2010:TDL

[744] Stanimire Tomov, Jack Dongarra, and Marc Baboulin. Towards dense linear algebra for hybrid GPU accelerated manycore systems. Parallel Computing, 36(5–6):232–240, June 2010. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Agullo:2011:FEA

[745] Emmanuel Agullo, Jack Dongarra, Rajib Nath, and Stanimire Tomov. A fully empirical autotuned dense QR factorization for multicore architectures. LAPACK Working Note 242, INRIA, ????, March 9, 2011. URL http:/

/www.netlib.org/lapack/lawnspdf/

lawn242.pdf. INRIA-7526.

Agullo:2011:QOM

[746] Emmanuel Agullo, Camille Coti, Thomas Herault, Julien Langou, Sylvain Peyronnet, Ala Rezmerita, Franck Cappello, and Jack Dongarra. QCG-OMPI: MPI applications on grids. Future Generation Computer Systems, 27 (4):357–369, April 2011. CODEN FG-SEVI. ISSN 0167-739X (print), 1872-7115 (electronic).

Anzt:2011:BAR

[747] Hartwig Anzt, Stanimire Tomov, Jack Dongarra, and Vincent Heuveline. A block-asynchronous relaxation method for graphics processing units. LAPACK Working Note 258, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, December

2011. URL http://www.netlib.org/
lapack/lawnspdf/lawn258.pdf. UT-

CS-11-687.

Anzt:2011:GAA

[748] Hartwig Anzt, Piotr Luszczek, Jack Dongarra, and Vincent Heuveline. GPU-accelerated asynchronous error correction for mixed precision iterative refinement. LAPACK Working Note 260, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, December

2011. URL http://www.netlib.org/
lapack/lawnspdf/lawn260.pdf. UT-

CS-11-690.

Baboulin:2011:ALS

[749] Marc Baboulin, Jack Dongarra, Julien Herrmann, and Stanimire Tomov. Accelerating linear system solutions using randomization techniques. LAPACK Working Note 246, INRIA, ????, May

15, 2011. URL http://www.netlib.
org/lapack/lawnspdf/lawn246.pdf.

INRIA RR-7616.

Baboulin:2011:PTS

[750] Marc Baboulin, Dulceneia Becker, and Jack Dongarra. A parallel tiled solver for dense symmetric indefinite systems on multicore architectures. LAPACK Working Note 261, INRIA (Institut National de Recherche en Informatique et en Automatique), Rocquencourt, France, December 2011. URL http:/

/www.netlib.org/lapack/lawnspdf/

lawn261.pdf. INRIA-7762.

Dongarra:2011:ANA

[751] Jack Dongarra, Mathieu Faverge, Hatem Ltaief, and Piotr Luszczek. Achieving numerical accuracy and high performance using recursive tile LU factorization. LAPACK Working Note 259, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, December

2011. URL http://www.netlib.org/
lapack/lawnspdf/lawn259.pdf. UT-

CS-11-688.

Dongarra:2011:F

[752] Jack Dongarra. Foreword, page ?? Volume 7 of Chapman and Hall/CRC computational science series [947], 2011. ISBN 1-4398-1192-X. LCCN QA76.88 .H34 2011.

Dongarra:2011:GEN

[753] Jack Dongarra and Bernard Tourancheau. Guest Editors’ note: special issue on clusters, clouds and grids for scientific computing. Parallel Processing Letters, 21(2):109, June 2011. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Dongarra:2011:HFA

[754] Jack Dongarra, Mathieu Faverge, Thomas Herault, Julien Langou, and Yves Robert. Hierarchical QR factorization algorithms for multi-core cluster systems. LAPACK Working Note 257, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, October 2011. URL http:/

/www.netlib.org/lapack/lawnspdf/

lawn257.pdf. UT-CS-11-684.

Dongarra:2011:IES

[755] Jack Dongarra, Pete Beckman, Terry Moore, Patrick Aerts, Giovanni Aloisio, Jean-Claude Andre, David Barkai, Jean-Yves Berthou, Taisuke Boku, Bertrand Braunschweig, Franck Cappello, Barbara Chapman, Xuebin Chi, Alok Choudhary, Sudip Dosanjh, Thom Dunning, Sandro Fiore, Al Geist, Bill Gropp, Robert Harrison, Mark Hereld, Michael Heroux, Adolfy Hoisie, Koh Hotta, Zhong Jin, Yutaka Ishikawa, Fred Johnson, Sanjay Kale, Richard Kenway, David Keyes, Bill Kramer, Jesus Labarta, Alain Lichnewsky, Thomas Lippert, Bob Lucas, Barney Maccabe, Satoshi Matsuoka, Paul Messina, Peter Michielse, Bernd Mohr, Matthias S. Mueller, Wolfgang E. Nagel, Hiroshi Nakashima, Michael E. Papka, Dan Reed, Mitsuhisa Sato, Ed Seidel, John Shalf, David Skinner, Marc Snir, Thomas Sterling, Rick Stevens, Fred Streitz, Bob Sugar, Shinji Sumimoto, William Tang, John Taylor, Rajeev Thakur, Anne Trefethen, Mateo Valero, Aad van der Steen, Jeffrey Vetter, Peg Williams, Robert Wisniewski, and Kathy Yelick. The International Exascale Software Project roadmap. The International Journal of High Performance Computing Applications, 25 (1):3–60, February 2011. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http://hpc.

sagepub.com/content/25/1/3.full.
pdf+html.

Dongarra:2011:SPW

[756] Jack Dongarra and Bernard Tourancheau. Selected papers of the Workshop on Clusters, Clouds and Grids for Scientific Computing (CCGSC). The International Journal of High Performance Computing Applications, 25 (3):259–260, August 2011. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

//hpc.sagepub.com/content/25/3/
259.full.pdf+html.

Du:2011:ABF

[757] Peng Du, Aurelien Bouteiller, George Bosilca, Thomas Herault, and Jack Dongarra. Algorithm-based fault tolerance for dense matrix factorizations. LAPACK Working Note 253, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, August 5,

2011. URL http://www.netlib.org/
lapack/lawnspdf/lawn253.pdf. UT-

CS-11-676.

Du:2011:HPL

[758] Peng Du, Piotr Luszczek, and Jack Dongarra. High performance linear system solver with resilience to multiple soft errors. LAPACK Working Note 256, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, October

2011. URL http://www.netlib.org/
lapack/lawnspdf/lawn256.pdf. UT-

CS-11-683.

Du:2011:SER

[759] Peng Du, Piotr Luszczek, Stanimire Tomov, and Jack Dongarra. Soft error resilient QR factorization for hybrid system. LAPACK Working Note 252, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, July 1,

2011. URL http://www.netlib.org/
lapack/lawnspdf/lawn252.pdf. UT-

CS-11-675.

Gustavson:2011:LCF

[760] Fred G. Gustavson, Jerzy Wásniewski, Jack J. Dongarra, José R. Herrero, and Julien Langou. Level-3 Cholesky factorization routines as part of many Cholesky algorithms. LAPACK Working Note 249, ????,

????, 2011. URL http://www.netlib.
org/lapack/lawnspdf/lawn249.pdf.

DTU/IMM-Technical-Report-2011-11, submitted at TOMS.

Haidar:2011:ADS

[761] Azzam Haidar, Hatem Ltaief, Asim YarKhan, and Jack Dongarra. Analysis of dynamically scheduled tile algorithms for dense linear algebra on multicore architectures. LAPACK Working Note 243, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, March 10, 2011. URL http://

www.netlib.org/lapack/lawnspdf/

lawn243.pdf. UT-CS-11-666. Submitted at Concurrency and Computations.

Haidar:2011:PRCa

[762] Azzam Haidar, Hatem Ltaief, and Jack Dongarra. Parallel reduction to condensed forms for symmetric eigenvalue problems using aggregated fine-grained and memory-aware kernels. LAPACK Working Note 254, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, August 5,

2011. URL http://www.netlib.org/
lapack/lawnspdf/lawn254.pdf. UT-

CS-11-677 Aug 5 2011.

Haidar:2011:PRCb

[763] Azzam Haidar, Hatem Ltaief, and Jack Dongarra. Parallel reduction to condensed forms for symmetric eigenvalue problems using aggregated fine-grained and memory-aware kernels. In Lathrop et al. [1136], pages 8:1–8:11. ISBN 1-4503-0771-X. LCCN ????

Jagode:2011:TBP

[764] Heike Jagode, Andreas Knüpfer, Jack Dongarra, Matthias Jurenz, Matthias S. Müller, and Wolfgang E. Nagel. Trace-based performance analysis for the petascale simulation code FLASH. The International Journal of High Performance Computing Applications, 25(4):428–439, November 2011. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL

http://hpc.sagepub.com/content/
25/4/428.full.pdf+html.

Kurzak:2011:AGF

[765] Jakub Kurzak, Stanimire Tomov, and Jack Dongarra. Autotuning GEMMs for Fermi. LAPACK Working Note 245, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, April 18,

2011. URL http://www.netlib.org/
lapack/lawnspdf/lawn245.pdf. UT-

CS-11-671. Submitted at SC11 November 12-18, 2011, Seattle, Washington, USA.

Ltaief:2011:HPB

[766] Hatem Ltaief, Piotr Luszczek, and Jack Dongarra. High performance bidiagonal reduction using tile algorithms on homogeneous multicore architectures. LAPACK Working Note 247, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, May 18,

2011. URL http://www.netlib.org/
lapack/lawnspdf/lawn247.pdf. UT-

CS-11-673. Submitted at TOMS.

Ltaief:2011:PHP

[767] Hatem Ltaief, Piotr Luszczek, and Jack Dongarra. Profiling high performance dense linear algebra algorithms on multicore architectures for power and energy efficiency. LAPACK Working Note 251, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, June 21,

2011. URL http://www.netlib.org/
lapack/lawnspdf/lawn251.pdf. UT-

CS-11-674.

Luszczek:2011:TST

[768] Piotr Luszczek, Hatem Ltaief, and Jack Dongarra. Two-stage tridiagonal reduction for dense symmetric matrices using tile algorithms on multicore architectures. LAPACK Working Note 244, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, April 18,

2011. URL http://www.netlib.org/
lapack/lawnspdf/lawn244.pdf. UT-

CS-11-670.

Nath:2011:OSD

[769] Rajib Nath, Stanimire Tomov, Tingxing “Tim” Dong, and Jack Dongarra. Optimizing symmetric dense matrixvector multiplication on GPUs. In Lathrop et al. [1136], pages 6:1–6:10. ISBN 1-4503-0771-X. LCCN ????

Song:2011:ESM

[770] Fengguang Song, Stanimire Tomov, and Jack Dongarra. Efficient support for matrix computations on heterogeneous multi-core and multi-GPU architectures. LAPACK Working Note 250, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, June 16,

2011. URL http://www.netlib.org/
lapack/lawnspdf/lawn250.pdf. UT-

CS-11-668.

Song:2011:STC

[771] Fengguang Song, Hatem Ltaief, Bilel Hadri, and Jack Dongarra. Scalable tile communication-avoiding QR factorization on multicore cluster systems. LAPACK Working Note 241, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, March 4,

2011. URL http://www.netlib.org/
lapack/lawnspdf/lawn241.pdf. UT-

CS-10-653. Published at SC’10.

Vetter:2011:KBH

[772] Jeffrey S. Vetter, Richard Glassbrook, Jack Dongarra, Karsten Schwan, Bruce Loftis, Stephen McNally, Jeremy Meredith, James Rogers, Philip Roth, Kyle Spafford, and Sudhakar Yalamanchili. Keeneland: Bringing heterogeneous GPU computing to the computational science community. Computing in Science and Engineering, 13(5):90– 95, September/October 2011. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Watkins:2011:FA

[773] David S. Watkins. Francis’s algorithm. American Mathematical Monthly, 118 (5):387–403, May 2011. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic). URL http:

//www.jstor.org/stable/info/10.
4169/amer.math.monthly.118.05.387.

White:2011:HPH

[774] J. B. White III and J. J. Dongarra. High-performance highresolution semi-Lagrangian tracer transport on a sphere. Journal of Computational Physics, 230(17): 6778–6799, July 20, 2011. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0021999111003123.

Bosilca:2012:DGD

[775] George Bosilca, Aurelien Bouteiller, Anthony Danalis, Thomas Herault, Pierre Lemarinier, and Jack Dongarra. DAGuE: a generic distributed DAG engine for High Performance Computing. Parallel Computing, 38(1–2):37– 51, January/February 2012. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167819111001347.

Bosilca:2012:DLA

[776] George Bosilca, Aurelien Bouteiller, Anthony Danalis, Thomas Herault, Piotr Luszczek, and Jack J. Dongarra. Dense linear algebra on distributed heterogeneous hardware with a symbolic DAG approach. LAPACK Working Note 264, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, January

2012. URL http://www.netlib.org/
lapack/lawnspdf/lawn264.pdf.

Bosilca:2012:UMA

[777] George Bosilca, Aurelien Bouteiller, Elisabeth Brunet, Franck Cappello, Jack Dongarra, Amina Guermouche, Thomas Herault, Yves Robert, Frederic Vivien, and Dounia Zaidouni. Unified model for assessing checkpointing protocols at extreme-scale. LA-PACK Working Note 269, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, June

2012. URL http://www.netlib.org/
lapack/lawnspdf/lawn269.pdf. UT-

CS-12-697.

Danalis:2012:BPH

[778] Anthony Danalis, Piotr Luszczek, Gabriel Marin, Jeffrey S. Vetter, and Jack Dongarra. Blackjack-Bench: portable hardware characterization. ACM SIGMETRICS Performance Evaluation Review, 40(2):74–79, September 2012. CODEN ???? ISSN 0163-5999 (print), 1557-9484 (electronic).

Dongarra:2012:ASC

[779] Jack Dongarra. Algorithmic and software challenges when moving towards exascale. In ????, editor, ATIP ’12: Proceedings of the ATIP/A*CRC Workshop on Accelerator Technologies for High-Performance Computing: Does Asia Lead the Way?, page ?? ACM Press, New York, NY 10036, USA, 2012. ISBN 1-4503-1644-1. LCCN ????

Dongarra:2012:HPC

[780] J. J. Dongarra and A. J. van der Steen. High-performance computing systems: Status and outlook. Acta Numerica, 21:379–474, 2012. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Dongarra:2012:LAL

[781] Jack Dongarra, James Demmel, Michael Heroux, and Jakub Kurzak. Linear algebra libraries for highperformance computing: Scientific computing with multicore and accelerators. Supercomputer ’2012 conference tutorial, November 10–16, 2012. ISBN 1-4673-0804-8. PDF with 76 lecture slides.

Dongarra:2012:RDC

[782] Jack Dongarra, Thomas Herault, and Yves Robert. Revisiting the double checkpointing algorithm. LAPACK Working Note 274, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, December

                                        2012. URL http://www.netlib.org/
                                        lapack/lawnspdf/lawn274.pdf. UT-

CS-13-705.

Du:2012:ABF

[783] Peng Du, Aurelien Bouteiller, George Bosilca, Thomas Herault, and Jack Dongarra. Algorithm-based fault tolerance for dense matrix factorizations. ACM SIGPLAN Notices, 47(8):225– 234, August 2012. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic). PPOPP ’12 conference proceedings.

Du:2012:COT

[784] Peng Du, Rick Weber, Piotr Luszczek, Stanimire Tomov, Gregory Peterson, and Jack Dongarra. From CUDA to OpenCL: Towards a performanceportable solution for multi-platform GPU programming. Parallel Computing, 38(8):391–407, August 2012.

CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

URL http://www.sciencedirect.
com/science/article/pii/S0167819111001335.

Du:2012:PGC

[785] Peng Du, Stanimire Tomov, and Jack Dongarra. Providing GPU capability to LU and QR within the ScaLA-PACK framework. LAPACK Working Note 272, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, Septem-

ber 2012. URL http://www.netlib.
org/lapack/lawnspdf/lawn272.pdf.

UT-CS-12-699.

Haidar:2012:ADS

[786] Azzam Haidar, Hatem Ltaief, Asim YarKhan, and Jack Dongarra. Analysis of dynamically scheduled tile algorithms for dense linear algebra on multicore architectures. Concurrency and Computation: Practice and Experience, 24(3):305–321, March 10, 2012. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Haidar:2012:THP

[787] Azzam Haidar, Hatem Ltaief, and Jack Dongarra. Toward a high performance tile divide and conquer algorithm for the dense symmetric eigenvalue problem. SIAM Journal on Scientific Computing, 34(6):C249–C274, ???? 2012. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Jia:2012:HFM

[788] Yulu Jia, Piotr Luszczek, and Jack Dongarra. Hybrid LU factorization on multi-GPU multi-core heterogeneous platforms. In ????, editor, ATIP ’12: Proceedings of the ATIP/A*CRC Workshop on Accelerator Technologies for High-Performance Computing: Does Asia Lead the Way?, page ?? ACM Press, New York, NY 10036, USA, 2012. ISBN 1-4503-1644-1. LCCN ????

Kurzak:2012:AGK

[789] Jakub Kurzak, Stanimire Tomov, and Jack Dongarra. Autotuning GEMM kernels for the Fermi GPU. IEEE Transactions on Parallel and Distributed Systems, 23(11):2045–2057, November 2012. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Kurzak:2012:FPP

[790] Jakub Kurzak, Piotr Luszczek, Mathieu Faverge, and Jack Dongarra. LU factorization with partial pivoting for a multi-CPU, multi-GPU shared memory system. LAPACK Working Note 266, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, April 2012. URL http:/

                                        /www.netlib.org/lapack/lawnspdf/
                                        lawn266.pdf.

Kurzak:2012:PRA

[791] Jakub Kurzak, Piotr Luszczek, Stanimire Tomov, and Jack Dongarra. Preliminary results of autotuning GEMM kernels for the NVIDIA Kepler Architecture GeForce GTX 680. LA-PACK Working Note 267, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, April

                                        2012. URL http://www.netlib.org/
                                        lapack/lawnspdf/lawn267.pdf.

Simon:2012:ISI

[792] Horst Simon, Jack Dongarra, and Hemant Shukla. Introduction to the Special Issue. The International Journal of High Performance Computing Applications, 26(4):335–336, November 2012. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL

http://hpc.sagepub.com/content/
26/4/335.full.pdf+html.

Vomel:2012:DCH

[793] Christof Vömel, Stanimire Tomov, and Jack Dongarra. Divide and conquer on hybrid GPU-accelerated multicore systems. SIAM Journal on Scientific Computing, 34(2):C70–C82, ???? 2012. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Anonymous:2013:CIF

[794] Anonymous. A conversation with ISC Fellow Jack Dongarra. Scientific Computing, April 10, 2013. CO-DEN SCHRCU. ISSN 1930-5753 (print), 1930-6156 (electronic). URL

http://www.scientificcomputing.
com/articles/2013/04/conversation-
isc-fellow-jack-dongarra.

Anzt:2013:BAR

[795] Hartwig Anzt, Stanimire Tomov, Jack Dongarra, and Vincent Heuveline. A block-asynchronous relaxation method for graphics processing units. Journal of Parallel and Distributed Computing, 73(12):1613– 1626, December 2013. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0743731513001147.

Aupy:2013:CSE

[796] Guillaume Aupy, Anne Benoit, Thomas Hérault, Yves Robert, Frédéric Vivien, and Dounia Zaidouni. On the combination of silent error detection and checkpointing. LAPACK Working Note 278, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, June

2013. URL http://www.netlib.org/
lapack/lawnspdf/lawn278.pdf. UT-

CS-13-710.

Aupy:2013:ISA

[797] Guillaume Aupy, Mathieu Faverge, Yves Robert, Jakub Kurzak, Piotr Luszczek, and Jack Dongarra. Implementing a systolic algorithm for QR factorization on multicore clusters with PaRSEC. LAPACK Working Note 277, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, May

2013. URL http://www.netlib.org/
lapack/lawnspdf/lawn277.pdf. UT-

CS-13-709.

Aupy:2013:OCP

[798] Guillaume Aupy, Anne Benoit, Thomas Herault, Yves Robert, and Jack Dongarra. Optimal checkpointing period: Time vs. energy. LAPACK Working Note 281, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, October 2013. URL http:/

/www.netlib.org/lapack/lawnspdf/

lawn281.pdf. UT-EECS-13-718.

Baboulin:2013:ALS

[799] Marc Baboulin, Jack Dongarra, Julien Herrmann, and Stanimire Tomov. Accelerating linear system solutions using randomization techniques. ACM Transactions on Mathematical Software, 39(2):8:1–8:13, February 2013.

CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Bland:2013:PFR

[800] Wesley Bland, Aurelien Bouteiller, Thomas Herault, George Bosilca, and Jack Dongarra. Post-failure recovery of MPI communication capability: Design and rationale. The International Journal of High Performance Computing Applications, 27 (3):244–254, August 2013. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

//hpc.sagepub.com/content/27/3/
244.full.pdf+html.

Bland:2013:SIP

[801] Wesley Bland, Peng Du, Aurelien Bouteiller, Thomas Herault, George Bosilca, and Jack J. Dongarra. Special issue papers: Extending the scope of the Checkpoint-on-Failure protocol for forward recovery in standard MPI. Concurrency and Computation: Practice and Experience, 25(17):2381– 2393, December 10, 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bosilca:2013:PEH

[802] George Bosilca, Aurelien Bouteiller, Anthony Danalis, Mathieu Faverge, Thomas Herault, and Jack J. Dongarra. PaRSEC: Exploiting heterogeneity to enhance scalability. Computing in Science and Engineering, 15 (6):36–45, November/December 2013. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Bouteiller:2013:CSC

[803] Aurelien Bouteiller, Thomas Herault, George Bosilca, and Jack J. Dongarra. Correlated set coordination in fault tolerant message logging protocols for many-core clusters. Concurrency and Computation: Practice and Experience, 25(4):572–585, ???? 2013. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Cao:2013:CHP

[804] Chongxiao Cao, Jack Dongarra, Peng Du, Mark Gates, Piotr Luszczek, and Stanimire Tomov. clMAGMA: High performance dense linear algebra with OpenCL. LAPACK Working Note 275, Department of Computer Science, University of Tennessee, Knoxville, TN, USA, March

2013. URL http://www.netlib.org/
lapack/lawnspdf/lawn275.pdf. UT-

CS-13-706.

Donfack:2013:AVP

[805] Simplice Donfack, Jack Dongarra, Mathieu Faverge, Mark Gates, Jakub Kurzak, Piotr Luszczek, and Ichitaro Yamazaki. On algorithmic variants of parallel Gaussian elimination: Comparison of implementations in terms of performance and numerical properties. LAPACK Working Note 280, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, July

2013. URL http://www.netlib.org/
lapack/lawnspdf/lawn280.pdf. UT-

CS-13-715.

Dongarra:2013:GEN

[806] Jack Dongarra and Bernard Tourancheau. Guest Editors’ note: Special issue on clusters, clouds, and data for scientific computing. Parallel Processing Letters, 23(2):1302001, 1, June 2013. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Dongarra:2013:HQF

[807] Jack Dongarra, Mathieu Faverge, Thomas Hérault, Mathias Jacquelin, Julien Langou, and Yves Robert. Hierarchical QR factorization algorithms for multi-core clusters. Parallel Computing, 39(4–5):212–232, April/May 2013. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167819113000100.

Dongarra:2013:IAS

[808] Jack Dongarra and Bernard Tourancheau. Introduction for August special issue CCDSC. The International Journal of High Performance Computing Applications, 27(3):231, August 2013. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL

http://hpc.sagepub.com/content/
27/3/231.full.pdf+html.

Du:2013:SER

[809] Peng Du, Piotr Luszczek, Stan Tomov, and Jack Dongarra. Soft error resilient QR factorization for hybrid system with GPGPU. Journal of Computational Science, 4(6): 457–464, November 2013. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S1877750313000161.

Faverge:2013:DHS

[810] Mathieu Faverge, Julien Herrmann, Julien Langou, Bradley Lowery, Yves Robert, and Jack Dongarra. Designing LU–QR hybrid solvers for performance and stability. LAPACK Working Note 282, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, October

                                        2013. URL http://www.netlib.org/
                                        lapack/lawnspdf/lawn282.pdf. UT-

EECS-13-719.

Gustavson:2013:LCF

[811] Fred G. Gustavson, Jerzy Waśniewski, Jack J. Dongarra, José R. Herrero, and Julien Langou. Level-3 Cholesky factorization routines improve performance of many Cholesky algorithms. ACM Transactions on Mathematical Software, 39(2):9:1–9:10, February 2013. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Haidar:2013:IPS

[812] Azzam Haidar, Piotr Luszczek, Jakub Kurzak, and Jack Dongarra. An improved parallel singular value algorithm and its implementation for multicore hardware. LAPACK Working Note 283, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA, October 2013. URL http:/

                                        /www.netlib.org/lapack/lawnspdf/

lawn283.pdf. UT-EECS-13-720.

Jia:2013:TER

[813] Yulu Jia, Piotr Luszczek, and Jack Dongarra. Transient error resilient Hessenberg reduction on GPU-based hybrid architectures. LAPACK Working Note 279, Department of Computer Science, University of Tennessee, Knoxville, Knoxville, TN 37996, USA,

June 2013. URL http://www.netlib.
org/lapack/lawnspdf/lawn279.pdf.

UT-CS-13-712.

Kurzak:2013:FPP

[814] Jakub Kurzak, Piotr Luszczek, Mathieu Faverge, and Jack Dongarra. LU factorization with partial pivoting for a multicore system with accelerators. IEEE Transactions on Parallel and Distributed Systems, 24(8):1613–1621, August 2013. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Li:2013:EWG

[815] Yinan Li, Asim YarKhan, Jack Dongarra, Keith Seymour, and Aurèlie Hurault. Enabling workflows in GridSolve: request sequencing and service trading. The Journal of Supercomputing, 64(3):1133–1152, June 2013. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic). URL http:

//link.springer.com/article/10.
1007/s11227-010-0549-1.

Ltaief:2013:HPB

[816] Hatem Ltaief, Piotr Luszczek, and Jack Dongarra. High-performance bidiagonal reduction using tile algorithms on homogeneous multicore architectures. ACM Transactions on Mathematical Software, 39(3):16:1–16:22, April 2013. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Ma:2013:KAT

[817] Teng Ma, George Bosilca, Aurelien Bouteiller, and Jack J. Dongarra. Kernel-assisted and topologyaware MPI collective communications on multicore/many-core platforms. Journal of Parallel and Distributed Computing, 73(7):1000– 1010, July 2013. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/S0743731513000166.

Baboulin:2014:EDR

[818] Marc Baboulin, Dulceneia Becker, George Bosilca, Anthony Danalis, and Jack Dongarra. An efficient distributed randomized algorithm for solving large dense symmetric indefinite linear systems. Parallel Computing, 40(7):213–223, July 2014. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167819113001488.

Ballard:2014:CAS

[819] Grey Ballard, Dulceneia Becker, James Demmel, Jack Dongarra, Alex Druinsky, Inon Peled, Oded Schwartz, Sivan Toledo, and Ichitaro Yamazaki. Communication-avoiding symmetricindefinite factorization. SIAM Journal on Matrix Analysis and Applications, 35(4):1364–1406, ???? 2014. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic).

Bosilca:2014:UMA

[820] George Bosilca, Aurélien Bouteiller, Elisabeth Brunet, Franck Cappello, Jack Dongarra, Amina Guermouche, Thomas Herault, Yves Robert, Frédéric Vivien, and Dounia Zaidouni. Unified model for assessing checkpointing protocols at extreme-scale. Concurrency and Computation: Practice and Experience, 26(17):2772–2791, December 10, 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Danalis:2014:BPH

[821] Anthony Danalis, Piotr Luszczek, Gabriel Marin, Jeffrey S. Vetter, and Jack Dongarra. BlackjackBench: Portable hardware characterization with automated results’ analysis. The Computer Journal, 57(7):1002– 1016, July 2014. CODEN CMPJA6. ISSN 0010-4620 (print), 1460-2067 (electronic). URL http://comjnl.

oxfordjournals.org/content/57/7/
1002.full.pdf+html.

Dongarra:2014:ANA

[822] Jack Dongarra, Mathieu Faverge, Hatem Ltaief, and Piotr Luszczek. Achieving numerical accuracy and high performance using recursive tile LU factorization with partial pivoting. Concurrency and Computation: Practice and Experience, 26(7):1408–1431, May 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dongarra:2014:AND

[823] Jack Dongarra, Mark Gates, Azzam Haidar, Jakub Kurzak, Piotr Luszczek, Stanimire Tomov, and Ichitaro Yamazaki. Accelerating numerical dense linear algebra calculations with GPUs. In Numerical Computations with GPUs, pages 3–28. Springer International Publishing, Cham, Switzerland, 2014.

Dongarra:2014:MDO

[824] Jack Dongarra, Azzam Haidar, Jakub Kurzak, Piotr Luszczek, Stanimire Tomov, and Asim YarKhan. Modeldriven one-sided factorizations on multicore accelerated systems. Supercomputing Frontiers and Innovations, 1(1): 85–115, ???? 2014. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.org/
                                        superfri/article/view/15.

Dongarra:2014:PHP

[825] Jack Dongarra, Mark Gates, Azzam Haidar, Yulu Jia, Khairul Kabir, Piotr Luszczek, and Stanimire Tomov. Portable HPC programming on Intel many-integrated-core hardware with MAGMA port to Xeon Phi. In Parallel processing and applied mathematics. Part I, volume 8384 of Lecture Notes in Computer Science, pages 571–581. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2014.

Haidar:2014:NHC

[826] Azzam Haidar, Stanimire Tomov, Jack Dongarra, Raffaele Solcà, and Thomas Schulthess. A novel hybrid CPU-GPU generalized eigensolver for electronic structure calculations based on finegrained memory aware tasks. The International Journal of High Performance Computing Applications, 28(2): 196–209, May 2014. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http://hpc.

                                        sagepub.com/content/28/2/196.

Luszczek:2014:LBD

[827] Piotr Luszczek, Jakub Kurzak, and Jack Dongarra. Looking back at dense linear algebra software. Journal of Parallel and Distributed Computing, 74 (7):2548–2560, July 2014. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0743731513002165.

Yamazaki:2014:DIL

[828] Ichitaro Yamazaki, Jakub Kurzak, Piotr Luszczek, and Jack Dongarra. Design and implementation of a large scale tree-based QR decomposition using a 3D virtual systolic array and a lightweight runtime. Parallel Processing Letters, 24(4):1442004, 23, December 2014. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Yamazaki:2014:TDS

[829] Ichitaro Yamazaki, Tingxing Dong, Raffaele Solcà, Stanimire Tomov, Jack Dongarra, and Thomas Schulthess. Tridiagonalization of a dense symmetric matrix on multiple GPUs and its application to symmetric eigenvalue problems. Concurrency and Computation: Practice and Experience, 26(16): 2652–2666, November 2014. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anzt:2015:AGB

[830] Hartwig Anzt, Stanimire Tomov, Piotr Luszczek, William Sawyer, and Jack Dongarra. Acceleration of GPU-based Krylov solvers via data transfer reduction. The International Journal of High Performance Computing Applications, 29(3):366–383, August 2015. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL

                                        http://journals.sagepub.com/doi/
                                        full/10.1177/1094342015580139.

Anzt:2015:EAM

[831] Hartwig Anzt, Blake Haugen, Jakub Kurzak, Piotr Luszczek, and Jack Dongarra. Experiences in autotuning matrix multiplication for energy minimization on GPUs. Concurrency and Computation: Practice and Experience, 27(17):5096–5113, December 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bouteiller:2015:ABF

[832] Aurelien Bouteiller, Thomas Herault, George Bosilca, Peng Du, and Jack Dongarra. Algorithm-based fault tolerance for dense matrix factorizations, multiple failures and accuracy. ACM Transactions on Parallel Computing (TOPC), 1(2):10:1–10:??, January 2015. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

Donfack:2015:SRD

[833] Simplice Donfack, Jack Dongarra, Mathieu Faverge, Mark Gates, Jakub Kurzak, Piotr Luszczek, and Ichitaro Yamazaki. A survey of recent developments in parallel implementations of Gaussian elimination. Concurrency and Computation: Practice and Experience, 27(5):1292–1309, April 10, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Dong:2015:FBG

[834] Tingzing Tim Dong, Stanimire Z. Tomov, Piotr R. Luszczek, and Jack J. Dongarra. A framework for batched and GPU-resident factorization algorithms applied to block Householder transformations. Technical report, Oak Ridge National Laboratory, Knoxville, TN, USA, January 1,

2015. URL https://www.osti.gov/
biblio/1261481-framework-batched-
gpu-resident-factorization-algorithms-
applied-block-householder-transformations.

Dongarra:2015:GEN

[835] Jack Dongarra and Bernard Tourancheau. Guest Editors’ note: Special issue on clusters, clouds and data for scientific computing. Parallel Processing Letters, 25(3):1502002:1– 1502002:2, September 2015. CODEN PPLTEE. ISSN 0129-6264 (print), 1793-642X (electronic).

Dongarra:2015:HPI

[836] Jack Dongarra, Mark Gates, Azzam Haidar, Yulu Jia, Khairul Kabir, Piotr Luszczek, and Stanimire Tomov. HPC programming on Intel Many-Integrated-Core hardware with MAGMA port to Xeon Phi. Scientific Programming, 2015(??):502593:1–502593:11, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

tronic). URL https://www.hindawi.
com/journals/sp/2015/502593/.

Dongarra:2015:PPM

[837] Jack Dongarra, M. Abalenkovs, A. Abdelfattah, M. Gates, A. Haidar, J. Kurzak, P. Luszczek, S. Tomov, I. Yamazaki, and A. YarKhan. Parallel programming models for dense linear algebra on heterogeneous systems. Supercomputing Frontiers and Innovations, 2(4):67–86, ???? 2015. CO-DEN ???? ISSN 2409-6008 (print), 2313-8734 (electronic). URL http:

                                        //superfri.org/superfri/article/
                                        view/90.

Faverge:2015:MLQ

[838] Mathieu Faverge, Julien Herrmann, Julien Langou, Bradley Lowery, Yves Robert, and Jack Dongarra. Mixing LU and QR factorization algorithms to design high-performance dense linear algebra solvers. Journal of Parallel and Distributed Computing, 85 (??):32–46, November 2015. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0743731515001057.

Haidar:2015:BMC

[839] Azzam Haidar, Tingxing Dong, Piotr Luszczek, Stanimire Tomov, and Jack Dongarra. Batched matrix computations on hardware accelerators based on GPUs. The International Journal of High Performance Computing Applications, 29 (2):193–208, May 2015. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

                                        //journals.sagepub.com/doi/full/
                                        10.1177/1094342014567546. See re-

traction notice [934].

Haidar:2015:FBG

[840] Azzam Haidar, Tingxing Tim Dong, Stanimire Tomov, Piotr Luszczek, and Jack Dongarra. A framework for batched and GPU-resident factorization algorithms applied to block Householder transformations. In High Performance Computing. ISC High Performance 2015, Lecture Notes in Computer Science, pages 31– 47. Springer International Publishing, Cham, Switzerland, 2015.

Haidar:2015:TBL

[841] Azzam Haidar, Tingxing Dong, Piotr Luszczek, Stanimire Tomov, and Jack Dongarra. Towards batched linear solvers on accelerated hardware platforms. ACM SIGPLAN Notices, 50(8): 261–262, August 2015. CODEN SIN-ODQ. ISSN 0362-1340 (print), 1523-2867 (print), 1558-1160 (electronic).

Reed:2015:ECB

[842] Daniel A. Reed and Jack Dongarra. Exascale computing and big data. Communications of the ACM, 58(7):56– 68, July 2015. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL http://cacm.

acm.org/magazines/2015/7/188732/
fulltext.

Song:2015:SAS

[843] Fengguang Song and Jack Dongarra. A scalable approach to solving dense linear algebra problems on hybrid CPU– GPU systems. Concurrency and Computation: Practice and Experience, 27 (14):3702–3723, September 25, 2015. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Strohmaier:2015:TLP

[844] Erich Strohmaier, Hans W. Meuer, Jack Dongarra, and Horst D. Simon. The TOP500 list and progress in highperformance computing. Computer, 48 (11):42–49, November 2015. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http:

//www.computer.org/csdl/mags/co/
2015/11/mco2015110042-abs.html.

Voevodin:2015:AOE

[845] Vladimir V. Voevodin, Alexander S. Antonov, and Jack Dongarra. Algo-Wiki: an open encyclopedia of parallel algorithmic features. Supercomputing Frontiers and Innovations, 2(1):4– 18, ???? 2015. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

                                        tronic). URL http://superfri.org/
                                        superfri/article/view/69.

Yamazaki:2015:CLR

[846] Ichitaro Yamazaki, Stanimire Tomov, and Jack Dongarra. Computing lowrank approximation of a dense matrix on multicore CPUs with a GPU and its application to solving a hierarchically semiseparable linear system of equations. Scientific Programming, 2015(??):246019:1–246019:17, ???? 2015. CODEN SCIPEV. ISSN 1058-9244 (print), 1875-919X (elec-

                                        tronic). URL https://www.hindawi.
                                        com/journals/sp/2015/246019/.

Yamazaki:2015:MPC

[847] Ichitaro Yamazaki, Stanimire Tomov, and Jack Dongarra. Mixed-precision Cholesky QR factorization and its case studies on multicore CPU with multiple GPUs. SIAM Journal on Scientific Computing, 37(3):C307–C330, ???? 2015. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Abdelfattah:2016:LAS

[848] A. Abdelfattah, H. Anzt, J. Dongarra, M. Gates, A. Haidar, J. Kurzak, P. Luszczek, S. Tomov, I. Yamazaki, and A. YarKhan. Linear algebra software for large-scale accelerated multicore computing. Acta Numerica, 25:1– 160, 2016. CODEN ANUMFU. ISSN 0962-4929 (print), 1474-0508 (electronic).

Abdelfattah:2016:POS

[849] Ahmad Abdelfattah, Hatem Ltaief, David Keyes, and Jack Dongarra. Performance optimization of sparse matrix-vector multiplication for multicomponent PDE-based applications using GPUs. Concurrency and Computation: Practice and Experience, 28(12): 3447–3465, August 25, 2016. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anzt:2016:DOI

[850] Hartwig Anzt, Edmond Chow, Daniel B. Szyld, and Jack Dongarra. Domain overlap for iterative sparse triangular solves on GPUs. In Software for exascale computing—SPPEXA 2013–2015, volume 113 of Lect. Notes Comput. Sci. Eng., pages 527–545. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016.

Anzt:2016:UIF

[851] Hartwig Anzt, Edmond Chow, Jens Saak, and Jack Dongarra. Updating incomplete factorization preconditioners for model order reduction. Numerical Algorithms, 73(3):611–630, November 2016. CODEN NUALEG. ISSN 1017-1398 (print), 1572-9265 (elec-

tronic). URL http://link.springer.
com/article/10.1007/s11075-016-
0110-2.

Baboulin:2016:DSI

[852] Marc Baboulin, Jack Dongarra, Adrien Rémy, Stanimire Tomov, and Ichitaro Yamazaki. Dense symmetric indefinite factorization on GPU accelerated architectures. In Parallel processing and applied mathematics. Part I, volume 9573 of Lecture Notes in Computer Science, pages 86–95. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016.

Dongarra:2016:HPC

[853] Jack Dongarra, Michael A. Heroux, and Piotr Luszczek. Highperformance conjugate-gradient benchmark: a new metric for ranking highperformance computing systems. The International Journal of High Performance Computing Applications, 30 (1):3–10, February 2016. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

//journals.sagepub.com/doi/full/
10.1177/1094342015593158.

Dongarra:2016:NMR

[854] Jack Dongarra, Michael A. Heroux, and Piotr Luszczek. A new metric for ranking high performance computing systems. National Science Review, ??(??):??, January 6, 2016. CO-DEN ???? ISSN 2053-714X (print), 2095-5138 (electronic). URL http:/

/nsr.oxfordjournals.org/content/
early/2016/01/06/nsr.nwv084.

Herrmann:2016:ACR

[855] Julien Herrmann, George Bosilca, Thomas Hérault, Loris Marchal, Yves Robert, and Jack Dongarra. Assessing the cost of redistribution followed by a computational kernel: Complexity and performance results. Parallel Computing, 52(??): 22–41, February 2016. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167819115001192.

Jagode:2016:ANC

[856] Heike Jagode, Anthony Danalis, George Bosilca, and Jack Dongarra. Accelerating NWChem coupled cluster through dataflow-based execution. In Parallel processing and applied mathematics. Part I, volume 9573 of Lecture Notes in Computer Science, pages 366–376. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016.

Kurzak:2016:ITB

[857] Jakub Kurzak, Hartwig Anzt, Mark Gates, and Jack Dongarra. Implementation and tuning of batched Cholesky factorization and solve for NVIDIA GPUs. IEEE Transactions on Parallel and Distributed Systems, 27(7): 2036–2048, July 2016. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL http://csdl.

computer.org/csdl/trans/td/2016/
07/07275187-abs.html.

Masliah:2016:HPM

[858] Ian Masliah, Ahmad Abdelfattah, A. Haidar, S. Tomov, Marc Baboulin, J. Falcou, and J. Dongarra. Highperformance matrix–matrix multiplications of very small matrices. In Euro-Par 2016: parallel processing, volume 9833 of Lecture Notes in Computer Science, pages 659–671. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2016.

Yamazaki:2016:SPV

[859] Ichitaro Yamazaki, Stanimire Tomov, and Jack Dongarra. Stability and performance of various singular value QR implementations on multicore CPU with a GPU. ACM Transactions on Mathematical Software, 43(2):10:1– 10:18, September 2016. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://dl.

                                        acm.org/citation.cfm?id=2898347.

Abdelfattah:2017:FCF

[860] Ahmad Abdelfattah, Azzam Haidar, Stanimire Tomov, and Jack Dongarra. Fast Cholesky factorization on GPUs for batch and native modes in MAGMA. Journal of Computational Science, 20:85–93, May 2017. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316305154.

Anzt:2017:PEE

[861] Hartwig Anzt, Stanimire Tomov, and Jack Dongarra. On the performance and energy efficiency of sparse linear algebra on GPUs. The International Journal of High Performance Computing Applications, 31(5): 375–390, September 2017. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

                                        //journals.sagepub.com/doi/full/
                                        10.1177/1094342016672081.

Anzt:2017:PKS

[862] Hartwig Anzt, Mark Gates, Jack Dongarra, Moritz Kreutzer, Gerhard Wellein, and Martin Köhler. Preconditioned Krylov solvers on GPUs. Parallel Computing, 68(??):

32–44, October 2017. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167819117300777.

Baboulin:2017:SDS

[863] Marc Baboulin, Jack Dongarra, Adrien Rémy, Stanimire Tomov, and Ichitaro Yamazaki. Solving dense symmetric indefinite systems using GPUs. Concurrency and Computation: Practice and Experience, 29(9):??, May 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Bell:2017:LBY

[864] Gordon Bell, David H. Bailey, Jack Dongarra, Alan H. Karp, and Kevin Walsh. A look back on 30 years of the Gordon Bell Prize. The International Journal of High Performance Computing Applications, 31(6): 469–484, November 2017. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

//journals.sagepub.com/doi/full/
10.1177/1094342017738610.

Dongarra:2017:ECR

[865] Jack Dongarra, Stanimire Tomov, Piotr Luszczek, Jakub Kurzak, Mark Gates, Ichitaro Yamazaki, Hartwig Anzt, Azzam Haidar, and Ahmad Abdelfattah. With extreme computing, the rules have changed. Computing in Science and Engineering, 19 (3):52–62, May/June 2017. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic). URL https:

//www.computer.org/csdl/mags/cs/
2017/03/mcs2017030052-abs.html.

Dongarra:2017:GEN

[866] Jack Dongarra and Bernard Tourancheau. Guest Editor’s note: Special issue on clusters, clouds and data for scientific computing. The International Journal of High Performance Computing Applications, 31 (1):3, January 2017. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

                                        //journals.sagepub.com/doi/full/
                                        10.1177/1094342015594514.

Haidar:2017:IHP

[867] Azzam Haidar, Panruo Wu, Stanimire Tomov, and Jack Dongarra. Investigating half precision arithmetic to accelerate dense linear system solvers. In Proceedings of the 8th Workshop on Latest Advances in Scalable Algorithms for Large-Scale Systems — ScalA’17. ACM Press, New York, NY 10036, USA, 2017. ISBN 1-4503-5125-5 (hardcover). LCCN ????

Kovalchuk:2017:DTC

[868] Sergey V. Kovalchuk, Tesfamariam M. Abuhay, Ilkay Altintas, Michael L. Norman, Michael H. Lees, Valeria V. Krzhizhanovskaya, Jack Dongarra, and Peter M. A. Sloot. Data through the Computational Lens. Journal of Computational Science, 20:81–84, May 2017. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750317305008.

Kurzak:2017:DIP

[869] Jakub Kurzak, Piotr Luszczek, Ichitaro Yamazaki, Yves Robert, and Jack Dongarra. Design and implementation of the PULSAR programming system for large scale computing. Supercomputing Frontiers and Innovations, 4(1):4– 26, ???? 2017. CODEN ???? ISSN 2409-6008 (print), 2313-8734 (elec-

tronic). URL http://superfri.org/
superfri/article/view/121.

Yamazaki:2017:NGR

[870] Ichitaro Yamazaki, Stanimire Tomov, and Jack Dongarra. Non-GPU-resident symmetric indefinite factorization. Concurrency and Computation: Practice and Experience, 29 (5):??, March 10, 2017. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

YarKhan:2017:PPN

[871] Asim YarKhan, Jakub Kurzak, Piotr Luszczek, and Jack Dongarra. Porting the PLASMA numerical library to the OpenMP standard. International Journal of Parallel Programming, 45(3):612–633, June 2017. CO-DEN IJPPE5. ISSN 0885-7458 (print), 1573-7640 (electronic).

Abdelfattah:2018:ADT

[872] Ahmad Abdelfattah, Azzam Haidar, Stanimire Tomov, and Jack Dongarra. Analysis and design techniques towards high-performance and energy-efficient dense linear solvers on GPUs. IEEE Transactions on Parallel and Distributed Systems, 29(12):2700– 2712, December 2018. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https://www.

computer.org/csdl/trans/td/2018/
12/08370686-abs.html.

Abdelfattah:2018:BOS

[873] Ahmad Abdelfattah, Azzam Haidar, Stanimire Tomov, and Jack Dongarra. Batched one-sided factorizations of tiny matrices using GPUs: Challenges and countermeasures. Journal of Computational Science, 26:226–236, May 2018. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750317311456.

Anzt:2018:ISA

[874] Hartwig Anzt, Thomas K. Huckle, Jürgen Bräckle, and Jack Dongarra. Incomplete sparse approximate inverses for parallel preconditioning. Parallel Computing, 71 (??):1–22, January 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S016781911730176X.

Anzt:2018:OPE

[875] Hartwig Anzt, Moritz Kreutzer, Eduardo Ponce, Gregory D. Peterson, Gerhard Wellein, and Jack Dongarra. Optimization and performance evaluation of the IDR iterative Krylov solver on GPUs. The International Journal of High Performance Computing Applications, 32(2):220–230, March 2018. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL

                                        http://journals.sagepub.com/doi/
                                        full/10.1177/1094342016646844.

Anzt:2018:PNP

[876] Hartwig Anzt, Edmond Chow, and Jack Dongarra. ParILUT — a new parallel threshold ILU factorization.

SIAM Journal on Scientific Computing, 40(4):C503–C519, ???? 2018. CO-DEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Asch:2018:BDE

[877] M. Asch, T. Moore, R. Badia, M. Beck, P. Beckman, T. Bidot, F. Bodin, F. Cappello, A. Choudhary, B. de Supinski, E. Deelman, J. Dongarra, A. Dubey, G. Fox, H. Fu, S. Girona, W. Gropp, M. Heroux, Y. Ishikawa, K. Keahey, D. Keyes, W. Kramer, J-F Lavignon, Y. Lu, S. Matsuoka, B. Mohr, D. Reed, S. Requena, J. Saltz, T. Schulthess, R. Stevens, M. Swany, A. Szalay, W. Tang, G. Varoquaux, J.-P. Vilotte, R. Wisniewski, Z. Xu, and I. Zacharov. Big data and extreme-scale computing. The International Journal of High Performance Computing Applications, 32(4):435–479, July 2018. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

//journals.sagepub.com/doi/full/
10.1177/1094342018778123.

Bosilca:2018:FDH

[878] George Bosilca, Aurelien Bouteiller, Amina Guermouche, Thomas Herault, Yves Robert, Pierre Sens, and Jack Dongarra. A failure detector for HPC platforms. The International Journal of High Performance Computing Applications, 32(1):139–158, January 2018. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL

http://journals.sagepub.com/doi/
full/10.1177/1094342017711505.

Chow:2018:UJI

[879] Edmond Chow, Hartwig Anzt, Jennifer Scott, and Jack Dongarra. Using Jacobi iterations and blocking for solving sparse triangular systems in incomplete factorization preconditioning. Journal of Parallel and Distributed Computing, 119(??): 219–230, September 2018. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0743731518303034.

Dong:2018:ASB

[880] Tingxing Dong, Azzam Haidar, Stanimire Tomov, and Jack Dongarra. Accelerating the SVD bi-diagonalization of a batch of small matrices using GPUs. Journal of Computational Science, 26:237–245, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S187775031731150X.

Dongarra:2018:GEN

[881] Jack Dongarra and Bernard Tourancheau. Guest editors’ note. The International Journal of High Performance Computing Applications, 32 (1):3, January 2018. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

                                        //journals.sagepub.com/doi/full/
                                        10.1177/1094342017745244.

Dongarra:2018:HPC

[882] Jack Dongarra, Michael A. Heroux, and Piotr Luszczek. The high-performance conjugate gradients benchmark. SIAM News, 51(1):12, 2018. ISSN 1557-9573.

Dongarra:2018:SVD

[883] Jack Dongarra, Mark Gates, Azzam Haidar, Jakub Kurzak, Piotr Luszczek, Stanimire Tomov, and Ichitaro Yamazaki. The singular value decomposition: anatomy of optimizing an algorithm for extreme scale. SIAM Review, 60(4):808–865, 2018. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Gates:2018:AST

[884] Mark Gates, Stanimire Tomov, and Jack Dongarra. Accelerating the SVD two stage bidiagonal reduction and divide and conquer using GPUs. Parallel Computing, 74:3–18, 2018. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167819117301758.

Haidar:2018:DFE

[885] Azzam Haidar, Ahmad Abdelfattah, Mawussi Zounon, Panruo Wu, Srikara Pranesh, Stanimire Tomov, and Jack Dongarra. The design of fast and energy-efficient linear solvers: on the potential of half-precision arithmetic and iterative refinement techniques. In Computational science—ICCS 2018. Part I, volume 10860 of Lecture Notes in Computer Science, pages 586–600. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2018. ISBN 3-319-93698-0; 3-319-93697-2.

Haidar:2018:GAH

[886] Azzam Haidar, Ahmad Abdelfattah, Mawussi Zounon, Stanimire Tomov, and Jack Dongarra. A guide for achieving high performance with very small matrices on GPU: A case study of batched LU and Cholesky factorizations. IEEE Transactions on Parallel and Distributed Systems, 29(5):

973–984, May 2018. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https://www.

                                        computer.org/csdl/trans/td/2018/
                                        05/08214236-abs.html.

Haidar:2018:HGT

[887] Azzam Haidar, Stanimire Tomov, Jack Dongarra, and Nicholas J. Higham. Harnessing GPU tensor cores for fast FP16 arithmetic to speed up mixedprecision iterative refinement solvers. In IEEE, editor, SC ’18 Proceedings of the International Conference for High Performance Computing, Networking, Storage, and Analysis, Dallas, Texas, November 11–16, 2018, pages 47:1–47:11. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2018. ISBN 1-5386-8384-9. LCCN

                                        ???? URL http://www.netlib.org/
                                        utk/people/JackDongarra/PAPERS/
                                        haidar_fp16_sc18.pdf;   https://
                                        dl.acm.org/citation.cfm?id=3291656.
                                        3291719.

Jagode:2018:ANC

[888] Heike Jagode, Anthony Danalis, and Jack Dongarra. Accelerating NWChem Coupled Cluster through dataflow-based execution. The International Journal of High Performance Computing Applications, 32 (4):540–551, July 2018. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL http:

                                        //journals.sagepub.com/doi/full/
                                        10.1177/1094342016672543.

Jagode:2018:EDP

[889] Heike Jagode, Anthony Danalis, Reazul Hoque, Mathieu Faverge, and Jack Dongarra. Evaluation of dataflow programming models for electronic structure theory. Concurrency and Computation: Practice and Experience, 30(17):e4490:1–e4490:??, September 10, 2018. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Kovalchuk:2018:ACS

[890] Sergey V. Kovalchuk, Valeria V. Krzhizhanovskaya, Petros Koumoutsakos, Eleni Chatzi, Michael H. Lees, Jack Dongarra, and Peter M. A. Sloot. The art of computational science: Bridging gaps — forming alloys. Journal of Computational Science, 26:190–192, May 2018. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750318304095.

Yamazaki:2018:SIL

[891] Ichitaro Yamazaki, Jakub Kurzak, Panruo Wu, Mawussi Zounon, and Jack Dongarra. Symmetric indefinite linear solver using OpenMP task on multicore architectures. IEEE Transactions on Parallel and Distributed Systems, 29(8):1879–1892, August 2018. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic). URL https://www.

computer.org/csdl/trans/td/2018/
08/08301559-abs.html.

Anzt:2019:APB

[892] Hartwig Anzt, Jack Dongarra, Goran Flegar, Nicholas J. Higham, and Enrique S. Quintana-Ortí. Adaptive precision in block-Jacobi preconditioning for iterative sparse linear system solvers. Concurrency and Computation: Practice and Experience, 31 (6):e4460:1–e4460:??, March 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Anzt:2019:FGB

[893] Hartwig Anzt, Jack Dongarra, and Enrique S. Quintana-Ortí. Fine-grained bit-flip protection for relaxation methods. Journal of Computational Science, 36:??, September 2019. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750316303891.

Anzt:2019:VSB

[894] Hartwig Anzt, Jack Dongarra, Goran Flegar, and Enrique S. Quintana-Ortí. Variable-size batched Gauss– Jordan elimination for block-Jacobi preconditioning on graphics processors. Parallel Computing, 81:131– 146, January 2019. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

                                        //www.sciencedirect.com/science/
                                        article/pii/S0167819117302107.

Dongarra:2019:GEN

[895] Jack Dongarra and Bernard Tourancheau. Guest editors’ note: Special issue on clusters, clouds, and data for scientific computing. The International Journal of High Performance Computing Applications, 33(6):1067– 1068, November 1, 2019. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https:

                                        //journals.sagepub.com/doi/full/
                                        10.1177/1094342019876834.

Dongarra:2019:HAM

[896] Jack Dongarra, Piotr Luszczek, and Yaohung (Mike) Tsai. HPL-AI mixed-precision benchmark. Web site, December 2019. URL https:

//bit.ly/hpl-ai/;   https://icl.
bitbucket.io/hpl-ai/;   https:/
/www.math.utah.edu/pub/tex/bib/
fparith.bib;   https://www.nextplatform.
com/2019/12/03/tweaked-math-libraries-
exploit-ai-hardware-for-traditional-
hpc/.

Dongarra:2019:PPL

[897] Jack Dongarra, Mark Gates, Azzam Haidar, Jakub Kurzak, Piotr Luszczek, Panruo Wu, Ichitaro Yamazaki, Asim Yarkhan, Maksims Abalenkovs, Negin Bagherpour, Sven Hammarling, Jakub Sístek, David Stevens, Mawussi Zounon, and Samuel D. Relton. PLASMA: Parallel linear algebra software for multicore using OpenMP. ACM Transactions on Mathematical Software, 45(2):16:1–16:35, April 2019. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL

https://dl.acm.org/citation.cfm?
id=3264491.

Dongarra:2019:RE

[898] J. Dongarra, S. Gottlieb, and W. T. C. Kramer. Race to exascale. Computing in Science and Engineering, 21(1): 4–5, January/February 2019. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366x (electronic).

Fevre:2019:CPR

[899] Valentin Le Fèvre, Thomas Herault, Yves Robert, Aurelien Bouteiller, Atsushi Hori, George Bosilca, and Jack Dongarra. Comparing the performance of rigid, moldable and gridshaped applications on failure-prone HPC platforms. Parallel Computing, 85(??):1–12, July 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167819118302230.

Haidar:2019:IPC

[900] Azzam Haidar, Heike Jagode, Phil Vaccaro, Asim YarKhan, Stanimire Tomov, and Jack Dongarra. Investigating power capping toward energy-efficient scientific applications. Concurrency and Computation: Practice and Experience, 31(6):e4485:1–e4485:??, March 25, 2019. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Jagode:2019:PSD

[901] Heike Jagode, Anthony Danalis, Hartwig Anzt, and Jack Dongarra. PAPI software-defined events for indepth performance analysis. The International Journal of High Performance Computing Applications, 33(6):1113– 1127, November 1, 2019. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https:

//journals.sagepub.com/doi/full/
10.1177/1094342019846287.

Kovalchuk:2019:SID

[902] Sergey V. Kovalchuk, Valeria V. Krzhizhanovskaya, Yong Shi, Haohuan Fua, Michael H. Lees, Jack Dongarra, and Peter M. A. Sloot. Science at the intersection of data, modelling, and computation. Journal of Computational Science, 34:117–119, May 2019.

CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750319305253.

Masliah:2019:AOT

[903] I. Masliah, A. Abdelfattah, A. Haidar, S. Tomov, M. Baboulin, J. Falcou, and J. Dongarra. Algorithms and optimization techniques for high-performance matrix–matrix multiplications of very small matrices. Parallel Computing, 81:1–21, January 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167819118301091.

Yamazaki:2019:DML

[904] Ichitaro Yamazaki, Akihiro Ida, Rio Yokota, and Jack Dongarra. Distributed-memory lattice H-matrix factorization. The International Journal of High Performance Computing Applications, 33(5):1046–1063, September 1, 2019. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https:/

/journals.sagepub.com/doi/full/
10.1177/1094342019861139.

Yamazaki:2019:PAO

[905] Ichitaro Yamazaki, Edmond Chow, Aurelien Bouteiller, and Jack Dongarra. Performance of asynchronous optimized Schwarz with one-sided communication. Parallel Computing, 86: 66–81, 2019. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic).

Zaitsev:2019:SLD

[906] D. Zaitsev, S. Tomov, and J. Dongarra.

Solving linear Diophantine systems on parallel architectures. IEEE Transactions on Parallel and Distributed Systems, 30(5):1158–1169, May 2019. CODEN ITDSEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Abdelfattah:2020:IBF

[907] Ahmad Abdelfattah, Stan Tomov, and Jack Dongarra. Investigating the benefit of FP16-enabled mixedprecision solvers for symmetric positive definite matrices using GPUs. In Krzhizhanovskaya et al. [1141], pages 237–250. ISBN 3-030-50416-6, 3-030-50417-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). URL

https://link.springer.com/book/
10.1007/978-3-030-50417-5.

Abdelfattah:2020:MMB

[908] Ahmad Abdelfattah, Stanimire Tomov, and Jack Dongarra. Matrix multiplication on batches of small matrices in half and half-complex precisions. Journal of Parallel and Distributed Computing, 145(??):188– 201, November 2020. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.sciencedirect.com/science/
article/pii/S0743731520303300.

Anonymous:2020:JDS

[909] Anonymous. Jack Dongarra selected to receive the 2020 IEEE Computer Society’s Computer Pioneer Award. University of Tennessee Knoxville Web site., February

4, 2020. URL https://www.eecs.
utk.edu/jack-dongarra-selected-
to-receive-the-2020-ieee-computer-

societys-computer-pioneer-award-
2/.

Anzt:2020:LBS

[910] Hartwig Anzt, Terry Cojean, Chen Yen-Chen, Jack Dongarra, Goran Flegar, Pratik Nayak, Stanimire Tomov, Yuhsiang M. Tsai, and Weichung Wang. Load-balancing sparse matrix vector product kernels on GPUs. ACM Transactions on Parallel Computing (TOPC), 7(1):2:1–2:26, April 2020. CODEN ???? ISSN 2329-4949 (print), 2329-4957 (electronic).

URL https://dl.acm.org/doi/abs/
10.1145/3380930.

Ayala:2020:HHE

[911] Alan Ayala, Stanimire Tomov, Azzam Haidar, and Jack Dongarra. heFFTe: highly efficient FFT for exascale. In Computational science— ICCS 2020. Part I, volume 12137 of Lecture Notes in Comput. Sci., pages 262–275. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2020. ISBN 3-030-50371-2; 3-030-50370-4.

Cardoso:2020:CSI

[912] Pedro J. S. Cardoso, João M. F. Rodrigues, Jânio Monteiro, Roberto Lam, Valeria V. Krzhizhanovskaya, Michael H. Lees, Jack Dongarra, and Peter M. A. Sloot. Computational science in the interconnected world: Selected papers from 2019 International Conference on Computational Science. Journal of Computational Science, 47:??, November 2020. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S1877750320305214.

Dongarra:2020:NAH

[913] Jack Dongarra, Laura Grigori, and Nicholas J. Higham. Numerical algorithms for high-performance computational science. Philosophical transactions of the Royal Society of London Series A, 378(2166):20190066:1– 20190066:18, 2020. CODEN PTR-MAD, PTMSFB. ISSN 1364-503X (print), 1471-2962 (electronic).

Farhan:2020:MTS

[914] Mohammed Al Farhan, Ahmad Abdelfattah, Stanimire Tomov, Mark Gates, Dalal Sukkari, Azzam Haidar, Robert Rosenberg, and Jack Dongarra. MAGMA templates for scalable linear algebra on emerging architectures. The International Journal of High Performance Computing Applications, 34(6): 645–658, November 1, 2020. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https:

                                        //journals.sagepub.com/doi/full/
                                        10.1177/1094342020938421.

Haidar:2020:MPI

[915] Azzam Haidar, Harun Bayraktar, Stanimire Tomov, Jack Dongarra, and Nicholas J. Higham. Mixed-precision iterative refinement using tensor cores on GPUs to accelerate solution of linear systems. Proceedings of the Royal Society of London. Series A, 476 (2243):20200110:1–20200110:30, 2020. CODEN PRLAAZ. ISSN 1364-5021 (print), 1471-2946 (electronic).

Lu:2020:RAC

[916] Yuechao Lu, Ichitaro Yamazaki, Fumihiko Ino, Yasuyuki Matsushita, Stanimire Tomov, and Jack Dongarra. Reducing the amount of out-of-core data access for GPU-accelerated randomized SVD. Concurrency and Computation: Practice and Experience, 32(19): e5754:1–e5754:??, October 10, 2020. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Abdelfattah:2021:SBB

[917] Ahmad Abdelfattah, Timothy Costa, Jack Dongarra, Mark Gates, Azzam Haidar, Sven Hammarling, Nicholas J. Higham, Jakub Kurzak, Piotr Luszczek, Stanimire Tomov, and Mawussi Zounon. A set of batched basic linear algebra subprograms and LAPACK routines. ACM Transactions on Mathematical Software, 47(3): 21:1–21:23, June 2021. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://dl.

acm.org/doi/10.1145/3431921.

Abdelfattah:2021:SNL

[918] Ahmad Abdelfattah, Hartwig Anzt, Erik G. Boman, Erin Carson, Terry Cojean, Jack Dongarra, Alyson Fox, Mark Gates, Nicholas J. Higham, Xiaoye S. Li, Jennifer Loe, Piotr Luszczek, Srikara Pranesh, Siva Rajamanickam, Tobias Ribizel, Barry F. Smith, Kasia Swirydowicz, Stephen Thomas, Stanimire Tomov, Yaohung M. Tsai, and Ulrike Meier Yang. A survey of numerical linear algebra methods utilizing mixed-precision arithmetic. The International Journal of High Performance Computing Applications, 35 (4):344–369, July 1, 2021. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https:

//journals.sagepub.com/doi/full/
10.1177/10943420211003313.

Dongarra:2021:TPM

[919] Jack Dongarra, Mark Gates, Piotr Luszczek, and Stanimire Tomov. Translational process: Mathematical software perspective. Journal of Computational Science, 52:??, May 2021. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

                                        //www.sciencedirect.com/science/
                                        article/pii/S1877750320305160.

Kolev:2021:EED

[920] Tzanio Kolev, Paul Fischer, Misun Min, Jack Dongarra, Jed Brown, Veselin Dobrev, Tim Warburton, Stanimire Tomov, Mark S. Shephard, Ahmad Abdelfattah, Valeria Barra, Natalie Beams, Jean-Sylvain Camier, Noel Chalmers, Yohann Dudouit, Ali Karakus, Ian Karlin, Stefan Kerkemeier, Yu-Hsiang Lan, David Medina, Elia Merzari, Aleksandr Obabko, Will Pazner, Thilina Rathnayake, Cameron W. Smith, Lukas Spies, Kasia Swirydowicz, Jeremy Thompson, Ananias Tomboulides, and Vladimir Tomov. Efficient exascale discretizations: High-order finite element methods. The International Journal of High Performance Computing Applications, 35(6): 527–552, November 1, 2021. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https:

                                        //journals.sagepub.com/doi/full/
                                        10.1177/10943420211020803.

Kovalchuk:2021:YCS

[921] Sergey V. Kovalchuk, Valeria V. Krzhizhanovskaya, Maciej Paszyński, Gábor Závodszky, Michael H. Lees, Jack Dongarra, and Peter M. A. Sloot. 20 years of computational science: Selected papers from 2020 International Conference on Computational Science. Journal of Computational Science, 53:??, July 2021. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:

//www.sciencedirect.com/science/
article/pii/S187775032100082X.

Abdulah:2022:AGM

[922] Sameh Abdulah, Qinglei Cao, Yu Pei, George Bosilca, Jack Dongarra, Marc G. Genton, David E. Keyes, Hatem Ltaief, and Ying Sun. Accelerating geostatistical modeling and prediction with mixed-precision computations: a high-productivity approach with PaRSEC. IEEE Transactions on Parallel and Distributed Systems, 33 (4):964–976, April 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Demmel:2022:PCEa

[923] James Demmel, Jack Dongarra, Mark Gates, Greg Henry, Julien Langou, Xiaoye Li, Piotr Luszczek, Weslley Pereira, Jason Riedy, and Cindy Rubio-González. Proposed consistent exception handling for the BLAS and LAPACK. arXiv.org, ??(??):92, July

19, 2022. URL https://arxiv.org/
abs/2207.09281.

Demmel:2022:PCEb

[924] James Demmel, Jack Dongarra, Mark Gates, Greg Henry, Julien Langou, Xiaoye Li, Piotr Luszczek, Weslley Pereira, Jason Riedy, and Cindy Rubio-González. Proposed consistent exception handling for the BLAS and LAPACK. In IEEE, editor, Correctness 2022: Sixth International Workshop on Software Correctness for HPC Applications: Held in conjunction with SC22: The International Conference for High Performance Computing, Networking, Storage and Analysis. Dallas, Texas, USA, November 13-18, 2022, page ?? IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2022. ISBN 1-66546-335-X. LCCN ????

Dongarra:2022:TLE

[925] Jack J. Dongarra. Turing lecture: The evolution of mathematical software. Communications of the ACM, 65 (12):66–72, December 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3554977.

Kovalchuk:2022:CSB

[926] Sergey V. Kovalchuk, Valeria V. Krzhizhanovskaya, Maciej Paszyński, Dieter Kranzlmüller, Jack Dongarra, and Peter M. A. Sloot. Computational science for a better future. Journal of Computational Science, 62:??, July 2022. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

                                        /www.sciencedirect.com/science/
                                        article/pii/S1877750322001351.

Lindquist:2022:ARG

[927] Neil Lindquist, Piotr Luszczek, and Jack Dongarra. Accelerating restarted GMRES with mixed precision arithmetic. IEEE Transactions on Parallel and Distributed Systems, 33(4): 1027–1037, April 2022. CODEN ITD-SEO. ISSN 1045-9219 (print), 1558-2183 (electronic).

Zhong:2022:ULV

[928] Dong Zhong, Qinglei Cao, George Bosilca, and Jack Dongarra. Using long vector extensions for MPI reductions. Parallel Computing, 109 (??):??, March 2022. CODEN PA-COEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S0167819121001137.

Dongarra:2023:GEN

[929] Jack Dongarra and Bernard Tourancheau. Guest editors note: Special issue on clusters, clouds, and data for scientific computing. The International Journal of High Performance Computing Applications, 37(3-4):211–212, July 1, 2023. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https://

journals.sagepub.com/doi/abs/10.
1177/10943420231180188.

Kovalchuk:2023:CP

[930] Sergey V. Kovalchuk, Clélia de Mulatier, Derek Groen, Maciej Paszyński, Valeria V. Krzhizhanovskaya, Jack Dongarra, and Peter M. A. Sloot. The computational planet. Journal of Computational Science, 72: ??, September 2023. CODEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL https:/

/www.sciencedirect.com/science/
article/pii/S187775032300162X.

Luszczek:2023:CMT

[931] Piotr Luszczek, Wissam M. Sid-Lakhdar, and Jack Dongarra. Combining multitask and transfer learning with deep Gaussian processes for autotuning-based performance engineering. The International Journal of High Performance Computing Applications, 37(3-4):229– 244, July 1, 2023. CODEN IH-PCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https://

journals.sagepub.com/doi/abs/10.
1177/10943420231166365.

Reed:2023:CAH

[932] Daniel Reed, Dennis Gannon, and Jack Dongarra. Contributed articles: HPC forecast: Cloudy and uncertain. Communications of the ACM, 66(2):82–90, February 2023. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

/dl.acm.org/doi/10.1145/3552309.

Abdelfattah:2024:MEE

[933] Ahmad Abdelfattah, Natalie Beams, Robert Carson, Pieter Ghysels, Tzanio Kolev, Thomas Stitt, Arturo Vargas, Stanimire Tomov, and Jack Dongarra. MAGMA: Enabling exascale performance with accelerated BLAS and LAPACK for diverse GPU architectures. The International Journal of High Performance Computing Applications, 38(5):468– 490, September 1, 2024. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https://

journals.sagepub.com/doi/abs/10.
1177/10943420241261960.

Anonymous:2024:RNA

[934] Anonymous. Retraction notice: Azzam Haidar and Tingxing Dong and Piotr Luszczek and Stanimire Tomov and Jack Dongarra, Batched matrix computations on hardware accelerators based on GPUs, Int. J. High Perform. Comput. Appl. 29(2) 193–208 (2015). The International Journal of High Performance Computing Applications, 38(4):377, July 1, 2024. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https://

journals.sagepub.com/doi/abs/10.
1177/10943420241257113. See [839].

Lindquist:2024:GRB

[935] Neil Lindquist, Piotr Luszczek, and Jack Dongarra. Generalizing random butterfly transforms to arbitrary matrix sizes. ACM Transactions on Mathematical Software, 50(4):26:1– 26:??, December 2024. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL https://dl.

acm.org/doi/10.1145/3699714.

Luszczek:2024:NES

[936] Piotr Luszczek, Anthony Castaldo, Yaohung M. Tsai, Daniel Mishler, and Jack Dongarra. Numerical eigen-spectrum slicing, accurate orthogonal eigen-basis, and mixedprecision eigenvalue refinement using OpenMP data-dependent tasks and accelerator offload. The International Journal of High Performance Computing Applications, 38(6):671– 691, November 1, 2024. CODEN IHPCFL. ISSN 1094-3420 (print), 1741-2846 (electronic). URL https://

journals.sagepub.com/doi/abs/10.
1177/10943420241281050.

Birkhoff:1980:BRC

[937] Garrett Birkhoff. Book review: Computational analysis with the HP 25 pocket calculator, by Peter Henrici, John Wiley, New York, 1977, 280 pp.; Compact numerical methods for computers: linear algebra and function minimization, by J. C. Nash, John Wiley & Sons, New York, 1979, x + 227 pp.; LINPACK: User’s guide, by J. J. Dongarra, J. R. Bunch, C. B. Moler, and G. W. Stewart, Society for Industrial and Applied Mathematics, Philadelphia, 1979 368 pp. Bulletin of the American Mathematical Society, 2(3):503–505, May 1980. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881X (elec-

tronic). URL https://www.ams.org/
journals/bull/1980-02-03/S0273-
0979-1980-14785-0/S0273-0979-1980-
14785-0.pdf.

Feldman:2019:TML

[938] Michael Feldman. Tweaked math libraries exploit AI hardware for traditional HPC. Web site., December 3, 2019. URL https:/

/bit.ly/hpl-ai/;   https://icl.
bitbucket.io/hpl-ai/;   https:/
/www.math.utah.edu/pub/tex/bib/
fparith.bib;   https://www.nextplatform.
com/2019/12/03/tweaked-math-libraries-
exploit-ai-hardware-for-traditional-

hpc/. Jack Dongarra describes work to benchmark, and utilize, 16-bit floatingpoint formats. The new benchmark is called HPL-AI.

Hammarling:2022:ICJ

[939] Sven Hammarling and Nicholas J. Higham. The influence and contribution of Jack Dongarra to numerical linear algebra. Computing in Science and Engineering, 24(4):6–11, July/August 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Hoefler:2024:XAS

[940] Torsten Hoefler, Marcin Copik, Pete Beckman, Andrew Jones, Ian Foster, Manish Parashar, Daniel Reed, Matthias Troyer, Thomas Schulthess, Daniel Ernst, and Jack Dongarra. XaaS: Acceleration as a service to enable productive high-performance cloud computing. Computing in Science and Engineering, 26(3):40– 51, July/September 2024. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Kovalchuk:2024:CCE

[941] Sergey V. Kovalchuk, Clélia de Mulatier, Valeria V. Krzhizhanovskaya, Jiří Mikyška, Maciej Paszyński, Jack Dongarra, and Peter M. A. Sloot. Computation at the cutting edge of science. Journal of Computational Science, 81(??):??, September 2024. CO-DEN ???? ISSN 1877-7503 (print), 1877-7511 (electronic). URL http:

//www.sciencedirect.com/science/
article/pii/S1877750324001728.

Parashar:2022:JDC

[942] Manish Parashar. Jack Dongarra: Catalyzing the transformation of highperformance computing. Computing in Science and Engineering, 24(4):4–5, July/August 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Vuduc:2022:JA

[943] Richard W. Vuduc. Jack, the autotuner. Computing in Science and Engineering, 24(4):24–27, July/August 2022. CODEN CSENFA. ISSN 1521-9615 (print), 1558-366X (electronic).

Urschel:2025:NSG

[944] John Urschel. Numerical stability in Gaussian elimination. Notices of the American Mathematical Society, 72(6): 605–613, June/July 2025. CODEN AMNOAN. ISSN 0002-9920 (print), 1088-9477 (electronic).

Rodrigue:1989:PPS

[945] G. Rodrigue, editor. Parallel processing for scientific computing. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1989. ISBN 0-89871-228-9. 428 pp. LCCN QA76.5 .S515 1987.

Sanders:2010:CEI

[946] Jason Sanders and Edward Kandrot. CUDA by Example: an Introduction to General-purpose GPU Programming. Addison-Wesley, Reading, MA, USA, 2010. ISBN 0-13-138768-5. xix + 290 pp. LCCN QA76.76.A65.

Hager:2011:IHP

[947] Georg Hager and Gerhard Wellein. Introduction to high performance computing for scientists and engineers, volume 7 of Chapman and Hall/CRC computational science series. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2011. ISBN 1-4398-1192-X. xxv + 330 + 4 pp. LCCN QA76.88 .H34 2011.

Buzbee:1978:PLW

[948] B. L. Buzbee and J. F. Morrison, editors. Proceedings of the 1978 LASL Workshop on Vector and Parallel Processors held at Los Alamos Scientific Laboratory, Los Alamos, New Mexico, September 20–22, 1978. Los Alamos Scientific Laboratory, Los Alamos, NM, USA, 1978. U.S. Scientific Laboratory, Los Alamos, NM Conference proceedings LA-7491-C.

Cowell:1984:SDM

[949] Wayne R. Cowell, editor. Sources and Development of Mathematical Software. Prentice-Hall Series in Computational Mathematics, Cleve Moler, Advisor. Prentice-Hall, Englewood Cliffs, NJ 07632, USA, 1984. ISBN 0-13-823501-5. xii + 404 pp. LCCN QA76.95 .S68 1984.

Dongarra:1984:IPS

[950] Jack Dongarra, Gene Golub, Jorge Moré, and Danny Sorensen, editors. Informal proceedings of the Symposium on Computational Mathematics — State of the Art: held at Argonne National Laboratory, September 20–21, 1984, in honor of James H. Wilkinson. Argonne National Laboratory, 9700 South Cass Avenue, Argonne, IL 60439-4801, USA, December 1984. LCCN QA297.S879 1984. Technical Report MCS-TM-42. It consists of copies of lecture slides from the ten symposium talks.

Glowinski:1984:CMA

[951] R. Glowinski and J.-L. Lions, editors. Computing Methods in Applied Sciences and Engineering, VI: Proceedings of the Sixth International Symposium on Computing Methods in Applied Sciences and Engineering, Versailles, France, December 12–16, 1983. North-Holland, Amsterdam, The Netherlands, 1984. ISBN 0-444-87597-2. LCCN QA297 .I57 1983.

Hwang:1985:PSC

[952] Kai Hwang, editor. Proceedings: 7th Symposium on Computer Arithmetic, June 4–6, 1985, University of Illinois, Urbana, Illinois. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1985. ISBN 0-8186-0632-0 (paperback), 0-8186-8632-4 (hard), 0-8186-4632-2 (microfiche). LCCN QA76.9.C62 S95 1985.

Bell:1986:DPC

[953] Alan G. Bell, editor. Digest of papers / Compcon 86, March 3–6, Spring; Thirty-first IEEE Computer Society International Conference, Cathedral Hill Hotel, San Francisco, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. CODEN PCICDQ. ISBN 0-8186-0692-4 (paperback), 0-8186-4692-6 (microfiche). LCCN QA75.5.C58 1986. IEEE Computer Society order number 692; IEEE catalog number 86CH2285-5.

Cullum:1986:LSE

[954] Jane Cullum and Ralph A. Willoughby, editors. Large Scale Eigenvalue Problems. Proceedings of the IBM European Institute Workshop. North-Hol-land, Amsterdam, The Netherlands, 1986. ISBN 0-444-70074-9. LCCN QA193 .I261 1985.

Feilmeier:1986:PCP

[955] M. Feilmeier, G. Joubert, and U. Schendel, editors. Parallel Computing 85: Proceedings of the Second International Conference on Parallel Computing, held at the Freie Universität Berlin, 23–25, September 1985. North-Holland, Amsterdam, The Netherlands, 1986. ISBN 0-444-70009-9. LCCN QA76.6 .I5471 1985.

Wouk:1986:NCE

[956] Arthur Wouk, editor. New computing environments: parallel, vector and systolic (Stanford University, Stanford, CA, November 7–9, 1984). Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1986. ISBN 0-89871-201-7. LCCN QA76.9.A73 N49 1986.

Anonymous:1987:ISS

[957] Anonymous, editor. International Seminar on Scientific Supercomputers. Inst. Nat. Rech. and Inf. Autom, Le Chesnay, France, 1987.

Jamieson:1987:CPA

[958] Leah H. Jamieson, Dennis B. Gannon, and Robert J. Douglass, editors. The Characteristics of Parallel Algorithms, volume 30 of MIT Press series in scientific computation. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-10036-3. LCCN QA76.6 .C42981 1987.

Houstis:1988:SIC

[959] E. N. Houstis, T. S. Papatheodorou, and C. D. Polychronopoulos, editors. Supercomputing: 1st International Conference, Athens, Greece, June 8–12, 1987: proceedings, volume 297 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. CODEN LNCSD9. ISBN 0-387-18991-2, 3-540-18991-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.297. The conference was organized and sponsored by the Computer Technology Institute (C.T.I.) of Greece.

IEEE:1988:PSN

[960] IEEE, editor. Proceedings, Supercomputing ’88: November 14–18, 1988, Orlando, Florida, volume 1. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1988. ISBN 0-8186-0882-X (v. 1; paper), 0-8186-8882-3 (v. 1; case), 0-8186-4882-1 (v. 1: microfiche) 0-8186-8923-4 (v. 2), 0-8186-5923-X (v. 2: microfiche), 0-8186-8923-4 (v. 2: case). LCCN QA76.5 .S894 1988. Two volumes. Available from IEEE Service Center (Catalog number 88CH2617-9), Piscataway, NJ, USA.

Paul:1988:PSC

[961] George Paul and George S. Almasi, editors. Parallel systems and computation: proceedings of the 1986 IBM Europe Institute—Seminar on Parallel Computing, Oberlech, Austria, August 11– 15, 1986. North-Holland, Amsterdam, The Netherlands, 1988. ISBN 0-444-70371-3. LCCN QA76.5 .I147 1986.

Schultz:1988:NAM

[962] Martin Schultz, editor. Numerical algorithms for modern parallel computer architectures (Minneapolis, MN, 1986– 87), volume 13 of The IMA volumes in mathematics and its applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-96733-8. LCCN QA76.5.N79 1988.

Carey:1989:PSM

[963] Graham F. Carey, editor. Parallel supercomputing: methods, algorithms and applications. Wiley series in parallel computing. Wiley, New York, NY, USA, 1989. ISBN 0-471-92436-9. x + 287 pp. LCCN M89.E02452; QA76.6.

Cosnard:1989:PDA

[964] Michel Cosnard, Y. Robert, P. Quinton, and M. Raynal, editors. Parallel & distributed algorithms: proceedings of the International Workshop on Parallel & Distributed Algorithms, Chateau de Bonas, Gers, France, 3–6 October, 1988. North-Holland, Amsterdam, The Netherlands, 1989. ISBN 0-444-87367-8. LCCN QA76.5 .I6191 1988.

Dongarra:1989:VPC

[965] Jack Dongarra, Iain S. Duff, Patrick W. Gaffney, and Sean StJ. McKee, editors. Vector and Parallel Computing: Issues in Applied Research and Development (Papers from LOEN II, a conference held June 6–10, 1988, in Tromsø, Norway). Ellis Horwood, New York, NY, USA, 1989. ISBN 0-7458-0756-9 (Ellis Horwood), 0-470-21571-2 (Halsted Press). LCCN QA76.5 .V398 1989. UK£29.95.

Evans:1989:PCM

[966] D. J. Evans and C. Sutti, editors. Parallel Computing. Methods, Algorithms and Applications Proceedings of the International Meeting on Parallel Computing. Adam Hilger, Bristol, UK, 1989. ISBN 0-85274-224-x. LCCN QA76.5.I5775 1988.

Wright:1989:ACA

[967] M. Wright, editor. Aspects of Computation on Asynchronous Parallel Processors. Proceedings of the IFIP WG 2.5 Working Conference, Stanford, CA, USA, 22–26 August, 1988. North-Hol-land, Amsterdam, The Netherlands, 1989. ISBN 0-444-87310-4. LCCN QA76.5 .I2775 1988.

Cox:1990:RNC

[968] M. G. Cox and S. Hammarling, editors. Reliable numerical computation. Oxford University Press, Walton Street, Oxford OX2 6DP, UK, 1990. ISBN 0-19-853564-3. LCCN QA297 .R435 1990. US$75.00. Based on papers from a conference in honour of the late James Hardy Wilkinson (died Sunday 5th October 1986) held at National Physical Laboratory, Teddington, Middlesex, UK, 8th–10th July 1987.

Dongarra:1990:PPS

[969] Jack Dongarra, Paul Messina, Danny C. Sorensen, and Robert G. Voigt, editors. Parallel Processing for Scientific Computing: Proceedings of the Fourth SIAM Conference on Parallel Processing for Scientific Computing, Chicago, Illinois, December 11–13, 1989. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1990. ISBN 0-89871-262-9. LCCN QA76.58.S55 1989.

IEEE:1990:PSN

[970] IEEE, editor. Proceedings, Supercomputing ’90: November 12–16, 1990, New York Hilton at Rockefeller Center, New York, New York. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-2056-0 (paperback) (IEEE Computer Society), 0-89791-412-0 (paperback) (ACM). LCCN QA 76.88 S87 1990. ACM order number 415903. IEEE Computer Society Press order number 2056. IEEE catalog number 90CH2916-5.

VanderSteen:1990:ESS

[971] Aad J. Van der Steen, editor. Evaluating supercomputers. Strategies for exploiting, evaluating and benchmarking computers with advanced architectures, volume 8 of Unicom applied information technology reports. Chapman and Hall, Ltd., London, UK, 1990. ISBN 0-412-37860-4, 0-442-31198-2 (U.S.). LCCN QA76.9.E94 E93 1990; QA76.88.E93 1990.

vanderVorst:1990:PAN

[972] Henk A. van der Vorst and Paul Van Dooren, editors. Parallel algorithms for numerical linear algebra, volume 1 of Advances in Parallel Computing. North-Holland, Amsterdam, The Netherlands, 1990. ISBN 0-444-88621-4. x + 330 pp. LCCN QA76.5 .P31458 1990.

Anonymous:1991:ISS

[973] Anonymous, editor. International Symposium on Supercomputing ’91. Amsterdam Universities Computing Centre (SARA), Amsterdam, The Netherlands, November 1991. CODEN SPCOEL. ISSN 0168-7875. Supercomputer, vol. 8, no. 6, Nov. (1991).

Anonymous:1991:PIS

[974] Anonymous, editor. Proceedings of the International Symposium on Supercomputing: Fukuoka, Japan, November 6–8, 1991. Kyushu University Press, Fukouka, Japan, 1991. ISBN 4-87378-284-8. LCCN QA76.88 .I587 1991.

Griffiths:1991:NAP

[975] D. F. Griffiths and G. A. Watson, editors. Numerical analysis, 1991: proceedings of the 14th Dundee Conference, June 1991, volume 260 of Pitman Res. Notes Math. Ser. Longman Scientific and Technical, Harlow, Essex, UK, 1991. ISBN 0-582-08908-5. LCCN QA297 .D85 1991.

IEE:1991:SIS

[976] IEE, editor. Second International Specialist Seminar on the Design and Application of Parallel Digital Processors (Conf. Publ. No. 334). IEE, London, UK, 1991. ISBN 0-85296-519-2. ISSN 0537-9989. LCCN QA76.58.I567 1991.

IEEE:1991:PSA

[977] IEEE, editor. Proceedings, Supercomputing ’91: Albuquerque, New Mexico, November 18–22, 1991. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-9158-1 (IEEE case), 0-8186-2158-3 (IEEE paper), 0-8186-6158-5 (IEEE microfiche), 0-89791-459-7 (ACM). LCCN QA76.5 .S894 1991. ACM order number 415913. IEEE Computer Society Press order number 2158. IEEE catalog number 91CH3058-5.

Stout:1991:SDM

[978] Quentin Stout and Michael Wolfe, editors. The Sixth Distributed Memory Computing Conference Proceedings: April 28–May 1, 1991, Portland, Oregon. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-2290-3 (paperback), 0-8186-2291-1 (fiche). LCCN QA76.5 .D58 1991.

Adeli:1992:SEA

[979] Hojjat Adeli, editor. Supercomputing in engineering analysis, volume 1 of New generation computing. Marcel Dekker, New York, NY, USA, 1992. ISBN 0-8247-8559-2. xi + 362 pp. LCCN TA345.S87; TA345 .S87 1991; TA345 .S87 1992.

Dongarra:1992:PFS

[980] Jack Dongarra, Ken Kennedy, Paul Messina, Danny C. Sorensen, and Robert G. Voigt, editors. Proceedings of the Fifth SIAM Conference on Parallel Processing for Scientific Computing (Houston, TX, March 25–27, 1991). Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1992. ISBN 0-89871-303-X. LCCN QA76.58 .P76 1992.

Griffiths:1992:NAP

[981] D. F. Griffiths and G. A. Watson, editors. Numerical analysis, 1991: proceedings of the 14th Dundee Conference, June 1991, volume 260 of Pitman Res. Notes Math. Ser. Longman Scientific and Technical, Harlow, Essex, UK, 1992. ISBN 0-582-08908-5. LCCN QA297.D85 1991.

IEEE:1992:SHP

[982] IEEE, editor. Scalable High Performance Computing Conference, SHPCC-92, April 26–29, 1992, Williamsburg, Virginia. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2775-1. LCCN QA76.76.A65 S33 1992.

Perrott:1992:SPC

[983] R. H. Perrott, editor. Software for parallel computers. Chapman and Hall, Ltd., London, UK, 1992. ISBN 0-412-39960-1. 350 pp. LCCN QA76.58.S63 1992.

Siegel:1992:FSF

[984] H. J. Siegel, editor. The Fourth Symposium on the Frontiers of Massively Parallel Computation: Frontiers ’92 / October 19–21, 1992, McLean, Virginia. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1992. ISBN 0-8186-2772-7 (hardback), 0-8186-2771-9 (microfiche). LCCN QA76.58 .S95 1992.

Dongarra:1993:CB

[985] Jack J. Dongarra and Wolfgang Gentzsch, editors. Computer Benchmarks, volume 8 of Advances in Parallel Computing. North-Holland, Amsterdam, The Netherlands, 1993. ISBN 0-444-81518-X. LCCN QA76.9.E94 C63 1993.

Dongarra:1993:ETP

[986] J. J. Dongarra and B. Tourancheau, editors. Environments and Tools for Parallel Scientific Computing (Saint Hilaire du Touvet, France, September 7– 8, 1992). North-Holland, Amsterdam, The Netherlands, 1993. ISBN 0-444-89963-4. LCCN QA76.642 .E58 1993.

Fincham:1993:CSO

[987] A. E. Fincham and B. Ford, editors. Parallel Computation Conference — September 1991, St. Catherine’s College, Oxford, UK, volume 46 of Institute of Mathematics and its Applications Conference Series. Clarendon Press, Oxford, UK, 1993. ISBN 0-19-853680-1. ISSN 0960-2526. LCCN QA76.58.P3755 1993.

Hoffmann:1993:PSA

[988] Geerd-R. Hoffmann and Tuomo Kauranne, editors. Parallel supercomputing in atmospheric science: proceedings of the fifth ECMWF Workshop on the Use of Parallel Processors in Meteorology, Reading UK, November 23–27, 1992. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 1993. ISBN 981-02-1429-4. LCCN QA76.58 E354 1992.

IEEE:1993:PSP

[989] IEEE, editor. Proceedings, Supercomputing ’93: Portland, Oregon, November 15–19, 1993. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4340-4 (paperback), 0-8186-4341-2 (microfiche), 0-8186-4342-0 (hardback), 0-8186-4346-3 (CD-ROM). ISSN 1063-9535. LCCN QA76.5 .S96 1993.

Kowalik:1993:SPC

[990] J. S. Kowalik and L. Grandinetti, editors. Software for Parallel Computation: Proceedings of the NATO Advanced Workshop on Software for Parallel Computation, held at Cetraro, Cosenza, Italy, June 22–26, 1992, volume 106 of NATO ASI Series F. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1993. ISBN 3-540-56451-9 (Berlin), 0-387-56451-9 (New York). LCCN QA76.58 .S629 1993.

Sincovec:1993:SCP

[991] R. F. Sincovec, editor. SIAM Conference on Parallel Processing for Scientific Computing (6th: 1993: Norfolk, VA, USA). Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1993. ISBN 0-89871-315-3. LCCN QA 76.58 S55 1993. Two volumes.

Anonymous:1994:HPC

[992] Anonymous, editor. High Performance Computing and Communications 1st Symposium – March 1994, Alexandria, VA, USA. ARPA-CSTO, Arlington, VA, USA, 1994.

Anonymous:1994:OON

[993] Anonymous, editor. Object oriented numerics: Annual conference: 2nd — April 1994, Sunriver, OR. RWS, Corvallis, OR, 1994.

Dongarra:1994:PSC

[994] J. J. Dongarra and Jerzy Waśniewski, editors. Parallel scientific computing: First International Workshop, PARA ’94, Lyngby, Denmark, June 20–23, 1994: proceedings, volume 879 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1994. CO-DEN LNCSD9. ISBN 3-540-58712-8 (Berlin), 0-387-58712-8 (New York). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1994. DM104.00. URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/tocs/t0879.htm;   http://www.
                                        springerlink.com/openurl.asp?genre=
                                        issue&issn=0302-9743&volume=879.

Dongarra:1994:PSW

[995] J. J. Dongarra and B. Tourancheau, editors. Proceedings of the Second Workshop on Environments and Tools for Parallel Scientific Computing, Townsend, TN, USA, May 25– 27, 1994, Proceedings of the Workshop on Environments and Tools for Parallel Scientific Computing. Society for Industrial and Applied Mathematics, Philadelphia, PA, USA, 1994. ISBN 0-89871-343-9. LCCN QA76.58.I568 1994.

Gilbert:1994:LMP

[996] John E. Gilbert and Donald Kershaw, editors. Large-scale matrix problems and the numerical solution of partial differential equations: 5th Summer school in numerical analysis — July 1992, Lancaster, UK, volume 3 of Advances in Numerical Analysis. Clarendon Press, Oxford, UK, 1994. ISBN 0-19-853463-9. LCCN QA374.L335 1994.

IEEE:1994:PSH

[997] IEEE, editor. Proceedings of the Scalable High-Performance Computing Conference, May 23–25, 1994, Knoxville, Tennessee. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5680-8, 0-8186-5681-6. LCCN QA76.58.S32 1994.

IEEE:1994:PSP

[998] IEEE, editor. Proceedings of the Scalable Parallel Libraries Conference, October 6–8, 1993, Mississippi State, Mississippi. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-4980-1 (paper), 0-8186-4981-X (microfiche). LCCN QA76.58 .S34 1993.

IEEE:1994:PTI

[999] IEEE, editor. Proceedings of the Third IEEE International Symposium on High Performance Distributed Computing, August 2-5, 1994, San Francisco, CA, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-6395-2, 0-8186-6396-0. LCCN QA76.9.D5I328 1994. IEEE catalog number 94TH0667-6.

Siegel:1994:PEI

[1000] Howard Jay Siegel, editor. Proceedings: Eighth International Parallel Processing Symposium, April 26–29, 1994, Cancun, Mexico. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1994. ISBN 0-8186-5602-6, 0-8186-5601-8. ISSN 1063-7133. LCCN QA 76.58 I56 1994.

Dongarra:1995:HPC

[1001] J. J. Dongarra et al., editors. High performance computing: technology, methods, and applications (Advanced workshop, June 1994, Cetraro, Italy), volume 10 of Advances in Parallel Computing. Elsevier, Amsterdam, The Netherlands, 1995. ISBN 0-444-82163-5. ISSN 0927-5452. LCCN QA76.88.H55 1995.

IEEE:1995:DPT

[1002] IEEE, editor. Digest of papers / the Twenty-fifth International Symposium on Fault-Tolerant Computing, June 27–30, 1995, Pasadena, California, FTCS 25th. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. CODEN DPFTDL. ISBN 0-8186-7079-7, 0-8186-7145-9. ISSN 0731-3071. LCCN QA 76.9 F38 I57 1995. IEEE catalog number 95CH35823.

IEEE:1995:FHC

[1003] IEEE, editor. Fourth heterogeneous computing workshop: April 25, 1995, Santa Barbara, CA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7121-1. LCCN QA76.9.D5 H48 1995.

IEEE:1995:PFI

[1004] IEEE, editor. Proceedings of the Fifth IEEE Computer Society Workshop on Future Trends of Distributed Computing Systems, August 28–30, 1995, Cheju Island, Korea. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7125-4. LCCN QA76.9.D5 I24 1995.

Karin:1995:PAI

[1005] Sidney Karin, editor. Proceedings of the 1995 ACM/IEEE Supercomputing Conference, December 3–8, 1995, San Diego Convention Center, San Diego, CA, USA. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-7803-3604-6, 0-89791-862-2, 0-7803-3605-4. ISSN 1063-9535. LCCN QA76.88 .S856

1995. URL http://www.supercomp.
org/sc95/proceedings/. ACM order

number 415932. IEEE Order Plan catalog number 95CB35990. These proceedings are not available in printed form.

However, they are available on the World Wide Web, and on CD-ROM, available from ACM (ACM Press order number 415952) and IEEE (IEEE Computer Society Press order number FW07435).

ACM:1996:SCP

[1006] ACM, editor. Supercomputing ’96 Conference Proceedings: November 17– 22, Pittsburgh, PA. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-89791-854-1. LCCN A76.88 .S8573

1996. URL http://www.supercomp.
org/sc96/proceedings/. ACM Order

Number: 415962, IEEE Computer Society Press Order Number: RS00126.

Bode:1996:PVM

[1007] Arndt Bode, J. Dongarra, T. Ludwig, and V. Sunderam, editors. Parallel virtual machine, EuroPVM ’96: third European PVM conference, Munich, Germany, October 7–9, 1996: proceedings, volume 1156 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61779-5. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E975 1996 Bar.

Bouge:1996:EPP

[1008] L. (Luc) Bouge, editor. Euro-Par’96: parallel processing: second International Euro-Par Conference, Lyon, France, August 26–29, 1996: proceedings, volume 1123, 1124 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61626-8 (vol. 1), 3-540-61627-6 (vol. 2). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.I554 1996.

Dongarra:1996:APC

[1009] J. J. Dongarra, Kaj Madsen, and Jerzy Waśniewski, editors. Applied parallel computing: computations in physics, chemistry, and engineering science: second international workshop, PARA ’95, Lyngby, Denmark, August 21– 24, 1995: proceedings, volume 1041 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. CODEN LNCSD9. ISBN 3-540-60902-4. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1995.

IEEE:1996:FSS

[1010] IEEE, editor. Frontiers ’96: The Sixth Symposium of Massively Parallel Computing: October 27–31, 1996, Annapolis, Maryland, Symposium on the Frontiers of Massively Parallel Computations. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7551-9 (paper), 0-8186-7553-5 (microfiche). ISSN 1088-4955. LCCN QA76.58 .S95 1996. IEEE Computer Society Press Order Number PR07551. IEEE order plan catalog number 96TB100062.

IEEE:1996:HCW

[1011] IEEE, editor. Heterogeneous computing workshop: 5th — April 1996, Honolulu, HI. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996.

Liddell:1996:HCN

[1012] Heather Mary Liddell, A. Colbrook, B. Hertzberger, and P. Sloot, editors. High-performance computing and networking: international conference and exhibition, HPCN EUROPE 1966, Brussels, Belgium, April 15–19, 1996: proceedings, volume 1067 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-61142-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .H52 1996.

Wasniewski:1996:APC

[1013] Jerzy Waśniewski, J. Dongarra, K. Madsen, and D. Olesen, editors. Applied parallel computing: industrialstrength computation and optimization: Third International Workshop, PARA 96, Lyngby, Denmark, August 18–21, 1996: proceedings, volume 1184 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1996. ISBN 3-540-62095-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P35 1996.

ACM:1997:SHP

[1014] ACM, editor. SC’97: High Performance Networking and Computing: Proceedings of the 1997 ACM/IEEE SC97 Conference: November 15–21, 1997, San Jose, California, USA. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-89791-985-8. LCCN QA76.9.A25 A265

1997. URL http://www.acm.org/
pubs/contents/proceedings/commsec/
266741/;   http://www.supercomp.
org/sc97/proceedings/.  ACM

SIGARCH order number 415972. IEEE Computer Society Press order number RS00160.

Anonymous:1997:VPC

[1015] Anonymous, editor. Vector and parallel computing: Workshop: 21st — March 1997, volume 11(1). SPEEDUP Society, Manno, Switzerland, 1997. ISSN 1421-6337.

Boisvert:1997:QNS

[1016] R. F. Boisvert, editor. Quality of numerical software: assessment and enhancement / proceedings of the IFIP TC2/WG2.5 Working Conference on the Quality of Numerical Software, Assessment and Enhancement, Oxford, United Kingdom, 8–12 July 1996. Chapman and Hall, Ltd., London, UK, 1997. ISBN 0-412-80530-8. LCCN QA297 .I35 1996.

Bubak:1997:RAP

[1017] Marian Bubak, J. J. Dongarra, and Jerzy Waśniewski, editors. Recent advances in parallel virtual machine and message passing interface: 4th European PVM/MPI user’s group meeting Cracow, Poland, November 3–5, 1997: proceedings, volume 1332 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CODEN LNCSD9. ISBN 3-540-63697-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58.E973

1997. URL http://link.springer-

                                        ny.com/link/service/series/0558/
                                        tocs/t1332.htm.

Dongarra:1997:PTW

[1018] J. J. Dongarra and Bernard Tourancheau, editors. Proceedings of the Third Workshop on Environments and Tools for Parallel Scientific Computing, volume 11(2) of International Journal of Supercomputer Applications and High Performance Computing. Sage Science Press, Thousand Oaks, CA, USA, 1997. LCCN QA 76.5 I55 v.11 no.2 1997.

Dongarra:1997:VPP

[1019] J. J. Dongarra and Jose M. L. M. Palma, editors. Vector and parallel processing—VECPAR ’96: Second International Conference on Vector and Parallel Processing-Systems and Applications, Porto, Portugal, September 25–27, 1996: Selected Papers, volume 1215 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1997. CO-DEN LNCSD9. ISBN 3-540-62828-

2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I552

                                        1996. URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        tocs/t1215.htm;   http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=1215.

Goscinski:1997:ICA

[1020] Andrzej Goscinski, Michael Hobbs, and Wanlei Zhou, editors. 1997 3rd International Conference on Algorithms and Architectures for Parallel Processing: ICA3PP/97: Melbourne, Australia, December 10–12, 1997, IN-TERNATIONAL CONFERENCE ON ALGORITHMS AND ARCHITEC-TURES FOR PARALLEL PROCESS-ING 1997; 3rd. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-7803-4229-1 (softbound), 0-7803-4230-5 (microfiche). LCCN QA76.58 .I528 1997.

Grandinetti:1997:HPC

[1021] Lucio Grandinetti, Janusz Kowalik, and Marian Vajtersic, editors. High performance computing: technology and applications: NATO Advanced Research Workshop — 24–26 June 1996, Cetraro, Italy, volume 30 of NATO ASI Series 3 High Technology. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 1997. ISBN 0-7923-4550-9. LCCN QA76.88 .A38 1997.

IEEE:1997:HPC

[1022] IEEE, editor. High performance computing on the information superhighway, HPC Asia ’97: proceedings, Seoul, Korea, April 28–May 2, 1997, HPC ASIA 1997. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7901-8, 0-8186-7902-6 (casebound) , 0-8186-7903-4 (microfiche). LCCN QA76.88 .H653 1997. IEEE Order Plan number 97TB100110.

IEEE:1997:PIC

[1023] IEEE, editor. Proceedings: 1997 International Conference on Parallel Architectures and Compilation Techniques: San Francisco, California, November 10–14, 1997. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997.

ISBN 0-8186-8090-3. LCCN QA76.58 .I5445 1997.

Sydow:1997:IWC

[1024] Achim Sydow, editor. 15th IMACS World Congress on Scientific Computation, Modelling and Applied Mathematics: Berlin, August 1997: proceedings, IMACS World Congress. Wissenschaft and Technik, Berlin, Germany, 1997. ISBN 3-89685-550-6 (set), 3-89685-551-4 (vol. 1), 3-89685-552-2 (vol. 2), 3-89685-553-0 (vol. 3),3-89685-554-9 (vol. 4), 3-89685-555-7 (vol. 5), 3-89685-556-5 (vol. 6). LCCN Q183.9 .I46 1997. In cooperation with R.-P. Schafer, W. Rufeger, and H. Lehmann.

Thiele:1997:IIC

[1025] L. Thiele, editor. IEEE International Conference on Application-Specific Systems, Architectures and Processors: proceedings, July 14–16, 1997: Zurich, Switzerland. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7959-X (casebound), 0-8186-7958-1, 0-8186-7960-3 (microfiche). ISSN 1063-6862. LCCN TK7874.6 .I58 1997.

ACM:1998:AWJ

[1026] ACM, editor. ACM 1998 Workshop on Java for High-Performance Network Computing, February 28 and March 1, 1998. ACM Press, New York, NY 10036, USA, 1998. ISBN ???? LCCN

                                        ???? URL http://www.cs.ucsb.edu/
                                        conferences/java98/program.html.

Possibly unpublished, except electronically.

ACM:1998:SHP

[1027] ACM, editor. SC’98: High Performance Networking and Computing: Proceedings of the 1998 ACM/IEEE SC98 Conference: Orange County Convention Center, Orlando, Florida, USA, November 7–13, 1998. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN ???? LCCN ???? URL http://www.

supercomp.org/sc98/papers/.

Alexandrov:1998:RAP

[1028] Vassil Alexandrov and J. J. Dongarra, editors. Recent advances in parallel virtual machine and message passing interface: 5th European PVM/MPI User’s Group Meeting, Liverpool, UK, September 7–9, 1998: Proceedings, volume 1497 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1998. CODEN LNCSD9. ISBN 3-540-65041-5 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1497. URL http://link.

springer-ny.com/link/service/series/
0558/tocs/t1497.htm;   http://www.
springerlink.com/openurl.asp?genre=
issue&issn=0302-9743&volume=1497.

Jointly sponsored by the Computer Science Dept., University of Liverpool and Oak Ridge National Laboratory.

Antonio:1998:SHC

[1029] John K. Antonio, editor. Proceedings: Seventh Heterogeneous Computing Workshop (HCW’98): March 30, 1998, Orlando, Florida, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-8365-1, 0-8186-8367-8 (microfiche). LCCN QA76.88 .H48 1998. IEEE catalog number 98EX126.

DHollander:1998:PCF

[1030] E. D’Hollander et al., editors. Parallel computing: fundamentals, applications, and new directions: Papers from ParCo97, held in Bonn, Germany, Sept. 19–22, 1997, volume 12 of Advances in Parallel Computing. Elsevier, Amsterdam, The Netherlands, 1998. ISBN 0-444-82882-6. LCCN QA76.58.P3795 1997.

IEEE:1998:PSI

[1031] IEEE, editor. Proceedings: the Seventh IEEE International Symposium on High Performance Distributed Computing, July 28–31, 1998, Chicago, Illinois. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1998. ISBN 0-8186-8579-4, 0-8186-8581-6. ISSN 1082-8907. LCCN QA76.9.D5 I157 1998. IEEE Order Plan Catalog Number 98TB100244. IEEE Computer Society Press order number PR08579.

Papailiou:1998:PFE

[1032] Kyriacos D. Papailiou, editor. Proceedings of the Fourth European Computational Fluid Dynamics Conference, 7–11 September 1998, Athens, Greece, volume 2. Wiley, New York, NY, USA, 1998. ISBN 0-471-98579-1 (vol. 1), 0-471-98580-5 (vol. 2). LCCN QA911 .E95 1998. Three volumes in two books.

Dongarra:1999:RAP

[1033] J. J. Dongarra, E. Luque, and Tomas Margalef, editors. Recent advances in parallel virtual machine and message passing interface: 6th European PVM/MPI Users’ Group Meeting, Barcelona, Spain, September 26– 29, 1999: Proceedings, volume 1697 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNCSD9. ISBN 3-540-66549-8 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E973

1999. URL http://link.springer-
ny.com/link/service/series/0558/
tocs/t1697.htm;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=1697.

Heath:1999:APP

[1034] Michael T. Heath, Abhiram Ranade, and Robert S. Schreiber, editors. Algorithms for parallel processing: Proceedings of the Workshop on Algorithms for Parallel Processing, held September 16–20, 1996, at the IMA, University of Minnesota, volume 105 of The IMA volumes in mathematics and its applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 0-387-98680-4. LCCN QA76.58 .A543 1999.

URL http://www.loc.gov/catdir/
enhancements/fy0817/98033425-t.
html.

Hernandez:1999:VPP

[1035] Vicente Hernandez, Jose M. L. M.

Palma, and J. J. Dongarra, editors. Vector and parallel process — VEC-PAR ’98: Third International Conference, Porto, Portugal, June 21–23, 1998: selected papers and invited talks, volume 1573 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. ISBN 3-540-66228-6 (softcover). LCCN QA76.58 .I552 1998 Bar.

Palma:1999:VPP

[1036] José M. L. M. Palma, J. J. Dongarra, and Vicente Hernández, editors. Vector and parallel processing — VEC-PAR ’98: Third International Conference, Porto, Portugal, June 21–23, 1998: selected papers and invited talks, volume 1573 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1999. CODEN LNCSD9. ISBN 3-540-66228-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.1573. URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/tocs/t1573.htm.

Webster:1999:WEE

[1037] John G. Webster, editor. Wiley encyclopedia of electrical and electronics engineering. Wiley, New York, NY, USA, 1999. ISBN 0-471-13946-7. vi + 758 pp. LCCN TK9 .W55 1999. 24 volumes.

ACM:2000:SHP

[1038] ACM, editor. SC2000: High Performance Networking and Computing. Dallas Convention Center, Dallas, TX, USA, November 4–10, 2000. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN ???? LCCN QA76.88. URL http://

                                        www.sc2000.org/proceedings/info/
                                        fp.pdf.

Dongarra:2000:RAP

[1039] J. J. Dongarra, Péter Kacsuk, and Norbert Podhorszki, editors. Recent Advances in Parallel Virtual Machine and Message Passing Interface: 7th European PVM/MPI Users’ Group Meeting, Balatonfüred, Hungary, September 10–13, 2000. Proceedings, volume 1908 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2000. CODEN LNCSD9. ISBN 3-540-41010-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E973

2000. URL http://link.springer-
ny.com/link/service/series/0558/
tocs/t1908.htm.

Sadayappan:2000:IWP

[1040] P. Sadayappan, editor. 2000 International Workshops on Parallel Processing: proceedings: 21–24 August, 2000, Toronto, Canada. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2000. ISBN 0-7695-0771-9, 0-7695-0773-5 (microfiche). LCCN QA76.58 .I575 2000. IEEE Computer Society Order Number PR00771.

Tentner:2000:PHP

[1041] Adrian M. Tentner, editor. Proceedings of the High Performance Computing Symposium: HPC 2000: 2000 Advanced Simulation Technologies Conference: Washington D.C., April 16– 20, 2000, Wyndham City Center Hotel. Society for Computer Simulation, San Diego, CA, USA, 2000. ISBN 1-56555-197-4. LCCN QA76.88 .H53 2000.

ACM:2001:PAJ

[1042] ACM, editor. Proceedings of the ACM 2001 Java Grande/ISCOPE Conference: Palo Alto, Calif., June 2–4, 2001. ACM Press, New York, NY 10036, USA, 2001. ISBN 1-58113-359-

6. LCCN QA76.9.O35 A26 2001.

ACM:2001:SHP

[1043] ACM, editor. SC2001: High Performance Networking and Computing. Denver, CO, November 10–16, 2001. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 1-58113-293-X. LCCN QA76.88 .S85 2001.

Alexandrov:2001:CSIa

[1044] Vassil Alexandrov, Jack J. Dongarra, Benjoe A. Juliano, René S. Renner, and C. J. Kenneth Tan, editors. Computational science — ICCS 2001: International Conference, San Francisco, CA, USA, May 28–30, 2001: proceedings, Part I, volume 2073 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNCSD9. ISBN 3-540-42232-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2001 v.1-2 (2001); QA267.A1 L43 no.2073-2074 Library has v.1-

2.  URL http://link.springer-
ny.com/link/service/series/0558/
tocs/t2073.htm;   http://link.springer-
ny.com/link/service/series/0558/
tocs/t2074.htm;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=2073.

Alexandrov:2001:CSIb

[1045] Vassil Alexandrov, Jack J. Dongarra, Benjoe A. Juliano, René S. Renner, and C. J. Kenneth Tan, editors. Computational science — ICCS 2001: International Conference, San Francisco, CA, USA, May 28–30, 2001: proceedings, Part II, volume 2074 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNCSD9. ISBN 3-540-42233-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2001 v.1-2 (2001); QA267.A1 L43 no.2073-2074 Library has v.1-

2.  URL http://link.springer-
ny.com/link/service/series/0558/
tocs/t2073.htm;   http://link.springer-
ny.com/link/service/series/0558/
tocs/t2074.htm;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=
0302-9743&volume=2074.

Boisvert:2001:ASS

[1046] Ronald F. Boisvert and Ping Tak Peter Tang, editors. The architecture of scientific software: IFIP TC2/WG2.5 Working Conference on the Architecture of Scientific Software, October 2– 4, 2000, Ottawa, Canada, volume 60 of IFIP. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 0-7923-7339-1. LCCN QA76.758 .I345 2000.

Cotronis:2001:RAP

[1047] Yiannis Cotronis and J. J. Dongarra, editors. Recent advances in parallel virtual machine and message passing interface: 8th European PVM/ MPI Users’ Group Meeting, Santorini/ Thera, Greece, September 23–26, 2001: Proceedings, volume 2131 of Lecture Notes in Computer Science and Lecture Notes in Artificial Intelligence. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNCSD9. ISBN 3-540-42609-4 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 E975 2001; QA267.A1 L43 no.2131. URL http://link.

                                        springer-ny.com/link/service/series/
                                        0558/tocs/t2131.htm;   http://www.
                                        springerlink.com/openurl.asp?genre=
                                        issue&issn=0302-9743&volume=2131.

IEEE:2001:IIS

[1048] IEEE, editor. IEEE International Symposium on Network Computing and Applications: NCA 2001: proceedings: 8–10 October, 2001, Cambridge, Massachusetts, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 0-7695-1432-4; 0-7695-1433-2 (case); 0-7695-1434-0 (microfiche). LCCN TK5105.5 .I323 2001.

Katz:2001:IIC

[1049] Daniel S. Katz, editor. 2001 IEEE International Conference on Cluster Computing: 8–11 October 2001, Newport Beach, California, USA: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 0-7695-1116-3; 0-7695-1117-1 (bookbroker); 0-7695-1118-X (microfiche). ISSN 0272-5428. LCCN QA76.58 .I38 2001.

Lee:2001:TAI

[1050] C. A. (Craig A.) Lee, editor. Third annual International Workshop on Active Middleware Services proceedings: 6 August 2001, San Francisco, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 0-7695-1528-2. LCCN QA76.76.M54 I58 2001. IEEE Computer Society Order Number PR01528.

Palma:2001:VPP

[1051] José M. L. M. Palma, Jack J. Dongarra, and Vicente Hernández, editors. Vector and parallel processing — VEC-PAR 2000: 4th International Conference, Porto, Portugal, June 21–23, 2000: selected papers and invited talks, volume 1981 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2001. CODEN LNCSD9. ISBN 3-540-41999-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I552

2000. URL http://link.springer-
ny.com/link/service/series/0558/
tocs/t1981.htm.

Sha:2001:PDC

[1052] E. Sha, editor. Parallel and distributed computing systems: proceedings of the ISCA 14th International Conference: Richardson, Texas, USA, August 8–10, 2001. International Society for Computers and Their Applications, Cary, NC, USA, 2001. ISBN 1-880843-39-0. LCCN QA76.58 .I5443 2001.

Tentner:2001:PHP

[1053] Adrian M. Tentner, editor. Proceedings of the High Performance Computing Symposium — HPC 2001: 2001 Advanced Simulation Technologies Conference, Seattle, Washington, April 22– 26, 2001, Renaissance Madison Hotel.

Simulation Councils, Inc., San Diego, CA, USA, 2001. ISBN 1-56555-237-7. LCCN ????

Abello:2002:HMD

[1054] James Abello, Panos M. Pardalos, and Mauricio G. C. Resende. Handbook of Massive Data Sets. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. ISBN 1-4020-0489-3. xii + 1223 pp. LCCN QA76.9.D3 H3474 2002.

Gropp:2002:PII

[1055] William Gropp, editor. Proceedings: 2002 IEEE International Conference on Cluster Computing, 23–26 September 2002, Chicago, Illinois. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1745-5. LCCN QA76.58 .I38 2002.

IEEE:2002:CIA

[1056] IEEE, editor. CCGrid2002: 2nd IEEE/ACM International Symposium on Cluster Computing and the Grid: proceedings: May 21–24, 2002, Berlin, Germany. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1582-7, 0-7695-1583-5 (bookbroker), 0-7695-1584-3 (microfiche). LCCN QA76.9.C58 I33 2002a. IEEE Computer Society order number PR01582.

IEEE:2002:HPI

[1057] IEEE, editor. HPDC-11 2002: proceedings: 11th IEEE International Symposium on High Performance Distributed Computing: 24–26 July, 2002, Edinburgh, Scotland. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1686-6; 0-7695-1688-2 (microfiche). LCCN QA76.9.D5I593 2002c.

IEEE:2002:PFA

[1058] IEEE, editor. Proceedings / Fourth Annual International Workshop on Active Middleware Services, AMS 2002: 23 July 2002, Edinburgh, United Kingdom. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1721-8, 0-7695-1723-4 (microfiche). LCCN QA76.76.M54 I58 2002.

IEEE:2002:PIP

[1059] IEEE, editor. Proceedings: 16th International Parallel & Distributed Processing Symposium: IPDPS 2002: 15–19 April, 2002, Ft. Lauderdale, Florida, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISBN 0-7695-1573-8, 0-7695-1574-6, 0-7695-1575-4. LCCN QA76.58.I583 2002.

Kranzlmuller:2002:RAP

[1060] Dieter Kranzlmüller, Peter Kacsuk, Jack Dongarra, and Jens Volkert, editors. Recent Advances in Parallel Virtual Machine and Message Passing Interface: 9th European PVM/ MPI Users’ Group Meeting, Linz, Austria, September 29–October 2, 2002. Proceedings, volume 2474 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNCSD9. ISBN 3-540-44296-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E975

2002. URL http://link.springer-
ny.com/link/service/series/0558/

tocs/t2474.htm. Also available via the World Wide Web.

Monien:2002:EPP

[1061] Burkhard Monien and Rainer Feldmann, editors. Euro-Par 2002, parallel processing: 8th International Euro-Par Conference, Paderborn, Germany, August 27–30, 2002: proceedings, volume 2400 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-44049-6 (softcover). LCCN QA76.58 .I553

2002. URL http://link.springer-
ny.com/link/service/series/0558/
tocs/t2400.htm.

Oldehoeft:2002:SIS

[1062] Rod Oldehoeft, editor. Special issue on software for high-performance systems: papers from the symposium of the Los Alamos Computer Science Institute, held in Santa Fe, NM, USA on October 15–18, 2001, volume 23(1) of The journal of supercomputing. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2002. CODEN JOSUED. ISSN 0920-8542 (print), 1573-0484 (electronic).

Parashar:2002:GCG

[1063] Manish Parashar, editor. Grid computing — GRID 2002: third international workshop, Baltimore, MD, USA, November 18, 2002: proceedings, volume 2536 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. ISBN 3-540-00133-6 (softcover). LCCN QA76.9.C58 G74

2002. URL http://link.springer-
ny.com/link/service/series/0558/

tocs/t2536.htm. Also available via the World Wide Web.

Sloot:2002:CSIa

[1064] Peter M. A. Sloot, C. J. Kenneth Tan, Jack J. Dongarra, and Alfons G. Hoekstra, editors. Computational Science–ICCS 2002: International Conference, Amsterdam, The Netherlands, April 21–24, 2002. Proceedings, Part I, volume 2329 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNCSD9. ISBN 3-540-43591-3 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2002 pt.1-3

(2002). URL http://link.springer-
ny.com/link/service/series/0558/
tocs/t2329.htm.

Sloot:2002:CSIb

[1065] P. M. A. Sloot, C. J. Kenneth Tan, J. J. Dongarra, and A. G. Hoekstra, editors. Computational Science–ICCS 2002: International Conference, Amsterdam, The Netherlands, April 21– 24, 2002. Proceedings, Part II, volume 2330 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNCSD9. ISBN 3-540-43593-X (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13

2002. URL http://link.springer-
ny.com/link/service/series/0558/
tocs/t2330.htm.

Sloot:2002:CSIc

[1066] P. M. A. Sloot, C. J. Kenneth Tan, J. J. Dongarra, and A. G. Hoekstra, editors. Computational Science–ICCS 2002: International Conference, Amsterdam, The Netherlands, April 21– 24, 2002. Proceedings, Part III, volume 2331 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNCSD9. ISBN 3-540-43594-8 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13

                                        2002. URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        tocs/t2331.htm.

Wyrzykowski:2002:PPA

[1067] Roman Wyrzykowski, Jack Dongarra, Marcin Paprzycki, and Jerzy Wasńiewski, editors. Parallel Processing and Applied Mathematics: 4th International Conference, PPAM 2001 Nalȩczów, Poland, September 9–12, 2001. Revised Papers, volume 2328 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2002. CODEN LNCSD9. ISBN 3-540-43792-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P69

                                        2001. URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        tocs/t2328.htm.

ACM:2003:CPI

[1068] ACM, editor. Conference proceedings of the 2003 International Conference on Supercomputing: June 23–26, 2003, San Francisco, California, USA. ACM Press, New York, NY 10036, USA, 2003. ISBN 1-58113-733-8. LCCN QA76.5 .I547 2003. ACM order number 415031.

Clematis:2003:EEC

[1069] Andrea Clematis, editor. Eleventh Euromicro Conference on Parallel, Distributed, and Network-Based Processing: proceedings: Genova, Italy, February 5–7, 2003. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 0-7695-1875-3. ISSN 1066-6192. LCCN QA76.58. IEEE Computer Society Order Number PR01875.

Dongarra:2003:RAP

[1070] Jack Dongarra, Domenico Laforenza, and Salvatore Orlando, editors. Recent advances in parallel virtual machine and message passing interface: 10th European PVM/MPI User’s group Meeting, Venice, Italy, September 29– October 2, 2003: Proceedings, volume 2840 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CO-DEN LNCSD9. ISBN 3-540-20149-1. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E973

2003. URL http://link.springer-
ny.com/link/service/series/0558/
tocs/t2840.htm.

Dongarra:2003:SPC

[1071] Jack Dongarra, Ian Foster, Geoffrey Fox, William Gropp, Ken Kennedy, Linda Torczon, and Andy White, editors. The Sourcebook of Parallel Computing. Morgan Kaufmann Publishers, San Francisco, CA, USA, 2003. ISBN 1-55860-871-0. xvi + 842 + 8 pp. LCCN QA76.58 S638 2003. US$59.95.

Gerndt:2003:PEI

[1072] Michael Gerndt, editor. Proceedings / Eighth International Workshop on High-Level Parallel Programming Models and Supportive Environments, 22 April 2003, Nice, France: held in conjunction with 17th International Parallel and Distributed Processing Symposium (IPDPS). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 0-7695-1880-X. LCCN QA76.642 .I586 2003.

IEEE:2003:CIA

[1073] IEEE, editor. CCGrid2003: 3rd IEEE/ACM International Symposium on Cluster Computing and the Grid: proceedings: Tokyo, Japan, 12–15 May, 2003. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 0-7695-1919-9. LCCN QA76.9.C58.

IEEE:2003:IPD

[1074] IEEE, editor. International Parallel and Distributed Processing Symposium: proceedings: April 22–26, 2003, Nice, France. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 0-7695-1926-1. ISSN 1530-2075. LCCN QA76.58.

IEEE:2003:PCI

[1075] IEEE, editor. Proceedings / CCGrid 2003, 3rd IEEE/ACM International Symposium on Cluster Computing and the Grid, Tokyo, Japan, 12–15 May 2003. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 0-7695-1919-9. LCCN ????

IEEE:2003:PIP

[1076] IEEE, editor. Proceedings: 17th International Parallel & Distributed Processing Symposium: IPDPS 2003: 22– 26 April, 2003, Nice, France. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 0-7695-1926-1. LCCN QA76.58 .M47 2003.

Kosch:2003:EPP

[1077] Harald Kosch, László Böszörményi, and Hermann Hellwagner, editors. Euro-Par 2003 parallel processing: 9th International Euro-Par Conference, Klagenfurt, Austria, August 26–29, 2003: proceedings, volume 2790 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. ISBN 3-540-40788-X. LCCN QA76.58 .I5443 2003.

Nabrzyski:2003:GRM

[1078] Jarek Nabrzyski, Jennifer M. Schopf, and Jan Weglarz, editors. Grid resource management: state of the art and future trends. Number ISOR 64 in International series in operations research & management science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2003. ISBN 1-4020-7575-8. xxi + 574 pp. LCCN QA76.9.C58 G78 2004.

Palma:2003:HPC

[1079] José M. L. M. Palma, Jack Dongarra, Vicente Hernández, and A. Augusto Sousa, editors. High performance computing for computational science, VECPAR 2002: 5th International Conference, Porto, Portugal, June 26–28, 2002: selected papers and invited talks, volume 2565 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNCSD9. ISBN 3-540-00852-7 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I552

2002. URL http://link.springer-
ny.com/link/service/series/0558/
tocs/t2565.htm.

Sloot:2003:CSIa

[1080] Peter M. A. Sloot, David Abramson, Alexander V. Bogdanov, Jack J. Dongarra, Albert Y. Zomaya, and Yuriy E. Gorbachev, editors. Computational science — ICCS 2003: International Conference, Melbourne, Australia and St. Petersburg, Russia, June 2–4, 2003: Proceedings, Part I, volume 2657 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNCSD9. ISBN 3-540-40194-6 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2003 pt.1-4

2003. URL http://link.springer-
ny.com/link/service/series/0558/
tocs/t2657.htm.

Sloot:2003:CSIb

[1081] Peter M. A. Sloot, David Abramson, Alexander V. Bogdanov, Jack J. Dongarra, Albert Y. Zomaya, and Yuriy E. Gorbachev, editors. Computational science — ICCS 2003: International Conference, Melbourne, Australia and St. Petersburg, Russia, June 2–4, 2003: Proceedings, Part II, volume 2658 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNCSD9. ISBN 3-540-40195-4 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2003 pt.1-4

2003. URL http://link.springer-
ny.com/link/service/series/0558/
tocs/t2658.htm.

Sloot:2003:CSIc

[1082] Peter M. A. Sloot, David Abramson, Alexander V. Bogdanov, Jack J. Dongarra, Albert Y. Zomaya, and Yuriy E. Gorbachev, editors. Computational science — ICCS 2003: International Conference, Melbourne, Australia and St. Petersburg, Russia, June 2–4, 2003: Proceedings, Part III, volume 2659 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNCSD9. ISBN 3-540-40196-2 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2003 pt.1-4

2003. URL http://link.springer-
ny.com/link/service/series/0558/
tocs/t2659.htm.

Sloot:2003:CSId

[1083] Peter M. A. Sloot, David Abramson, Alexander V. Bogdanov, Jack J. Dongarra, Albert Y. Zomaya, and Yuriy E. Gorbachev, editors. Computational science — ICCS 2003: International Conference, Melbourne, Australia and St. Petersburg, Russia, June 2–4, 2003: Proceedings, Part IV, volume 2660 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2003. CODEN LNCSD9. ISBN 3-540-40197-0 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13 2003 pt.1-4

                                        2003. URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        tocs/t2660.htm.

Bozdogan:2004:EMP

[1084] Hamparsum Bozdogan, editor. Statistical Data Mining and Knowledge Discovery. CRC Press, 2000 N.W. Corporate Blvd., Boca Raton, FL 33431-9868, USA, 2004. ISBN 1-58488-344-8. 588 pp. LCCN QA76.9.D343 S685 2004.

Bubak:2004:CSIa

[1085] Marian Bubak, Geert Dick van Albada, Peter M. A. Sloot, and Jack J. Dongarra, editors. Computational Science–ICCS 2004: 4th International Conference Kraków, Poland, June 6– 9, 2004 Proceedings, Part I, volume 3036 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CO-DEN LNCSD9. ISBN 3-540-22114-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13

                                        2004. URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        tocs/t3036.htm.

Bubak:2004:CSIb

[1086] Marian Bubak, Geert Dick van Albada, Peter M. A. Sloot, and Jack J. Dongarra, editors. Computational Science–ICCS 2004: 4th International Conference Kraków, Poland, June 6– 9, 2004 Proceedings, Part II, volume 3037 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CO-DEN LNCSD9. ISBN 3-540-22115-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13

2004. URL http://link.springer-
ny.com/link/service/series/0558/
tocs/t3037.htm.

Bubak:2004:CSIc

[1087] Marian Bubak, Geert Dick van Albada, Peter M. A. Sloot, and Jack J. Dongarra, editors. Computational Science–ICCS 2004: 4th International Conference Kraków, Poland, June 6– 9, 2004 Proceedings, Part III, volume 3038 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CO-DEN LNCSD9. ISBN 3-540-22116-6. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13

2004. URL http://link.springer-
ny.com/link/service/series/0558/
tocs/t3038.htm.

Bubak:2004:CSId

[1088] Marian Bubak, Geert Dick van Albada, Peter M. A. Sloot, and Jack J. Dongarra, editors. Computational Science–ICCS 2004: 4th International Conference Kraków, Poland, June 6– 9, 2004 Proceedings, Part IV, volume 3039 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CO-DEN LNCSD9. ISBN 3-540-22129-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5 .I13

2004. URL http://link.springer-
ny.com/link/service/series/0558/
tocs/t3039.htm.

Eigenmann:2004:IIC

[1089] Rudolf Eigenmann, editor. ICPP 2004: 2004 International Conference on Parallel Processing: proceedings: 15–18 August, 2004, Montreal, Quebec, Canada. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7695-2197-5. LCCN QA76.6 .I548 2004.

IEEE:2004:CII

[1090] IEEE, editor. CCGrid 2004: 2004 IEEE International Symposium on Cluster Computing and the Grid: April 19–22, 2004, Chicago, IL. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7803-8430-X. LCCN QA76.9.C58 I42 2004. IEEE catalog number 04EX836.

IEEE:2004:IPD

[1091] IEEE, editor. 18th International Parallel and Distributed Processing Symposium: Santa Fe, New Mexico, April 26– 30, 2004: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7695-2132-0. LCCN QA76.58.

IEEE:2004:SIC

[1092] IEEE, editor. Seventh International Conference on High Performance Computing and Grid in Asia Pacific Region proceedings: Omiya Sonic City, Tokyo, Japan, 20–22 July, 2004. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7695-2138-X. LCCN QA76.88.

Kranzlmuller:2004:RAP

[1093] Dieter Kranzlmüller, Péter Kacsuk, and Jack J. Dongarra, editors. Recent advances in parallel virtual machine and message passing interface: 11th European PVM/MPI Users’ Group Meeting, Budapest, Hungary, September 19–22, 2004: proceedings, volume 3241 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CODEN LNCSD9. ISBN 3-540-23163-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E973 2004. URL http:

//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&
volume=3241;   http://www.springerlink.
com/openurl.asp?genre=volume&id=
doi:10.1007/b100820.

Sprague:2004:PAH

[1094] Ralph H. Sprague, editor. Proceedings of the 37th Annual Hawaii International Conference on System Sciences abstracts and CD-ROM of full papers: 5–8 January, 2004, Big Island, Hawaii. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2004. ISBN 0-7695-2056-1. LCCN Q350. HICSS-37, Hilton Waikoloa Village, Big Island, Hawaii.

Wyrzykowski:2004:PPA

[1095] Roman Wyrzykowski, Jack Dongarra, Marcin Paprzycki, and Jerzy Waśniewski, editors. Parallel Processing and Applied Mathematics: 5th International Conference, PPAM 2003, Czȩstochowa, Poland, September 7–10, 2003: Revised Papers, volume 3019 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CO-DEN LNCSD9. ISBN 3-540-21946-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .P69

                                        2003. URL http://link.springer-
                                        ny.com/link/service/series/0558/
                                        tocs/t3019.htm;   http://www.springerlink.
                                        com/openurl.asp?genre=issue&issn=
                                        0302-9743&volume=3019;   http://
                                        www.springerlink.com/openurl.asp?
                                        genre=volume&id=doi:10.1007/b97218.

Dayde:2005:HPC

[1096] Michel Daydé, Jack J. Dongarra, Vicente Hernández, and José M. L. M. Palma, editors. High Performance Computing for Computational Science: VECPAR 2004: 6th International Conference Valencia, Spain, June 28– 30, 2004 Revised Selected and Invited Papers, volume 3402 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNCSD9. ISBN 3-540-25424-2. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .V43 2004. URL http:

                                        //www.springerlink.com/openurl.
                                        asp?genre=issue&issn=0302-9743&
                                        volume=3402;   http://www.springerlink.
                                        com/openurl.asp?genre=volume&id=
                                        doi:10.1007/b106965.

DiMartino:2005:RAP

[1097] Beniamino Di Martino, Dieter Kranzlmüller, and J. J. Dongarra, editors. Recent advances in parallel virtual machine and message passing interface: 12th European PVM/MPI User’s Group Meeting, Sorrento, Italy, September 18–21, 2005: proceedings, volume 3666 of Lecture Notes in Computer Science. Springer-Ver-lag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNCSD9. ISBN 3-540-29009-5 (paperback). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .E973

2005. URL http://springerlink.
metapress.com/openurl.asp?genre=
issue&issn=0302-9743&volume=3666.

IEEE:2005:IPD

[1098] IEEE, editor. 19th International Parallel and Distributed Processing Symposium: proceedings: April 4–8, 2005, Denver, Colorado. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2005. ISBN 0-7695-2312-9. LCCN QA76.58 .I583 2005. IEEE Computer Society Order Number P2312.

Pan:2005:PDP

[1099] Yi Pan, Daoxu Chen, Minyi Guo, Jiannong Cao, and Jack J. Dongarra, editors. Parallel and distributed processing and applications: third international symposium, ISPA 2005, Nanjing, China, November 2–5, 2005: proceedings, volume 3758 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNCSD9. ISBN 3-540-29769-3. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58 .I88 2005. URL http:

//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&
volume=3758.

Sunderam:2005:CSIa

[1100] Vaidy S. Sunderam, Geert Dick van Albada, Peter M. A. Sloot, and Jack J. Dongarra, editors. Computational science: ICCS 2005: 5th international conference, Atlanta, GA, USA, May 22–25, 2005, proceedings, Part I, volume 3514 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNCSD9. ISBN 3-540-26032-3 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA75.5

.I13 2005. URL http://www.loc.
gov/catdir/enhancements/fy0663/
2005925759-d.html;   http://www.
springerlink.com/openurl.asp?genre=
issue&issn=0302-9743&volume=3514;
 http://www.springerlink.com/openurl.
asp?genre=volume&id=doi:10.1007/
b136570.

Sunderam:2005:CSIb

[1101] Vaidy S. Sunderam, Geert Dick van Albada, Peter M. A. Sloot, and Jack J. Dongarra, editors. Computational science: ICCS 2005: 5th international conference, Atlanta, GA, USA, May 22–25, 2005, proceedings, Part II, volume 3515 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNCSD9. ISBN 3-540-26043-9. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&
volume=3515;   http://www.springerlink.
com/openurl.asp?genre=volume&id=
doi:10.1007/b136571.

Sunderam:2005:CSIc

[1102] Vaidy S. Sunderam, Geert Dick van Albada, Peter M. A. Sloot, and Jack J. Dongarra, editors. Computational science: ICCS 2005: 5th international conference, Atlanta, GA, USA, May 22–25, 2005, proceedings, Part III, volume 3516 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNCSD9. ISBN 3-540-26044-7. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&
volume=3516;   http://www.springerlink.
com/openurl.asp?genre=volume&id=
doi:10.1007/b136575.

Yang:2005:HPC

[1103] Laurence Tianruo Yang, Omer F.

Rana, Beniamino Di Martino, and Jack Dongarra, editors. High performance computing and communications: first international conference, HPCC 2005, Sorrento, Italy, September 21–23, 2005: proceedings, volume 3726 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2005. CODEN LNCSD9. ISBN 3-540-29031-1 (softcover). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.88 .H655 2005. URL http:

//www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&
volume=3726.

ACM:2006:SCH

[1104] ACM, editor. SC’06: Conference on High Performance Networking and Computing: proceedings of the 2006 ACM/IEEE conference on Supercomputing, November 11–17, 2006, Tampa Convention Center, Tampa, Florida, USA. ACM Press, New York, NY 10036, USA, 2006. ISBN 0-7695-2700-

0. LCCN QA76.5 .P742 2006. Contains one CD-ROM.

Alexandrov:2006:CSIb

[1105] Vassil N. Alexandrov, Geert Dick van Albada, Peter M. A. Sloot, and Jack Dongarra, editors. Computational Science – ICCS 2006: 6th International Conference, Reading, UK, May 28– 31, 2006. Proceedings, Part II, volume 3992 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNCSD9. ISBN 3-540-34381-4 (print), 3-540-34382-2 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        LCCN ????  URL http://www.
                                        springerlink.com/content/978-3-
                                        540-34382-0.

Alexandrov:2006:CSIc

[1106] Vassil N. Alexandrov, Geert Dick van Albada, Peter M. A. Sloot, and Jack Dongarra, editors. Computational Science – ICCS 2006: 6th International Conference, Reading, UK, May 28– 31, 2006. Proceedings, Part III, volume 3993 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNCSD9. ISBN 3-540-34383-0 (print), 3-540-34384-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        LCCN ????  URL http://www.
                                        springerlink.com/content/978-3-
                                        540-34384-4.

Alexandrov:2006:CSId

[1107] Vassil N. Alexandrov, Geert Dick van Albada, Peter M. A. Sloot, and Jack Dongarra, editors. Computational Science – ICCS 2006: 6th International Conference, Reading, UK, May 28– 31, 2006, Proceedings, Part IV, volume 3994 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNCSD9. ISBN 3-540-34385-7 (print), 3-540-34386-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ????  URL http://www.
springerlink.com/content/978-3-
540-34386-8.

Dongarra:2006:APC

[1108] Jack Dongarra, Kaj Madsen, and Jerzy Wasniewski, editors. Applied Parallel Computing: State of the Art in Scientific Computing. 7th International Workshop, PARA 2004, Lyngby, Denmark, June 20–23, 2004. Revised Selected Papers, volume 3732 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNCSD9. ISBN 3-540-29067-2 (softcover), 3-540-33498-X. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76.58

.P353 2004. URL http://www.loc.
gov/catdir/enhancements/fy0663/
2006920921-d.html;   http://www.
springerlink.com/openurl.asp?genre=
issue&issn=0302-9743&volume=3732.

Guo:2006:PDP

[1109] Minyi Guo, Laurence T. Yang, Beniamino Di Martino, Hans P. Zima, Jack Dongarra, and Feilong Tang, editors. Parallel and Distributed Processing and Applications: 4th International Symposium, ISPA 2006, Sorrento, Italy, December 4–6, 2006. Proceedings, volume 4330 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CO-DEN LNCSD9. ISBN 3-540-68067-5 (print), 3-540-68070-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.com/content/
                                        978-3-540-68070-3.

Mohr:2006:RAP

[1110] Bernd Mohr, Jesper Larsson Träff, Joachim Worringen, and Jack Dongarra, editors. Recent Advances in Parallel Virtual Machine and Message Passing Interface: 13th European PVM/MPI User’s Group Meeting Bonn, Germany, September 17– 20, 2006 Proceedings, volume 4192 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. CODEN LNCSD9. ISBN 3-540-39110-X (print), 3-540-39112-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        LCCN ????  URL http://www.
                                        springerlink.com/content/978-3-
                                        540-39112-8.

Shahzadeh-Fazeli:2006:ECN

[1111] I. Lirkov, S. Margenov, and J. Wasniewski, editors. Eigenvalue computation with NetSolve global computing system, volume 3743. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-31994-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ????

Wyrzykowski:2006:PPA

[1112] Roman Wyrzykowski, Jack Dongarra, and Norbert Meyer, editors. Parallel Processing and Applied Mathematics: 6th International Conference, PPAM 2005, Poznan, Poland, September 11– 14, 2005, Revised Selected Papers, volume 3911 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2006. ISBN 3-540-34141-2 (softcover). LCCN QA76.58

.P69 2005. URL http://www.loc.
gov/catdir/enhancements/fy0661/
2006925464-d.html;   http://www.
springerlink.com/openurl.asp?genre=
issue&issn=0302-9743&volume=3911.

Cappello:2007:RAP

[1113] Franck Cappello, Thomas Herault, and Jack Dongarra, editors. Recent Advances in Parallel Virtual Machine and Message Passing Interface: 14th European PVM/MPI User’s Group Meeting, Paris, France, September 30 — October 3, 2007. Proceedings, volume 4757 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNCSD9. ISBN 3-540-75415-6 (print), 3-540-75416-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ????  URL http://www.
springerlink.com/content/978-3-
540-75416-9.

Oram:2007:BC

[1114] Andrew Oram and Greg Wilson, editors. Beautiful Code. Theory in practice. O’Reilly & Associates, Inc., 103a Morris Street, Sebastopol, CA 95472, USA, Tel: +1 707 829 0515, and 90 Sherman Street, Cambridge, MA 02140, USA, Tel: +1 617 354 5800, 2007. ISBN 0-596-51004-7 (paperback). xxi + 593 pp. LCCN QA76.758 .B428 2007; QA76.758 .B43 2007; QA76.758 .B48 2007. URL http:/

/proquest.safaribooksonline.com/

                                        9780596510046;   http://www.oreilly.
                                        com/catalog/9780596510046.

Hogben:2007:HLA

[1115] Leslie Hogben, editor. Handbook of Linear Algebra. Discrete Mathematics and its Applications (Boca Raton). Chapman and Hall/CRC, Boca Raton, FL, USA, 2007. ISBN 1-58488-510-6 (hardcover), 1-4200-1057-3 (e-book). xxx + 1370 pp. LCCN QA184.2 .H36

                                        2007. URL http://www.crcnetbase.
                                        com/isbn/9781420010572;   http:/
                                        /www.crcnetbase.com/isbn/9781584885108;
                                         http://www.loc.gov/catdir/enhancements/
                                        fy0647/2006045491-d.html. Asso-

ciate editors: Richard Brualdi, Anne

Greenbaum and Roy Mathias.

Kaagstrom:2007:APC

[1116] Bo Kågström, Erik Elmroth, Jack Dongarra, and Jerzy Waśniewski, editors. Applied Parallel Computing. State of the Art in Scientific Computing: 8th International Workshop, PARA 2006, Umeå, Sweden, June 18– 21, 2006, Revised Selected Papers, volume 4699 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNCSD9. ISBN 3-540-75754-6 (print), 3-540-75755-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        LCCN ????  URL http://www.
                                        springerlink.com/content/978-3-
                                        540-75755-9.

Shi:2007:CSIa

[1117] Yong Shi, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational Science – ICCS 2007: 7th International Conference, Beijing, China, May 27 — 30, 2007, Proceedings, Part I, volume 4487 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNCSD9. ISBN 3-540-72583-0 (print), 3-540-72584-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ????  URL http://www.
springerlink.com/content/978-3-
540-72584-8.

Shi:2007:CSIb

[1118] Yong Shi, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational Science – ICCS 2007: 7th International Conference, Beijing, China, May 27 — 30, 2007, Proceedings, Part II, volume 4488 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNCSD9. ISBN 3-540-72585-7 (print), 3-540-72586-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ????  URL http://www.
springerlink.com/content/978-3-
540-72586-2.

Shi:2007:CSIc

[1119] Yong Shi, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational Science – ICCS 2007: 7th International Conference, Beijing, China, May 27 — 30, 2007, Proceedings, Part III, volume 4489 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNCSD9. ISBN 3-540-72587-3 (print), 3-540-72588-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        LCCN ????  URL http://www.
                                        springerlink.com/content/978-3-
                                        540-72588-6.

Shi:2007:CSId

[1120] Yong Shi, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational Science – ICCS 2007: 7th International Conference, Beijing, China, May 27 — 30, 2007, Proceedings, Part IV, volume 4490 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2007. CODEN LNCSD9. ISBN 3-540-72589-X (print), 3-540-72590-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        LCCN ????  URL http://www.
                                        springerlink.com/content/978-3-
                                        540-72590-9.

Bubak:2008:CSIa

[1121] Marian Bubak, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational Science – ICCS 2008: 8th International Conference, Kraków, Poland, June 23– 25, 2008, Proceedings, Part I, volume 5101 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNCSD9. ISBN 3-540-69383-1 (print), 3-540-69384-X (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

                                        LCCN ????  URL http://www.
                                        springerlink.com/content/978-3-
                                        540-69384-0.

Bubak:2008:CSIb

[1122] Marian Bubak, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational Science – ICCS 2008: 8th International Conference, Kraków, Poland, June 23– 25, 2008, Proceedings, Part II, volume 5102 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNCSD9. ISBN 3-540-69386-6 (print), 3-540-69387-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ????  URL http://www.
springerlink.com/content/978-3-
540-69387-1.

Bubak:2008:CSIc

[1123] Marian Bubak, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational science – ICCS 2008. 8th international conference, Kraków, Poland, June 23– 25, 2008. Proceedings, Part III, volume 5103 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008.

Chatterjee:2008:PPA

[1124] Siddhartha Chatterjee, editor. PPoPP ’08: proceedings of the 2008 ACM SIG-PLAN Symposium on Principles and Practice of Parallel Programming: Salt Lake City, Utah, USA, February 20– 23, 2008. ACM Press, New York, NY 10036, USA, 2008. ISBN 1-59593-795-1. LCCN QA76.642 .A27 2008.

Dongarra:2008:DHP

[1125] M. V. Zelkowitz, editor. DARPX’s HPCS program: History, models, tools, languages, volume 72 of Advances in Computers. Academic Press, New York, NY, USA, 2008. ISBN 0-12-374411-3. ISSN 0065-2458. LCCN ????

Lastovetsky:2008:RAP

[1126] Alexey Lastovetsky, Tahar Kechadi, and Jack Dongarra, editors. Recent Advances in Parallel Virtual Machine and Message Passing Interface: 15th European PVM/MPI Users’ Group Meeting, Dublin, Ireland, September 7–10, 2008. Proceedings, volume 5205 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNCSD9. ISBN 3-540-87474-7 (print), 3-540-87475-5 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/
978-3-540-87475-1.

Wyrzykowski:2008:PPA

[1127] Roman Wyrzykowski, Jack Dongarra, Konrad Karczewski, and Jerzy Wasniewski, editors. Parallel Processing and Applied Mathematics: 7th International Conference, PPAM 2007, Gdansk, Poland, September 9–12, 2007 Revised Selected Papers, volume 4967 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2008. CODEN LNCSD9. ISBN 3-540-68105-1 (print), 3-540-68111-6 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ????  URL http://www.
springerlink.com/content/978-3-
540-68111-3.

Allen:2009:CSIa

[1128] Gabrielle Allen, Jarosław Nabrzyski, Edward Seidel, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational science – ICCS 2009. 9th international conference Baton Rouge, LA, USA, May 25–27, 2009. Proceedings, Part I, volume 5544 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009.

Allen:2009:CSIb

[1129] Gabrielle Allen, Jarosław Nabrzyski, Edward Seidel, Geert Dick van Albada, Jack Dongarra, and Peter M. A. Sloot, editors. Computational science – ICCS 2009. 9th international conference Baton Rouge, LA, USA, May 25–27, 2009. Proceedings, Part II, volume 5545 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009.

Ropo:2009:RAP

[1130] Matti Ropo, Jan Westerholm, and Jack Dongarra, editors. Recent Advances in Parallel Virtual Machine and Message Passing Interface: 16th European PVM/MPI Users’ Group Meeting, Espoo, Finland, September 7–10, 2009. Proceedings, volume 5759 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2009. CODEN LNCSD9. ISBN 3-642-03769-0 (print), 3-642-03770-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

//www.springerlink.com/content/
978-3-642-03770-2.

Bultheel:2010:BNA

[1131] Adhemar Bultheel and Ronald Cools, editors. The birth of numerical analysis. World Scientific Publishing Co., Singapore; Philadelphia, PA, USA; River Edge, NJ, USA, 2010. ISBN 981-283-625-X. LCCN QA297 .B54 2010.

Keller:2010:RAM

[1132] Rainer Keller, Edgar Gabriel, Michael Resch, and Jack Dongarra, editors. Recent Advances in the Message Passing Interface: 17th European MPI Users’ Group Meeting, EuroMPI 2010, Stuttgart, Germany, September 12– 15, 2010. Proceedings, volume 6305 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNCSD9. ISBN 3-642-15645-2 (print), 3-642-15646-0 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ????  URL http://www.
springerlink.com/content/978-3-
642-15646-5.

Wyrzykowski:2010:PPAa

[1133] Roman Wyrzykowski, Jack Dongarra, Konrad Karczewski, and Jerzy Wasniewski, editors. Parallel Processing and Applied Mathematics: 8th International Conference, PPAM 2009, Wroclaw, Poland, September 13–16, 2009. Revised Selected Papers, Part I, volume 6067 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNCSD9. ISBN 3-642-14389-X (print), 3-642-14390-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ????  URL http://www.
springerlink.com/content/978-3-
642-14390-8.

Wyrzykowski:2010:PPAb

[1134] Roman Wyrzykowski, Jack Dongarra, Konrad Karczewski, and Jerzy Wasniewski, editors. Parallel Processing and Applied Mathematics: 8th International Conference, PPAM 2009, Wroclaw, Poland, September 13–16, 2009, Revised Selected Papers, Part II, volume 6068 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2010. CODEN LNCSD9. ISBN 3-642-14402-0 (print), 3-642-14403-9 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ????  URL http://www.
springerlink.com/content/978-3-
642-14403-5.

Cotronis:2011:RAM

[1135] Yiannis Cotronis, Anthony Danalis, Dimitrios S. Nikolopoulos, and Jack Dongarra, editors. Recent Advances in the Message Passing Interface: 18th European MPI Users’ Group Meeting, EuroMPI 2011, Santorini, Greece, September 18–21, 2011. Proceedings, volume 6960 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2011. CODEN LNCSD9. ISBN 3-642-24448-3 (print), 3-642-24449-1 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ????  URL http://www.
springerlink.com/content/978-3-
642-24449-0.

Lathrop:2011:SPI

[1136] Scott Lathrop, Jim Costa, and William Kramer, editors. SC’11: Proceedings of 2011 International Conference for High Performance Computing, Networking, Storage and Analysis, Seattle, WA, November 12–18 2011. ACM Press and IEEE Computer Society Press, New York, NY 10036, USA and 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2011. ISBN 1-4503-0771-X. LCCN ????

Hollingsworth:2012:SPI

[1137] Jeffrey Hollingsworth, editor. SC ’12: Proceedings of the International Conference on High Performance Computing, Networking, Storage and Analysis, Salt Lake Convention Center, Salt Lake City, UT, USA, November 10–16, 2012. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2012. ISBN 1-4673-0804-8.

Traff:2012:RAM

[1138] Jesper Larsson Träff, Siegfried Benkner, and Jack J. Dongarra, editors. Recent Advances in the Message Passing Interface: 19th European MPI Users’ Group Meeting, EuroMPI 2012, Vienna, Austria, September 23–26, 2012. Proceedings, volume 7490 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNCSD9. ISBN 3-642-33517-9 (print), 3-642-33518-7 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN ???? URL http:

                                        //www.springerlink.com/content/
                                        978-3-642-33518-1.

Wyrzykowski:2012:PPAa

[1139] Roman Wyrzykowski, Jack Dongarra, Konrad Karczewski, and Jerzy Waśniewski, editors. Parallel Processing and Applied Mathematics: 9th International Conference, PPAM 2011, Torun, Poland, September 11–14, 2011. Revised Selected Papers, Part I, volume 7203 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNCSD9. ISBN 3-642-31463-5 (print), 3-642-31464-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ????  URL http://www.
springerlink.com/content/978-3-
642-31464-3.

Wyrzykowski:2012:PPAb

[1140] Roman Wyrzykowski, Jack Dongarra, Konrad Karczewski, and Jerzy Waśniewski, editors. Parallel Processing and Applied Mathematics: 9th International Conference, PPAM 2011, Torun, Poland, September 11–14, 2011. Revised Selected Papers, Part II, volume 7204 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. CODEN LNCSD9. ISBN 3-642-31499-6 (print), 3-642-31500-3 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic).

LCCN ????  URL http://www.
springerlink.com/content/978-3-
642-31500-8.

Krzhizhanovskaya:2020:CSI

[1141] Valeria V. Krzhizhanovskaya, Gábor Závodszky, Michael H. Lees, Jack J. Dongarra, Peter M. A. Sloot, Sérgio Brissos, and João Teixeira, editors. Computational Science — ICCS 2020 20th International Conference, Amsterdam, The Netherlands, June 3– 5, 2020, Proceedings, Part II, volume 12138 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2020. ISBN 3-030-50416-6, 3-030-50417-4 (e-book). ISSN 0302-9743 (print), 1611-3349 (electronic). URL https://link.

                                        springer.com/book/10.1007/978-3-
                                        030-50417-5.

Wyrzykowski:2020:PPA

[1142] Roman Wyrzykowski, Ewa Deelman, Jack Dongarra, and Konrad Karczewski, editors. Parallel Processing and Applied Mathematics: 13th International Conference, PPAM 2019, Bialystok, Poland, September 8–11, 2019, Revised Selected Papers, Part I, Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2020. ISBN 3-030-43229-7. ISSN 0302-9743 (print), 1611-3349 (electronic).

Anonymous:2022:ATA

[1143] Anonymous. ACM Turing Award honors Jack Dongarra for pioneering concepts and methods which resulted in world-changing computations: Dongarra’s algorithms and software fueled the growth of high-performance computing and had significant impacts in many areas of computational science from AI to computer graphics. ACM Web site, March 29,

                                        2022. URL https://awards.acm.
                                        org/about/2021-turing.

Savage:2022:NAI

[1144] Neil Savage. News: Always improving performance. Communications of the ACM, 65(6):16–18, June 2022. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic). URL https:/

                                        /dl.acm.org/doi/10.1145/3530689.

Wyrzykowski:2023:PPA

[1145] Roman Wyrzykowski, Jack Dongarra, Ewa Deelman, and Konrad Karczewski, editors. Parallel Processing and Applied Mathematics: 14th International Conference, PPAM 2022, Gdansk, Poland, September 11–14, 2022, Revised Selected Papers, Part II, volume 13827 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2023. ISBN 3-031-30441-1, 3-031-30444-6, 3-031-30445-4 (e-book). LCCN Q334-342.

URL https://link.springer.com/
book/10.1007/978-3-031-30445-3.

Anonymous:1995:BRB

[1146] Anonymous. Book reviews: 26. Barrett, Berry, Chan, Demmel, Donato, Dongarra, Eijkhout, Pozo, Romine, and van der Vorst. Mathematics of Computation, 64(211):??, July 1995. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).